A Bibliography of Publications in IEEE Computer, 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

21 December 2023 Version 2.47

Title word cross-reference

$100 [2272]. $20 [1508, 1783]. 3 [62, 194, 406, 408, 461, 496, 649, 661, 718, 927, 986, 1197, 1246, 1270, 1512, 1515, 1516, 1809, 2022, 2052, 2055, 2132, 2345, 2368]. $300 [2568]. 4 [407, 2140]. 6 [2672]. $99 [80].

-D [986, 2368].

.NET [450, 753, 871, 954, 964, 1048].

000 [1244].

1-GHz [985]. 1.5 [1007]. 10 [1007]. 10-Gigabit [658, 1007]. 10x [1106]. 1220 [1729]. 1471 [364]. 15 [43, 579]. 1599 [2507].

1B [1095].

2 [48, 122, 426]. 2.0 [244, 267, 1106, 1652, 2020, 2117, 2131, 2157, 2291, 2302, 2564, 2680]. 20 [520]. 20-GHz [520]. 2000 [80, 84, 85, 143, 172, 191, 267, 485]. 2001 [308, 357, 449, 485]. 2003 [871, 1033]. 2004 [1244]. 2005 [1346]. 2006 [1654, 1677, 1896, 1942]. 2008 [2612]. 2018 [2604]. 20th [1988]. 20th-Century [1988]. 21st [349, 865, 2258]. 24-Hour [2460]. 254 [2445]. 2nd [806].

3 [165]. 3.0 [2203, 2467]. 31 [331]. 32-Bit [778]. 37 [1381]. 3G [164, 547].

4 [244]. 4.0 [694, 897]. 4758 [483]. 4GL [205].

5 [122, 694]. 5.6 [244]. 50th [2433].

6 [224, 1106]. 6.0 [244]. 60th [1836]. 64 [69]. 64-bit [985].

7 [447].

802.11-Based [942]. 802.11n [2580]. 802.1X [2232]. 8th [532].

Abaco [406]. Absolutely [1503]. Abstraction [904, 980, 1472]. Absurdity [56]. Abuse [436]. Academia [161, 1017, 1603]. Academic [171, 622, 2503, 2575]. Academics [357]. Academy [373]. Accelerate [975, 1966]. Accelerates [874]. Accelerating [375, 2260]. Acceleration [65]. Accelerators [96]. Acceptance [112, 412]. Accepted [279]. Access [143, 304, 331, 807, 840, 864, 907, 1121, 1540, 2070, 2603]. Accessibility [832, 860, 1281, 2139, 2475]. Accessible [747, 1134, 1737]. Accessing [1605]. Accomplishments [516]. According [2213, 2445]. Accreditation [2682]. Accurately [1941]. Achievable [283, 341]. achievement [937]. Achievements [75, 308, 423, 669, 961]. Achievers [1565]. Achieving [1201, 1721, 1964]. Acing [204]. Acoustic [1903]. Acquisition [2249]. Acronyms [788]. Across [550, 1758]. Actel [736]. Acting [2252]. Action [2010]. Active [49, 169, 400, 1947, 2012]. Active-Content [169]. ActiveCampus [1298]. Actively [1384]. ActiveState [244, 267, 426]. Activism [389]. Activities [43, 99]. Activity [1411, 2271]. Ad [156, 397, 912, 1120, 1121, 1123–1126]. Adam [2608]. adapt [952]. Adaptable [71]. Adaptation [1295, 2313, 2651, 2652]. Adaptec [1007]. Adapter [605]. Adaptive [534, 994, 1071, 1165, 1168, 1230, 1672, 1864, 1940, 2486]. Add [974]. Added [448, 2411].

Adding [593, 2345]. Additions [432]. Address [8, 444, 465, 733, 1030, 1541, 2104, 2353, 2611]. Addressing [2368]. Adds [49, 244, 649, 829, 915, 1033]. Adequate [721]. Administration [406]. Adobe [672]. Adopt [112]. Adopts [2475]. Ads [1903, 2264]. Advance [8, 423, 691, 1383]. Advanced [422, 426, 626, 876, 2287, 2358]. Advances [7, 104, 118, 145, 174, 212, 822, 916, 2051]. Advancing [1024]. Advantage [94]. Adventure [1107]. Adventures [314]. advertisements [2219]. Advertiser [18, 47, 78, 310, 404]. Advertiser/Product [18, 47, 78, 310, 404]. Advertising [910, 2106]. Aerospace [319]. Aestiva [361]. Aether [426]. Afar [123]. Affect [2617]. Afflicts [130]. Affordable [49, 2626]. after [172]. Again [89, 319, 1142]. against [44, 996]. Age [347, 401, 796, 991, 1039, 1130, 1593, 1688, 1885, 2359, 2408]. Agenda [1630]. Agent [42, 243, 330, 344, 375, 418, 420, 421, 731, 1112, 1155]. Agent-Based [421]. Agents [31, 771, 943, 1133, 1560, 1778]. Agere [672]. Aggregation [2543]. Agile [470, 513, 551, 673, 701, 930, 933, 935, 936, 970, 986, 1080, 1308, 1338, 2629]. Agility [931, 2109, 2597]. Aging [1187]. Ago [1092, 1137, 1182, 1223, 1266, 1311, 1360, 1382, 1403, 1423, 1442, 1467, 1487, 1509, 1531, 1551, 1576, 1598, 1623, 1644, 1667, 1689, 1714, 1736, 1759, 1782, 1807, 1832, 1856, 1881, 1910, 1931, 1957, 1979, 2003, 2026, 2047, 2068, 2092, 2116, 2138, 2162, 2184, 2208, 2231, 2275, 2297, 2321, 2342, 2365, 2385, 2406, 2428, 2453, 2472, 2492, 2516, 2537, 2558, 2579, 2601, 2624, 2645, 2667, 2688]. Ahead [152, 196, 394, 507, 826, 1051, 1468, 2580]. AI [205, 322, 422, 436, 498, 952, 1008, 1088, 2529, 2573]. Aid [651]. Aim [343, 747]. Aims [1738]. Air [1017]. AireWeb [534]. Airline [941]. AIX [487]. Ajax [1554, 1731]. Ajax-Style [1731]. Akyildiz [669]. Alert [1274]. Alfred [937]. Algebra [2683]. Algebraic [2661]. Algorithm [79, 1168]. Algorithms [186, 312, 618, 686, 735, 777, 873, 964, 2194, 2437]. Alignment [1604, 2666]. Alive [147]. Allen [1077]. Alliance [546, 2238]. Along [885]. also [1894]. Altera [807, 985]. Alternative [417, 2282]. Alternatives [743]. Altia [605]. AMC [165]. AMD [985]. Amdahl [2327, 2432]. Amend [2462]. Amendment [99, 555, 691, 893]. Amendments [161]. America [1872]. American [224]. Ames [14]. Among [2687]. Amphibious [1916]. Ampro [627]. AMS [852]. Analog [852]. Analysis [99, 182, 224, 260, 291, 361, 534, 583, 671, 778, 829, 928, 954, 1023, 1054, 1079, 1296, 1369, 2054, 2261]. Analytics [2474]. Analyze [594]. Analyzer [649, 851, 874, 1054]. analyzes [985]. Analyzing [970, 1607, 2654]. Anatomy [563]. Anchors [784]. Ancient [2647]. Andhra [929]. Animal [824]. Animation [296, 885, 1365]. Animator [239]. Annie [2067]. Anniversary [1836, 2433]. Annotating [420]. Announce [455, 1677]. Announces [851, 915, 963, 1007, 1033, 1054, 1654, 1723, 2626]. Annual [240, 263, 531, 804, 1075, 1343, 1613, 1896, 2174, 2439, 2706]. Anonymity [140]. Anonymization [2668]. ANSI [2604]. ANSI/CEA [2604]. ANSI/CEA-2018 [2604]. Answer [1005]. Answers [2525]. Anthrax [520]. Antikythera [2279]. Antivirus [2005]. ANTS [869]. Any [180]. Anymore [1558]. Anytime [53, 54]. Anywhere [53, 54]. AOL [84, 85]. Apart [48]. API [52, 484, 1079]. API-Level [484]. APIs [138, 753]. Apollo [2434]. App [279, 829]. Appealing [1080]. AppForce [1054]. AppForge [583]. Apple [186, 384, 637, 694, 985]. Apples [48]. Appliance [103, 143, 452, 1079]. Appliances [21, 412]. Applicants [331, 579]. Application [49, 187, 223, 251, 291, 313, 335, 361, 384, 406, 426, 435, 447, 467, 472, 478, 558, 583, 597, 605, 694, 736, 963, 1054, 1516, 2322, 2610, 2647]. application-layer [1054]. Application-Monitoring [736]. Application-Specific [2610]. Applications [42, 59, 72, 80, 93, 110, 129, 187, 383, 586, 605, 666, 672, 694, 806, 873, 874, 876, 1053, 1232, 1329, 1439, 1554, 1563, 1582, 1649, 1653, 1935, 2052, 2122, 2343, 2369, 2414, 2564, 2674]. Applied [583, 964, 1686]. Applying [809, 1008, 2302]. Appointments [161, 381]. Approach [82, 389, 532, 562, 588, 635, 682, 735, 747, 774, 777, 873, 891, 896, 907, 920, 941, 957, 975, 1034, 1043, 1044, 1080, 1584, 1699, 1857, 1958, 2075, 2223, 2455, 2538, 2545, 2572, 2581, 2669]. Approaches [51, 864]. Approves [99, 161, 555, 637, 691, 893, 982]. Approximation [1148]. Apps [144, 224, 915]. Apptivity [244]. Aprisa [605]. AQUA [1916]. Arc [267]. Archaic [2194]. Archibald [1622]. Architect [937]. Architecting [1107, 1447, 1474]. Architectural [350, 1003, 2083, 2374, 2574, 2652]. Architecture [5, 20, 95, 97–99, 103, 125, 243, 303, 331, 356, 364, 491, 553, 626, 669, 754, 777, 848, 976, 1107, 1208, 1295, 1449, 1494, 1631, 1865, 1887, 2316, 2351, 2400, 2476, 2583, 2676, 2704]. Architecture-Based [1295]. Architecture-Centric [2400]. Architectures [144, 338, 504, 877, 1142, 1229, 1329, 1367, 1913, 1988, 2149, 2191, 2281, 2347, 2671, 2691]. Archive [205, 234, 1068]. Archiving [467]. Area [1122, 1557, 2063, 2073, 2139, 2459, 2547]. Argument [290]. Aria [420]. Arithmetic [806, 858, 2039]. Arm [243, 915, 1494, 2277, 2620]. Armando [893]. Armenia [2115]. Army [1872, 2448]. Array [1370, 2391]. Arrived [1578]. Arrives [1950]. Arrogance [1060]. Art [24, 234, 331, 498, 648, 867, 1001, 1243, 2144]. Artemis [504]. Arthur [801]. Article [9, 33, 60, 90, 113, 131, 153, 176, 195, 213, 233, 255, 280, 300, 323, 345, 371, 393, 415, 437, 456, 476, 499, 521, 545, 568, 590, 612, 634, 656, 679, 704, 722, 742, 762, 787, 812, 836, 859, 880, 903, 923, 947, 969, 992, 1016, 1040, 1061, 1087, 1113, 1135, 1158, 1179, 1199, 1220, 1241, 1264, 1286, 1309, 1331, 1355, 1379, 1400, 1421, 1441, 1464, 1485, 1507, 1529, 1550, 1573, 1594, 1619, 1641, 1664, 1687, 1711, 1733, 1756, 1779, 1804, 1829, 1854, 1876, 1905, 1929, 1954, 1976]. Article [2000, 2024, 2045, 2066, 2089, 2113, 2135, 2159, 2183, 2205, 2228, 2250, 2273, 2294, 2319, 2340, 2363, 2382, 2404, 2424]. Artifact [963]. Artifacts [1409]. Artificial [20, 416, 909, 1060]. Artist [1745]. Artist-Technologist [1745]. ASC [2238]. Ascend [447]. Asian [279]. Asim [577]. Asking [543]. ASP [143]. Aspect [343]. Aspect-Oriented [343]. Aspects [1367, 1472]. Aspiration [1107]. Assembled [1367]. Assembly [20, 281, 686]. Assessing [312, 625, 1728, 2595]. Assessment [384, 1293, 1985, 2191, 2280]. Asset [80]. Assets [21]. Assimilation [2436]. Assistant [1007, 1590]. Assisted [317, 927]. Assistive [1775, 2336]. ASSL [2572]. Associates [205, 291, 583]. Association [1561]. Associations [1740]. Assumptions [27]. Assurance [218, 526, 616, 800, 1504, 2240, 2347, 2348, 2350, 2352, 2544]. Assuring [2351]. astral [702]. Astrology [544]. Astrophysicist [99]. ATA [1007]. Atinav [649]. ATM [189, 754]. Atomic [851]. Atoms [108, 268]. Attached [103, 649]. Attachments [1995]. Attack [1268, 2005, 2368]. Attacker [1939]. Attacks [88, 484, 535, 793, 2004, 2034, 2562, 2661]. Attaway [331]. Attending [2227]. Attention [1218, 2422]. Attribute [2543]. Auction [1293]. Auctions [856]. Audience [2274]. Audio [746]. Audit [2062]. Auditions [952]. Auditory [1274].

Augmented [596, 2689]. Augmented-Reality [596]. Authenticated [1164]. Authentication [219, 261, 715, 1139, 1578, 1658, 1676, 1995, 2034, 2239]. Authoring [187, 467, 473]. Authority [785]. Authorizing [2217]. AutoCAD [649]. AutoCAD-to-PDF [649]. Automata [2191]. Automate [706, 885]. Automated [500, 619, 736, 940, 2030, 2158, 2444]. Automatic [235]. Automatically [729]. Automating [1651, 2329, 2534]. Automation [356, 482, 2512]. Automobile [928]. Automotive [554, 2122, 2704]. Autonomic [571, 821, 2572, 2650]. Autonomic-Computing [571]. Autonomous [1887, 1888, 1892, 1916]. Autonomy [2435]. Autostereoscopic [1515]. AV [2314]. Available [75, 99, 579, 733, 893, 940, 1106]. Avalanche [1925]. Avoiding [246]. Award [99, 240, 308, 331, 485, 507, 532, 669, 1701, 1942, 2612, 2698]. Awarded [264, 622]. Awards [44, 99, 139, 140, 220, 331, 357, 485, 779, 937, 961, 1565, 1746, 2483, 2591, 2657]. Aware [439, 750, 864, 898, 1000, 1069, 1074, 1130, 1261, 1408, 1582, 1583, 1672, 2170, 2239]. Awareness [397, 1305, 2322, 2485]. Awari [1023]. Away [2568].

B2B [152]. Back

[146, 539, 571, 694, 792, 965, 1142, 1200]. Back-Office [694]. Backbone [383, 520]. Bad [1268, 1537, 1831, 2154, 2666]. Bag [1100]. Balance [933]. Balanced [452]. Balancing [626, 815, 970]. Ballistic [975]. Banerjee [485]. Banking [2562]. Bar [1284]. Barbacci [1746]. Barbarians [188]. Barrier [453, 648]. Barriers [341]. Based [80, 160, 165, 177, 187, 204, 238, 245, 262, 291, 376, 384, 387, 406, 421, 462, 536, 558, 640, 731, 735, 773, 778, 866, 870, 873, 921, 942, 986, 1007, 1033, 1054, 1121, 1166, 1232, 1295, 1296, 1318, 1378, 1650, 1668, 1811, 1886, 1891, 1964, 2013, 2034, 2051, 2075, 2263, 2283, 2284, 2359, 2414, 2523, 2532, 2581, 2604, 2649, 2676]. Basic [406, 435, 467, 2249]. Basics [626]. Basis [291, 406]. Bathroom [2296]. Batteries [1042]. Battery [438, 472, 495, 1074, 1583]. Battle [710, 884, 925, 2277, 2299, 2386, 2602]. Battling [767]. BCI [2392]. BCIs [2391]. Be [89, 122, 279, 322, 538, 547, 747, 818, 968, 1514, 2190, 2256, 2473, 2499]. BEA [1079]. Beacon [489]. Beans [2137]. Beating [604]. Beautiful [1352]. Become [31, 252]. Becomes [1188, 1489, 1660]. Becoming [617]. Beeline [1599]. Been [790]. BEES [1272]. Before [48, 79]. Begin [112, 370, 592, 975]. Begins [232, 308, 571, 2140, 2323]. Behavior [158, 204, 1668, 2571]. Behavior-Based [204, 1668]. Behavioral [103, 186, 772, 2531]. Behind [172, 852, 1053, 1381, 2314]. Beholder [2143]. Being [1855, 1890]. Belt [170]. Benchmark [845]. Benchmarketing [1063]. Benchmarking [844, 1037]. Benchmarks [915, 2511]. Benefits [588, 607, 663, 1855]. Benjamin [2612]. BeOS [122]. Bernoulli [905]. Best [226, 622, 964, 1111, 1730, 1738, 1978, 2030, 2593]. Better [243, 254, 347, 814, 819, 873, 973, 995, 1062, 1642, 1915]. Between [389, 701, 904, 2588]. Beyond [11, 22, 569, 632, 718, 823, 1146, 1270, 1281, 1459, 1527, 1533, 1852, 1913, 2501, 2509, 2584, 2639, 2674, 2701]. Bharat [99]. Bhargava [99]. Big [127, 547, 641, 659, 1399, 1787, 1981]. Billboard [2345]. Billboard-Viewership [2345]. Billion [21, 1142]. Billion-Transistor [1142]. Binaries [2328]. Binary [25, 26, 53, 54, 63, 65, 68, 70, 71, 84, 85, 108, 126, 148, 170, 192, 209, 229, 1425, 2039]. Binding [406, 672]. BioconX [447]. BioID [39]. Bioinformatic [688]. Bioinformatics [660, 685]. Bioinspired [1272]. Biological [735, 1407]. Biologically [2529]. Biology [102, 1030, 1186]. Biomedical [2261, 2415]. Biometric [38–41, 52, 109, 447, 1658].

Biometrics [36, 1107, 1178]. Biometry [290]. Biomimetic [1273]. Biomolecular [1966]. BioSig [688, 2392]. Bipolar [2582]. Bird’s [230]. Bird’s-Eye [230]. Birth [449]. Bit [778, 985]. Bite [2468]. Bits [11, 108]. Black [1402]. BlackBerry [649]. Blaxxun [267]. Blind [794]. Blue [852, 937, 1044]. Blueprint [684]. Bluestone [187]. Bluetooth [244, 510, 649, 675, 736]. Blunders [700, 743, 763]. Board [99, 161, 464, 517, 555, 691, 893, 982, 1894, 1980, 2462]. Body [578, 2052, 2459]. Bolsters [299]. Books [20, 48, 79, 102, 123, 144, 164, 186, 204, 223, 243, 268, 290, 312, 334, 360, 383, 405, 427, 446, 468, 488, 511, 535, 702, 1575, 2481]. Bookshelf [559, 582, 604, 626, 648, 671, 693, 716, 735, 754, 777, 806, 830, 852, 873, 896, 916, 941, 964, 986, 1008, 1034, 1053, 1080, 1107, 1153, 1171, 1194, 1213, 1234, 1257, 1278, 1302, 1323, 1347, 1373, 1393, 1416, 1434, 1458, 1481, 1502, 1523, 1545, 1568, 1589, 1611, 1634, 1657, 1704, 1725, 1748, 1771, 1797, 1822, 1845, 1867, 1898, 1923, 1945, 1971, 2017, 2038, 2060, 2079, 2107, 2128, 2153, 2176, 2198, 2220, 2242, 2265, 2285, 2312, 2334, 2356, 2377, 2398, 2417, 2438, 2482, 2506, 2526, 2549, 2570, 2590, 2614, 2635, 2656, 2678, 2699]. Boolean [2282]. Boomer [127]. Boost [593, 726, 1715, 2603]. Boosting [529, 1989, 2547]. Boosts [344, 975]. Booth [99, 485]. Boots [1304]. Bootups [2561]. Borg [86]. Borland [244, 267, 361, 426, 558]. Born [585]. Borrows [102]. Both [212, 646, 1738]. Bots [1361]. Bottle [21]. Bottlenecks [341, 391, 613, 2694]. Bounce [861]. Bouncing [904]. Bound [600]. Boundaries [2046, 2588]. Boundary [1939]. Bounds [346]. Bourne [164]. Box [146, 284, 1144, 1806]. Boxes [1304, 1402]. Boy [223]. Boys [2067]. BPC [1681]. Braille [794]. Brain [224, 653, 698, 1911, 2393–2395]. Brain-Computer [1911, 2393–2395]. Brains [824]. Brainstorming [1160].

Branching [728, 815]. Brand [777]. Brand-Name [777]. Brave [2646]. Brazil [2307]. Breadth [2683]. Breaking [648]. Breakthrough [498, 747, 2561, 2626]. breakthroughs [660]. Breath [1017]. Breeze [672]. Brewing [2210]. Bridge [341, 467, 485, 650, 778]. Bridging [2316]. Brief [932, 1717]. Briefs [8, 32, 59, 89, 112, 130, 152, 175, 212, 232, 254, 279, 299, 322, 344, 370, 392, 414, 436, 455, 475, 498, 520, 546, 571, 593, 614, 637, 659– 661, 682, 706, 747, 767, 794, 819, 841, 864, 885, 908, 927, 952, 975, 1021, 1067, 1095, 1118, 1141, 1162, 1185, 1203, 1226, 1247, 1269, 1292, 1314, 1337, 1363, 1385, 1406, 1427, 1446, 1470, 1491, 1513, 1535, 1556, 1579, 1601, 1626, 1647, 1670, 1692, 1762, 1786, 1810, 1835, 1859, 1884, 1912, 1934, 1960, 1982, 2006, 2028, 2050, 2071, 2095, 2118, 2141, 2165, 2345, 2368, 2388, 2409]. Briefs [726, 996, 1044, 2188, 2211, 2234, 2278, 2324, 2431, 2455, 2475, 2495, 2519, 2540, 2561, 2581]. Bright [658, 711]. Brightest [964]. Brilliant [2449]. Brinch [669]. Bring [127, 726]. Bringing [398, 852, 991]. Bristol [291]. Broadband [151, 232, 754, 908, 1201, 2031]. Broadcast [209, 595, 711, 2586]. Broadcasting [2416]. Broadcom [244, 1007]. Broadening [2443, 2533]. Broke [99]. Broker [1984]. Browser [49, 187, 1007]. browser-based [1007]. Browsing [428, 518, 1809]. Bruce [240]. BSquare [224]. Bud [401]. Buffer [916]. Bug [534, 583]. Bug-Tracking [534]. Bugs [806, 1053, 1952, 2695]. Build [80, 103, 307, 650, 2343]. Building [58, 144, 178, 243, 299, 427, 483, 555, 558, 754, 813, 814, 852, 995, 1008, 1320, 1357, 1417, 1554, 1577, 1606, 1744, 2020, 2291, 2501, 2604]. Built [1388]. Built-In [1388]. Bull [765, 975]. Bulldogs [316]. Bullet [1119]. Burden [2429]. Bus [344, 985, 1980, 2200]. Business [31, 102, 144, 152, 186, 217, 291, 335, 406, 470, 627, 641, 680, 694, 862, 873, 920, 1111, 1174, 1177, 1411, 1808, 1833, 1839, 2062, 2083, 2108, 2189, 2225, 2335]. Business-to-Business [217]. Businesses [144]. Busy [1045, 1140]. Buy [56]. Buzz [1505]. Buzzword [432]. Bylaw [982]. Bylaws [99, 161, 555, 691, 893, 2462].

C

[1, 97, 123, 291, 384, 406, 426, 627, 1033, 2348]. C# [625, 1008]. C. [485]. C/C [426, 1033]. C3 [315]. Cable [244]. Cache [262, 1121, 1168]. Cache-Based [1121]. caches [668]. Caching [512, 1987, 2603]. CAD [165, 267, 1106]. CAD-UL [165]. Cadena [1650]. caGrid [2415]. Calculator [904]. Calendar [17, 45, 76, 100, 121, 141, 162, 184, 202, 221, 241, 265, 288, 309, 332, 358, 380, 402, 424, 445, 466, 486, 508, 556, 580, 602, 623, 645, 670, 713, 734, 756, 780, 803, 827, 850, 872, 894, 917, 938, 962, 983, 1006, 1031, 1052, 1078, 1103, 1128, 1150, 1170, 1192, 1211, 1233, 1255, 1280, 1300, 1325, 1344, 1371, 1395, 1414, 1435, 1456, 1478, 1498, 1521, 1543, 1566, 1587, 1610, 1633, 1655, 1678, 1702, 1724, 1747, 1770, 1794, 1823, 1847, 1869, 1895, 1920, 1943, 1968, 1992, 2015, 2036]. Calendar [2059, 2081, 2105, 2126, 2151, 2173, 2197, 2222, 2244, 2267, 2289, 2311, 2333, 2355, 2376, 2397, 2419, 2442, 2463, 2484, 2505, 2528, 2569, 2592, 2613, 2634, 2658]. Calender [692]. Call [17, 45, 76, 100, 121, 141, 162, 184, 202, 221, 241, 265, 288, 309, 332, 340, 358, 380, 402, 424, 445, 466, 486, 508, 556, 580, 602, 623, 645, 670, 692, 713, 734, 756, 780, 803, 827, 850, 872, 894, 917, 938, 962, 983, 1006, 1031, 1052, 1078, 1103, 1128, 1150, 1170, 1192, 1211, 1233, 1255, 1280, 1300, 1325, 1344, 1371, 1395, 1414, 1435, 1456, 1478, 1498, 1521, 1543, 1566, 1587, 1610, 1633, 1655, 1678, 1702, 1724, 1747, 1770, 1794, 1823, 1847, 1869, 1895, 1920, 1943, 1968, 1992]. Call [44, 140, 389, 610, 1548, 2015, 2036, 2059, 2081, 2105, 2126, 2151, 2173, 2197, 2222, 2244, 2267, 2289, 2311, 2333, 2355, 2376, 2397, 2419, 2442, 2463, 2484, 2505, 2528, 2569, 2592, 2613, 2634, 2658]. Calling [2111]. CALM [1650]. Cam [123]. Came [132]. Camera [1342, 1789]. Cameras [730, 1653, 1788]. Camino [2025]. Campus [712]. Can [34, 160, 273, 778, 1067, 1510, 1570, 1720, 1766, 1841, 2190, 2460, 2499, 2628]. Canada [75]. Canadian [871]. CANcentrate [2547]. Candidates [161, 201, 357, 465, 733, 893, 1005, 1541, 1819, 2104, 2353, 2611]. Can’t [56, 272, 405]. Capabilities [2414]. Capability [34, 916, 1630, 2552]. Capacity [406, 726, 975, 1050]. Cape [627]. Capture [224, 2093]. Captured [1735]. Caravela [2013]. Carbon [2681]. Card [794, 2217]. Cardiff [467]. Cards [487, 726]. Care [551, 928]. Career [19, 46, 77, 101, 142, 163, 185, 203, 222, 242, 266, 289, 311, 333, 359, 382, 403, 425, 509, 533, 557, 581, 603, 624, 647, 695, 714, 737, 757, 781, 805, 828, 853, 875, 895, 919, 939, 984, 1009, 1032, 1055, 1081, 1105, 1151, 1172, 1193, 1214, 1237, 1256, 1276, 1301, 1326, 1349, 1372, 1396, 1412, 1433, 1454, 1479, 1500, 1522, 1542, 1564, 1585, 1614, 1656, 1680, 1700, 1726, 1749, 1772, 1795, 1820, 1844, 1897, 1922, 1944, 1970, 1990, 2035, 2057, 2082, 2106, 2127, 2152, 2175, 2195, 2219]. Career [6, 99, 2241, 2286, 2309, 2331]. Careers [1129]. Carnegie [989]. Carnivore [385]. Carriers [1384]. Carver [1077]. CASA [1864]. Case [260, 866, 902, 1146, 1471, 1606, 1681, 1740, 2029, 2054, 2167, 2238, 2446, 2650]. Catalog [871]. Catalyst [874]. Causes [593]. CaveUT [1438]. CD [622]. CD-ROM [622]. CDIDC [871]. CE [510, 1033]. CE-based [1033]. CEA-2018 [2604]. Ceiling [1095]. Celebrates [2433]. Celebrating [1621]. Celebrity [791]. Cell [927, 2031, 2209]. Cells [1042, 1389]. Cellular [199, 1468, 2191]. Celluloid [629]. Center [14, 103, 1318, 1319]. Centered [505, 1305, 2008, 2214, 2485]. Centers [1317, 2194, 2374]. Central [736]. Centric [152, 290, 1742, 2235, 2400]. Century [349, 865, 1988, 2236, 2258]. Certificates [290]. Certification [177, 331, 458, 622, 663, 664, 723, 2223, 2292]. Certified [1030]. Cetus [2693]. Cg [851]. Chadless [341]. Chads [341]. Chain [606]. Chair [893]. Chairs [1062]. Chalking [726]. Challenge [178, 208, 312, 413, 454, 548, 887, 951, 1708, 1840, 1842, 1885, 2373, 2451]. challenged [80]. Challenges [51, 155, 174, 235, 247, 306, 369, 394, 474, 479, 494, 584, 706, 831, 910, 960, 976, 1186, 1316, 1368, 1387, 1752, 1838, 1935, 1936, 2094, 2117, 2144, 2166, 2307, 2369, 2374, 2389, 2459, 2477, 2502, 2544]. Challenging [1062]. Chamois [640]. Champ [79]. Champion [852]. Chang [801]. Change [31, 107, 167, 256, 930, 1651, 2042, 2202, 2571]. Change-Tolerant [2042]. Changes [127, 171, 212, 507, 608, 982, 2625]. Changing [55, 109, 117, 475, 2163, 2189, 2456, 2539, 2540]. Channel [2589]. Channels [49, 1217]. Chaos [341, 1549]. Chaotic [1288]. Character [2578]. Characteristics [2370]. Chat [320]. Chattels [837]. Cheaper [498]. Checker [1053]. Checking [1, 218]. Chess [852]. Chicken [702, 2203]. Chief [3, 16, 161, 287, 308, 381, 579, 691]. Child [268]. Children [747]. Chimera [2088]. China [572, 712, 1271, 1765]. Chip [130, 204, 231, 232, 244, 354, 392, 637, 706, 726, 841, 927, 952, 967, 1007, 1145, 1362, 1444, 1495, 1581, 1827, 1989, 2277, 2368, 2386, 2494]. Chipmaking [1981]. Chips [23, 89, 194, 279, 341, 498, 520, 552, 593, 659, 726, 794, 817, 883, 994, 1093, 1109, 1291, 1405, 1493, 2094, 2345]. Chipsets [765]. Choosing [698]. Choreographing [767]. Choreography [1026]. Chosen [801]. Chronicles [1628]. Cinema [629]. Cinematography [1791]. Circuitry [1962]. Circuits [1148, 2111, 2328]. Cisco [212, 777]. CiSE [264]. City [1813, 1814]. Civil [1178]. Clarification [193, 472, 970, 1062]. Clarity [858, 2683]. Clash [246, 754]. Class [350, 530, 2002]. Classes [736]. Classification [238]. Classified [2106, 2264]. Classroom [5, 336, 864, 972, 1218, 2101, 2134]. Claude [357]. Clean [2187]. Cleaning [1034]. Clear [627]. Clearly [896]. ClearMail [1106]. ClearSight [1054]. Click [2061, 2364]. Client [2007, 2201, 2465, 2562]. Client-Server [2007]. Client-Side [2562]. Clients [560, 1674, 1800]. Climate [1177, 1358]. Climbing [5]. Clinical [1694]. Cloaking [468]. clock [794]. Clockless [1405]. Cloning [205, 487]. Close [500, 1669, 2474]. Closing [1013]. Clothing [469]. Cloud [2454, 2596, 2618]. Clustering [690, 1165]. CMMI [916]. Cobol [89, 205, 335]. Cocke [712]. Cockpit [2130]. CoCreate [1033]. Codagen [851]. Code [5, 49, 122, 187, 205, 253, 267, 283, 335, 534, 571, 583, 605, 941, 963, 1232, 1754, 2238, 2444, 2567]. CodeFutures [1106]. CodeGuide [778]. Codemesh [384, 467]. Codesign [82, 783, 888, 892, 1390, 2486]. CodeWarrior [143, 244, 778]. CodeWeavers [940]. Coding [248, 451, 452]. Cogito [123]. Cognition [1693]. Cognitive [710, 763, 784, 2247]. ColdFusion [672]. Collaboration [335, 1033, 1327, 2044, 2551]. Collaborative [66, 1792, 2335, 2500]. Colleague [240]. Collection [227, 304, 669, 779, 940, 2659]. College [2245]. Collide [108]. Color [2256]. Column [904]. Combinatorial [96, 2615]. Combine [1738]. Combines [59]. Combining [1102]. Come [57, 1268, 1518, 1885]. Comeback [186, 298]. Comedians [1532]. Comes [249, 1130, 1691, 2359, 2408]. Comfort [1880]. Coming [120, 450, 691, 899, 1688]. Commandments [1627]. Comments [5, 248]. Commerce [30, 143, 215– 217, 269, 586, 608, 845, 1350, 1527, 1715]. Commercial [843, 846, 1070, 2654]. Commitment [779, 2676]. Commitment-Based [2676]. Committee [381]. Committee-Sponsored [381]. Commodity [2032]. Common [1021, 2618].

Communicating [282]. Communication [79, 152, 164, 461, 731, 898, 996, 1000, 1408, 1669, 1940, 2009, 2301]. Communications [102, 104, 145, 189, 228, 267–269, 317, 363, 407, 452, 467, 482, 536, 675, 718, 726, 759, 808, 825, 856, 898, 942, 964, 987, 1021, 1037, 1085, 1124, 1189, 1215, 1225, 1526, 1722, 2300, 2314, 2494]. Communities [421, 550, 599, 1728, 2423, 2593]. Community [622, 987, 1298, 1606, 1985, 2075, 2245]. Community-Based [2075]. Community-Oriented [1298]. Companies [175, 377, 1021, 1833]. Company [44, 661, 1660, 2368]. Comparative [1369]. Compares [75]. Comparing [189]. Comparison [214, 913, 2655]. Compatible [694]. Compendium [626]. Compete [161, 871, 1766]. Competisoft [2119]. Competition [43, 240, 308, 443, 507, 546, 601, 703, 972, 1063]. Competitive [2192]. Compilation [915]. Compiler [64, 97, 98, 335, 467, 627, 941, 1107, 2693]. Compilers [626]. Complaints [904]. Complex [934, 1248, 1365, 1989, 2498, 2517]. Complex-Event [2517]. Complexity [343, 626, 1553, 1914, 2179, 2400]. Compliant [335, 915]. Component [2, 73, 74, 171, 238, 458, 558, 640, 873, 940, 1650, 2149]. Component-Based [238, 640, 873, 1650]. Componentization [290, 1763]. Components [132, 193, 375, 447, 873, 1054, 2553]. Composing [502, 1230]. Compositions [2350]. Compound [1270]. Compression [236, 885]. Computation [447, 648, 1408, 1743, 1791]. Computational [312, 687, 776, 1186, 1636, 1787, 1788, 1790, 2660, 2700]. Computations [1525, 1965, 2694]. Computer [3, 5, 20, 55, 86, 99, 103, 120, 123, 144, 182, 200, 212, 251, 264, 271, 278, 282, 291, 299, 339, 356, 360, 383, 422, 423, 464, 485, 532, 538, 572, 576, 598, 622, 626, 669, 681, 691, 705, 712, 727, 732, 733, 758, 767, 777, 801, 806, 824, 848, 852, 864, 873, 927, 976, 982, 986, 997, 1004, 1030, 1047, 1076, 1083, 1102, 1115, 1134, 1175, 1206, 1275, 1315, 1345, 1346, 1517, 1544, 1547, 1686, 1818, 1828, 1836, 1907, 1911, 1942, 2103, 2266, 2279, 2283, 2307, 2361, 2375, 2393– 2395, 2433, 2443, 2496, 2587, 2633, 2637]. Computer [16, 43, 75, 99, 120, 139, 161, 172, 201, 220, 240, 256, 287, 308, 331, 357, 381, 384, 401, 423, 443, 444, 465, 485, 494, 507, 520, 532, 555, 579, 583, 601, 622, 646, 669, 691, 712, 733, 755, 777, 779, 802, 814, 826, 849, 871, 893, 918, 961, 982, 1005, 1030, 1051, 1077, 1089, 1090, 1104, 1127, 1149, 1169, 1191, 1210, 1236, 1277, 1299, 1324, 1334, 1370, 1394, 1413, 1437, 1455, 1477, 1499, 1520, 1541, 1565, 1586, 1608, 1609, 1723, 1793, 1896, 2104, 2172, 2174, 2186, 2450, 2469, 2489, 2555, 2611, 2621, 2642, 2664, 2684]. Computer [263, 264, 531, 532, 691, 802, 804, 1075, 1077, 1632, 1701, 1769, 1868, 1893, 1894, 1921, 1967, 1991, 2014, 2037, 2058, 2080, 2125, 2150, 2171, 2196, 2221, 2243, 2288, 2332, 2353, 2354, 2396, 2418, 2440, 2441, 2462, 2483, 2504, 2527, 2548, 2568, 2591, 2657, 2697]. Computer-Assisted [927]. Computer-Generated [727, 1517]. computerized [794]. Computers [5, 10, 87, 122, 405, 553, 567, 609, 610, 632, 654, 677, 683, 726, 729, 825, 927, 1044, 1095, 1340, 1473, 1663, 2065, 2182, 2672]. Computing [20, 21, 80, 94, 105, 146, 164, 166, 188, 225, 247, 290, 295, 312, 330, 346, 362, 442, 475, 536, 540, 541, 698, 700, 707, 708, 719, 750, 763, 806, 821, 833, 871, 899, 901, 918, 922, 924, 943, 944, 967, 988, 997, 1013, 1024, 1034, 1046, 1080, 1083, 1086, 1109, 1130, 1157, 1231, 1298, 1318, 1320, 1367, 1368, 1582, 1685, 1811, 1816, 1928, 1933, 1961, 1964, 2008, 2018, 2023, 2167, 2204, 2214, 2218, 2258, 2262, 2326, 2374, 2402, 2414, 2421, 2443, 2446, 2447, 2454, 2456, 2476, 2485, 2529, 2530, 2533, 2552, 2571, 2583, 2650, 2663, 2670, 2675]. Computing [35, 50, 93, 109, 134, 193, 208, 271, 308, 314, 331, 344, 361, 381, 401, 408, 439, 444, 532, 538, 561, 562, 565, 571, 646, 676, 725, 728, 739, 758, 783, 795, 809, 854, 855, 863–865, 877, 910, 911, 916, 921, 928, 937, 954, 977, 980, 989, 998, 1003, 1015, 1021, 1023, 1034, 1036, 1038, 1045, 1056, 1068, 1069, 1080, 1084, 1307, 1316, 1354, 1462, 1580, 1631, 1812, 1813, 2013, 2144, 2158, 2170, 2202, 2310, 2366, 2389, 2401, 2432, 2458, 2473, 2487, 2496, 2508, 2521, 2531, 2541, 2562, 2572, 2575, 2596, 2618, 2671, 2682]. Computing/Enigma [290]. Compuware [715, 736, 829, 851, 963, 1007]. Concept [49, 550, 1811]. Concept-Based [1811]. Concepts [1538, 2145]. Conceptual [527, 2133]. Concern [279, 1625]. Concerning [1875]. Concerns [232, 465, 479, 565, 654, 733, 906, 1336, 1541, 1639, 2104, 2344, 2345, 2353, 2611]. Concurrency [2695]. Concurrent [186, 384, 1053]. Conduct [44, 140, 1572]. Conference [532, 2288]. Conferences [120, 381, 555, 646, 826, 918, 1219, 1723, 2227]. Conferencing [1383]. Conficker [2560]. Confidence [1695]. Configurable [23, 93, 94, 833, 1495]. Configuration [103, 896, 1209]. Conflicting [815]. Confrontational [452]. Confusion [2064]. Connect [1065]. Connected [1312, 1606, 2700]. Connection [16, 43, 75, 99, 120, 139, 161, 201, 220, 240, 264, 287, 308, 331, 357, 381, 401, 423, 444, 485, 507, 532, 555, 579, 601, 622, 646, 669, 691, 712, 733, 755, 779, 802, 826, 849, 871, 893, 918, 961, 982, 1005, 1030, 1051, 1077, 1104, 1127, 1149, 1169, 1191, 1210, 1236, 1277, 1299, 1324, 1394, 1413, 1437, 1455, 1477, 1499, 1520, 1793, 1921, 1967, 1991, 2014, 2037, 2058, 2080, 2104, 2125, 2150, 2196, 2221, 2243, 2332, 2354, 2396, 2418, 2441, 2462, 2483, 2504, 2527, 2548, 2568, 2591]. Connections [571]. Connectivity [614, 1067, 1101]. Conquering [2179]. Conscious [262]. Consciousness [416, 2065]. considers [952]. Consistency [1987]. Console [2602]. Constrained [2044]. Constraint [350]. Construction [175]. Consumer [73, 74, 1948]. Consumption [1021, 1072]. Contemporary [964]. Content [169, 187, 237, 426, 438, 480, 743, 960, 1482, 1705, 1865, 2051, 2235, 2371, 2423, 2446]. Content-Based [2051]. Content-Centric [2235]. Contest [264]. Context [440, 740, 898, 1000, 1261, 1812]. Context-Aware [898, 1000, 1261]. Context-Sensitive [440]. Contextual [1890]. continue [927]. Continuity [627]. Continuous [459, 667, 2662]. Contracts [1048, 2149, 2335]. Contribute [802]. Contributions [1345]. Control [32, 102, 114, 224, 291, 896, 1625, 2240, 2695]. Controller [2547]. Controllers [2602]. Controlling [80, 2070, 2091]. Controls [1]. Controversial [475, 1021, 1095]. Controversy [49, 212, 520, 593, 1760]. Convenience [632]. Converged [1577]. Converging [1825]. Conversation [2091]. conversion [1007]. Converting [122, 205]. Cooler [347]. Coolest [764]. Cools [436]. Cooper [1346]. Cooperation [2284]. Cooperative [125, 303, 869, 1121, 1987]. Coordination [42, 418]. Coping [92, 1163, 1696]. Coplink [594]. Copper [1119]. Coprocessor [483]. Copy [481, 1764]. Copyright [279, 986, 991]. Copyrights [565, 986]. Copywrongs [986]. CORBA [136, 754, 1927]. Core [223, 351, 754, 1053, 2432, 2691]. Corel [940]. Corporate [636, 1094]. Corporation [291, 357]. Corps [299]. Correction [28, 29, 87]. Cosimulation [890, 915]. Cost [71, 102, 752, 1033, 2315, 2521]. Cost-Effective [102]. Costs [927]. COTS [167, 376, 378, 387, 411, 868]. COTS-Based [376, 387]. Could [89, 112, 455, 593, 794, 975, 2561, 2581, 2603]. Count [322]. Counters [829]. Counting [2137]. Counts [2527]. Coupled [2413]. Coupled-Simulation [2413]. Course [783, 896, 1999, 2318]. Courses [601, 712, 871]. Court [996]. Courtship [316]. Cover [686–690, 709– 711, 729, 730, 749–751, 769–776, 796–800, 820– 825, 844–848, 866–868, 888–892, 911, 912, 931– 936, 956–960, 978, 979, 999–1003, 1025– 1028, 1047, 1048, 1070–1074, 2391, 2412, 2434, 2456, 2478, 2498, 2522, 2543, 2564, 2576, 2585]. Coverage [1586, 1868, 2418]. CPE [1563]. CPU [154, 1109, 2521]. CPU/FPGA [1109]. CPU2000 [154]. CPUs [546]. CPXe [1027]. CRA [331, 2527]. Craft [193, 296, 2293]. Crafting [490]. Crash [84, 85, 299, 928]. Crash-Proof [299]. Crawlers [996]. Cray [308, 355, 937, 2698]. Create [129, 149, 546, 879]. Creates [498, 2117]. Creating [171, 243, 864, 1398, 1590, 2361, 2481]. Creation [187, 426]. Creative [2690]. Creativity [308, 986]. Credibility [1767]. Credit [2468]. Crime [896, 1167]. Crimes [182]. Criminal [594, 896]. Criminal-Justice [594]. Crisis [1262]. Critic [25, 26, 53, 54, 84, 85, 108, 126, 148, 170, 192, 209, 229, 230]. Critical [175, 735, 1005, 1047, 1238, 1625, 1722, 1754, 2062, 2280, 2445, 2638]. Critique [452]. Cross [291, 583, 597, 605, 1123, 1837, 1940]. Cross-Application [597]. Cross-Layer [1940]. Cross-Layering [1123]. Cross-Platform [291, 583, 605, 1837]. CrossOver [940]. Crossroads [922]. Crouching [471]. Crowdsourcing [2422]. CROWN [2348]. CROWN-C [2348]. Crucial [210]. Crunch [2468]. Crusoe [59]. cryptographic [957]. Cryptography [27, 1034, 1592, 1937]. Cryptology [626]. CS [555, 2450, 2469, 2489, 2513, 2535, 2555, 2577, 2598]. CSIDC [161, 172, 308, 357, 423, 507, 691, 703, 972, 1244, 1508, 1677, 1783]. Cube [2661]. Cube-Type [2661]. Cubic [282]. Cubic-Millimeter [282]. Cues [1903]. Culler [308]. Cultural [1813, 2194]. Culture [754, 1743, 2112]. Cultures [457]. Curb [1117]. Curl [467]. Currency [1358].

Current [1073, 1450, 2108]. Curricula [401]. Curriculum [733, 1254]. Curse [362, 1419]. Curve [1643]. Curves [873]. Custom [49, 729, 964]. Customer [168, 307, 977]. Customizable [49]. Customizing [186]. Cut [927, 1764]. Cuts [975]. Cutting [2585]. Cutting-Edge [2585]. Cyber [299, 535, 2508]. Cyber-physical [2508]. Cyberattack [344, 1584]. Cybercourt [414]. Cyberinfrastructure [1864, 2212]. Cyberprivacy [292]. Cybersecurity [2346]. Cyberspace [140, 419, 810, 965]. Cybervote [322]. Cycle [159, 1058, 2072, 2148, 2170, 2486]. Cynicism [515].

D [62, 194, 406, 408, 496, 649, 661, 927, 986, 1197, 1246, 1512, 1515, 1516, 1809, 2022, 2052, 2055, 2132, 2345, 2368, 2672]. Dahl [755]. Daily [105, 677]. DakNet [1101]. Dance [316]. Dancing [702]. Dangerous [8]. Dark [538]. DARPA [864, 1885]. Dashboard [49]. Data [21, 48, 102, 122, 144, 168, 227, 262, 290, 304, 307, 327, 331, 361, 399, 400, 406, 455, 468, 488, 523, 526– 528, 532, 594, 613, 615, 636, 672, 682, 706, 753, 775, 860, 864, 896, 928, 981, 1008, 1034, 1045, 1049, 1095, 1106, 1121, 1167, 1198, 1317– 1319, 1604, 1675, 1755, 1853, 1984–1987, 2139, 2254, 2258, 2259, 2262, 2317, 2367, 2373, 2374, 2429, 2436, 2456, 2480, 2605, 2609, 2623, 2659]. data-access [864]. Data-Binding [406, 672]. Data-Centric [290]. Data-Intensive [2258, 2262, 2456]. Data-Sharing [2373]. Database [187, 197, 224, 243, 267, 467, 524, 525, 605, 694, 829, 896, 907, 941, 1080, 1106, 1459, 2133]. database-access [907]. Databases [57, 173, 305, 399, 841, 1108, 1224, 1863]. Datasets [2261, 2263]. David [532]. Davidson [331]. DaVinci [2123]. Dawn [348]. Days [701]. DBA [1008]. Dea [2403]. Dead [295, 709]. Deaf [794, 2336]. Deal [1978]. Death [791]. Debate [56, 663, 664, 723, 931]. Debt [1902]. Debues [940]. Debug [788]. Debugger [510, 715, 736]. Debugging [500, 715, 806, 1033, 1989, 2329]. Debunking [1727]. Debuts [672]. Decade [888, 1701, 2443]. Decadent [860]. Decades [1377]. Deciphering [2077]. Decision [523, 525, 829, 1862, 1908]. Decision-Support [829]. Decisions [971, 2567]. Deck [1062]. Declarative [2631]. Decomposition [530]. Decompositions [2391]. Decoy [370]. Deduplication [2429]. Deep [760, 852, 964, 1044, 2257, 2380]. Deepen [1108]. Defeated [852]. Defect [49, 293]. Defect-Tracking [49]. Defective [1696]. Defending [810, 970]. defends [1079]. Defense [829, 945, 1240, 1440]. Defined [2124]. Defining [13, 186, 401, 1088]. Definitional [2525]. Degree [1263]. Delays [522, 566]. Deliberate [130]. Delight [806]. Deliver [111, 1313]. Delivers [807]. Delivery [243, 447]. Dell [223]. Demand [379, 459, 460, 1260, 1482]. Dementia [2009]. Democracy [587, 610]. Demos [622]. Denial [88, 751]. Denial-of-Service [88]. Denneau [937]. Density [94]. Departing [2441]. Departments [55, 2382, 2404, 2424, 2470, 2490, 2513, 2535, 2556, 2577]. Dependability [1320, 2202]. Dependencies [2565]. Deploying [2305]. Deployment [426, 534, 2102]. Depth [896, 941, 2683]. Derek [2207]. Derived [1333]. Dertouzos [485]. Dervishes [963]. Descrambling [48]. Describes [331]. Description [771, 2704]. Design [43, 122, 144, 149, 164, 218, 232, 240, 243, 291, 306, 308, 328, 335, 337, 350, 426, 443, 447, 502, 503, 505–507, 583, 605, 606, 703, 706, 770, 778, 807, 829, 874, 887, 889, 915, 921, 940, 941, 964, 967, 972, 985, 986, 988, 1008, 1013, 1018, 1054, 1056, 1062, 1074, 1080, 1091, 1123, 1145, 1175, 1387, 1388, 1452, 1495, 1553, 1583, 1631, 1649, 1743, 1824, 1902, 2056, 2083, 2121, 2122, 2303, 2304, 2536, 2554, 2631, 2673].

Designer [852]. Designers [652]. Designing [48, 286, 527, 696, 729, 731, 848, 852, 1047, 1274, 1409, 1815, 1817, 1865, 1914, 2571]. Designs [103, 627, 890, 1389, 1639, 1675]. Desk [143, 558]. Desktop [212, 940, 1404, 1974]. Despair [1713]. Detailed [1034]. Details [366]. Detecting [1903, 2655]. Detection [583, 896, 1251, 1584, 2177]. Determining [440, 1354, 2237]. Deterministic [643]. Detroit [108]. Dev [244]. Develop [370, 422, 2647]. Developer [313, 331, 487, 694, 915, 1080, 1504]. Developer-Focused [1504]. Developers [986, 1030, 1460, 1646, 1741, 1932]. Developing [177, 224, 386, 596, 725, 754, 873, 934, 964, 1053, 1101, 1471, 1496, 1649, 1754, 1811, 1901, 2022, 2214, 2248, 2298, 2389, 2603, 2659]. Development [1, 79, 80, 91, 114, 133, 143, 144, 186, 205, 231, 267, 291, 313, 335, 361, 375, 384, 395, 406, 426, 435, 447, 467, 470, 510, 513, 534, 558, 588, 605, 610, 627, 639, 649, 736, 743, 753, 795, 807, 815, 851, 852, 873, 915, 929, 930, 932, 936, 941, 963, 970, 986, 1007, 1154, 1196, 1271, 1338, 1376, 1534, 1538, 1539, 1629, 1650, 1652, 1731, 1744, 1837, 1872, 1953, 2235, 2238, 2301–2303]. Developmental [1637]. Developments [593, 1383]. Develops [510, 794, 927, 2345]. Device [313, 778, 874, 1067, 1341, 1695, 1698, 1699, 1730, 1995]. Devices [80, 198, 253, 368, 398, 510, 560, 749, 1033, 1042, 1133, 1259, 1339, 1368, 1431, 1673, 1696, 1722, 1966, 2123, 2280, 2282, 2318, 2393]. Devil [366]. Devour [385]. Diagnosing [1218]. Diagram [1740]. Dialog [619, 1248, 2636]. Dialogs [1560]. Dialogues [916]. Dials [49]. Diary [2011]. Dictator [2536]. Dictionary [123, 224, 2034]. Didn’t [633]. Difference [2236]. Different [610, 763, 990, 1855]. Digital [49, 50, 80, 108, 204, 290, 302, 306, 312, 324, 341, 438, 462, 472, 480, 495, 629, 631, 637, 653, 669, 715, 761, 854, 873, 885, 956, 960, 1034, 1053, 1086, 1262, 1285, 1336, 1342, 1353, 1389, 1398, 1482, 1497, 1571, 1593, 1602, 1603, 1661, 1690, 1975, 2101, 2123, 2191, 2193, 2212, 2230, 2271, 2313, 2416, 2468, 2510, 2520, 2530, 2563, 2565, 2641]. DigiTerra [694]. Dilemmas [2461]. Dimension [593]. Dimensions [726, 1813]. Dinosaurs [2384]. Dip [127]. Direction [1538]. Directions [14, 347, 2502]. Director [99, 172, 357, 532, 893]. Director-Elect [99, 357, 893]. Directorie [86]. Dirty [1930]. Disabilities [2018]. Disability [1281]. disabled [747]. Disappear [1340]. disaster [703]. disaster-response [703]. Disasters [1008]. Discipline [931, 2023, 2121, 2290, 2575]. Disciplines [444, 541]. Disciplinesand [1096]. Discovery [455, 463, 675, 1088, 2411, 2465, 2616]. Discrepancy [693]. Discrete [447]. Discrete-Event [447]. Discryption [2043]. Discussion [1062]. Disk [49, 205, 487, 1072, 2043]. Disk-Cloning [487]. Disks [400]. Disorders [1218]. Display [736, 839, 852, 927, 1067, 1514, 1517, 1518, 2055]. Displays [326, 1515, 1516, 1519, 1672, 1709, 2012, 2626]. Dispute [637]. Dissemination [1190, 1947]. Distance [712, 871, 879, 2341]. Distinguished [75, 308, 485]. Distinguishing [389, 701]. Distributed [72, 82, 129, 134, 138, 165, 188, 276, 294, 352, 361, 529, 532, 550, 665, 706, 725, 754, 911, 937, 1110, 1190, 1251, 1453, 1540, 1582, 1653, 1938, 1948, 2013, 2414, 2524]. Distributed-Computing [725]. Distribution [267, 583, 2371]. Diversity [191, 1947, 2443]. Divide [341, 761, 1086, 1758]. Divisible [914, 1088]. Division [99, 357, 815, 893]. DNA [689, 874]. Do [152, 405, 546, 664, 745, 1308, 1941, 1996, 2182, 2445]. DO-254 [2445]. Document [267, 325, 467, 1166, 2370]. Documents [975, 2647]. Dodging [701]. Does [86, 212, 2381]. Doesn’t [207, 633]. Dog [816]. Doing [492, 678, 2226]. DOM [299]. Domain [243, 279, 329, 502, 892, 1607, 2004, 2368]. Domain-Oriented [243]. Domain-Specific [502, 892]. Dominant [1489]. Donates [849]. Done [1334]. Don’t [381, 697, 2585]. Doomed [481]. Doughnut [1755]. Down [130, 562, 689, 752, 852, 1902, 1933, 2181]. downloads [975]. Downturn [475, 953]. Draft [733]. Drafting [873]. Dragon [572]. Dramatic [2625]. drastic [952]. Draw [2538]. Dream [283, 341]. Dreams [1860]. Drift [2582]. Drinking [2384]. Drive [231, 855, 1066, 1889, 2083]. Driven [852, 933, 1411, 1538, 1560, 1648, 1649, 1651, 1652, 2638, 2662]. Driver [503, 511, 736, 852, 963, 1888, 2012]. Drives [50, 1445]. Driving [752, 967, 2019]. Drop [392, 1021, 2266]. Droplet [291]. Drops [1095]. DRPM [1072]. Drummer [610]. DSL [748]. DSP [829, 915]. DSPs [546]. Dual [842]. Due [43, 579, 849]. During [133, 2443]. Dust [282, 1937]. DVD [661]. dye [952]. Dyes [1021]. Dynamic [70, 752, 873, 1232, 1365, 1752, 1932, 2268, 2328, 2478, 2627, 2651]. Dynamically [1071].

E-Business [152, 291, 335, 641]. E-Commerce [30, 143, 215, 217, 608, 845, 1715]. E-Communities [421]. E-Emperor [469]. E-Government [303, 1874]. E-Grind [48]. E-Health [1862]. E-mail [7, 130, 414, 1095, 1578, 1792, 1936, 2589]. E-Mailing [2115]. E-Micropayments [1245]. E-Paper [2256]. E-Passports [2129]. E-Process [218]. e-Science [2351, 2411, 2413, 2415, 2416]. E-Sex [608]. E-Shopping [130]. E-Wallet [219]. E.T. [702]. E3 [449]. Each [2213]. EAL4 [857]. Early [349]. Earned [970]. Earned-Value [970]. Earnings [75]. Ears [249]. Earth [764]. Easier [2338, 2519]. Easy [2215]. eBook [715]. ECBS [610]. Eclectic [2636]. Eclipse [1489]. Economic [389, 1983]. Economically [738]. Economics [748, 1219]. Economy [124, 312, 392, 1378, 1784]. Edge [1229, 1617, 2551, 2585]. Editing [267, 2368]. edition [806]. Editor [3, 16, 49, 122, 143, 161, 187, 287, 308, 381, 520, 534, 579, 955, 1514, 1557, 1648, 1694, 1741, 1787, 2073, 2120, 2206, 2229, 2251, 2295, 2390, 2411, 2490, 2584, 2622, 2685]. Editor-in-Chief [16, 161, 287, 381, 579]. Editors [36, 134, 215, 374, 439, 459, 479, 685, 691, 768, 843, 887, 930, 998, 1024, 1069, 1120, 1205, 1249, 1317, 1448, 1493, 1602, 1812, 1961, 2008, 2097, 2433, 2441, 2497, 2542, 2563]. EDK [807]. Educating [1741]. Education [475, 485, 893, 961, 996, 1462, 1475, 1476, 1742, 1743, 1928, 2101, 2212]. Educational [5, 99, 240, 264, 543, 610, 1057, 2617]. Edward [331, 2405]. Effective [102, 1080]. Effects [5, 631, 997, 2584, 2585, 2587]. Efficiency [2168]. Efficient [379, 668, 754, 1122, 1145, 2326, 2368, 2432, 2675]. Effort [325, 1760]. Efforts [747, 1534]. eHelp [187]. EIC [55, 273, 541, 790, 813, 1089, 2582]. Einstein [27]. EJB [361]. Ektron [1007]. Elect [99, 201, 357, 465, 733, 801, 893, 1005, 1076, 1346, 1541, 1609, 1819, 1894, 2104, 2171, 2353, 2440, 2611, 2697]. Election [200, 308, 579, 691, 732, 1004, 1544, 1818, 2103, 2633]. Elections [2375]. Electoral [318]. Electricity [1930]. Electronic [164, 438, 554, 889, 1067, 1475, 1527, 2085, 2481]. Electronics [73, 74, 506, 824, 1926, 1948]. Electrophysiologically [598]. Elects [250, 517]. Elements [916]. Elgar [2405]. Elicits [5, 515]. Eliminate [2561]. Eliminating [2695]. Elite [2040]. Elliott [275]. Elsewhere [2450, 2469, 2489, 2513, 2535, 2555, 2577, 2598]. Elusive [806]. Embarcadero [605]. Embedded [57, 119, 165, 196, 197, 205, 223, 291, 294, 313, 327, 337, 356, 361, 384, 406, 426, 447, 479, 484, 487, 497, 504, 510, 558, 561, 583, 605, 606, 627, 652, 698, 715, 730, 739, 777, 778, 783, 807, 833, 877, 887, 892, 915, 921, 967, 1013, 1054, 1056, 1109, 1156, 1239, 1366, 1390, 1391, 1476, 1629, 1639, 1653, 1775, 2022, 2120–2122, 2202, 2326, 2421, 2508, 2522, 2523, 2552, 2638, 2696]. Embedded-System [361]. Embedded-Systems [504]. Embody [339]. Embrace [2323]. Embracing [1988]. EMC [357]. Emerge [614]. Emergence [269]. Emergency [1768]. Emerges [629]. Emerging [4, 52, 134, 407, 685, 1430, 2280, 2281, 2467]. EmmaSoft [205, 487]. Emperor [469]. Empire [146]. Empirical [214, 2007]. Empirix [361, 534, 851]. Employee [724]. Employment [475, 1100, 1316]. Empowered [2304]. Empowering [1282]. Empty [1806, 2364]. Emulation [874, 890]. emWare [778]. Enable [2168]. Enabled [190]. Enablers [2131]. Enables [952, 2690]. Enabling [303, 672, 1176, 1638, 1671, 2369]. Encoder [122]. Encoding [2507]. Encounters [341]. Encouraging [2169]. Encrypted [1328]. Encryption [174, 232, 254, 711, 747, 819, 1958, 2043, 2359, 2519, 2661]. End [148, 212, 224, 341, 391, 541, 629, 644, 1097, 1229, 1596, 1750, 2148, 2263]. End-to-End [644, 2148, 2263]. Ender [1397]. Ends [1760]. Enemies [1764]. Energy [304, 322, 1070, 1074, 1122, 1321, 1583, 1672, 2167, 2168, 2368, 2432, 2675]. Energy-Adaptive [1672]. Energy-Aware [1074, 1583, 1672]. Energy-Efficiency [2168]. Energy-Efficient [1122, 2368, 2432, 2675]. Energy-Proportional [2167]. Enforcement [278]. Engagement [2099]. Engel [485, 1076]. Engine [109, 243, 267, 829, 1306, 1438, 2031]. Engineer [5, 120, 794, 1730]. Engineering [23, 27, 82, 83, 102, 115, 123, 125, 133, 161, 168, 206, 247, 294, 308, 331, 337, 373, 374, 376, 386, 389, 410, 430, 472, 530, 532, 539, 540, 579, 589, 621, 640, 654, 655, 669, 701, 733, 774, 830, 866, 959, 971, 993, 996, 1053, 1107, 1181, 1315, 1375, 1472, 1563, 1630, 1640, 1648, 1651, 1694, 1729, 1825, 1839, 1842, 2021, 2109, 2237, 2247, 2292, 2308, 2461, 2496, 2563, 2567, 2572, 2609, 2638, 2662, 2677, 2695]. Engineers [75, 669, 735, 1265, 1402, 1575, 1941]. Engines [13, 477, 543, 885, 1404, 1559, 1751, 1798, 2074]. English [1803]. Enhance [158]. Enhanced [675, 736]. Enhancement [605]. Enhancements [2426]. Enhances [851, 963, 1007, 1054, 1079]. Enhancing [421, 1215, 1473, 1849, 2101, 2180]. Enigma [290]. Enough [1094, 2053]. Enriching [2701]. Enrollments [2266]. Ensembles [1341]. Ensuring [1698, 2564]. Enter [572]. Enterasys [143]. Enterprise [103, 426, 558, 605, 874, 1107, 1204, 1534, 1980, 2021, 2084, 2146, 2147, 2157, 2420, 2559]. Enterprises [1107, 2248]. Enters [1036]. Entertain [2447]. Entertaining [989]. Entertainment [50, 105, 146, 188, 198, 225, 271, 314, 362, 408, 538, 642, 676, 719, 758, 809, 854, 899, 943, 960, 989, 1036, 1083, 1366, 1497, 1660, 2402, 2447, 2487, 2531]. Entry [507]. Environment [1, 62, 165, 291, 335, 361, 384, 406, 426, 447, 467, 605, 649, 778, 851, 889, 963, 1002, 1007, 1428, 2013, 2031]. Environmental [1250, 2595]. Environments [37, 502, 536, 869, 1011, 1110, 1126, 1409, 1453, 1474, 1649, 1752, 1892, 1987, 2044, 2410, 2691]. EPIC [35]. Epidemic [1190]. Epidemics [1039, 1207]. Epidemiology [2700]. Equal [207]. Equals [229]. Equation [389]. Equator [447]. Era [232, 348, 1495, 2042, 2096, 2327, 2432, 2680]. Ergo [123]. Errata [367]. Erratum [566]. Erroneous [27]. Error [1, 102, 471, 1388, 2272]. Error-Control [102]. Errors [1209]. Escher [1972]. Essential [5, 941, 1053]. Establishes [996].

Estimation [80, 1391, 2696]. Ethernet [89, 392, 658, 1007]. Ethical [2461]. Ethics [2338, 2563, 2567]. Etiquette [2564]. European [2362]. EuroSoft [205, 487]. Evaluating [38, 378, 843, 1497, 2349]. Evaluation [411, 845, 857, 868, 976, 997, 1722, 2096, 2502]. Evaluations [34]. Even [806, 1260, 1642]. Event [447, 2517]. Events [68, 381]. EventWeb [2214]. Every [322, 864]. Everybody [2373]. Everyday [2571]. Evidential [312]. Evitability [61]. Evolution [429, 457, 512, 689, 800, 1651, 1886, 2142, 2145, 2193, 2400, 2532]. Evolutionary [648, 916, 943, 964, 2427]. Evolves [59, 89, 1340]. Evolving [964, 1090, 1745, 2023]. Examination [2677]. Examining [2166]. Example [716, 852, 1048, 1763]. Examples [1167]. Exascale [2672]. Excel [447, 487, 534]. Excellence [99, 622, 754, 2657]. Exchange [1676]. Executable [294]. Executing [600]. Execution [69]. Executive [172, 532, 2287]. Exhibition [286]. Exhibitions [1683]. ExoLab [406]. Expand [59, 89]. Expanded [669]. Expanding [439, 554, 1080, 1342, 2575]. expands [907]. Expect [1841]. Expectations [24]. Experience [941, 1674, 2180, 2306]. Experiences [286, 520, 660, 2305]. Experimental [1067]. Experimentation [123]. Experiments [1298, 2110]. Expert [601, 779, 1608, 1893, 2172, 2441]. Expertise [139, 264, 532, 1999]. Experts [2396]. Expiration [232]. Explain [158]. Explained [896]. Explanatory [1560]. Explicit [319]. Explicitly [35, 2692]. Exploding [2407]. Exploiting [354]. Exploration [822, 1630, 1861]. Exploratory [1034]. Explore [918]. Exploring [211, 504, 690, 1001, 1272]. Expression [1878]. Extended [510]. Extending [1174, 1988, 2432]. Extends [143, 820, 829, 851, 963, 1007]. Extensible [1963]. Extension [244, 313, 963]. Extensions [583, 934]. Extracting [2573]. Extreme [715, 1093, 2670, 2671, 2675]. Extreme-Scale [2670, 2671, 2675]. Eye [230, 2199]. Eyes [249, 1850, 2143, 2401, 2585]. Eyes-Free [2401].

F [669]. Fabric [1010, 1021, 1827, 2194, 2366]. Fabrics [1037]. Face [37, 277, 546, 2539]. Face-to-Face [277]. Facebook [2405]. Faces [474, 1690, 2209, 2625, 2646]. FaceTime [753]. Facilitating [1814, 2102]. Facing [413, 494]. Factor [513, 672, 1871]. Factorization [2606]. Factory [467, 2460]. Facts [830, 2522]. Faculty [75, 331, 871, 2443]. Failure [207]. Failures [2566]. Fair [357, 982, 1769, 2504]. Fairies [1663]. Fairness [632]. Fairs [882, 1488]. Fake [370, 2655]. Fall [48, 1723, 2557]. Fallacies [830]. False [830]. Family [1007, 2600]. Fans [1282]. Fantasy [2588]. Farewell [229]. Farmer [1663]. Farms [2203]. Fashioning [166]. Fast [362, 975]. Faster [8, 498, 571, 613, 614, 883]. Father [2048]. Fault [386, 447, 1722, 1927]. Fault-Tolerant [1927]. Faultless [2628]. Favorite [149]. FCD [588]. Fear [451]. Feature [686–690, 709–711, 729, 731, 749, 750, 752, 769– 776, 796–800, 820–825, 844–848, 866–870, 889– 892, 911–914, 931–936, 956–960, 978–981, 999– 1003, 1025–1029, 1047, 1048, 1070–1074, 2373, 2391, 2412, 2434, 2437, 2456, 2478, 2480, 2498, 2522, 2524, 2543, 2547, 2564, 2585, 2589]. Feature [730, 751, 888, 1049, 1050]. Features [871]. Federal [41]. Federated [907, 1616, 2543, 2544]. Feedback [930, 970, 2074, 2240]. Feel [80, 254]. Feeling [2087]. Fellow [43, 240, 579, 849]. Fellows [75, 308, 579, 849, 1654]. Females [2533]. Femtocells [2323]. Fernbach [99, 331]. Fervor [2427]. Feudalism [365]. Few [32, 68]. Fewer [793]. FGPA [736]. Fi [951, 1351, 1492, 1834, 1997, 2300]. Fiber [165, 927, 1044]. Fiber-Optic [165]. Fiction [10, 87]. Field [134, 331, 1669, 1677, 1784, 1860]. Fields [646, 1790, 2018]. Fight [706, 996, 1432, 1959]. Fighting [414, 535, 860, 896, 1039, 1293, 1952]. Figures [2341, 2522]. File [188, 276, 649, 2139]. Files [122]. FileStream [205]. Filtering [1792]. finalizes [1007]. Finally [434, 519, 1313, 1445, 1950, 2232, 2276]. Finals [161, 423, 691]. Financial [183, 1563]. Find [1559]. Finding [806, 1044]. Findings [646]. Fingerprint [715]. Fingerprinting [615, 2702]. Fingertip [2069]. Firewall [627, 1209]. Firewalls [945, 970]. FireWire [510]. Firms [1766]. First [49, 99, 144, 350, 510, 778, 896, 915, 1054, 1508, 1840, 2368]. First-Class [350]. Fishing [927]. Fitness [1569]. Fittest [2629]. Five [1053, 2131]. Fixed [151, 908, 1231]. Fixing [48, 79]. Flash [778, 1624, 1645]. Flash-Based [778]. Flaws [149]. Flex [1389]. Flexibility [503, 974]. Flexible [735, 2029, 2455, 2626]. Flick [2701]. Flight [225]. Flock [1254]. Fluke [1]. Fly [175, 1834]. Flynn [106]. Focus [99, 109, 230, 322, 507]. Focused [1504, 2223]. Focuses [308]. Follies [655, 723]. Food [27]. Fookes [122]. Fooling [109]. Footprint [2681]. Force [122, 851, 2161]. Forces [444]. Forecasting [1864]. Forensics [278]. Foresight [970]. Forgetting [259]. Forging [285]. Fork [112]. form [778]. Formal [378, 891, 1056, 1627, 1699, 2629]. Formalization [1828]. Format [205]. Formats [1690]. Forms [1800]. Forte [447, 649]. Forum [2287]. Forward [274, 539, 563, 1064]. Forward-Looking [563]. fossil [727]. Fostering [1002]. Fosters [170]. Foundation [166, 467, 849, 1034, 2568]. Founder [99]. Foundries [348]. Four [264, 2054, 2245]. Four-Year [2245]. Fourth [940]. Fourth-Generation [940]. Foxes [296].

FPGA [807, 1109, 1964, 1966, 2328].

FPGA-Based [1964]. FPGAs [778, 1873]. Fractal [1973]. Fragility [1091]. Fragmented [171]. FrameMaker [672]. Frames [1466]. Framework [80, 137, 187, 406, 427, 577, 597, 640, 869, 1167, 1315, 1391, 1963, 2436]. Frameworks [217, 429, 432, 1376]. Fraud [1293, 2061]. Fray [673, 701]. Free [1, 109, 122, 267, 467, 662, 726, 807, 874, 1112, 2392, 2401]. FreeBSD [187]. Freedom [140, 701]. Freeman [99]. Fresh [1017]. Friend [1730]. Friendly [1778]. Frontier [216, 1426]. Frontiers [281, 1557]. Fuel [1042]. Full [575, 1632, 2148]. Fully [21]. Fun [1134, 2225]. Function [530]. Function-Class [530]. Fundamental [107]. Fundamentals [102, 852]. Funding [357, 893]. Funds [893]. Funny [126, 149]. Future [10, 11, 20, 36, 86, 87, 109, 110, 127, 215, 408, 414, 435, 493, 494, 542, 658, 711, 792, 813, 840, 942, 978, 989, 1104, 1225, 1288, 1357, 1428, 1450, 1538, 1621, 1684, 1694, 1774, 1860, 1879, 1909, 1936, 1950, 2019, 2049, 2209, 2277, 2357, 2435, 2473, 2522, 2559]. Fuzzy [896].

G [407, 461, 534, 694, 718, 1106, 1270, 2140]. G3 [228]. Gain [127, 497]. Gaining [114, 434]. Gains [8, 906, 1555]. Game [130, 271, 422, 538, 604, 610, 1080, 1175, 1397, 1438, 1505, 1741, 1743, 1744, 1994, 2531, 2602]. Game-Console [2602]. Gamers [232, 1080]. Games [206, 230, 490, 952, 1536, 1742, 1753, 1901, 2087, 2313, 2587]. Gaming [243, 449, 719, 818, 1593, 1785, 1872]. Gap [1013, 1294, 2316]. gaps [1015]. Gardens [2402]. Garp [97]. Gates [188, 766]. Gateway [753]. Gateways [1865, 2412]. Gathering [186]. Gator [1410]. Gauges [49]. Gave [132]. Gears [1267]. Gender [648, 1080, 1870]. Gene [937]. General [1167, 1525]. General-Purpose [1525]. Generated [727, 1517, 2446]. Generates [279]. Generation [89, 112, 180, 236, 237, 335, 363, 369, 534, 667, 940, 941, 1037, 1046, 1083, 1741, 2097, 2132, 2291, 2369, 2372, 2552]. Generator [605]. Generic [72, 137, 596, 1517]. Generics [2610]. Genetic [245, 2437]. GENI [1824]. Genius [127, 1528]. Genome [686]. genomes [689]. Genre [1526]. Geolocation [317]. Geometric [873]. Geophysical [2259]. George [1622]. Georgie [2320]. Geospatial [2414]. Gerald [485]. Gesture [1290, 2701]. Gesture-Recognition [1290]. Get [8, 192, 209, 276, 551, 883, 1334, 1716, 1833, 2368]. Gets [1095]. Getting [204, 254, 806, 1044, 1980, 2134]. GHz [520, 985]. Giants [1766]. Gigabit [658, 1007, 1208]. gInstall [915]. Give [1140]. Gives [112, 2568]. Glass [1463, 2130]. Glen [308]. Global [210, 565, 707, 879, 929, 1014, 1766, 2303, 2682]. Globalization [1774]. Globally [2252]. Glue [1184]. glut [706]. Go [614, 999, 1093, 1469, 1996]. Goals [357, 815, 2083, 2465]. Goes [108, 344, 414, 1511, 1668]. Going [79, 841, 1146]. Good [2, 29, 127, 389, 664, 1463, 1537, 1904, 2154]. Goode [331, 485, 669]. Google [1660]. Googlewhacking [614]. Got [571, 614]. Governance [1005, 1819]. Government [149, 302, 303, 305, 306, 1602, 1603, 1605, 1607, 1874, 2554]. Governors [464, 691, 893, 982, 1894, 2462]. GPS [317, 628]. GPU [1036, 1062]. GPUs [1949]. Graduates [570]. Grand [1708, 1840, 1842, 1885, 2307]. Granddaddy [86]. Graphical [257, 327, 682, 963, 2072]. Graphics [103, 194, 519, 605, 622, 635, 649, 758, 806, 873, 940, 986, 1102, 1306, 1525, 1533, 1799, 2386, 2538, 2586]. Graphics-Chip [2386]. Graphics-Hardware [2538]. Graphing [426]. Gratitude [1878]. Graveyard [2460]. Great [86, 388, 467, 700, 702, 743, 790, 1758]. Great-Granddaddy [86]. Greed [86, 837].

Green [384, 510, 2464, 2511, 2550, 2595]. Green500 [2169]. Greener [2550]. Grid [665, 1430, 2348, 2413, 2457, 2476]. Grids [823]. Grind [48]. Grooving [20]. Ground [99, 192, 434, 861, 1887, 2299]. Group [130, 406, 952, 1124, 2368]. Groups [322, 1111]. Grove [99]. Grow [592]. Grows [1624, 2620]. Growth [22, 164, 1267, 1926, 2517]. Guest [36, 134, 215, 374, 439, 459, 479, 685, 768, 843, 887, 930, 955, 998, 1024, 1069, 1120, 1205, 1249, 1317, 1448, 1493, 1514, 1602, 1648, 1694, 1741, 1787, 1812, 1961, 2008, 2097, 2120, 2390, 2411, 2433, 2497, 2542, 2563, 2584]. GUI [720]. Guide [440, 806, 852, 885, 941, 1008, 1107]. Guided [1697]. Guidelines [48, 860, 916]. Guilt [431]. Guises [2616].

H [485, 1095]. H-1B [1095]. Habits [130]. Hacker [344, 806]. Hackers [370, 604, 1426, 1490, 1716]. Hacking [710, 763, 1177]. Haeberer [893]. Half [777, 1826]. Half-Million [1826]. Hand [8, 342, 344, 1781]. Handbook [941, 1107]. Handheld [8, 368, 414, 571, 998, 1002, 1003, 1339, 1473]. Handheld-to-Handheld [414]. Handhelds [231, 232, 583, 817, 999, 2345]. Handle [167]. Handling [1034]. Hands [897]. HandTalk [2336]. Handwriting [715, 2100]. Hansen [669]. Happen [1510]. Happened [173]. Haptic [897, 1044]. Haptics [719, 747]. Hard [270, 1066, 1445, 1958, 2043]. Hard-Disk [2043]. Hard-Line [1958]. Hardware [23, 224, 426, 735, 806, 886, 888, 915, 1022, 1304, 1390, 1858, 1962, 1965, 2111, 2262, 2486, 2538, 2552]. Hardware-in-the-Loop [915]. Hardware/Software [888, 1390, 2486]. Hardwire [392]. Harm [664]. Harmonization [2064]. Harnessing [2193]. Harold [99, 401]. Harry [485, 2320]. Haul [2493]. HAVi [198]. HCI [1870]. HD [1852, 2430]. HDL [103]. Headaches [452].

Heal [705, 841]. Healing [179]. Health [5, 1728, 1862, 2284]. Healthcare [526, 1085, 1694, 1984, 2052]. Healthwear [1188]. Hear [746, 1200]. Heart [172]. Heat [758]. Heats [2430]. Heavy [103]. Hedgehog [941]. Hedgehogs [296]. Helix [874]. Hell [1155]. Helm [186]. Help [5, 103, 143, 152, 187, 341, 558, 726, 864, 996, 1021, 1067, 1117, 2581]. Helps [130]. Hennage [532]. Hennessy [120, 669]. Herbert [99]. Here [1578, 2357, 2364]. Hero [2002]. Heterogeneous [1581]. Hewlett [308]. Hidden [471, 1048]. Hiding [57]. Hierarchical [690]. Hierarchies [2692]. High [103, 199, 204, 220, 224, 278, 308, 357, 386, 423, 475, 666, 712, 715, 817, 937, 953, 980, 1008, 1011, 1067, 1143, 1205, 1207, 1228, 1283, 1291, 1419, 1582, 1695, 1888, 1961, 1962, 1964, 2218, 2262, 2347, 2348, 2350, 2352, 2386, 2668]. High-Assurance [2347, 2348, 2350, 2352]. High-Confidence [1695]. High-End [224]. High-Level [204, 386, 980, 1962]. High-Performance [199, 308, 475, 666, 937, 1011, 1582, 1961, 2218, 2262]. High-Powered [817]. High-Rise [1291]. High-Speed [103, 715, 1205, 1207, 1228, 1283, 1888]. High-Stakes [2386]. High-Tech [278, 953, 1008, 1067]. Higher [1475, 1928]. Highest [1424]. Highlights [893, 2450, 2469, 2489, 2555, 2621, 2642, 2664, 2684]. Hijacks [2368]. Hills [384, 510]. Hindsight [1330]. Hints [652]. Hist [475]. Historic [432]. History [68, 223, 295, 932, 941, 1001, 2204]. Hits [49, 212, 571]. Hobby [149]. Hoc [156, 397, 912, 1120, 1121, 1123–1126]. Hog [322]. Hold [254, 2493]. Holds [948, 996]. Hole [5, 1755]. Holistic [2012, 2704]. Hollywood [585]. Holmes [276]. Holograms [1102]. holographic [661]. Holography [498, 1517]. Home [198, 224, 244, 480, 482, 676, 702, 708, 723, 2011, 2271]. Homeland [651]. HomePlug [2314]. Homes [2650]. Honest [588]. Honor [99, 961, 2483, 2687]. Honored [240, 1746].

Honors [172, 669, 1077, 1769, 2396]. Hope [1713, 2247]. Hopes [927]. Horizon [2471]. Horizons [439, 740]. Horse [538]. Hosaka [1942]. Hospitals [1000]. Hot [279, 436]. Hour [2460, 2521]. House [1410, 2534]. Household [652]. How-To [2086]. HP [2626]. HPC [915]. HTML [59]. HTMT [475]. Human [127, 152, 328, 588, 1275, 1472, 1817, 1871, 1889, 1911, 2008, 2010, 2214, 2283, 2403]. Human-Centered [2008, 2214]. Human-Computer [1275, 2283]. Humans [389, 2705]. Humble [588]. Humility [431]. Humming [885]. Hunting [1053]. Hurdle [453]. Hurdles [1364]. Hurt [152]. Hurts [130, 637]. Hybrid [530, 1109]. Hydrologic [2436]. Hype [569, 632, 909, 2247]. Hyperformix [447]. Hyperlinking [441].

I-Logix [361]. I/O [391, 950]. IA [69]. IA-64 [69]. IAB [2582]. Ian [669]. Ibarra [485]. IBM [483, 487, 548, 571, 767, 864, 937]. IC [165]. ICDE [532]. Icebergs [1160]. iCompression [122]. ICT [285, 2302, 2306, 2595]. ICT4D [2302]. IDCs [1540]. IDE [49, 335, 447, 467, 605, 627, 694, 715, 736, 778, 874, 884, 1033, 1106, 1489]. Idea [2, 29, 2449]. Ideas [107, 1463, 2644]. Identification [36, 39, 599]. Identify [2647]. Identifying [428, 2511, 2694]. Identity [109, 1616, 2359, 2542, 2543, 2545, 2546]. Identity-Based [2359]. IDL [122]. IEEE [99, 201, 240, 251, 308, 357, 364, 423, 444, 465, 532, 579, 669, 691, 712, 732, 733, 765, 785, 801, 849, 893, 982, 1004, 1005, 1030, 1076, 1328, 1541, 1544, 1591, 1654, 1729, 1818, 1819, 1836, 1872, 2103, 2104, 2232, 2353, 2375, 2507, 2568, 2580, 2611, 2633]. IETF [927]. If [132, 633]. IFIP [893]. Ignites [520]. Ignorance [1060]. IGSTK [1697]. IHMC [284]. III [244]. IKIWISI [167]. Iliffe [331]. ILOG [558]. ILP [574]. IM [753]. Image [95, 234, 361, 671, 1007, 1697, 1788, 2051, 2261]. Image-Guided [1697]. Images [420, 1067, 1095]. Imagination [1735]. Imaging [688, 963, 1027, 1790]. Imagining [1813]. Immersidata [2054]. Immersion [279, 2053]. Immersive [1008, 1106]. iMouse [2033]. Impact [5, 399, 534, 1926, 2201, 2237, 2412, 2553, 2566, 2595]. Impasse [1429]. Imperceptible [1217]. Impetus [290]. Implementation [1107, 1204]. Implementing [197, 442, 542, 964, 1107, 2686]. Implications [1861]. Implicit [2074, 2619]. Importance [728, 1890]. Important [116, 660]. Improve [109, 144, 160, 455, 747, 755, 885, 908, 1785, 2345, 2494]. Improved [1624]. Improvement [27, 916, 1119, 2119]. Improvements [81, 682]. Improving [186, 223, 243, 414, 717, 832, 1572, 2139, 2284]. Improvisation [2099]. IMS [1577]. IMSI [1106]. In-Depth [896, 941]. In-Flight [225]. In-House [2534]. Inaugural [43]. Inboxes [1558]. Incarnations [2073]. Incentive [2284]. Incentive-Based [2284]. Incentives [171]. InCert [447]. Inclusive [1080]. Incompatible [1690]. Inconsistency [91]. Incorporating [1999]. Increased [601, 893]. Increases [357]. Increasing [2018, 2533]. Incremental [932]. Index [18, 47, 78, 263, 310, 404, 531, 804, 1075, 1343, 1613, 1896, 2174, 2439, 2706]. India [1369]. Indian [1766]. Indispensable [806]. Individualized [1862]. Industry [6, 30, 52, 55, 57, 86, 110, 128, 150, 173, 194, 211, 231, 252, 270, 277, 297, 319, 320, 342, 368, 390, 412, 473, 518, 572, 680, 705, 725, 745, 765, 792, 817, 839, 842, 862, 883, 906, 925, 941, 950, 973, 994, 1017, 1019, 1042, 1065, 1093, 1116, 1369, 1444, 1765, 1926, 2209, 2323, 2366, 2386, 2407, 2429, 2473, 2493, 2503, 2517, 2538, 2559, 2636]. Industry-Academic [2503]. Inevitability [2703]. Inevitable [2072]. Inflation [837]. Inflection [25, 26, 53, 54]. Influence [44, 86, 271, 1160, 1418]. Inform [1618]. Informatics [1842]. Information [11, 80, 178, 181, 259, 271, 417, 616, 654, 707, 727, 753, 779, 800, 991, 1014, 1039, 1058, 1190, 1198, 1559, 1605– 1607, 1673, 1925, 2282, 2290, 2301, 2330, 2497– 2500, 2502, 2566]. Information-Intensive [178]. Information-Processing [2282]. Information-Seeking [2497, 2498, 2502]. Information-Sharing [417]. Informix [267]. Infragistics [940]. Infrastructure [58, 179, 210, 261, 306, 576, 1164, 1184, 1231, 1295, 1453, 1516, 1933, 2317, 2416, 2518, 2693]. Infrastructures [2457]. Ingredient [5]. Inhomogeneous [2330]. initial [985]. Initiative [279, 322]. Ink [2099, 2101]. Inner [1814]. Inner-City [1814]. Innovation [105, 164, 220, 229, 470, 788, 881, 964, 1643, 2379, 2510, 2639, 2680]. Innovative [240, 286, 314, 2337]. Inprise [244]. Inprise/Borland [244]. Input [2100]. Ins [950, 1080]. Insecure [674, 1227]. Insightful [654]. Inspectable [2653]. Inspection [2257]. Inspections [160]. Inspector [1]. Inspired [2529]. InstallAnywhere [694, 1106]. Installation [205, 487, 829, 2056]. Installations [80]. Installer [291]. Installs [103]. InstallShield [291, 487, 778, 829]. Instant [130, 320, 571, 862, 904, 1019, 1490]. Instantiation [335]. Instantiations [583]. Institute [546]. Instruction [35, 352, 601, 1494, 1988]. Instruction-Level [352]. Instruments [487]. Integrated [23, 82, 125, 168, 247, 294, 337, 384, 386, 430, 510, 539, 672, 754, 774, 778, 889, 1034, 1865, 2033, 2191, 2436]. Integrates [228]. Integrating [912, 916, 1407, 1571, 1743]. Integration [122, 384, 665, 694, 753, 916, 1028, 1604, 1984, 2021]. Integrity [169, 807]. Intel [232, 299, 520, 820, 982, 1021, 1033, 1451, 1769, 2277, 2504]. Intellectual [30, 114, 329, 956, 986]. Intellectual-Property [30]. Intelligence [312, 680, 770, 772, 775, 776, 909, 1008, 1053, 1060, 1359, 1808, 1828, 1998, 2108]. Intelligent [204, 227, 559, 736, 778, 873, 943, 964, 1560, 1882, 2012]. Intelligentsia [916].

IntelliJ [874]. Intensive [178, 1842, 2258, 2262, 2456]. Interacting [1709]. Interaction [749, 1275, 1340, 2056, 2283, 2507, 2619, 2701]. Interactions [1571]. Interactive [62, 143, 182, 267, 598, 622, 649, 699, 771, 901, 1044, 1095, 1102, 1674, 2011, 2099, 2101]. Interactively [690]. Interactivity [1131]. Interconnect [2672]. Interconnection [1428, 2374]. Interdomain [777]. Interest [171]. Interesting [1596, 2160]. Interface [321, 328, 411, 583, 596, 852, 1333, 1390, 2393, 2394]. Interfaces [189, 326–328, 462, 798, 852, 1010, 1131, 1672, 1705, 1811, 1911, 2069, 2395, 2401, 2604]. Interfacing [290]. Interim [8]. Internal [2043]. International [165, 357, 406, 532, 1219, 2248, 2302]. Internet [8, 22, 42, 51, 80, 81, 88, 110–112, 124, 152, 155, 165, 168, 169, 174, 180, 191, 208, 224, 245, 254, 292, 322, 336, 361, 374, 377, 385, 398, 412, 428, 436, 452, 455, 511, 518, 610, 614, 637, 699, 738, 749, 759, 831, 840, 844, 874, 911, 975, 1027, 1049, 1189, 1205, 1228, 1281, 1293, 1317, 1333, 1335, 1386, 1429, 1549, 1915, 2187, 2215, 2343, 2368, 2423, 2518, 2562, 2581, 2603]. Internet-Addressing [2368]. Internet-Based [165]. Internet-challenged [80]. Internet-Derived [1333]. Internet-Driver [511]. Internet-Traffic [455]. Internet2 [520]. InterNetwork [406]. Interns [2679]. Interoperability [1024, 1140, 1336, 1638, 1718]. Interoperable [2542, 2545]. Interpolation [735]. Interruptions [1875]. Interspace [550]. Interview [106]. Interviewing [916]. Intervoice [1054]. Intranet [874, 1028]. Intro [144]. Introduces [384, 406, 426, 605, 649, 851, 985]. Introducing [364, 935, 2287, 2582]. Introduction [36, 38, 134, 215, 374, 439, 459, 479, 685, 768, 843, 887, 930, 955, 986, 998, 1024, 1069, 1120, 1205, 1249, 1317, 1448, 1493, 1514, 1602, 1628, 1648, 1694, 1741, 1787, 1812, 1836, 1961, 2008, 2097, 2120, 2390, 2411, 2433, 2497, 2542, 2563, 2584]. Introductory [2318]. Introspection [2464]. Intruders [766]. Intrusion [2177]. Intrusions [767]. Intuitive [291]. Invasive [681, 723]. Inventing [1187]. Invention [411]. Inventor [401]. Investigating [312]. Investigative [896]. Invests [357]. Invisible [247, 855, 901, 944, 988, 1038, 1084, 1816, 2401, 2446, 2485, 2530, 2571]. Involvement [357]. Iomegas [649]. IONA [426]. Iopsis [447]. IP [348, 644, 754, 842, 881, 912, 921, 964, 1140]. IP-based [921]. iPhones [2040]. iPlanet [361]. IPsec [1208]. IPv6 [279, 434, 510, 840, 2477]. Iris [40]. Irish [2308]. Iron [362]. IronGrid [940, 1007]. Isaak [2440]. Isolate [1939]. Isolation [2032]. ISPs [21]. Issue [1015, 1836, 1838]. Issues [20, 201, 238, 407, 632, 686, 1005, 1228, 1819, 2200, 2361]. Italian [303]. Iterative [932, 1537]. Itsy [346]. IV&V [86]. IVR [963].

J [702]. J2EE [694, 829, 1007, 1376]. J2EE-Compatible [694]. J2ME [649, 1935]. Jabberwocky [290]. Jack [240]. JafSoft [122]. JAR [205]. Jargon [860]. Java [1, 32, 48, 80, 92, 135, 149, 175, 267, 291, 312, 335, 384, 406, 414, 426, 447, 467, 510, 583, 605, 649, 654, 657, 672, 736, 743, 753, 777, 778, 807, 829, 851, 874, 884, 885, 915, 940, 964, 1007, 1008, 1011, 1053, 1054, 1079, 1163, 1489]. Java-Based [406]. Java-Compliant [335]. Java/C [384]. JBuilder [244]. JCanvas [335]. JDBC [1007]. JDO [605]. JetBrains [874]. Jini [147, 708]. Jn [224]. Jn-Software [224]. Job [79, 436, 605]. Jobs [741, 788]. Johan [755]. John [331, 712, 2230]. Johnny [272, 296]. Jolly [2230]. Joseph [240]. Journals [1030]. Joy [1880]. JProbe [487]. Judging [1488]. Juggernaut [312]. Jump [232]. Jumps [1335]. Junction [122].

Jungle [1802]. Juniper [777]. junk [975]. Just [64, 320, 709, 1019, 1202, 1558]. Just-in-Time [64]. Justice [594]. JVM [915].

Kanai [485, 937, 2612]. Kasturi [1894]. Kaydara [694]. KDE [267]. Keep [268, 296, 433, 885]. Keepin [433]. Keeping [20, 21, 117, 296, 588]. Kemma [143]. kernel [1033]. kernel-level [1033]. Key [89, 164, 201, 236, 252, 331, 593, 918, 1077, 1132, 1592, 1676, 2317]. Key-Exchange [1676]. Key-Management [2317]. Keyframe [462]. Keyframe-Based [462]. Keying [2272]. keys [957]. Keyword [2501]. Kid [186]. Kids [1606]. Kilby [240]. Killer [279, 2112]. Killers [2158]. Kinks [204]. Kit [1, 122, 244, 313, 406, 426, 510, 534, 753, 778, 807, 2486]. Know [79, 148, 212, 541, 1750, 2274]. Knowing [123]. Knowledge [5, 116, 297, 329, 463, 578, 582, 620, 640, 773, 784, 1039, 1378, 1765, 2213, 2460, 2501, 2573]. Knowledge-Based [1378]. Known [2364, 2384, 2405, 2427, 2452, 2471, 2491, 2515, 2536, 2557, 2578]. Koala [73, 74]. Kristen [755]. Kudos [904]. Kunii [99]. KWIC [295].

L [99, 165, 485]. L-3 [165]. L. [99]. Lab [1461, 1547, 1802, 2596]. Labor [815]. Lalit [99]. LAMP [1534]. LAN [672, 819]. Land [1666, 2171]. Landscape [685, 2120]. Language [102, 406, 447, 453, 735, 771, 811, 980, 1043, 1560, 1831, 1962, 2010, 2177, 2434, 2631, 2690]. Languages [214, 249, 1157, 1932]. LANscape [211]. Laptop [1997]. Large [123, 400, 618, 1338, 1709, 2102, 2238, 2261, 2673]. Large-Scale [123, 400, 2102, 2238, 2673]. Large-Vocabulary [618]. Laser [370]. Last [126, 151, 1739]. Lasts [1837]. Latency [355]. Later [1627]. Latest [605]. Laugh [126].

Launches [344, 384, 605, 715, 897, 915, 1033, 2310]. Law [22, 278, 475, 611, 654, 739, 820, 1073, 2327, 2432, 2620]. Laws [226]. Lawson [401]. Lay [1666]. Layer [1054, 1706, 1940]. Layered [1631]. Layering [1123]. Laying [1034]. Layton [649]. Lead [579, 1864]. Leader [230]. Leaders [43, 331]. Leadership [250, 2451]. Leading [669]. Leads [915]. Leakage [1073, 2367]. Leaks [79]. Lean [986]. Leap [20]. Leaps [1739]. Learn [2074]. Learned [999, 1843, 2110, 2434]. Learning [336, 712, 716, 774, 871, 879, 1029, 1110, 1472, 1473, 1891, 2099, 2605, 2617]. Least [1826]. Leave [99, 905]. Lecture [139, 755]. LegacyJ [335]. Legal [584]. Legislation [41]. Lengths [1093]. Less [2182]. Lesson [295, 941]. Lessons [707, 929, 999, 1115, 1532, 2110, 2389, 2434]. Let [139, 746, 1200, 2447]. Lets [747]. Letters [5, 27, 56, 86, 109, 127, 149, 171, 193, 210, 230, 249, 276, 296, 319, 341, 366, 389, 411, 432, 452, 472, 495, 515, 543, 565, 588, 610, 632, 654, 701, 723, 743, 763, 788, 815, 837, 860, 881, 904, 948, 970, 1017, 1062, 1088, 1114, 1136, 1159, 1180, 1221, 1242, 1287, 1310, 1332, 1356, 1380, 1401, 1422, 1465, 1486, 1530, 1552, 1574, 1595, 1620, 1665, 1712, 1734, 1757, 1780, 1805, 1830, 1877, 1906, 1955, 1977, 2001, 2090, 2114, 2136, 2206, 2229, 2251, 2295, 2383, 2425, 2470, 2490, 2514, 2556, 2599, 2622, 2643, 2665, 2685]. Level [12, 204, 352, 386, 426, 484, 505, 921, 980, 1033, 1339, 1389, 1861, 1962, 1999, 2181, 2466, 2631]. Leverage [1424]. Leveraging [4, 91, 1432]. Liable [674, 1227]. Liberated [2190]. Liberties [1178]. Libraries [779]. Library [291, 406, 669, 963, 1033, 1563, 2212, 2392]. Licensing [637]. Lies [493, 1468]. Life [20, 105, 159, 677, 684, 2148, 2170, 2372, 2486, 2571]. Life-Cycle [2148, 2486]. Lifestyle [127]. Lifetime [331]. Light [455, 682, 1790]. Lights [1534]. Lightweight [597, 1857]. Like [79]. Likely [744]. Limit [225].

Lindows.com [1033]. Line [20, 735, 1650, 1958]. Lineo [313]. Lines [2268]. Linking [1839]. Links [102]. Linux [80, 112, 164, 187, 204, 212, 267, 361, 384, 390, 406, 426, 447, 538, 548, 583, 585, 610, 694, 778, 782, 851, 897, 985, 1033, 1054]. Lips [1021]. List [293, 387, 1364, 2169]. Little [659, 1107]. Live [1161]. Liver [392]. Living [1661, 2132, 2256]. Load [361, 406, 649, 914, 1088]. Load-Scheduling [1088]. Load-Testing [361, 649]. Local [950, 1386]. Locality [2480]. Localization [1252]. Locally [2252]. Location [397, 414, 439, 440, 628, 750, 1084, 1130, 1305, 2473]. Location-Aware [439, 750, 1130]. Location-Tracking [628]. Locations [441]. Log [1411]. Logic [205, 806, 1053, 2280, 2282]. Logix [361]. Long [240, 279, 396, 2230, 2493]. Long-Term [396]. Longer [320, 2620, 2689]. Look [80, 826, 941, 1021, 1034, 1051, 1238, 2187, 2474]. Looking [539, 563, 813, 1064, 1357, 1463, 1621]. Looks [3]. Loop [915]. Lost [2129, 2420]. Lot [344]. Lotus [175, 331, 467]. Love [130, 1831]. Low [71, 156, 351, 817, 1033, 2315]. Low-Cost [1033, 2315]. Low-Power [351, 817]. Lowdown [1291]. LRU [1168]. Ltd. [291]. Luddism [543]. Lunch [109]. LVDS [778]. LynuxWorks [583].

M [99, 223, 269, 485]. M-Commerce [269]. M-CORE [223]. Mac [244]. Machine [223, 794, 816, 941, 1029, 1267, 1359, 1450, 1452, 1555, 1911, 2541, 2605]. Machine-Learning [1029]. Machine-to-Machine [1267]. Machines [389, 559, 844, 1449]. Macraigor [715]. Macromedia [672, 694, 753]. Made [161, 249, 381, 2426, 2686]. Madrid [2449]. Magazines [507, 1051, 1586, 1868, 2418, 2450, 2469, 2489, 2555, 2621, 2642, 2664, 2684]. Magic [361, 1304, 2098]. Magnetic [175]. Magnets [1067]. Mail [7, 130, 414, 975, 1095, 1578, 1792, 1936, 2589].

Mailbox [731]. Mailbox-Based [731]. Mailing [2115]. Mainstream [150, 617, 1036, 1668]. Maintaining [1358]. Maintenance [1941, 2532]. Make [115, 287, 519, 1345, 1599, 1600, 1720, 1882, 2254]. Makers [392, 841, 1444, 2602]. Makes [298, 2322, 2519]. Making [171, 262, 322, 482, 754, 873, 952, 1034, 1044, 1134, 1147, 1208, 1512, 1642, 1681, 1737, 1778, 1808, 1816, 1851, 2182, 2236, 2338, 2567]. Malicious [253, 571, 1361]. Malware [2276]. Mambo [702]. Man [816, 1112, 1862, 2618]. Manage [322, 2652]. Managed [2692]. Management [1, 24, 83, 102, 103, 122, 167, 207, 224, 246, 267, 293, 297, 325, 338, 361, 387, 429, 446, 447, 534, 620, 649, 672, 715, 754, 773, 778, 807, 896, 964, 970, 993, 1053, 1070, 1080, 1082, 1132, 1259, 1321, 1540, 1616, 1938, 1985, 2019, 2147, 2317, 2457, 2478, 2520, 2542, 2543, 2546, 2640]. Manager [80, 963]. Managers [109, 164, 296, 2201]. Manages [2407]. Managing [5, 48, 169, 305, 438, 501, 630, 718, 878, 1154, 1662, 1936, 2084, 2679]. Manet [1767]. Manifesto [515]. Manufacturable [1519]. Many [32, 128, 877, 2432, 2616, 2691]. Many-Core [2432, 2691]. Map [1539]. Maple [447]. Mapping [62, 187]. Maps [357, 2181]. March [43, 302, 579, 849]. Marginalize [1308]. Mark [622]. Market [8, 436, 765, 883, 925, 926, 1080, 1645]. Marketing [993, 1008]. Marketplace [613, 1065, 1139, 1224, 1475, 2386, 2644]. Markets [512]. Marking [1939, 2557]. Markup [471, 515]. Mars [683, 1272, 2435]. Mash [2226]. Mashups [2551]. Masses [224, 636]. Massive [234, 2263, 2270]. Massively [899, 2371]. Masters [48]. Match [791]. Matching [735, 2133, 2257]. Mate [172]. Materials [240, 392, 2561]. Math [986]. Math-Based [986]. Mathematical [447, 986, 1777]. Mathematics [1686]. MathType [605]. MathWorks [291, 313, 963]. Matrix [1965, 2606]. Matter [249, 1147, 1483]. Maturing [171]. Maturity [916, 1985, 2420]. May [89, 192, 212, 254, 322, 331, 747, 1144]. McDowell [331]. McMaster [172]. MDA [1539]. MDD [406]. Me [838]. Mead [1077]. Mean [2094]. Meaning [2653]. Meaningful [1297]. Measure [158]. Measurement [165, 487, 578, 672, 807]. measures [952]. Measuring [154, 1985, 2421]. Mechanic [244]. Mechanics [589]. Mechanism [2279]. Mechanisms [644, 2562]. Medal [99, 1077]. Media [322, 459, 460, 463, 546, 648, 667, 715, 916, 1106, 2270, 2416]. Mediated [1028]. Medical [1188, 1695, 1696, 1698, 1699, 2269]. Meet [20, 174, 824, 1316, 1911]. Meeting [312, 454, 579, 635]. Meets [699, 1017, 1073, 1359, 1948]. Mega [855]. Mega-Utilities [855]. Meld [23]. Mellon [989]. Melting [2588]. Member [139]. Members [161, 172, 250, 517, 601, 801, 1076, 1289, 1508, 1894, 2375, 2440]. Membership [251, 287, 308, 357]. Memory [414, 574, 747, 996, 1088, 1624, 1891, 2280, 2692]. Memory-Based [1891]. MEMS [152, 753]. Menagerie [455]. Mentor [532]. Mercedes [56]. Mercury [649]. Merging [411, 727]. Merit [507]. Merwin [264, 331, 485, 507]. Mesh [1351, 2672]. Mesh/Torus [2672]. Message [3, 4, 55, 273, 274, 494, 516, 541, 542, 789, 790, 813, 814, 1064, 1089, 1090, 2178, 2426, 2451, 2582]. Messaging [130, 320, 571, 862, 904, 1019, 1490]. MetaCreation [1]. Metadata [1057, 1068]. Metalanguage [472]. Metamodeling [1650]. Metamorphic [1143]. Metamorphosis [144]. Metaphor [904]. MetaWare [426]. Metering [1164]. Method [331, 530]. Methodology [2346, 2654]. Methods [223, 307, 551, 673, 701, 933, 936, 970, 1029, 1056, 1308, 1627, 1699]. Metric [997]. Metrics [376, 532, 610, 1053, 2168]. Metropolis [889]. Metropolitan [1351]. Metropolitan-Scale [1351]. Metrowerk [627]. Metrowerks [143, 205, 778, 985, 1033, 1054]. MFC [49, 187]. Michael [99, 275, 485]. Michigan [414]. Micro [175, 1381]. Micro-37 [1381]. Microarchitecture [351]. Micropayment [1951]. Micropayments [1245]. Microprocessor [455]. Microscope [851]. Microsoft [224, 455, 558, 672, 871, 954, 1007]. Microsystems [361, 583]. Microtest [165, 224]. Middleware [80, 187, 330, 394, 432, 868, 1184, 2348, 2389]. Midlife [1262]. MigraTEC [406]. Migrating [936]. Migration [406]. Mile [151, 1739]. Millennium [3, 6, 15, 154, 274, 292]. Millimeter [282]. Million [1826]. Millions [130, 546]. Mills [669]. MIMO [1225]. Mind [653, 710]. Miniaturization [548]. Minimizing [1615]. Mining [307, 488, 582, 636, 746, 775, 896, 1034, 1045, 1167, 1986, 2194, 2594, 2605, 2609]. Minolta [267]. Minorities [2533]. Mirrors [1519, 2653]. Misguided [89]. Misleading [1447]. Miss [16, 381]. Missed [56]. Misses [611]. Mission [735, 893, 1099, 2290]. Mission-Critical [735]. Misspent [701]. Misunderstood [89]. Miura [2698]. Mix [745]. Mixed [231, 809, 852, 1100]. Mixed-Signal [231, 852]. Mixed-Technology [852]. MKS [534, 829]. MMX [67]. Mobile [21, 42, 59, 157, 164, 216, 244, 253, 330, 346, 361, 397, 398, 406, 444, 461, 467, 510, 518, 583, 586, 592, 642, 649, 657, 731, 747, 759, 808, 840, 841, 885, 912, 1000, 1033, 1042, 1054, 1123– 1126, 1133, 1184, 1195, 1231, 1232, 1350, 1426, 1439, 1511, 1600, 1673, 1691, 1705, 1940, 1987, 2033, 2077, 2124, 2131, 2164, 2180, 2226, 2276, 2277, 2299, 2318, 2358, 2408, 2446, 2473, 2559, 2701]. Mobile-Agent [42]. Mobile-Code [1232]. mobile-platform [1054]. Model [73, 74, 159, 218, 246, 361, 379, 406, 478, 505, 577, 847, 870, 881, 1053, 1459, 1482, 1648, 1649, 1651, 1652, 1801, 1972, 2235, 2240, 2330, 2437, 2486, 2523, 2638, 2649, 2662]. Model-Based [2523, 2649]. Model-Centered [505]. Model-Clash [246]. Model-Driven [1648, 1649, 1651, 1652, 2638, 2662]. Modeler [447, 649]. Modeling [137, 335, 426, 447, 487, 488, 576, 641, 852, 873, 1029, 1074, 1248, 1275, 1297, 1365, 1863, 2436, 2479, 2498, 2673]. Models [42, 527, 728, 916, 1109, 1871, 2168, 2420, 2531, 2648–2653]. Modem [244, 874]. Modern [401, 1034, 2200, 2411]. Modular [164]. Module [627, 715]. Modules [1999]. Molecular [12, 1080]. Molecule [1862]. Moment [1789]. Monash [1]. Money [126, 149, 1082, 2076]. Mongering [632]. Monitor [839]. Monitoring [372, 597, 736, 829, 1122, 1411, 2269, 2284]. Monitors [411, 963, 1450, 1452]. Monotonic [1296]. Months [120]. Monty [937]. Mood [312]. Moore [22, 820, 1073]. Moronic [1008]. Morph [226]. Most [249, 287, 744, 798, 806]. Mote [1254]. Motif [49]. Motion [2093, 2602]. Motion-Sensitive [2602]. MotionBuilder [694]. Motivation [1983]. Motivator [2076]. Motocoder [915]. Motorola [778, 874, 915]. Mount [694]. Move [21, 680, 1224, 1597, 2254]. Moves [253, 1858, 2093]. Moving [539, 657, 964, 1067, 1852, 2189, 2233]. MPEG [122, 637]. MPEG-2 [122]. MPLS [1183]. MpSoC [891]. MTA [355]. Much [637, 2053]. Multiagent [223]. Multicast [586, 777, 1126]. Multicasting [777]. Multicore [1444, 2032, 2094, 2260, 2327, 2421, 2691]. Multicores [2693]. Multidimensional [524]. Multiengineering [506]. Multifactor [2545]. Multigate [841]. MultiGen [649]. MultiGen-Paradigm [649]. Multilevel [896]. Multimedia [620, 682, 956, 1033, 1215, 1474, 1540, 2011, 2696]. Multimodal [39, 1275]. Multiplayer [2371]. Multiple [245, 856, 1183, 2094, 2498]. Multiprocessing [354]. Multiprocessor [847, 890, 1493, 1496]. Multiprocessors [574, 1581, 2692]. Multiscale [1863, 1864]. Multistandard [2124]. Multiterabit [363]. Multithreaded [353, 367]. Multithreading [355]. Multitier [1320]. Multiuser [899]. Multiway [2391]. Murphy [611, 654]. Museum [286, 498, 1683]. Museums [854]. Music [678, 864, 996, 1044, 1591, 1690, 2368, 2507, 2641]. Music-Editing [2368]. Musical [1029]. Musician [1997]. Musings [743]. Must [109]. Must-Read [109]. Mutation [626]. Mutations [245]. MX [672, 753]. My [144, 344, 838, 1021, 1907, 2681]. Mystery [632]. Myths [452].

’n [567]. Naked [702]. Nam [239]. Name [777, 1682, 2004]. Named [75, 99, 120, 161, 308, 331, 357, 669, 893, 1346, 1894, 1942, 2440]. Names [279, 849, 2368]. namespace [963]. Nanocomputing [978]. Nanoelectronics [281, 1913]. Nanoscale [1387, 2280]. Nanotech [546, 996]. Nanotechnogy [851]. Nanotechnology [1849]. NAREGI [2413]. Narrowly [128]. NASA [14, 436, 822, 1099, 1630, 2433]. National [161, 261, 487, 778, 1077, 2212]. National-Scale [261]. Nations [1101]. Native [65]. Natural [619, 1560, 2177, 2619]. Natural-Language [1560, 2177]. Nature [1973, 2161]. Navigating [1812]. Navigation [550, 1768, 1892]. NC [1508, 1783]. NCSU [2596]. Near [611, 1669]. Near-Field [1669]. Nears [341]. Necessary [171]. Necessity [632]. Need [50, 409, 1144, 2021]. Needs [452, 635, 1732, 2487]. Needy [1067]. Negative [5]. Negatives [830]. Negotiate [377]. Negotiation [2479]. Neighborhoods [1814]. Nerd [1727]. Nerve [344]. Net [8, 32, 89, 129]. NetAdvantage [940]. Netscout [963]. Network [1, 21, 32, 79, 80, 103, 118, 130, 165, 187, 224, 239, 267, 322, 406, 570, 649, 715, 718, 767, 864, 963, 1067, 1079, 1123, 1161, 1165, 1177, 1183, 1253, 1307, 1361, 1425, 1431, 1668, 1865, 2164, 2448, 2702].

Network-based [80]. Network-Security [570]. Networked [102, 501, 666, 996, 1259, 1935]. Networking [20, 58, 104, 145, 156, 175, 198, 244, 277, 370, 417, 708, 726, 883, 1814, 2224, 2581]. Networks [48, 125, 143, 157, 363, 370, 397, 407, 455, 546, 552, 593, 674, 715, 751, 754, 769, 777, 873, 912, 916, 956, 987, 1035, 1046, 1050, 1054, 1120– 1122, 1124, 1125, 1207, 1249– 1251, 1351, 1384, 1432, 1492, 1561, 1577, 1592, 1698, 1768, 1792, 1802, 2070, 2139, 2291, 2374, 2399, 2437, 2459, 2494, 2547]. NetworkView [187]. Neural [896, 1165]. Neuroscience [1359]. NeuVis [335]. Never [192]. Newest [224]. News [7, 8, 31, 32, 58, 59, 88, 89, 111, 112, 129, 130, 151, 152, 174, 175, 212, 232, 253, 254, 278, 279, 298, 299, 321, 322, 343, 344, 369, 370, 391, 392, 413, 414, 436, 455, 474, 475, 498, 520, 546, 547, 571, 593, 614, 637, 659– 661, 681, 682, 706, 726, 746, 747, 755, 766, 767, 794, 818, 819, 841, 864, 885, 907, 908, 926, 927, 952, 974, 975, 996, 1020, 1021, 1067, 1094, 1095, 1118, 1141, 1162, 1185, 1203, 1226, 1269, 1363, 1385, 1470, 2367, 2387, 2430, 2454, 2474, 2494, 2515, 2518, 2539, 2560, 2580]. News [519, 840, 863, 884, 951, 995, 1043, 1066, 1247, 1292, 1314, 1337, 1406, 1427, 1445, 1446, 1491, 1513, 1535, 1556, 1579, 1601, 1626, 1647, 1670, 1692, 1717, 1762, 1786, 1810, 1835, 1859, 1884, 1912, 1934, 1960, 1982, 2006, 2028, 2050, 2071, 2095, 2118, 2141, 2154, 2165, 2188, 2211, 2234, 2278, 2324, 2345, 2368, 2388, 2408, 2409, 2431, 2455, 2475, 2495, 2519, 2540, 2561, 2581]. NEXIQ [736]. Next [15, 25, 26, 89, 112, 180, 236, 237, 363, 369, 534, 547, 641, 642, 1037, 1046, 1083, 1339, 1426, 1741, 1787, 1981, 2097, 2291, 2299, 2302, 2369, 2631]. Next-Generation [89, 112, 180, 236, 237, 363, 369, 534, 1037, 1046, 1083, 2291, 2369]. Night [2623]. Nimble [1997]. Nine [1216]. NIST [2029, 2084]. No [69, 109, 320, 2339, 2689]. Nobel [27, 240]. Nobody [132]. Nodes [2552]. Noise [2358].

Nomadic [347]. Nominations [43, 579, 849, 1701]. Nominees [464]. Noninvasive [2391]. Nonlinear [534]. nonuniform [668, 735]. Norman [99]. Notable [68]. Note [229]. Notebook [1765]. Notes [331]. Notion [319]. Novadigm [534]. Novel [2013]. Novell [224, 649]. Novelty [2689]. Novice [205]. NQL [447]. number [702]. Numbers [777, 2654]. Numerical [406, 1460]. Numerics [406, 426, 753]. Numina [1473]. Nurturers [1561]. Nurturing [1083]. NuSphere [426]. Nvidia [851]. Nygaard [755].

O [391, 950]. O. [702]. Obelix [168].

Obituary [485]. Object

[2, 123, 134, 138, 173, 187]. Object-Oriented [173]. ObjectFX [291]. Objects [20, 668, 2617]. Obligation [318]. Obligations [1580]. Observations [654, 2007, 2023]. Observing [941]. Obvious [318]. OC [829]. OFDM [792]. Off [192, 279, 1335, 2232]. Offbeat [941]. Offer [331, 601, 646, 1067, 1095]. Offered [99]. Offers [8, 175, 291, 406, 601, 712, 779, 874, 963, 1007, 1106, 1370, 1632]. Office [224, 464, 694, 940, 1054]. Officers [517]. Offices [161]. Offline [662]. Offshore [1263]. Offshore-Resistant [1263]. Ohio [240]. Oklahoma [1769]. OLAP [2263]. Old [56, 102, 546]. Older [2009]. Ole [755]. Ole-Johan [755]. Olympic [206, 230]. Omnicore [406, 778, 915]. On-Chip [2494]. On-Demand [1482]. On-Line [735]. On-the-Fly [175]. Once [436, 2087]. Once-Hot [436]. OnCore [205, 487]. One [21, 650, 852]. Ongoing [302, 2510]. Online [252, 331, 385, 490, 601, 605, 662, 721, 779, 856, 885, 952, 1327, 2110, 2181, 2298, 2371, 2531, 2594, 2703]. Onloading [1319]. onto [1362]. Ontology [773, 774, 1028]. Ontology-Based [773]. Ontology-Mediated [1028]. OO [102, 527]. Open [16, 191, 227, 254, 287, 348, 395, 406, 467, 537, 563, 591, 672, 788, 958, 1106, 1224, 1238, 1420, 1460, 1496, 1697, 1728, 1752, 1838, 1983, 2031, 2142, 2209, 2392, 2476, 2539, 2659]. Open-Secret [1420]. Open-Source [406]. Open-World [1838]. OpenGL [558, 986, 1007]. Openness [409]. OpenOffice.org [672, 1054]. OpenPath [467]. Opens [30, 122]. Operating [165, 384, 1720]. Operation [2393]. Operational [754, 2370]. Opinion [1418]. Opinions [2594]. Opportunities [19, 46, 63, 77, 101, 142, 163, 185, 203, 222, 242, 266, 289, 311, 333, 359, 382, 403, 425, 509, 533, 557, 581, 603, 624, 647, 695, 714, 737, 757, 781, 805, 828, 853, 875, 895, 919, 939, 984, 1009, 1032, 1055, 1081, 1105, 1151, 1172, 1193, 1214, 1237, 1256, 1276, 1301, 1326, 1349, 1372, 1396, 1412, 1433, 1454, 1479, 1500, 1522, 1542, 1564, 1585, 1614, 1656, 1680, 1700, 1726, 1749, 1772, 1795, 1820, 1844, 1897, 1922, 1944, 1970, 1990, 2035, 2057, 2082, 2106, 2127, 2152, 2175, 2219]. Opportunities [75, 910, 1580, 2195, 2241, 2286, 2309, 2331, 2378, 2459, 2530]. Opportunity [591, 601]. Optic [165]. Optical [8, 58, 152, 392, 498, 571, 593, 615, 794, 883, 1021, 1044, 1067, 1102, 1384, 2055]. Optical-Networking [58]. Optics [754, 1738]. Optimistic [529]. Optimization [49, 896, 1389]. Optimizations [941, 2168]. Optimizing [528, 2607]. Option [308]. Options [411, 797, 881, 1675]. Oracle [267, 694, 807, 1008]. Oranges [48]. Orchestration [1026]. Orders [1443]. Organically [1080]. organisms [683]. Organization [232, 365, 599, 935, 1112, 2368, 2457]. Organizational [2686]. Organizations [1338]. Organizers [1677]. Orientation [2146]. Oriented [1, 103, 173, 243, 343, 672, 1298, 1320, 1685, 2144, 2147, 2149, 2347–2349, 2351, 2551, 2676]. Originality [2512, 2534]. OriginLab [426].

Origins [86, 432]. Oscar [485]. OSDL [694]. OSs [497, 925]. Other [691]. Our [55, 149, 516, 542, 677, 1642, 1732, 1879, 2067]. Ourselves [564]. Outcomes [2617]. Outdated [945]. Outperforming [1168]. Outposts [1956]. Outs [950, 1080]. Outsider [639]. Outsourcing [1080, 1547, 1629, 2163, 2201, 2540]. Outstanding [961]. Overclocking [1138]. Overcoming [1429]. Overheads [784]. Overlay [1126]. Overload [1936]. Overview [136, 395, 1249]. Own [604]. Oxymoron [591, 1684]. Oz [20]. Ozzie [331].

P2P [417, 455, 498]. PA [69]. PA-RISC [69]. Pacific [851]. Package [313]. Packet [363, 2257]. Packeteer [736, 874]. PACT [627]. Paessier [205]. Page [226, 2405]. Pages [1386, 2086]. Paid [1776]. Painting [1]. Pair [315, 1727]. Palm [267, 715]. PalmOS [8]. Palmtop [212]. Panda [1079]. Panda3D [1306]. Panoramic [1095]. Paper [622, 1067, 2098, 2256]. Papers [340]. Paradigm [268, 460, 552, 649, 865, 1156, 2283, 2456]. Paradigms [478, 2072]. Paradox [324, 2410]. Parallel [35, 103, 268, 627, 908, 1023, 1563, 2238, 2674]. Parallelism [319, 1494]. Parallelization [1900]. Parameter [2696]. Parameters [2259]. Parascript [715]. Parasitic [219]. ParaSoft [1, 143, 736, 1079, 1106]. Parents [747]. Parents-to-Be [747]. Park [314]. Parnas [2631]. Parsing [2370]. Part [1751, 1798]. Participants [240]. Participation [507, 601, 1548, 2018, 2443, 2533]. Partnership [779]. Partnerships [755]. Party [56]. Passages [1107]. Passé [2534]. Passive [1384, 2055, 2702]. PassMark [224]. Passport [954]. Passports [2129]. Password [406, 682, 2034]. Password-Based [2034]. Past [493, 813, 1357, 1621, 2435, 2443]. Paste [1764]. Pastures [2550]. Patch [103, 1216, 1260]. Patchwork [224]. Patent [193, 232]. patented [927]. Patents [61, 127, 152, 565, 745]. Path [30]. Paths [6, 158]. Patient [372, 1698]. Patnaik [99]. Pattern [99, 638, 735]. Pattern-Matching [735]. Patterns [144, 186, 428, 988, 1008, 2525, 2574]. Paving [1104]. Paying [1902]. Payments [808, 2217]. PC [56, 223, 267, 436, 496, 760, 846, 972, 985, 1033, 1106, 1765, 2097, 2100, 2102]. PCs [546]. PDAs [605]. PDF [649]. PE [2677]. Pecking [1443]. Pedagogical [2102]. Peer [277, 344, 1094, 2255, 2300]. Peer-to-Peer [277, 344, 1094, 2255, 2300]. Pencel [583]. Penny [981]. Pentium [244, 1628]. People [109, 249, 513, 609, 1045, 1378, 2009, 2018]. PeopleWeb [2078]. Perception [1887]. Perfect [243]. Perforce [851]. Performance [133, 154, 199, 224, 308, 341, 344, 475, 510, 528, 529, 573, 577, 583, 605, 666, 724, 891, 896, 937, 940, 963, 1007, 1011, 1033, 1497, 1582, 1961, 1964, 2096, 2218, 2262, 2370, 2421, 2654, 2673, 2694]. Performing [851]. perimeters [1079]. Period [2190]. Periodicals [220, 444]. Perl [244, 426]. Persistence [301, 335, 366, 605, 1106]. Persistent [490, 899]. Persistent-State [490]. Personal [40, 325, 560, 944, 963, 1095, 1215, 2663]. Personalizing [2075]. PersonalJava [143]. Perspective [160, 562, 740, 782, 832, 878, 920, 948, 966, 970, 1011, 2279, 2448]. Perspective-Based [160]. Perspectives [683, 684, 696, 727, 748, 842, 865, 886, 909, 929, 953, 976, 1022, 1058, 1096, 1097, 1110, 1983, 2380, 2410, 2432, 2465, 2477, 2509, 2520, 2553, 2596]. Perspicuous [902]. Pervasive [444, 536, 708, 829, 865, 944, 1034, 1085, 1410, 1815, 1816]. Pervasiveness [2605]. Petascale [1636, 2674]. Peter [106]. Petersburg [264]. Phase [2302]. PhDs [2527]. Phenomenon [411]. Phenomics [688]. Philosophy [791, 837]. Phishing [1469]. Phone [702, 925, 1691, 1833, 2209]. Phones [657, 927, 994, 1426, 1600, 2124, 2180, 2389]. Photography [637, 1787]. Phrenology [544]. Phylogenetics [687]. Physical [441, 1569, 1571, 1580, 2508]. Physics [240, 702, 1785, 1799]. Physiome [1863]. Piano [996]. Piano-to-Piano [996]. Picks [254]. PICO [729]. PicoRadio [156]. Picture [1399]. Picturing [819]. Pinch [2701]. Pinning [130]. PINs [2217]. Pioneer [99, 319, 401, 485, 1942]. Pioneering [319, 669]. PipeRench [98]. Piracy [392, 495, 955, 959, 996]. Piranha [354]. Pitches [595]. Pitfalls [720, 1652]. Pixels [21, 1518, 2051]. PKI [252, 709]. Place [342, 1187]. Places [2056]. Plagiarism [706]. Plagues [130]. Plan [357, 933, 1067, 1868, 2418]. Plan-Driven [933]. Plane [702, 948]. Planet [384]. Planning [600, 873, 1817, 1887]. Plans [164, 444]. Platform [175, 267, 291, 313, 337, 361, 384, 426, 510, 534, 575, 583, 605, 915, 958, 985, 1054, 1496, 1837, 2283, 2298]. Platform-as-a-Service [2298]. Platforms [143]. Play [21, 283, 1077, 1742, 2452, 2578]. Play-Centric [1742]. Play-In [283]. Player [1645]. Playing [2192]. PlayStation2 [49]. PLD [985]. Plot [2420]. Plug [21]. Plug-and-Play [21]. Podcasts [1848]. Poet [187]. Point [25, 26, 53, 54, 390, 860, 1015, 1066, 1290, 1424]. Pointer [262, 401]. Pointer-Based [262]. Points [411]. Poised [2517]. PolarLake [605]. Policy [44, 1232, 2147]. Policy-Based [1232]. Policy-Oriented [2147]. Polish [443]. Political [968]. Politics [13, 109, 2452]. PONs [2493]. Pool [2192]. Pool-Playing [2192]. Poopers [56]. Pop [1115]. Popkin [447]. Popular [130, 798, 1106]. Popularity [870, 1624]. popularity-based [870]. Porn [436].

Porn-Seeking [436]. Portability [268]. Portable [534, 2254]. Portal [80, 331, 2056]. Portals [558, 642]. Portfolio [807]. POS [189]. Positions [16, 287]. Positively [1503]. Positives [830]. Possibility [416]. Possible [1243]. Post [99]. Postsilicon [2329]. Potential [32]. Power [152, 156, 175, 350, 351, 659, 744, 817, 1069, 1072, 1073, 1253, 1259, 1321, 1391, 1582, 1617, 1754, 1799, 2640]. Power-Aware [1069, 1582]. Powered [21, 144, 817]. Powering [1933]. Powerline [2314]. PowerPC [165, 406]. PowerPoint [1240]. Powers [314, 2498]. PowerSteering [313]. Poznan [1244]. Practical [119, 123, 334, 553, 735, 777, 852, 1034, 1107, 1729, 2283]. Practice [735, 896, 936, 1034, 1985, 2237, 2583]. Practices [226, 707, 708, 728, 795, 910, 928, 954, 977, 997, 1023, 1045, 1046, 1068, 1111, 1155, 2030, 2102, 2389, 2476, 2496, 2521, 2541, 2562, 2583, 2593]. Practitioner [941]. Pradesh [929]. Pragmatism [2325]. Praise [1710, 2325]. Precautionary [2565]. Precision [1710]. Precollege [99]. Predestination [1996]. Predict [1296, 1941]. Predicting [86]. prediction [870]. Predictive [2240]. Preemptive [643]. prefetching [870]. Premia’a [187]. Prepare [43]. Presence [1019]. Present [369, 2435]. Presentation [204, 1532]. Presenter [2101]. Presenters [755]. Presents [240, 2591]. Preservation [324]. Preserving [1986]. President [4, 120, 161, 201, 274, 465, 494, 516, 542, 669, 733, 789, 801, 814, 1005, 1064, 1076, 1090, 1346, 1541, 1609, 1819, 1894, 2104, 2171, 2353, 2426, 2440, 2451, 2611, 2697]. President-Elect [201, 465, 733, 801, 1005, 1076, 1346, 1541, 1609, 1819, 1894, 2104, 2171, 2353, 2440, 2611, 2697]. Presidential [1077]. Presidents [1894]. Preventing [959, 2034]. Prevention [1]. Prevents [2367]. preview [1079]. Pricing [51, 81]. Pride [431]. Prime [297, 676, 701, 747, 2454]. Primer [735, 852, 1053, 2476]. primitives [1033]. Princeton [467]. Principle [2565]. Principles [243, 835, 873, 881, 896, 1824, 2360, 2631]. Print [1781]. Printed [327]. Printer [279]. Prioritized [1126]. Priority [2567]. PrismTech [605]. Prithviraj [485]. Privacy [44, 232, 372, 385, 454, 543, 701, 721, 955, 1012, 1035, 1084, 1752, 1986, 2345, 2564, 2627]. Privacy-Preserving [1986]. Private [127, 1189, 2164]. Private-to-Private [1189]. Prize [27, 172, 240, 331, 622, 712, 982, 1508, 1783]. Prizes [220, 2504]. Proactive [165]. Probability [27, 806]. Probe [510]. Prober [753]. Problem [946, 1088, 1308, 1716, 1719, 1951, 1953]. Problems [32, 96, 614, 806, 1140, 2496]. Process [27, 32, 177, 218, 308, 378, 411, 641, 720, 885, 916, 935, 1918, 2062, 2119, 2235, 2616]. Processes [109, 1174, 1196, 2148]. Processing [11, 95, 199, 268, 352, 400, 908, 1071, 1148, 1785, 2177, 2282, 2328, 2391, 2517]. Processor [59, 341, 627, 807, 1071, 1161, 1533, 2654]. Processors [165, 205, 349, 353, 367, 487, 571, 574, 592, 833, 979, 1283, 1444, 1495, 1525, 1873, 2032, 2260, 2494, 2696]. ProDA [2263]. Product [18, 47, 78, 143, 310, 404, 544, 916, 985, 1007, 1650, 2223, 2268, 2610]. Product-Focused [2223]. Product-Line [1650]. Production [1765, 2586]. Productivity [124, 335, 672]. Products [21, 49, 80, 103, 122, 143, 165, 187, 205, 224, 244, 267, 291, 313, 335, 361, 378, 384, 406, 426, 447, 467, 487, 510, 534, 558, 583, 605, 627, 649, 672, 694, 715, 736, 753, 778, 807, 829, 851, 874, 897, 915, 940, 963, 985, 1007, 1033, 1054, 1079, 1106, 1152, 1173, 1202, 1212, 1235, 1258, 1279, 1303, 1322, 1327, 1348, 1370, 1374, 1392, 1415, 1436, 1457, 1480, 1501, 1524, 1546, 1567, 1588, 1612, 1635, 1679, 1703, 1773, 1796, 1821, 1846, 1866, 1899, 1924, 1946, 1969, 1993, 2016, 2189, 2426]. Products’ [467]. Profession [166, 171, 193, 206, 248, 272, 295, 318, 339, 365, 388, 410, 431, 493, 514, 540, 543, 564, 565, 579, 587, 677, 700, 721, 741, 761, 786, 788, 811, 835, 858, 879, 902, 922, 946, 948, 968, 991, 1014, 1039, 1060, 1086, 1112, 1157, 1399, 1732, 1928, 2112, 2158, 2204, 2381, 2403, 2423, 2449, 2468, 2488, 2512, 2534, 2554, 2575, 2597]. Professional [44, 116, 140, 331, 534, 588, 723, 778, 1572, 1710, 2461]. Professionalism [622]. Professionals [941, 968, 1030, 2663]. Professor [264]. Professors [706]. Profile [2668]. Profiles [307]. Profiling [66, 143, 291, 829]. Profitable [537]. Profiting [448]. Program [109, 143, 158, 212, 272, 498, 622, 1263, 1744, 1841]. Programmable [979, 1410, 1431, 1483]. Programmers [89, 122, 128, 187, 205, 368, 1402, 2094, 2381]. Programming [79, 92, 102, 103, 138, 149, 205, 214, 257, 268, 291, 296, 315, 343, 426, 609, 701, 715, 851, 873, 902, 1109, 1134, 1143, 1727, 1843, 1932, 2072, 2190, 2293, 2318, 2325, 2534, 2690, 2692]. Programs [139, 186, 248, 555, 755, 826, 1008, 1104, 1316, 1632, 2630]. Progress [244, 2426, 2686]. Project [27, 80, 102, 122, 207, 304, 315, 446, 571, 703, 747, 1473, 1863, 2119, 2201]. Projectors [1366]. Projects [5, 600, 849, 934, 1067, 2238]. Prominence [906]. Promise [190, 659, 996, 1015, 1021, 1161, 2218, 2493]. Promises [8, 155, 475, 498, 520, 747, 1652]. Proof [299]. Propagating [1947]. Property [30, 956, 986, 1198]. Proponents [1021, 2255]. Proportional [2167]. Proposal [520, 615, 1021, 1044, 1293]. Proposed [81, 982, 1049, 1600]. Proprietary [1974]. Pros [651]. Prospects [1100, 2124]. Prosthetic [2393]. Protect [219, 615]. Protecting [372, 438, 480, 565, 956, 957, 960, 1012, 1398, 2070, 2317]. Protection [7, 30, 392, 481, 495, 648, 743, 996, 1084, 2562].

Proteins [2458]. Protocol [59, 112, 122, 152, 614, 1283]. Protocol-Related [614]. Protocols [529, 731, 2661]. Prototype [2315]. Prototypes [2530]. Provider [478]. Provides [331, 622]. Providing [2076]. Provision [1050]. ProxySource [335]. PTC [963]. Public [40, 127, 1132, 1418, 1592, 1764, 1816, 2056, 2199, 2567, 2627]. Public-Key [1592]. Publications [99, 691, 755, 826]. Publisher [2487]. Publishing [715, 1059, 1077, 1572]. Pulling [1765]. Pun [56]. punch [726]. Punches [164]. Purpose [1525, 1753]. Pursue [1384]. Push [115, 392, 819]. Pushing [21, 274]. Put [864]. Puts [103, 230, 862, 864, 2368]. Putting [693, 852, 948, 970]. Puzzling [2496]. Pyramid [873]. Python [267, 335].

Q [771]. Q3 [2288]. QNX [715]. QoS

[644, 1940, 2240, 2350].

QoS-Reconfigurable [2350]. Quagmire [429, 432]. Qualifies [5]. Qualifying [1740]. Quality [127, 377, 384, 526, 544, 578, 885, 1119, 2088]. Quantifying [1674]. Quantitative [777, 1209, 2259]. Quantum [11, 20, 290, 553, 589, 632, 654, 1368, 1631]. Quasireform [1803]. Quell [130]. QuERIES [2346]. Query [1918]. Querying [2261]. Quest [1054]. Question [5, 1460]. Questions [32, 86, 543, 654, 743, 1005, 1819]. Queuing [806, 847]. Quipunet [99]. Quo [1873]. Quotas [1333].

R [308, 2403]. R.C. [1622]. Rack [694]. Rack-Mount [694]. Rackmount [103]. RAD [583, 897]. Radiation [5, 1251]. Radio [209, 1417, 2124]. Radios [974]. RadView [313]. Rages [2386]. RAID [1007]. Rails [1646]. Rainbow [1295]. Raise [632]. Raises [232, 906]. Raising [1284]. RAM [175, 613, 996]. Ramamoorthy [485]. Random [223, 544, 655, 678, 689, 702, 724, 744, 764, 791, 816, 838, 861, 882, 905, 924, 949, 971, 993, 1018, 1041, 1063, 1091]. Range [89, 396]. Rapid [167, 335, 361, 426]. RapidStream [143]. Rasterization [103]. Rate [1296]. Rational [243, 267, 558, 605, 753]. Rationality [1285]. Raymond [331]. Raytheon [406]. Razor [1147]. Reach [637, 1174, 1342]. Reaches [1066]. Reaching [175, 1879]. Reactive [1365]. Read [109, 1021, 1575]. Reading [99, 160, 286, 555, 691, 893, 982, 2481]. Readings [161]. Ready [297, 319, 496, 551, 676, 701, 1313, 2454]. Real [20, 134–136, 138, 143, 313, 416, 467, 643, 649, 941, 1206, 1353, 1721, 1864, 2025, 2134, 2260, 2325, 2381, 2521]. Real-Life [20]. Real-Time [134–136, 138, 143, 313, 467, 643, 649, 1721, 1864, 2260]. Real-World [941]. Realistic [1365, 2481]. Reality [452, 596, 809, 876, 909, 1438, 1536, 1737, 1927, 2053, 2395, 2588, 2689]. Realization [2147]. Realizing [607]. Really [2, 29, 547, 2454]. Realm [1602]. Realms [2093]. RealNetworks [874]. Rearranging [1062]. Reasoning [123, 334]. Reasons [914]. Recall [2390]. ReCANcentrate [2547]. Receive [331, 893]. Receiver [204]. Receives [99, 240, 331, 532, 669, 712, 1077]. Recipes [671, 1661]. Recognition [37, 321, 617, 618, 715, 926, 1290, 1858, 2100, 2271]. Recognize [220, 485, 937, 2657]. Recognized [99, 240, 357, 669]. Recognizes [264, 308, 532, 669, 1030, 1608, 2172, 2441]. Recommendation [1038, 1715]. Recommender [2606]. Recompilation [69]. Reconfigurable [96, 98, 980, 1961, 1963, 1965, 2218, 2350]. Reconfiguration [1232, 2552]. Reconfiguring [994]. Reconnaissance [2702]. Record [851]. Record-Performing [851]. Recovery [1320]. Recovery-Oriented [1320]. Redefining [6, 788, 1788]. Reduce [451, 784]. Reducing [522, 566, 795, 1072, 2429]. Reduction [229, 293, 2358]. Redux [1397, 2529, 2573]. Reengineering [1915]. Refactoring [1902, 2480]. Reference [122, 366, 510, 778, 874, 1053, 1466]. References [301]. Refind [1559]. Reflect [55]. Reflection [1472]. Reflections [256]. Refocusing [1871]. Reform [206, 318, 341]. Regions [1471]. Registration [785]. Regulation [748]. Regulations [1607]. Rehabilitate [2255]. Rehabilitation [2394]. Reinforce [779]. Reinventing [21]. Reisman [2697]. Rejuvenate [1952]. Related [614]. Relational [1459]. relationship [825]. Relationships [2503, 2636]. Release [694]. Released [267]. Releases [143, 267, 361, 384, 406, 447, 467, 534, 583, 736, 829, 851, 874, 940, 963, 1007, 1054, 1079, 1106, 2288, 2368]. Relevant [2076]. Reliability [157, 258, 319, 806, 977, 1639]. Reliable [258, 460, 553, 1099, 1145, 1720, 2524]. Relighting [1791]. Remarks [515]. Reminder [357, 1732]. Remote [224, 372, 406, 2269]. Remote-Patient [372]. Removable [981]. Remove [1952]. Renaissance [170, 1448]. Renew [56]. Repair [2329]. Replace [1042, 2460]. Replacement [668, 1168]. Replacing [1974, 2129]. Replicate [1952]. Reply [2339]. Reply-to-All [2339]. Report [172, 183, 250, 517, 801, 1076, 1289, 1508, 2375, 2426, 2440]. Reporting [736]. Reports [224]. Representative [587]. Requests [392]. Requirements [160, 167, 186, 338, 867, 873, 904, 1504, 1851, 2532, 2632]. Requirements-Based [2532]. Requires [296]. Research [14, 99, 122, 148, 171, 175, 291, 301, 392, 532, 570, 605, 646, 731, 752, 820, 848, 869, 870, 881, 893, 913, 914, 941, 967, 980, 981, 1029, 1049, 1050, 1058, 1407, 1590, 1602, 1603, 2098, 2144, 2237, 2281, 2306, 2307, 2373, 2378, 2392, 2415, 2437, 2480, 2524, 2547, 2581, 2589]. Researcher [864, 1044]. Researchers [370, 392, 414, 726, 794, 885, 908, 1067, 1882, 2187, 2483, 2647]. Reservations [941]. Resilience [1388]. Resistance [475]. Resistant [1263]. Resolving [1951]. Resource [99, 361, 753, 779, 1448, 2044]. Resource-Constrained [2044]. Resources [5, 137, 331, 610, 718, 1057, 1071]. Respect [1555, 2597]. Respectable [115]. Respond [1819]. Responds [565]. Response [210, 355, 703, 2589]. Responsibly [811]. Resting [709]. Restructuring [1316]. Results [183, 690, 1197, 2076]. Retail [1671]. Rethinking [319, 761, 1101, 1452]. Retires [172]. Retrieval [236, 620, 2051, 2259]. Retrieving [420]. Retrograde [1023]. Retry [1952]. Returns [186]. Reusable [1295, 1693]. Reuse [243, 648, 2650]. Reusing [2170]. Revamping [21]. Reveals [436]. Reverse [540, 959, 996]. Reversing [2582]. Review [733, 789, 1699, 2444, 2583]. Reviewer [264, 532, 940]. Reviewers [802, 1077, 1345, 1608, 1893, 2172, 2441]. Reviewing [516]. Reviews [724, 940, 1018, 1062]. Revised [507, 601, 881, 1053, 2636]. Revises [406]. Revising [835]. Revisions [109]. Revisited [1097]. Revisiting [1592]. Revitalization [2185]. Revitalizing [854, 1462]. Revive [1116]. revoke [957]. Revolution [132, 193, 321, 899, 1096, 1640, 1659, 2008]. Revolutionize [637]. Revolutionized [127]. Revolutionizing [1799]. Rewards [423]. Rework [1537]. RFC [123]. RFID [1176, 1671, 1761, 2369]. Rich [334, 1554, 2343]. Richard [485]. Ride [790, 1646]. RidgeRun [426]. Right [79, 319, 543]. Rights [44, 842, 881, 1336, 2520]. Rigorous [149, 868]. RIM [649]. Rip [88]. Ripe [1827]. RISC [69]. Rise [871, 986, 1291, 1959]. Rises [1645]. Risk [229, 544, 610, 735, 837, 933, 1082, 1353, 2084]. Risk-Based [735]. Risks [259, 630, 1615]. Road [112, 341, 1539, 2437, 2550, 2580, 2696].

Road-Traffic [2696]. Roadmap [549, 941, 1098, 1204]. roaming [951]. Robbery [388]. Robert [1077]. Robonaut [1364]. Robot [144, 1916, 2192, 2608, 2620]. Robotic [204, 822, 2393]. Robotics [1476]. Robots [1273]. Robust [1049, 1388]. Robustness [2547]. Rock [567]. Rocket [2207]. Roger [2230]. Rogue [187, 1033]. Role [149, 1014, 1068, 1077, 1354]. Roll [567]. Rolling [164]. Rollover [32]. ROM [622]. Romanian [703]. Roomware [1340]. Root [1716, 2004]. Root-Server [2004]. Rose [940]. Rosetta [505, 2466]. Rough [152]. Router [8, 510, 1939, 2284]. Routing [777, 963, 1125, 2518, 2625]. Rover [750]. Rovers [2435]. RSA [807]. RSIM [574]. RTF [122]. RTSJ [915]. RTSJ-Compliant [915]. Ruby [1646]. Rule [920, 1622]. Rules [507, 806, 1754]. Ruling [279]. Run.time [2648, 2651]. Running [144]. Runtime [2649, 2650, 2652, 2653]. Rural [2305]. Russian [982]. Rust [170]. Rust-Belt [170].

S [99, 331, 583, 698, 937, 1156]. S2ESC [1377]. Safeguards [721]. Safety [1047, 1698, 1754, 1801, 2200, 2445, 2638]. Safety-Critical [1047, 1754, 2445, 2638]. Sakamura [532]. Salaries [871]. Salary [75, 331]. SALT [1054]. SALT-based [1054]. Salute [275]. Sampling [735]. Sanctions [632]. Sanctum [361]. Sandia [331]. Sands [224]. Satellite [209, 268, 2259]. Satellites [436]. Saves [1216, 1260]. Saving [806]. Says [1021]. Scalability [1721]. Scalable [460, 738, 750, 986, 2371]. Scale [123, 261, 400, 1274, 1351, 1570, 2102, 2238, 2670, 2671, 2673, 2675]. Scales [2498]. Scaling [418, 1050, 1229, 1431, 1861]. Scan [752]. Scanning [2052, 2471]. Scans [829]. ScanSoft [1033]. Scare [520]. Scatterers [2055]. SCE [109]. Scenario [771, 1155]. Scenarios [283]. Scenes [1381]. Schatz [485]. Schedulability [1296]. Schedule [579]. Scheduler [605]. Scheduling [363, 643, 1088]. Schema [1008, 2133]. Schemas [1008]. schematics [963]. Scheme [731, 873, 2603]. Schemes [2292]. Schism [1686]. Scholarship [264, 331, 507, 622, 893]. School [220, 357, 423, 712, 999]. Schools [254]. Science [10, 87, 357, 589, 605, 791, 806, 837, 882, 982, 1041, 1053, 1097, 1175, 1315, 1354, 1462, 1488, 1686, 1769, 1784, 1861, 1919, 2212, 2266, 2279, 2307, 2351, 2361, 2411– 2413, 2415, 2416, 2443, 2448, 2504, 2640, 2675]. Science-Philosophy [791]. Scientific [271, 301, 331, 1053, 1068, 1582, 1584, 1811, 1861, 2166, 2238]. Scientist [2207, 2608]. Scientists [1083]. SCinet [666]. SCM [144, 728, 815, 851]. Scob [1268]. Screen [224]. Scripting [447, 2041, 2325, 2674]. SCSI [270]. SCTP [1049]. SDE [778]. SDH [754]. SDH/SONET [754]. SDK [224, 291, 313, 715]. SDL [257]. Sea [1044]. Seamless [1231]. Search [13, 109, 245, 370, 477, 543, 608, 735, 768, 831, 885, 913, 973, 1008, 1197, 1386, 1404, 1511, 1557, 1559, 1560, 1562, 1751, 1776, 1798, 2073– 2077, 2501, 2553, 2574, 2669]. Searches [168, 237, 414, 1882]. Searching [5, 1246, 1561, 2027, 2260, 2525]. Seas [152]. Seat [852]. Secnes [172]. Second [21, 240, 267, 667, 2372]. Second-Generation [667]. Secret [1420]. Secrets [1676]. Secure [482, 483, 754, 1020, 1094, 1202, 1208, 1600, 1706, 1720, 1761, 2043, 2215, 2330]. Securing [1205, 1492, 1874, 1935, 2518, 2632]. Security [1, 130, 143, 165, 210, 252, 290, 306, 322, 344, 361, 406, 413, 447, 477, 479, 536, 543, 570, 584, 591, 614, 630, 651, 654, 674, 717, 753, 760, 807, 810, 819, 834, 857, 892, 896, 900, 945, 954, 990, 1030, 1035, 1082, 1095, 1106, 1125, 1166, 1177, 1202, 1206, 1228, 1239, 1304, 1361, 1469, 1548, 1615, 1625, 1637, 1668, 1722, 1752, 1801, 1849, 1865, 1871, 1915, 2029, 2084, 2117, 2157, 2202, 2344, 2410, 2444, 2448, 2477, 2574]. Security-Critical [1722]. Security-Safety [1801]. See [564, 697]. Seek [391, 507].

Seeking [436, 973, 2497–2500, 2502]. Seeks [130, 240, 308, 331, 546, 579, 755, 864, 1701]. Selected [43, 308, 691]. Selecting [197]. Self [179, 281, 599, 864, 1295, 1367, 1662, 1985]. Self-Adaptation [1295]. Self-Assembled [1367]. Self-Assembly [281]. Self-Assessment [1985]. self-aware [864]. Self-Healing [179]. Self-Managing [1662]. Self-Organization [599]. Semantic [56, 505, 772, 1044, 1562, 2051, 2156, 2203, 2261, 2362, 2380]. Semantics [1297]. Semiconductor [778]. Semiconductors [549, 1098]. sending [975]. Sensation [719]. Sense [1851]. Sensing [125, 1176, 1583, 2012, 2607]. Sensitive [440, 2602]. Sensor [751, 1035, 1122, 1249–1251, 1253, 1592, 1768, 1802, 2033, 2063, 2399, 2459, 2552]. SensorMap [2063]. Sensors [1254]. Sensory [1217]. Sentient [442]. Sequence [686]. Sequences [735]. Serial [1007]. Serious [209]. Server [187, 205, 224, 267, 335, 467, 667, 694, 846, 874, 915, 1321, 1777, 2004, 2007]. Server-Based [187]. Servers [558, 649, 782, 843, 844, 851, 1070, 1072, 1319]. Service [25, 26, 88, 220, 478, 485, 543, 675, 751, 779, 961, 1025, 1685, 1839, 1919, 1980, 2019, 2143, 2144, 2146, 2149, 2178, 2298, 2347– 2349, 2351, 2381, 2465, 2551, 2575, 2676]. Service-Oriented [1685, 2144, 2149, 2347– 2349, 2351, 2551, 2676]. Services [51, 81, 426, 448, 569, 627, 665, 694, 718, 740, 767, 841, 900, 927, 964, 966, 985, 1015, 1020, 1024, 1026–1028, 1033, 1050, 1106, 1166, 1270, 1313, 1369, 1370, 1407, 1431, 1570, 1659, 1718, 1784, 1874, 1940, 1973, 2189, 2350, 2352, 2426, 2446, 2476, 2479]. Session [2239]. Session-Aware [2239]. Set [49, 1494, 1586, 1988]. Setting [1377]. Setup [915]. Seven [214, 528, 700, 743, 2360]. Sex [608]. SGDL [406]. SGI [558, 753, 851, 915, 1007]. Shades [2511]. Shannon [357, 2056]. Shape [149, 1246, 1879, 2140]. Shaping [186, 989, 2212]. Share [201, 2373]. Shared [574, 1453, 1676]. Shared-Memory [574]. Sharing [188, 276, 399, 417, 2373]. Sharp [927]. Sheet [56]. Shift [32, 268, 1932, 2270, 2460]. Shifting [389, 478]. Shine [139]. Ships [384, 426, 447, 467, 487, 510, 534, 694, 807, 829, 1033]. Shooter [1252]. Shopbots [31]. Shopping [130]. ’Short [396, 1077, 1364, 2178]. Short-Range [396]. Shortage [1975]. Should [44, 968, 1575]. Showcase [139, 779]. Shrinking [1445]. Side [152, 631, 2562]. Sight [614]. Sign [1268]. Signal [231, 807, 852, 1140, 2391]. Signal-Processing [2391]. Signals [489]. SIISAM [2330]. Silas [736]. Silicon [778, 829, 1147, 1229, 1738, 1913, 2170]. Silo [2639]. Simics [575]. SimpleScalar [576]. Simplifies [152]. Simplifying [144, 304, 1059, 1132, 1731, 2705]. Simpson [702]. Simulate [105]. Simulating [574, 846]. Simulation [118, 165, 199, 285, 386, 447, 573, 575, 777, 795, 847, 1053, 1536, 1638, 1860, 2413]. Simulations [1767, 1817, 1966]. Sing [1254]. Single [354, 1007, 2366]. Single-Chip [354, 1007]. SIP [254]. Sirius [209]. Site [344, 696, 2310]. Sites [152, 964, 1095]. Sitraka [487]. Situ [1891]. Six [691]. Sizing [1777]. Sketch [2098]. Sketching [328]. Skies [2632]. Skill [1888]. Skills [741, 788, 1975, 2249, 2663]. Skin [406, 760]. Skin-Deep [760]. Sky [225]. Slate [1632, 2187, 2288]. Slaves [48]. Slide [1622]. Slightly [871]. Slow [883]. Slump [436]. Small [5, 27, 80, 103, 152, 546, 749, 1245, 1352, 2248]. Small-Project [27]. Smalltalk [312]. Smart [37, 282, 635, 680, 730, 794, 925, 940, 1131, 1366, 1409, 1410, 1653, 1800, 2389, 2650]. Smart-Phone [925]. Smarter [796, 873]. Smarts [1994]. Smoke [1519]. SMS [2178]. Sniff [807]. Snooping [593]. SOA [2145, 2246]. SoC [510, 552, 795, 890, 921, 1013]. Social [584, 769, 770, 1294, 1432, 1439, 1814, 1914, 2041, 2156, 2194, 2224, 2254, 2378, 2499, 2566]. Social-Mobile [1439]. Social-Technical [1294]. Sociality [2641]. Socially [1408]. Society [16, 43, 75, 99, 120, 139, 161, 172, 200, 201, 220, 240, 250, 251, 256, 264, 287, 308, 331, 357, 381, 401, 410, 423, 443, 444, 464, 465, 472, 485, 517, 520, 654, 732, 733, 755, 802, 814, 849, 871, 918, 961, 982, 1004, 1005, 1014, 1076, 1077, 1090, 1127, 1149, 1169, 1191, 1210, 1236, 1277, 1299, 1324, 1334, 1346, 1370, 1394, 1413, 1437, 1455, 1477, 1499, 1544, 1609, 1642, 1746, 1818, 1836, 1894, 2103, 2104, 2171, 2375, 2418, 2426, 2440, 2450, 2469, 2489, 2555, 2611, 2621, 2642, 2664, 2684, 2697]. Society [485, 494, 507, 532, 555, 579, 601, 622, 646, 669, 691, 712, 779, 801, 826, 893, 982, 1030, 1051, 1104, 1520, 1541, 1586, 1632, 1701, 1723, 1769, 1793, 1868, 1921, 1967, 1991, 2014, 2037, 2058, 2080, 2125, 2150, 2196, 2221, 2243, 2288, 2310, 2332, 2353, 2354, 2396, 2418, 2441, 2462, 2483, 2504, 2527, 2548, 2568, 2591, 2633, 2657, 2677]. Society-Sponsored [99]. Sociotechnical [2589]. Sockets [1706]. SoCs [1496]. Soft [164, 874, 1388]. Soft-Error [1388]. Softcoded [638]. Softech [467]. Softening [886]. SoftQuad [187, 426]. Softswitches [32]. SoftUDC [1318]. Software [1, 24, 61, 83, 114, 122, 143, 165, 167, 177, 190, 205, 207, 224, 244, 246, 264, 291, 293, 313, 315, 319, 335, 361, 377, 384, 387, 406, 426, 447, 467, 470, 475, 487, 491, 510, 513, 530, 534, 605, 627, 639, 649, 664, 673, 701, 723, 736, 753, 778, 807, 815, 829, 866, 874, 940, 1022, 1024, 1106, 1119, 1164, 1318, 1378, 1390, 1460, 1472, 1646, 1662, 1697, 1707, 1708, 1755, 1766, 1837, 1838, 1842, 1870, 1914, 1941, 1974, 2005, 2021, 2031, 2088, 2189, 2236, 2268, 2337, 2445, 2461, 2522, 2532, 2567, 2583, 2615]. Software [5, 23, 27, 32, 34, 75, 79, 91, 107, 115, 116, 119, 123, 144, 147, 159, 186, 193, 196, 223, 227, 254, 308, 331, 338, 343, 364, 373, 374, 376, 378, 411, 429, 457, 458, 532, 537, 543, 578, 588, 610, 621, 630, 648, 654, 663, 669, 681, 687, 720, 735, 743, 806, 830, 941, 974, 975, 995, 1003, 1021, 1025, 1053, 1080, 1230, 1338, 1539, 1629, 1631, 1640, 1694–1696, 1730, 1778, 1825, 1858, 1917, 1953, 1983, 2119, 2124, 2142, 2223, 2237, 2292, 2298, 2308, 2316, 2349, 2360, 2368, 2392, 2400, 2486, 2523, 2563, 2572, 2574, 2581, 2638]. Software [73, 74, 122, 127, 143, 165, 187, 224, 258, 267, 335, 361, 395, 406, 426, 447, 467, 487, 534, 558, 605, 627, 685, 694, 753, 819, 829, 851, 888, 929, 930, 963, 977, 986, 1030, 1054, 1062, 1106, 1417, 1840, 1926, 2230, 2609, 2677]. Software-Based [1318, 2581]. Software-Defined [2124]. Software-Intensive [1842]. SOI [807]. SOISIC [807]. Solaris [384, 406]. Solaris-to-Linux [406]. Solid [1368]. Solid-State [1368]. Solitary [1528]. Solution [8, 1253, 2596]. Solutions [777, 778, 1012, 1936, 2369]. Solvers [1308]. Solving [1023]. Some [248, 1167, 2361]. SONET [21, 754]. Songs [1115, 1880]. Sonic [95]. Sonice [941]. Soon [691, 2256]. Sooner [1851]. Soothsayers [1265]. Sophisticated [80, 103]. Sophomore [783]. Sorry [1618]. Sort [763]. Source [49, 122, 187, 254, 395, 406, 467, 537, 591, 672, 874, 1106, 1224, 1238, 1697, 1728, 1983, 2031, 2142, 2392, 2539, 2659, 2693]. Source-to-Source [2693]. Sourcing [707]. Space [822, 1410, 1812]. Spaces [2051]. Spam [952, 1017, 1117, 1294, 1432, 1440, 1558, 1792, 1959, 2589]. Spammers [975]. Sparks [1336, 1625, 2344, 2345]. Sparrows [2557]. Sparse [1965]. Speaks [1139]. Spec [112, 154, 605]. Special [1836]. Specific [502, 892, 2610]. Specification [135, 136, 717, 735, 1007, 1166, 1927]. Specifications [294, 1718, 2475]. Specifying [735, 2352]. specimens [727]. Speck [1937]. SPECpower [2511]. Specs [944, 1146]. Spector [937]. Spectrum [254]. Speculative [353, 367, 1900]. Speech [140, 268, 321, 617, 618, 662, 672, 701, 926, 1858, 2669]. Speech-Enabling [672]. SpeechStudio [807]. Speed [50, 103, 180, 626, 682, 715, 1205, 1207, 1228, 1283, 1425, 1888]. Speeding [223, 908, 1148, 1283]. Speeds [1095]. SpeedStep [244]. Speedware [361]. Spell [1]. Spell-Checking [1]. Spelling [1803]. Spiders [913]. Spiderweb [246]. Spielberg [1942]. SPIN [1053, 2368]. Spoken [619, 1248]. Sponsored [99, 381, 2076, 2553]. Sponsors [381, 2504]. Sports [1282]. Spring [1713]. Spyglass [143]. SQL [224]. Squeezing [1362]. SRC [205]. SRM [2407]. SSL [2239]. SSL/TLS [2239]. St. [264]. Stage [175]. Staged [159]. Stages [2303]. Stakeholder [1983]. Stakes [2386]. Standard [8, 52, 254, 364, 386, 546, 563, 1049, 1328, 1591, 1600, 2368]. Standardization [881, 1760, 1774, 2155, 2466]. Standardized [936, 1684, 2704]. Standards [32, 52, 56, 106, 127, 130, 191, 226, 270, 316, 409, 520, 563, 651, 697, 745, 755, 785, 788, 834, 842, 852, 900, 927, 970, 1010, 1057, 1111, 1117, 1377, 1430, 1825, 1863, 1872, 1904, 2029, 2064, 2084, 2109, 2200, 2248, 2337, 2379, 2414, 2420, 2466, 2507, 2593]. Standards-Based [2414]. Stanford [120, 669]. Star [585, 702, 1008]. Starbase [267, 829]. StarCore [510]. Stardock [406]. Start [79, 2455]. Start-up [2455]. Starts [1246, 2049]. Startup [537, 706]. Stashing [1853]. State [240, 264, 331, 490, 648, 966, 1368, 1508, 1783, 2108, 2144, 2583]. State-of-the-Art [331]. Statement [893]. Static [291, 1073]. Statistical [1043, 1555, 1605]. Statistics [588, 806]. StatSoft [406]. Status [942]. STC [99]. Stealing [2694]. Stealth [829]. Steering [312]. Step [641, 708, 1787]. Stephen [331]. Steps [528, 642, 1020, 1840, 1919]. Stereo [2022]. Stereo-Vision [2022]. StereoGraphic [313]. Stereoscopic [313]. Stereotype [1727]. Steven [937]. Stibitz [1622]. Still [270, 1082]. Stimulate [788]. Stiquito [1476]. Stitching [2366]. Stolen [2129]. Stone [99]. Stop [952, 1445, 2111]. Stopping [766, 1017]. Storage [102, 103, 181, 234, 614, 649, 661, 706, 726, 796– 799, 834, 860, 963, 975, 981, 1328, 1548, 1883, 1948, 1995, 2043, 2322, 2407, 2429, 2524]. Stored [2317]. Stores [2216]. Stories [1222]. Storing [455]. Storm [2210]. Story [588]. Strangest [1917]. Strategic [357, 516, 1902]. Strategies [418, 723, 788, 815, 1080, 2346, 2686]. Stream [682, 979, 2013]. Stream-Based [2013]. Streaming [322, 461, 463, 546, 908, 1106]. Streaming-Media [463, 546, 1106]. Streamlines [829]. Streams [150, 1903]. Street [2181]. Streets [212]. Strengthening [102]. Stress [92, 149, 205]. Stretching [346]. StrikeIron [985]. Strikes [146, 436]. String [2257, 2260]. Strings [735]. Strong [1826, 2519]. Stronger [2004]. Structural [1607, 2400]. Structure [329]. Structures [262, 807]. Struggle [412, 452]. Student [240, 357, 423, 601, 691, 712, 1473]. Students [139, 149, 172, 264, 331, 443, 507, 864, 893]. Studies [2054]. Studio [224, 487, 940]. Study [206, 260, 866, 1209, 1765, 2238]. Studying [688]. Stuff [546, 1245]. Stupidity [1008]. Style [56, 916, 1029, 1731]. Stymied [322]. Substance [32]. Substitutes [2335]. Success [555, 802, 1043, 1646, 1669]. Successful [1053, 1062, 1080]. Successfully [1629]. Sudden [1959]. Suit [862]. Suite [291, 361, 447, 467, 558, 649, 672, 694, 736, 1054]. Sum [123]. Summaries [9, 33, 60, 90, 113, 131, 153, 176, 195, 213, 233, 255, 280, 300, 323, 345, 371, 393, 415, 437, 456, 476, 499, 521, 545, 568, 590, 612, 634, 656, 679, 704, 722, 742, 762, 787, 812, 836, 859, 880, 903, 923, 947, 969, 992, 1016, 1040, 1061, 1087, 1113, 1135, 1158, 1179, 1199, 1220, 1241, 1264, 1286, 1309, 1331, 1355, 1379, 1400, 1421, 1441, 1464, 1485, 1507, 1529, 1550, 1573, 1594, 1619, 1641, 1664, 1687, 1711, 1733, 1756, 1779, 1804, 1829, 1854, 1876, 1905, 1929, 1954, 1976]. Summaries [2000, 2024, 2045, 2066, 2089, 2113, 2135, 2159, 2183, 2205, 2228, 2250, 2273, 2294, 2319, 2340, 2363, 2382, 2404, 2424]. Summarization [235]. Summer [1723]. Sun [32, 361, 455, 885]. Supercomputers [129, 1195, 1362, 2368]. Supercomputing [1116, 1963, 2169, 2455]. superscalar [844]. Supersystems [1375]. Supplement [622]. Support [72, 122, 143, 505, 510, 523, 525, 586, 610, 829, 908, 1003, 1021, 1033, 1862, 1872, 1989, 2009, 2012, 2148, 2413, 2497, 2502, 2651]. Supporting [491, 1892, 2044, 2249, 2544, 2653]. Supports [156, 205, 874, 2677]. Surface [2525]. Surfaces [873]. Surgery [1697]. Surveillance [1653, 2033, 2627]. Survey [75, 331, 436, 797, 2527]. Surveying [798, 964]. Survivability [157]. Survivable [178, 181]. Survival [2486, 2629]. Survive [171]. Surviving [953]. SuSE [406]. SUSSMicroTec [753]. Sustainability [2305]. Sustainable [427, 1181, 2169, 2361]. SVG [986]. Swarm [1998]. Swarms [1998]. Sweat [546, 1245]. Swift [109, 2546]. Switch [1007, 1010, 1037]. Switches [498]. Sybase [897, 1033]. Syclo [510]. Symantec [627]. Symbian [627]. Symbiotic [825, 1002]. Symbolic [1591]. Symposium [532]. Synchronization [1675]. synchronize [794]. Synergy [1901, 2636]. Synthetic [62]. System [39, 40, 48, 103, 169, 187, 197, 204, 206, 223, 228, 243, 244, 267, 313, 361, 384, 406, 442, 447, 529, 575, 576, 621, 627, 649, 654, 665, 688, 698, 705, 715, 753, 754, 807, 829, 851, 852, 877, 878, 887, 889, 915, 921, 967, 972, 1074, 1106, 1145, 1204, 1239, 1274, 1388, 1418, 1420, 1497, 1583, 1637, 1639, 1861, 1999, 2004, 2009, 2022, 2031, 2033, 2122, 2214, 2263, 2305, 2348, 2466, 2541, 2566, 2669]. System-Chip [967]. System-Level [921, 1861]. System-on-Chip [1145]. SystemC [890]. Systems [38, 82, 83, 123, 178, 179, 204, 238, 296, 339, 387, 421, 461, 484, 498, 501, 504– 506, 523, 525, 554, 559, 562, 583, 598, 606, 696, 735, 754, 777, 782, 807, 832, 852, 878, 920, 944, 964, 966, 1011, 1034, 1038, 1047, 1053, 1058, 1110, 1154, 1186, 1187, 1190, 1196, 1227, 1307, 1321, 1365, 1390, 1476, 1484, 1493, 1606, 1615, 1625, 1636, 1694, 1695, 1707, 1721, 1729, 1815, 1825, 1842, 1918, 1938, 1984, 1989, 2007, 2021, 2042, 2121, 2247, 2290, 2366, 2380, 2394, 2434, 2445, 2465, 2497, 2498, 2502, 2509, 2542, 2553, 2596, 2606, 2619, 2628, 2673]. Systems [122, 181, 205, 224, 236, 291, 294, 337, 376, 426, 487, 510, 652, 672, 685, 715, 730, 740, 829, 847, 873, 937, 963, 997, 1034, 1079, 1391, 1580, 1720, 1919, 1986, 2012, 2120, 2330, 2350, 2373, 2508]. Systems-Level [505]. Systems-on-Chips [1493]. Syware [605].

T. [275]. Tablet [972, 2097, 2100, 2102]. Tablet-PC [972]. Tablets [2099]. Tackle [1404]. Tackles [613]. Tag [981]. Takada [532]. Take [1246, 1503, 2140, 2232, 2474, 2546]. Takes [172, 279, 343, 344, 2668, 2669]. Taking [12, 254, 682, 887, 965, 1020, 1339, 1533, 1802, 1958, 2181, 2631]. Taks [103]. Taks-Oriented [103]. Tale [1096]. Talents [485]. Talk [344]. Talking [1528]. Talks [299]. Taming [1022]. Tampering [959]. Tamperproof [208]. tank [819]. Tapping [360]. Target [498, 1490, 2040]. targets [819]. Tascom [143]. Task [444, 2604]. Task-Based [2604]. Tasking [406]. Tasks [643, 1941]. Taulbee [331]. Taxonomy [1307, 1538]. Taylor [485]. TC [893]. TC-2 [893]. TCP [754, 1319]. TCP/IP [754]. TCPA [717]. teach [1041]. Teachers [240]. Teaches [149]. Teaching [336, 779, 1175, 1474, 1843, 2142, 2703]. Team [109, 1244, 1508, 1783]. Teaming [79]. Teams [43, 133, 161, 308, 423, 607, 691, 985, 993]. TEAtime [1146]. Tech [278, 392, 436, 475, 953, 1008, 1067, 1410, 1419, 1505, 2040]. Technical [44, 140, 204, 313, 357, 381, 444, 584, 960, 961, 1053, 1294, 1358, 2591]. Technique [819, 975, 1095, 1939, 2519]. Techniques [204, 370, 908, 1080, 2606]. Technological [743, 1012]. Technologies [4, 7, 8, 107, 147, 165, 190, 226– 228, 315, 342, 384, 391, 447, 455, 537, 560, 565, 605, 614, 627, 673, 699, 720, 736, 738, 754, 784, 831, 839, 851, 855, 876, 926, 927, 965, 973, 981, 1012, 1015, 1054, 1059, 1108, 1272, 1469, 1775, 2262, 2301, 2378, 2400, 2422, 2445, 2467, 2486, 2510, 2532, 2551, 2572, 2594]. Technologist [1745]. Technology [2, 7, 20, 31, 48, 58, 88, 105, 111, 123, 129, 151, 152, 174, 232, 253, 259, 278, 285, 298, 321, 343, 344, 369, 391, 392, 413, 453, 454, 461, 474, 496, 510, 514, 520, 525, 547, 549, 571, 593, 614, 628, 631, 637, 651, 678, 681, 682, 706, 707, 746, 766, 818, 852, 854, 868, 907, 926, 967, 974, 1020, 1094, 1098, 1267, 1282, 1285, 1290, 1336, 1358, 1418, 1468, 1471, 1512, 1577, 1593, 1599, 1715, 1848, 1853, 1883, 1901, 1975, 2170, 2255, 2287, 2298, 2303, 2314, 2367, 2387, 2430, 2454, 2474, 2494, 2518, 2539, 2560, 2580, 2641, 2668]. Technology [12, 41, 93, 147, 171, 224, 291, 341, 481, 519, 524, 747, 779, 794, 840, 863, 884, 885, 908, 951, 952, 995, 1019, 1043, 1044, 1065, 1066, 1077, 1188, 1364, 1445, 1450, 1451, 1511, 1517, 1701, 1739, 1972, 2097, 2123, 2336, 2345, 2408]. Technorealism [835, 881]. Technosoft [1]. Ted [654]. Teen [982, 1769]. Teenagers [2216]. Tektronix [736, 807]. Tele [279]. Tele-Immersion [279]. Telecom [122, 605, 1384]. Telecommunications [511, 2145]. Telelogic [143, 384, 807]. Telemedicine [2305]. Telephones [249, 873]. Telephony [254, 807, 1140, 1335, 2358]. Telescope [767]. Television [595]. Tempest [473]. Temporal [1367, 2485, 2616]. Ten [120, 161, 423, 691, 914, 1627]. TeraGrid [823, 2412]. Term [388, 396]. Termination [207]. Terminology [970]. Terraforming [419]. Terrain [649, 1252]. Terrorism [514, 543]. Test [165, 205, 224, 534, 633, 694, 736, 752, 852, 1079, 1387, 1461, 1538, 2007, 2509, 2630]. Test-Driven [852, 1538]. Testable [1693]. Testbed [666]. Tester [165, 583, 672]. Testing [102, 165, 313, 331, 361, 406, 626, 644, 649, 720, 735, 736, 753, 1034, 1053, 1888, 2315, 2349, 2360, 2531, 2615]. Tests [852]. Text [122, 143, 236, 625, 626, 1008, 2100]. Texts [735, 779]. th [2403]. Thanks [230, 1893]. Theater [676, 723]. Theft [109]. Their [201, 604, 747, 975, 1161, 1785, 2412]. theKompany.com [335]. Them [541]. Theme [314]. Themselves [841, 2630]. Theoretic [1062]. Theoretically [334]. Theory [27, 223, 258, 319, 693, 735, 914, 1034]. Theosophist [2296]. There [109, 650, 1142, 2485, 2488]. Theremin [838]. Thesaurus [1]. These [381, 2048]. Thick [560]. Thieves [2687]. Thin [1674]. ThinAirApp [467]. Thing [764, 1917, 1981]. Things [15, 48, 1268, 1312, 2399, 2444]. Think [15, 149, 819]. Thinking [284, 878, 986, 2252, 2639, 2660]. Third [274, 593, 2552]. Third-Generation [2552]. Those [1200]. Thought [27, 389]. Thoughts [276]. Threads [1163, 1719]. Threat [191, 477, 837, 1178]. Threaten [614, 1361]. Threatens [986]. Threats [793]. Three [228, 726, 913, 1096, 1377, 1764, 2488]. throughout [75]. Thunderbolts [701]. Thwarts [2004]. Thyself [705]. Tiger [2308]. Time [64, 84, 85, 134–136, 138, 143, 297, 313, 319, 374, 467, 600, 610, 643, 649, 676, 701, 795, 806, 1216, 1260, 1405, 1503, 1597, 1721, 1827, 1864, 2046, 2067, 2111, 2260, 2276, 2454]. Time-Bound [600]. Timeliness [2351]. TimeSys [897, 915]. Timothy [240]. Tiny [299, 726]. TLA [735]. TLS [2239]. Today [56, 147, 798, 1358, 1784, 1883, 1996, 2164]. Tofu [2672]. Together [273, 338, 1334]. Told [20]. Tolerant [1927, 2042]. Tome [1053]. Tomorrow [268, 1883]. Tone [777].

Tonic [627]. Too [128]. Tool

[1, 49, 122, 165, 187, 205, 224, 291, 313, 334, 361, 406, 426, 447, 467, 487, 534, 558, 583, 672, 694, 715, 736, 753, 778, 807, 829, 851, 897, 915, 928, 940, 985, 1007, 1033, 1054, 1079, 1106]. Toolbox [806]. toolboxes [963]. Toolkit [122, 165, 224, 244, 291, 447, 735, 986, 1697]. toolkits [1033]. Tools [1, 28, 102, 122, 149, 165, 199, 205, 267, 335, 406, 426, 467, 473, 573, 583, 627, 672, 715, 735, 806, 807, 915, 949, 963, 964, 986, 995, 1054, 1631, 1863, 2072, 2444, 2445, 2532, 2655, 2691, 2702]. Toolset [467, 694, 1033]. Top [120, 172, 293, 381, 387, 562, 982, 1565, 2396, 2483, 2515]. Top-Down [562]. Topics [918]. Topologies [799]. Torus [2672]. Tosiyasu [99]. Total [187, 2390]. Total-e-Server [187]. Touch [127, 747, 1044, 2069]. Tougher [109]. Tour [440]. Touting [588]. Town [102, 1978]. Toyland [89]. Toys [1131]. TPC [845]. TPC-W [845]. trace [957]. trace-and-revoke [957]. Traceability [2030]. Traces [2649]. Tracing [432]. Track [21, 99, 497]. Tracking [49, 53, 54, 268, 414, 534, 595, 628]. Trade [193, 949]. Tradeoffs [1447]. Trades [741, 788]. Traditional [646]. Traffic [322, 455, 1425, 2696]. Tragedies [842]. Trail [2560]. Trails [1261]. Trained [128]. Training [75, 149, 212, 296, 616, 701]. Transactional [335]. Transactions [308, 444]. Transforming [1334, 1539]. Transient [1995]. Transim [165]. Transistor [1142, 1389]. Transistor-Level [1389]. Transistors [299, 841]. Transitioning [1972, 2245]. Translates [1043]. Translating [67]. Translation [8, 63, 68, 70, 71, 175, 361, 453, 1555, 2328]. Translational [2415]. Translator [649]. Transmeta [59]. Transmission [1095]. Transparent [69, 70, 605]. Transport [1049]. Traps [432]. Trash [2364]. Travel [1222]. Traveling [702]. Travels [151]. Trek [702]. Trends [6, 22, 30, 57, 110, 128, 150, 173, 194, 211, 231, 252, 277, 297, 320, 342, 368, 390, 412, 428, 473, 475, 518, 680, 705, 725, 745, 765, 792, 817, 839, 862, 883, 906, 925, 950, 970, 973, 994, 1019, 1042, 1065, 1093, 1116, 1444, 1450, 2077, 2366, 2386, 2407, 2429, 2473, 2493, 2517, 2538, 2559, 2682]. Trendy [452]. Triage [867, 904]. Trial [702]. Tricks [2005]. Trident [1962]. Tries [453, 885]. Troubling [411]. True [830]. Trust [34, 536, 1938, 2335, 2478, 2479, 2564, 2585, 2632]. Trust-Based [536]. Trusted [298, 760, 863, 958, 2542]. Trusting [1288]. Trustworthy [863, 1707, 2541]. Truth [27, 858, 2683]. Try [2255]. Tube [1750]. Tuna [927]. Tune [79]. Tune-up [79]. Tuning [337, 896, 1071]. TurboPower [583]. Turing [2509]. Turn [392, 498, 841, 1444]. Turning [390, 1025, 1066, 1506, 2216]. Turtles [316]. Tutorial [331, 621, 735]. Tutorials [622]. TV [143, 699, 1691, 1833]. Tweaking [896]. Twisty [1107]. Two [161, 322, 381, 444, 747, 897, 926, 1067]. Type [2661]. Types [1183]. Typical [1147].

U [164]. U.S. [254]. Ubicomp [1615, 2315]. Ubiquitous [449, 988, 1038, 1176, 1298, 2637, 2660]. UCITA [89, 212, 475]. Ugly [1537]. UGVs [1886]. UL [165]. Ulead [534]. Ultimate [1514, 1518, 1590]. Ultra [156]. Ultra-Low [156]. Ultralow [1253]. Ultralow-Power [1253]. Ultrawideband [1065]. UML [59, 137, 143, 144, 257, 426, 447, 851, 916, 1296, 1652]. UML-Based [1296]. UML-Powered [144]. Unauthorized [807]. Unavoidable [588]. Unborn [747]. Uncanny [916]. Uncertainty [2345]. Uncloaking [468]. Unconventional [1143]. Uncovering [27, 152, 723, 1048]. Undergraduate [240, 893, 1254]. Underground [546]. Underlying [1848]. Underrepresented [2533]. Understanding [204, 522, 566, 857, 2098, 2403]. Underway [507]. unemployment [953]. Unicode [946]. Unified [103, 223, 1801]. Unify [864, 1183]. Unifying [83]. Union [2362]. UniPress [558]. Unique [291]. Uniqueness [1684]. Unit [736, 1021]. Units [1161, 1693]. Universal [767, 884, 996, 1512, 2434]. Universities [871]. University [1, 120, 879, 1244, 1547, 2245, 2636]. University-Industry [2636]. UNIX [164, 384]. UNIX-Based [384]. Unmanned [1885, 1891]. Unsafe [180, 2155]. Unspoken [1640]. Untrustworthy [1659]. Unveils [59, 291, 361, 627, 649, 661, 672, 2455]. Update [267]. Updates [267, 291, 335, 361, 406, 426, 447, 467, 510, 534, 583, 627, 649, 860, 1033, 1054, 1106]. Upgrade [299, 558, 747, 794, 1785]. Upgrades [361, 426, 487, 520, 534, 558, 583, 605, 649, 736, 778, 807, 829, 851, 874, 897, 915, 963, 985, 1007, 1054]. Upper [8]. UQBT [71]. Urban [1252, 1274, 1812, 1813, 1815, 1817, 1892, 2194]. Urban-Scale [1274]. URL [588]. US$18 [1244]. US$2K [846]. US$2M [846]. US$425 [849]. US-Centric [152]. Usability [491, 696]. Usage [177]. Usage-Based [177]. USB [122, 224, 874, 2049]. Use [40, 436, 706, 914, 1094, 1740, 2519]. Used [392]. Useful [1808, 2322]. Usefulness [1330]. Usenet [21]. User [103, 326, 327, 462, 597, 635, 941, 1305, 1672, 1674, 1705, 1778, 1811, 2180, 2239, 2371, 2401, 2446, 2485, 2604]. User-Centered [1305, 2485]. User-Friendly [1778]. User-Generated [2446]. Users [109, 546, 952, 1106, 1165, 1599, 2474]. Uses [351, 637, 897]. Using [105, 158, 168, 218, 227, 257, 307, 322, 341, 422, 440, 537, 594, 677, 749, 782, 784, 811, 927, 933, 934, 952, 996, 1008, 1029, 1067, 1183, 1296, 1431, 1525, 1559, 1560, 1607, 1649, 1652, 1676, 1792, 1872, 1903, 1966, 2055, 2123, 2257, 2259, 2284, 2437, 2525, 2567, 2617, 2649, 2652, 2673]. Utilities [855]. Utility [122, 187, 224, 1318, 1482, 2374]. Utopia [416, 452, 778]. Utopia-LVDS [778]. UWB [1760].

V [485]. V8 [829]. Vacuum [753]. Vadis [1873]. Vain [831]. Validation [672]. Valis [689]. Valuable [254]. Value [124, 448, 866, 970, 1333, 2379, 2411]. Value-Added [448]. Value-Based [866]. Values [339, 638, 1622, 1817]. Vanity [431]. Variability [2650]. Variable [401, 1999]. Variable-Expertise-Level [1999]. Variety [485]. Vector [519, 986]. VeGame [1001]. Vehicles [1885, 1887, 1888, 1891]. Vehicular [1350]. Vendors [112, 370, 392, 819, 908, 1093, 1785, 1959, 2345, 2538]. Venice [1001]. VentureCom [313]. Venus [683]. Veridicom [715]. Verification [103, 119, 467, 891, 1107, 1841, 2523, 2545]. Verified [1708, 1840]. Vernon [485]. Version [244, 778]. Versus [837, 881, 986, 1062, 1800, 2402, 2629]. Vertical [913]. Very [749, 1045, 1596, 2248]. Vforce [1963]. VHDL [852]. VHDL-AMS [852]. via [2429]. Viable [2520]. Vice [161, 1894]. Video [95, 150, 342, 379, 408, 462, 534, 885, 908, 1645, 1903, 2123, 2283, 2291, 2430]. Video-Based [2283]. Video-on-Demand [379]. Videogames [2395]. VideoSoft [1]. vie [926]. View [230, 279, 396, 411, 639, 977, 1637]. Viewer [267]. Viewership [2345]. Viewing [1095]. Viewpoints [743]. Views [201]. VIII [99, 357, 893]. VIM [1022]. VINT [239]. Viosoft [447]. Virtual [129, 365, 506, 607, 767, 782, 876, 916, 1218, 1438, 1449, 1450, 1452, 1453, 1461, 1536, 1569, 1683, 1737, 1791, 2053, 2132, 2164, 2372, 2395, 2410, 2491, 2586, 2596]. Virtualization [940, 1429, 1448, 1451, 1857, 2344]. Virtually [2357]. Viru [130]. Virus [436, 793]. Viruses [8, 369, 498, 975]. Visa [392]. Visas [1095]. Visibility [165]. Visible [1816]. VisiComp [736]. Vision [542, 821, 1684, 1886, 1892, 1907, 2022]. Vision-Based [1886]. Visionary [308]. Visitor [1763]. VisLab [1886]. Visual [224, 237, 335, 406, 426, 435, 467, 753, 1536, 1903, 2027, 2474, 2584, 2585, 2587]. Visualization [239, 271, 291, 313, 406, 753, 1106, 2532, 2690]. Visualize [726, 2652]. Visualizing [329, 1197, 1673]. Vital [143, 187, 1345]. Vitality [893]. VM [874]. VMware [874, 940]. Vocabulary [618]. VoD [699]. Voice [112, 190, 952, 964, 1139]. Voice-Enabled [190]. VoiceGenie [753, 963, 1054]. VoiceXML [753, 1054]. VoIP [244, 851, 1271, 2559]. Volume [263, 531, 733, 804, 1075, 1896, 2174, 2407]. Volumetric [326, 411, 1516]. Voted [1076, 1609, 1894, 2171, 2697]. Votes [2462]. Voting [1353, 1418, 1420, 2085, 2215, 2541]. VPN [627, 1208]. VRML [876]. Vulnerabilities [501]. Vulnerability [260].

W [331, 401, 532, 801, 845]. W3C

[520, 1044, 2475]. Wah [2612]. walk [689]. Walled [2402]. Wallet [219, 2530]. Walls [2491]. Want [1996]. Wanted [368, 570]. Wants [299, 706]. WAP [111, 192, 210]. War [446, 726]. Warehouse [528, 1675]. Warehouses [144, 523, 527]. Warehousing [526, 1008]. Warhol [739]. Warner [84, 85]. Warp [2328]. Wars [130, 758, 793]. Watch [22, 51, 81, 124, 169, 208, 245, 292, 336, 385, 428, 548]. Water [2607]. Waterloo [447]. Watermarking [2510]. Wave [187, 1033, 1104, 1909]. Wavelengths [455]. Wavelet [122, 2263]. Wavelet-Based [2263]. Waves [5]. Way [452, 650, 1290]. Ways [2343]. WDM [754, 2493]. WDM-PONs [2493]. WDM/Optics [754]. Wearable [1046, 1188]. Weather [1864]. Weaving [338]. Web [1, 13, 48, 56, 59, 80, 112, 117, 123, 144, 152, 170, 190, 210, 226, 267, 291, 301, 361, 366, 406, 426, 435, 467, 512, 558, 566, 599, 605, 627, 668, 672, 694, 738, 745, 769, 772, 774– 776, 807, 832, 860, 870, 964–966, 985, 986, 1008, 1012, 1015, 1020, 1024, 1033, 1050, 1059, 1106, 1108, 1164, 1166, 1197, 1313, 1383, 1407, 1411, 1549, 1559, 1560, 1562, 1570, 1645, 1659, 1751, 1798, 1809, 1973, 2020, 2041, 2117, 2156, 2157, 2203, 2233, 2362, 2378, 2380, 2422, 2465, 2467, 2525, 2536, 2551, 2564, 2607, 2680]. Web [109, 234, 237, 254, 428, 441, 473, 519, 522, 560, 569, 608, 696, 699, 740, 749, 767, 768, 770, 784, 788, 831, 841, 876, 900, 973, 996, 1026– 1028, 1044, 1095, 1165, 1554, 1718, 1731, 1800, 1882, 1940, 2027, 2075, 2086, 2350, 2408, 2475, 2479, 2510, 2573, 2594, 2637]. Web-Accessibility [2475]. Web-Application [435]. Web-Authoring [473]. Web-Based [291, 406, 558]. Web-Caching [512]. Web-Driven [1560]. Web-Log-Driven [1411]. Web-Oriented [1]. Web-Video [1645]. Web3 [2203]. Weblogs [1059]. Webs [2063]. Webservers [2240]. Websigns [441]. Websites [2655]. Weide [240]. Weighing [44]. Weird [702]. Welcome [63, 2600]. Welcomes [532]. Well [147]. Wellness [1187]. Were [2488]. West [2402]. Westbridge [807]. Weyuker [1746]. Whatever [173]. Wheel [1506]. Where [20, 79, 212, 489, 1082, 1518, 1911, 1996]. Whirling [963]. White [224]. Whither [739]. Who [5, 127, 230, 360, 674, 681, 1227, 2487]. Wi [951, 1351, 1492, 1834, 1997, 2300]. Wi-Fi [951, 1351, 1492, 1834, 1997, 2300]. Wide [1370, 2063]. Wide-Area [2063]. Widen [507]. Widgets [2048]. Wikis [2213]. Wild [194, 2315, 2402]. WildPackets [649]. Wiley [779]. Will [111, 126, 127, 152, 252, 279, 385, 436, 519, 538, 546, 547, 637, 658, 664, 726, 779, 818, 879, 927, 1042, 1065, 1117, 1140, 1161, 1290, 1334, 1425, 1445, 1514, 1518, 1600, 1646, 1715, 1834, 2202, 2232, 2473]. William [308, 485]. Williams [1609]. WiMax [1201, 2299]. Win [172, 443, 507]. WinDK [244]. Windows [80, 143, 191, 260, 267, 291, 406, 510, 558, 605, 736, 857, 1033]. Winner [99, 308, 331, 401, 669, 818, 1419]. Winners [99, 357, 972, 1942]. Winning [133, 446]. Wins [264, 703, 972, 982, 1244, 1508, 1783, 2612, 2698]. Winston [801]. Winter [1713]. Wipe [2187]. Wired [104]. Wireless [21, 32, 111, 145, 151, 155– 157, 211, 279, 313, 321, 336, 396– 398, 406, 407, 413, 426, 448, 467, 510, 534, 546, 547, 560, 583, 614, 637, 649, 672, 718, 726, 792, 818, 819, 840, 871, 908, 910, 974, 987, 990, 994, 1067, 1084, 1133, 1184, 1201, 1225, 1253, 1270, 1592, 1698, 1739, 1768, 2033, 2049, 2140, 2284, 2299, 2305, 2323, 2399, 2430, 2661]. Wires [1335]. Wisdom [768]. WiseNET [1253]. Withdraws [32]. within [480]. Without [103]. Wizard [20]. Wizards [829]. WLAN [765]. WLANs [942]. Women [308, 532, 893, 1080]. Wonder [152, 223]. Words [204]. Work [273, 331, 414, 633, 693, 869, 873, 1765, 1949, 2399, 2444, 2694]. Work-Stealing [2694]. Workflow [49]. Workflows [2166]. Workhorse [270]. Working [186, 223, 243, 270, 1883, 2002]. workload [997]. Workloads [843, 848]. Works [669, 767, 1044]. Workspaces [901]. Workstation [897]. World [56, 161, 194, 290, 423, 427, 691, 700, 786, 852, 871, 893, 941, 1090, 1095, 1181, 1206, 1262, 1334, 1838, 1850, 2163, 2164, 2361, 2364, 2384, 2405, 2427, 2452, 2471, 2491, 2515, 2536, 2540, 2557, 2563, 2565, 2573, 2578, 2603, 2646, 2659, 2700]. Worlds [490, 899, 1398, 1569, 1738, 2132, 2371, 2372]. Worm [130, 1207, 2210, 2560]. Worry [2276]. Worst [1146, 1155, 1510]. Worst-Case [1146]. Would [653, 885, 1503]. Writer [1107]. Writing [10, 87, 357, 873]. Wrong [319, 389, 611, 654].

X [146, 876]. X-Box [146]. X-VRML [876]. XAWare [753]. XDK [267]. Xerox [927]. Xilinx [778, 807, 829, 915]. Xiran [1106]. XM [209]. XML [49, 59, 80, 103, 106, 122, 164, 187, 299, 384, 399, 426, 474, 583, 605, 672, 754, 807, 900, 906, 970, 1079, 1166, 1425, 2370, 2381]. XML-Based [80, 1166]. XML-Oriented [672]. XP [558, 934]. XScale [426].

Y2K [32]. Yao [937]. Year [789, 1596, 1908, 2160, 2185, 2245]. Years [172, 256, 1008, 1092, 1137, 1182, 1223, 1266, 1311, 1360, 1382, 1403, 1423, 1442, 1467, 1487, 1509, 1531, 1551, 1576, 1598, 1623, 1627, 1644, 1667, 1689, 1714, 1736, 1759, 1782, 1807, 1832, 1856, 1881, 1910, 1931, 1957, 1979, 2003, 2026, 2047, 2068, 2092, 2116, 2138, 2162, 2184, 2208, 2231, 2275, 2297, 2321, 2342, 2365, 2385, 2406, 2428, 2453, 2472, 2492, 2516, 2537, 2558, 2579, 2601, 2624, 2645, 2667, 2688]. Yellow [1386]. Yes [2628]. YesSoftware [335, 605]. Yima [667]. You’re [79].

Z [937]. Zachary [240]. Zayante [510]. Zen [1138]. Zero [325, 534, 694, 1106]. Zero-Effort [325]. ZigBee [1599]. Zondigo [406]. Zork [1107].

References

Kroeker:1999:NTSf

[1] Kirk L. Kroeker. New tools: Software development: Monash University’s free Java development environment; ParaSoft’s error prevention tool; Technosoft’s C++ development kit; VideoSoft’s spell-checking and thesaurus controls. network development: Network Security’s security management tool; Fluke’s network inspector; MetaCreation’s Web-oriented painting tool. Computer, 32(11):120–122, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/

co/books/co1999/pdf/ry120.pdf. See

correction [28].

Meyer:1999:COTc

[2] Bertrand Meyer. Component and object technology: A really good idea. Computer, 32(12):144–147, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rz144.pdf. See

correction [29].

Aylor:2000:ECM

[3] James H. Aylor. Editor in Chief’s message: Computer looks to the new millennium. Computer, 33(1):4, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r1004.pdf.

Pollock:2000:PML

[4] Guylaine M. Pollock. President’s message: Leveraging emerging technologies. Computer, 33(1):6–7, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r1006.pdf.

Lawson:2000:LCA

[5] Harold ‘Bud’ Lawson, Dave Fleck, David Manthey, Barry G. Brown, Don Gotterbarn, Keith Miller, Simon Rogerson, Asoke K. Bhattacharya, Craig E. Groeschel, Roger Sorenson, Neville Holmes, Larry Bernstein, and Ted Lewis. Letters: Computer architecture is climbing out of the hole; help for managing small projects; code elicits comments and a question; who qualifies as a software engineer; negative impact of classroom computers; knowledge is the essential ingredient; searching for educational resources; health effects of radiation waves. Computer, 33(1):8–11, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r1008.pdf.

Edwards:2000:ITR

[6] John Edwards. Industry trends: Redefining IT career paths for the new millennium. Computer, 33(1):16– 20, January 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2000/pdf/
r1016.pdf.

Ortiz:2000:TNM

[7] Sixto Ortiz Jr. Technology news: Email protection advances with new technologies. Computer, 33(1):21– 23, January 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2000/pdf/
r1021.pdf.

Lear:2000:NBP

[8] Anne C. Lear. News briefs: PalmOS gains upper hand in handheld market; Internet viruses get more dangerous; standard offers interim net address solution; optical router promises faster net; translation technologies advance. Computer, 33(1):24–26, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r1024.pdf.

Anonymous:2000:ASa

[9] Anonymous. Article summaries. Computer, 33(1):27–28, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r1027.pdf.

Post:2000:WFC

[10] Jonathan Vos Post and Kirk L. Kroeker. Writing the future: Computers in science fiction. Computer, 33(1):29–37, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r1029.pdf;
                                        http://www.computer.org/computer/
                                        co2000/r1029abs.htm. See correction

[87].

Steane:2000:BBF

[11] Andrew M. Steane and Eleanor G. Rieffel. Beyond bits: The future of quantum information processing. Computer, 33(1):38–45, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r1038.pdf;
                                        http://www.computer.org/computer/
                                        co2000/r1038abs.htm.

Peterson:2000:TTM

[12] Christine Peterson. Taking technology to the molecular level. Computer, 33(1):46–53, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r1046.pdf;
                                        http://www.computer.org/computer/
                                        co2000/r1046abs.htm.

Introna:2000:DWP

[13] Lucas Introna and Helen Nissenbaum. Defining the Web: The politics of search engines. Computer, 33(1): 54–62, January 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2000/pdf/
r1054.pdf;   http://www.computer.
org/computer/co2000/r1054abs.htm.

Cooke:2000:NDN

[14] Daniel Cooke and Scott Hamilton. New directions at NASA Ames Research Center. Computer, 33(1):63–71, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r1063.pdf;
http://www.computer.org/computer/
co2000/r1063abs.htm.

Wilson:2000:TTT

[15] Janet Wilson. Toward things that think for the next millennium. Computer, 33(1):72–76, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r1072.pdf;
http://www.computer.org/computer/
co2000/r1072abs.htm.

Piner:2000:CSCa

[16] Mary-Louise G. Piner. Computer Society connection: What not to miss in 2000: Editor-in-Chief positions open for 2001. Computer, 33(1):77–82, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r1077.pdf.

Anonymous:2000:CCa

[17] Anonymous. Call and calendar. Computer, 33(1):82–84, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r1082.pdf.

Anonymous:2000:APIa

[18] Anonymous. Advertiser/product index. Computer, 33(1):85, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r1082.pdf.

Anonymous:2000:COa

[19] Anonymous. Career opportunities. Computer, 33(1):86–112, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r1086.pdf.

Lutz:2000:NBQ

[20] Michael J. Lutz. New books: A quantum leap in computing; the future of Oz, as told by the wizard; keeping the assembly line grooving; where real-life and artificial objects meet; networking technology; issues in computer architecture. Computer, 33(1):113, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r1113.pdf.

Kroeker:2000:PMC

[21] Kirk L. Kroeker. Products: The move to computing appliances; plug-and-play ISPs; fully powered mobile appliances; revamping wireless computing. reinventing the network: Usenet in a bottle; pushing data over Sonet. this and that: Keeping track of assets; one billion pixels per second. Computer, 33(1):114–116, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r1114.pdf;
http://www.3dfx.com/.

Roberts:2000:IWB

[22] Lawrence G. Roberts. Internet watch: Beyond Moore’s law: Internet growth trends. Computer, 33(1):117–119, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2000/pdf/r1117.pdf.

Sipper:2000:IEC

[23] Moshe Sipper and Eduardo Sanchez. Integrated engineering: Configurable chips meld software and hardware. Computer, 33(1):120–121, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r1120.pdf.

Boehm:2000:SMAa

[24] Barry Boehm. Software management: The art of expectations management. Computer, 33(1):122–124, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r1122.pdf. Ded-

icated to the memory of Tom Bauer.

Lewis:2000:BCS

[25] Ted Lewis. Binary critic: Service: The next inflection point. Computer, 33(1):128, 126–127, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r1128.pdf.

News:2000:BCS

[26] Ted Lewis. Binary critic: Service: The next inflection point. Computer, 33(1):128, 126–127, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r1128.pdf.

MacAndrew:2000:LPT

[27] Tim MacAndrew, Robert H. Norman, Jeff Templon, Kevin W. Wall, Shari Lawrence Pfleeger, Joseph C. Sligo, Christopher Jack, and Terry Ritter. Letters: Probability theory and software engineering; food for thought; uncovering erroneous assumptions; Einstein’s Nobel Prize; small-project process improvement; the truth about cryptography. Computer, 33(2):4, 6–8, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r2004.pdf.

Anonymous:2000:CNT

[28] Anonymous. Correction: New tools. Computer, 33(2):7, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r2004.pdf. See

[1].

Anonymous:2000:CRG

[29] Anonymous. Correction: A really good idea. Computer, 33(2):7, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2004.pdf. See

[2].

Lawton:2000:ITI

[30] George Lawton. Industry trends: Intellectual-property protection opens path for E-commerce. Computer, 33 (2):14–17, February 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2000/pdf/
r2014.pdf.

Clark:2000:TNS

[31] David Clark. Technology news: Shopbots become agents for business change. Computer, 33(2):18–21, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2000/pdf/r2018.pdf.

Lear:2000:NBY

[32] Anne C. Lear. News briefs: Y2K rollover: Few problems, many questions; Sun withdraws Java from standards process; softswitches shift network control to software; wireless net has more potential than substance. Computer, 33(2):22–24, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2022.pdf.

Anonymous:2000:ASb

[33] Anonymous. Article summaries. Computer, 33(2):25–26, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r2025.pdf.

OConnell:2000:CYT

[34] Emilie O’Connell and Hossein Saiedian. Can you trust software capability evaluations? Computer, 33(2):28– 35, February 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2000/pdf/
                                        r2028.pdf;   http://www.computer.
                                        org/computer/co2000/r2028abs.htm.

Schlansker:2000:EEP

[35] Michael S. Schlansker and B. Ramakrishna Rau Cover. EPIC: Explicitly parallel instruction computing. Computer, 33(2):37–45, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r2037.pdf;
                                        http://www.computer.org/computer/
                                        co2000/r2037abs.htm.

Pankanti:2000:GEI

[36] Sharath Pankanti, Ruud M. Bolle, and Anil Jain. Guest Editors’ introduction: Biometrics: The future of identification. Computer, 33(2):46–49, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r2046.pdf.

Pentland:2000:FRS

[37] Alex (Sandy) Pentland and Tanzeem Choudhury. Face recognition for smart environments. Computer, 33(2):50– 55, February 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2000/pdf/
r2050.pdf;   http://www.computer.
org/computer/co2000/r2050abs.htm.

Phillips:2000:IEB

[38] P. Jonathon Phillips, Alvin Martin, C. L. Wilson, and Mark Przybocki. An introduction to evaluating biometric systems. Computer, 33(2):56–63, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2056.pdf;
http://www.computer.org/computer/
co2000/r2056abs.htm.

Frischholz:2000:BMB

[39] Robert W. Frischholz and Ulrich Dieckmann. BioID: A multimodal biometric identification system. Computer, 33(2):64–68, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2064.pdf;
http://www.computer.org/computer/
co2000/r2064abs.htm.

Negin:2000:IBS

[40] Michael Negin, Thomas A. Chmielewski, Jr., Marcos Salganicoff, Theodore A. Camus, Ulf M. Cahn von Seelen, Péter L. Venetianer, and Guanghua G. Zhang. An iris biometric system for public and personal use. Computer, 33(2):70–75, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2070.pdf;
http://www.computer.org/computer/
co2000/r2070abs.htm.

Wayman:2000:FBT

[41] James L. Wayman. Federal biometric technology legislation. Computer, 33(2):76–80, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2076.pdf;
http://www.computer.org/computer/
co2000/r2076abs.htm.

Cabri:2000:MAC

[42] Giacomo Cabri, Letizia Leonardi, and Franco Zambonelli. Mobile-agent coordination models for Internet applications. Computer, 33(2):82–89, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2000/pdf/r2082.pdf.

Piner:2000:CSCb

[43] Mary-Louise G. Piner. Computer Society connection: Inaugural design competition teams selected; Fellow nominations due 15 March; Society leaders prepare for 2000 activities. Computer, 33(2):90–92, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2090.pdf.

Bowyer:2000:YCS

[44] Kevin W. Bowyer. Your call: Should professional conduct influence technical awards; weighing privacy rights against company policy. Computer, 33(2):93–95, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2093.pdf.

Anonymous:2000:CCb

[45] Anonymous. Call and calendar. Computer, 33(2):96–100, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2096.pdf.

Anonymous:2000:COb

[46] Anonymous. Career opportunities. Computer, 33(2):102–117, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2102.pdf.

Anonymous:2000:APIb

[47] Anonymous. Advertiser/product index. Computer, 33(2):118, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2119.pdf.

Lutz:2000:NBMa

[48] Michael J. Lutz. New books: Masters of technology or slaves to the Egrind?; managing apples and oranges; descrambling data on the web; designing system networks; fixing things before things fall apart; Java 2 guidelines. Computer, 33(2):119, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2119.pdf.

Kroeker:2000:PDD

[49] Kirk L. Kroeker. Products: Digital dashboard concept now affordable; defect-tracking tool adds custom workflow; active application source code browser for MFC; customizable set of dials and gauges for Motif; first IDE for PlayStation2; new XML editor hits channels; the disk optimization controversy. Computer, 33(2):120–122, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2120.pdf.

Macedonia:2000:ECW

[50] Michael Macedonia. Entertainment computing: Why digital entertainment drives the need for speed. Computer, 33(2):124–127, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2124.pdf.

McKnight:2000:IWPa

[51] Lee W. McKnight and Jahangir Boroumand. Internet watch: Pricing Internet services: Approaches and challenges. Computer, 33(2):128– 129, February 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2000/pdf/
r2128.pdf.

Tilton:2000:SEB

[52] Catherine J. Tilton. Standards: An emerging biometric API industry standard. Computer, 33(2):130–132, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2000/pdf/r2130.pdf.

Lewis:2000:BCT

[53] Ted Lewis. Binary critic: Tracking the ‘anywhere, anytime’ inflection point. Computer, 33 (2):136, 134–135, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2136.pdf.

News:2000:BCT

[54] Ted Lewis. Binary critic: Tracking the ‘anywhere, anytime’ inflection point. Computer, 33 (2):136, 134–135, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2136.pdf.

Aylor:2000:EMN

[55] James H. Aylor. EIC message: New Computer departments reflect our changing industry. Computer, 33(3):4, 6, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3004.pdf.

Donne:2000:LSS

[56] Werner Donné, P. M. Marden, Jr., E. V. Munson, Jack M. Nilles, Sebastian Tyrrell, Bertrand Meyer, David Wall, Tom Dodington, and Ted Lewis. Letters: Style sheet standards renew old debate; A missed pun; an absurdity in today’s world; semantic Web party poopers; why can’t I buy a Mercedes PC? Computer, 33(3):8–10, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3008.pdf.

Ortiz:2000:ITE

[57] Sixto Ortiz Jr. Industry trends: Embedded databases come out of hiding. Computer, 33(3):16–19, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3016.pdf.

Edwards:2000:TNB

[58] John Edwards. Technology news: Building the optical-networking infrastructure. Computer, 33(3):20–23, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2000/pdf/r3020.pdf.

Lear:2000:NBT

[59] Anne C. Lear. News briefs: Transmeta unveils Crusoe mobile processor; new Web protocol combines HTML and XML; as UML evolves, its applications expand. Computer, 33(3):24–26, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3024.pdf.

Anonymous:2000:ASc

[60] Anonymous. Article summaries. Computer, 33(3):27–28, March 2000.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3027.pdf.

Holmes:2000:ESP

[61] W. Neville Holmes. The evitability of software patents. Computer, 33(3):30–34, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3030.pdf;
http://www.computer.org/computer/
co2000/r3030abs.htm.

Ladner:2000:MIS

[62] Roy Ladner, Mahdi Abdelguerfi, and Kevin Shaw Cover. 3D mapping of an interactive synthetic environment. Computer, 33(3):35–39, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3035.pdf;
http://www.computer.org/computer/
co2000/r3035abs.htm.

Altman:2000:WOB

[63] Erik R. Altman, David Kaeli, and Yaron Sheffer. Welcome to the opportunities of binary translation. Computer, 33(3):40–45, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3040.pdf.

Moon:2000:JTC

[64] SooMook Moon and Kemal Ebcioğlu. A just-in-time compiler. Computer, 33(3):41, March 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2000/pdf/
                                        r3040.pdf.

Bala:2000:NBA

[65] Vasanth Bala, Evely Duesterwald, and Sanjeev Banerjia. Native binary acceleration. Computer, 33(3):42, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wu:2000:CP

[66] Youfeng Wu. Collaborative profiling. Computer, 33(3):42, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r3040.pdf.

Drongowski:2000:TM

[67] Paul J. Drongowski. Translating MMX. Computer, 33(3):43, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r3040.pdf.

Sites:2000:FNE

[68] Richard L. Sites. A few notable events in the history of binary translation. Computer, 33(3):44, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zheng:2000:PRI

[69] Cindy Zheng and Carol Thompson. PA-RISC to IA-64: Transparent execution, no recompilation. Computer, 33(3):47–52, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r3047.pdf;

http://www.computer.org/computer/
co2000/r3047abs.htm.

Gschwind:2000:DTB

[70] Michael Gschwind, Erik R. Altman, Sumedh Sathaye, Paul Ledak, and David Appenzeller. Dynamic and transparent binary translation. Computer, 33(3):54–59, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3054.pdf;
http://www.computer.org/computer/
co2000/r3054abs.htm.

Cifuentes:2000:UAB

[71] Cristina Cifuentes and Mike Van Emmerik. UQBT: Adaptable binary translation at low cost. Computer, 33(3):60–66, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3060.pdf;
http://www.computer.org/computer/
co2000/r3060abs.htm.

Bacon:2000:GSD

[72] Jean Bacon, Ken Moody, John Bates, Richard Hayton, Chaoying Ma, Andrew McNeil, Oliver Seidel, and Mark Spiteri. Generic support for distributed applications. Computer, 33(3):68–76, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3068.pdf;
http://www.computer.org/computer/
co2000/r3068abs.htm.

vanOmmering:2000:KCMa

[73] Rob van Ommering, Frank van der Linden, Jeff Kramer, and Jeff Magee. The Koala component model for consumer electronics software. Computer, 33(3):78–85, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r3078.pdf;
                                        http://www.computer.org/computer/
                                        co2000/r3078abs.htm.

vanOmmering:2000:KCMb

[74] Rob van Ommering, Frank van der Linden, Jeff Kramer, and Jeff Magee. The Koala component model for consumer electronics software. Computer, 33(3):78–85, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r3078.pdf;
                                        http://www.computer.org/computer/
                                        co2000/r3078abs.htm.

Piner:2000:CSCc

[75] Mary-Louise G. Piner. Computer Society connection: New training opportunities available for software engineers; Fellows named for distinguished achievements; salary survey compares faculty earnings throughout US and canada. Computer, 33(3):86–89, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r3086.pdf.

Anonymous:2000:CCc

[76] Anonymous. Call and calendar. Computer, 33(3):90–93, March 2000.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3090.pdf.

Anonymous:2000:COc

[77] Anonymous. Career opportunities. Computer, 33(3):94–102, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3094.pdf.

Anonymous:2000:APIc

[78] Anonymous. Advertiser/product index. Computer, 33(3):103, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3104.pdf.

Lutz:2000:NBTa

[79] Michael J. Lutz. New books: Teaming up for software development; fixing leaks in your network; know where you’re going before you start; communication network tune-up; programming like a champ; the right algorithm for the job. Computer, 33(3):104, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3104.pdf.

Kroeker:2000:PCL

[80] Kirk L. Kroeker. Products: Controlling the look and feel of Windows; sophisticated project estimation for $99; Web applications for the Internet-challenged; network-based digital asset manager; Linux middleware framework; XML-based information portal; Java development for small computing devices; build installations for Windows 2000. Computer, 33(3):105–107, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r3105.pdf;
                                        http://www.stardock.net/.

McKnight:2000:IWPb

[81] Lee W. McKnight and Jahangir Boroumand. Internet watch: Pricing Internet services: Proposed improvements. Computer, 33(3):108–109, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co2000/pdf/r3108.pdf.

Sarjoughian:2000:IEC

[82] Hessam S. Sarjoughian, Bernard P. Zeigler, and Daryl R. Hild. Integrated engineering: A codesign approach for distributed systems. Computer, 33(3):110–113, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r3110.pdf.

Boehm:2000:SMU

[83] Barry Boehm. Software management: Unifying software engineering and systems engineering. Computer, 33(3):114–116, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r3114.pdf.

Lewis:2000:BCA

[84] Ted Lewis. Binary critic: AOL, Time Warner, and the Crash of 2000. Computer, 33(3):120, 118–119, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3120.pdf.

News:2000:BCA

[85] Ted Lewis. Binary critic: AOL, Time Warner, and the Crash of 2000. Computer, 33(3):120, 118–119, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r3120.pdf.

Parhami:2000:LDG

[86] Behrooz Parhami, Ted Lewis, Bruno Hanzen, Ted Lewis, Mike Pelley, A. D. Arthur, Chris Plambeck, J. Vos Post, K. L. Kroeker, Andrew E. Jove, Jr., J. Vos Post, and K. L. Kroeker. Letters: Does greed influence the computer industry? directories’ great-granddaddy; IV&V questions; Borg’s origins; predicting the future. Computer, 33(4):4, 6–7, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r4004.pdf.

Anonymous:2000:CWF

[87] Anonymous. Correction: Writing the Future: Computers in Science Fiction. Computer, 33(4):7, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r4012.pdf. See

Garber:2000:TND

[88] Lee Garber. Technology news: Denialof-service attacks rip the Internet. Computer, 33(4):12–17, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r4012.pdf.

Lear:2000:NBU

[89] Anne C. Lear. News briefs: UCITA: Misguided or misunderstood? chips in toyland; Cobol programmers could be key to new IT; Ethernet evolves again, may expand range; what’s next for the next-generation net? Computer, 33(4):18–21, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r4018.pdf.

Anonymous:2000:ASd

[90] Anonymous. Article summaries. Computer, 33(4):22–23, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r4022.pdf.

Nuseibeh:2000:LIS

[91] Bashar Nuseibeh, Steve Easterbrook, and Alessandra Russo. Leveraging inconsistency in software development. Computer, 33(4):24–29, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r4024.pdf;
                                        http://www.computer.org/computer/
                                        co2000/r4024abs.htm.

Alexander:2000:CJP

[92] Roger T. Alexander, James M. Bieman, and John Viega. Coping with Java programming stress. Computer, 33(4):30–38, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r4030.pdf;
http://www.computer.org/computer/
co2000/r4030abs.htm.

Vemuri:2000:CCT

[93] Ranga R. Vemuri and Randolph E. Harr. Configurable computing: Technology and applications. Computer, 33 (4):39–40, April 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2000/pdf/
r4039.pdf.

DeHon:2000:DAC

[94] André DeHon. The density advantage of configurable computing. Computer, 33(4):41–49, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r4041.pdf;
http://www.computer.org/computer/
co2000/r4041abs.htm.

Haynes:2000:VIP

[95] Simon D. Haynes, John Stone, Peter Y. K. Cheung, and Wayne Luk. Video image processing with the Sonic architecture. Computer, 33(4):50–57, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r4050.pdf;

                                        http://www.computer.org/computer/
                                        co2000/r4050abs.htm.

Platzner:2000:RAC

[96] Marco Platzner. Reconfigurable accelerators for combinatorial problems. Computer, 33(4):58–60, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r4058.pdf;
                                        http://www.computer.org/computer/
                                        co2000/r4058abs.htm.

Callahan:2000:GAC

[97] Timothy J. Callahan, John R. Hauser, and John Wawrzynek. The Garp architecture and C compiler. Computer, 33(4):62–69, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r4062.pdf;
                                        http://www.computer.org/computer/
                                        co2000/r4062abs.htm.

Goldstein:2000:PRA

[98] Seth Copen Goldstein, Herman Schmit, Mihai Budiu, Srihari Cadambi, Matt Moe, and R. Reed Taylor. PipeRench: A reconfigurable architecture and compiler. Computer, 33(4):70–77, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r4070.pdf;
                                        http://www.computer.org/computer/
                                        co2000/r4070abs.htm.

Piner:2000:CSCd

[99] Mary-Louise G. Piner. Computer Society connection: Focus on awards: Computer pioneer broke ground in architecture and pattern analysis (Herbert Freeman); astrophysicist named Fernbach winner (Michael L. Norman); Booth winners: Educational awards: Tosiyasu L. Kunii, Harold S. Stone; Grove named IEEE Medal of Honor winner; excellence in research recognized: Lalit M. Patnaik, Bharat Bhargava; Quipunet founder receives first precollege award; new career resource available from the Computer Society; Society-sponsored track offered at STC; Board approves reading of bylaws amendment on publications activities board; IEEE Division VIII Director-Elect to leave post. Computer, 33(4):80–84, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r4080.pdf.

Anonymous:2000:CCd

[100] Anonymous. Call and calendar. Computer, 33(4):86–88, 90, 92, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r4086.pdf.

Anonymous:2000:COd

[101] Anonymous. Career opportunities. Computer, 33(4):94–101, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lutz:2000:NBTb

[102] Michael J. Lutz. New books: Tools for OO testing; this old networked town; programming fundamentals, language by language; error-control in communications engineering; business borrows from biology; cost-effective data storage; strengthening the project management links. Computer, 33(4):103, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r4103.pdf.

Kroeker:2000:PBV

[103] Kirk L. Kroeker. Products: Behavioral verification for HDL designs; build patch installs without heavy programming; new network attached storage appliance; parallel rasterization architecture for high-speed graphics; small rackmount computer; XML for a unified enterprise; sophisticated configuration management; taksoriented help system puts user at center. Computer, 33(4):104–106, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r4104.pdf;
                                        http://www.3dlabs.com/.

Varshney:2000:CRAa

[104] Upkar Varshney. Communications: Recent advances in wired networking. Computer, 33(4):107–109, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r4107.pdf.

Macedonia:2000:ECU

[105] Michael Macedonia. Entertainment computing: Using technology and innovation to simulate daily life. Computer, 33(4):110–112, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r4110.pdf.

Wiggins:2000:SXI

[106] Richard Wiggins. Standards: XML: An interview with Peter Flynn. Computer, 33(4):113–114, 116, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r4113.pdf.

Lutz:2000:STF

[107] Michael J. Lutz. Software technologies: Fundamental ideas and change. Computer, 33(4):115–116, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r4115.pdf.

Lewis:2000:BCWa

[108] Ted Lewis. Binary critic: When atoms and bits collide: Detroit goes digital. Computer, 33(4):120, 118–119, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r4120.pdf.

Pittman:2000:LTN

[109] Tom Pittman, Roger Coombs, William D. Jackson, Eric A. Weiss, Jonathan Vos Post, Kirk L. Kroeker, George Mc-Kee, Paul Losleben, Sharath Pankanti, Ruud M. Bolle, Anil Jain, David Cox, Mark D. Levin, Roy Kimbrell, Borisk Mutafelija, Emilie O’Connell, and Hossein Saiedian. Letters: There’s no free lunch for Web users; search engine politics; the future of computing; not so swift; biometric identity theft; changing the SCE focus from processes to people; A must-read for program managers; fooling the team is tougher now; revisions improve SCE processes. Computer, 33(5):4, 9–11, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r5004.pdf.

Hellman:2000:ITF

[110] Reed Hellman. Industry trends: The future of the OS for Internet applications. Computer, 33(5):12–15, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r5012.pdf.

Leavitt:2000:TNW

[111] Neal Leavitt. Technology news: Will WAP deliver the wireless Internet? Computer, 33(5):16–20, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r5016.pdf.

Lear:2000:NBV

[112] Anne C. Lear. News briefs: Vendors begin to adopt ‘next-generation’ Internet protocol; road to Linux acceptance could fork; new spec gives voice to the Web. Computer, 33(5):21–23, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r5021.pdf.

Anonymous:2000:ASe

[113] Anonymous. Article summaries. Computer, 33(5):24–25, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r5024.pdf.

Boehm:2000:GIC

[114] Barry Boehm and Victor R. Basili. Gaining intellectual control of software development. Computer, 33(5):27–33, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2000/pdf/r5027.pdf;
http://www.computer.org/computer/
co2000/r5027abs.htm.

Pour:2000:PMS

[115] Gilda Pour, Martin L. Griss, and Michael Lutz. The push to make software engineering respectable. Computer, 33(5):35–43, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

dlib.computer.org/co/books/co2000/
pdf/r5035.pdf;   http://www.computer.
org/computer/co2000/r5035abs.htm.

Lethbridge:2000:WKI

[116] Timothy C. Lethbridge. What knowledge is important to a software professional? Computer, 33(5):44–50, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r5044.pdf;
http://www.computer.org/computer/
co2000/r5044abs.htm.

Brewington:2000:KCW

[117] Brian E. Brewington and George Cybenko. Keeping up with the changing Web. Computer, 33(5):52–58, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r5052.pdf;
http://www.computer.org/computer/
co2000/r5052abs.htm.

Breslau:2000:ANS

[118] Lee Breslau, Deborah Estrin, Kevin Fall, Sally Floyd, John Heidemann, Ahmed Helmy, Polly Huang, Steven McCanne, Kannan Varadhan, Ya Xu, and Haobo Yu. Advances in network simulation. Computer, 33(5):59–67, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r5059.pdf;
http://www.computer.org/computer/
co2000/r5059abs.htm.

Staunstrup:2000:PVE

[119] Jørgen Staunstrup, Henrik Reif Andersen, Henrik Hulgaard, Jørn Lind-Nielsen, Kim G. Larsen, Gerd Behrmann, Kåre Kristoffersen, Arne Skou, Henrik Leerberg, and Niels Bo Theilgaard. Practical verification of embedded software. Computer, 33 (5):68–75, May 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2000/pdf/
r5068.pdf;   http://www.computer.
org/computer/co2000/r5068abs.htm.

Piner:2000:CSCe

[120] Mary-Louise G. Piner. Computer Society connection: Top ten conferences for the coming months; computer engineer Hennessy named Stanford University president. Computer, 33(5):76–80, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r5076.pdf.

Anonymous:2000:CCe

[121] Anonymous. Call and calendar. Computer, 33(5):82–86, 88, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r5082.pdf.

Kroeker:2000:PIM

[122] Kirk L. Kroeker. Products: iCompression’s MPEG-2 encoder reference design kit for USB; free project management support tools; Force Computers’ protocol software for telecom integration; Data Junction’s free XML tool; Fookes Software’s text editor for programmers; Research Systems’ IDL wavelet toolkit; Be opens source code to BeOS 5; JafSoft’s utility for converting text files into RTF. Computer, 33(5):98–100, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r5098.pdf.

Lutz:2000:NBCa

[123] Michael J. Lutz. New books: Cogito ergo Web cam sum: Knowing from afar; software engineering experimentation; object technology and largescale systems; reasoning with your computer; the RFC dictionary; practical C++. Computer, 33(5):101, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r5101.pdf.

Barua:2000:IWV

[124] Anitesh Barua, Andrew B. Whinston, and Fang Yin. Internet watch: Value and productivity in the Internet economy. Computer, 33(5):102–105, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r5102.pdf.

Agre:2000:IEI

[125] Jon Agre and Loren Clare. Integrated engineering: An integrated architecture for cooperative sensing networks. Computer, 33(5):106–108, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r5106.pdf.

Lewis:2000:BCWb

[126] Ted Lewis. Binary critic: Why funny money will have the last laugh. Computer, 33(5):112, 110–111, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r5112.pdf.

Zeliff:2000:LSP

[127] Bob Zeliff, Neville Holmes, Peggy Dymond, Ted Lewis, K. V. K. K. Prasad, Norman R. Scott, Jonathan Vos Post, and Kirk L. Kroeker. Letters: Software patents: Public good or private gain? boomers’ lifestyle changes will bring the big dip; quality standards with a human touch; the genius who revolutionized the future. Computer, 33(6):4, 6–7, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r6004.pdf.

Clark:2000:ITT

[128] David Clark. Industry trends: Are too many programmers too narrowly trained? Computer, 33(6):12–15, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r6012.pdf.

Lawton:2000:TND

[129] George Lawton. Technology news: Distributed net applications create virtual supercomputers. Computer, 33 (6):16–20, June 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2000/pdf/
r6016.pdf.

Lear:2000:NBL

[130] Anne C. Lear. News briefs: Love hurts: New E-mail worm afflicts millions; standards group seeks to quell instant messaging wars; deliberate ‘virus’ plagues popular game; new chip helps with network security; pinning down E-shopping habits. Computer, 33(6):22–24, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r6022.pdf.

Anonymous:2000:ASf

[131] Anonymous. Article summaries. Computer, 33(6):26–27, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r6026.pdf.

Maurer:2000:CWI

[132] Peter M. Maurer. Components: What if they gave a revolution and nobody came? Computer, 33(6):28–34, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r6028.pdf;
                                        http://www.computer.org/computer/
                                        co2000/r6028abs.htm.

Cover:2000:WTP

[133] Robert S. Oshana Cover. Winning teams: Performance engineering during development. Computer, 33(6):36–44, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co2000/pdf/r6036.pdf;
                                         http://www.computer.org/computer/
                                        co2000/r6036abs.htm.

Shokri:2000:GEI

[134] Eltefaat Shokri and Phillip Sheu. Guest Editors’ introduction: Real-time distributed object computing: An emerging field. Computer, 33(6):45–46, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r6045.pdf.

Bollella:2000:RTS

[135] Greg Bollella and James Gosling. The real-time specification for Java. Computer, 33(6):47–54, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r6047.pdf;
                                        http://www.computer.org/computer/
                                        co2000/r6047abs.htm.

Schmidt:2000:ORT

[136] Douglas C. Schmidt and Fred Kuhns. An overview of the real-time CORBA specification. Computer, 33(6):56–63, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2000/pdf/r6056.pdf;
 http://www.computer.org/computer/
co2000/r6056abs.htm.

Selic:2000:GFM

[137] Bran Selic. A generic framework for modeling resources with UML. Computer, 33(6):64–69, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r6064.pdf;
http://www.computer.org/computer/
co2000/r6064abs.htm.

Kim:2000:ART

[138] K. H. (Kane) Kim. APIs for realtime distributed object programming. Computer, 33(6):72–80, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r6072.pdf;
http://www.computer.org/computer/
co2000/r6072abs.htm.

Piner:2000:CSCf

[139] Mary-Louise G. Piner. Computer Society connection: Awards programs let Computer Society students shine: Lecture programs showcase member expertise. Computer, 33(6):82–84, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r6082.pdf.

Boyer:2000:YCA

[140] Kevin W. Boyer. Your call: Anonymity and freedom of speech in cyberspace; professional conduct and technical awards. Computer, 33(6):85–86, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r6085.pdf.

Anonymous:2000:CCf

[141] Anonymous. Call and calendar. Computer, 33(6):87–89, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r6087.pdf.

Anonymous:2000:COe

[142] Anonymous. Career opportunities. Computer, 33(6):90–94, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r6090.pdf.

Kroeker:2000:PEN

[143] Kirk L. Kroeker. Products: Enterasys Networks’ E-commerce access platforms; Tascom Software’s ASP editor; Vital’s text editor for program development; RapidStream’s security appliance; Kemma Software’s help desk software; Telelogic’s real-time UML profiling software; ParaSoft extends product support to Windows 2000; Spyglass’ interactive TV software; Metrowerks releases CodeWarrior with PersonalJava support. Computer, 33(6):96–98, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r6096.pdf.

Lutz:2000:NBS

[144] Michael J. Lutz. New books: Simplifying with design patterns; my first robot; improve software development with SCM; running businesses with data warehouses; UML-powered Web apps; an intro to building computer architectures; business metamorphosis. Computer, 33(6):99, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r6100.pdf.

Varshney:2000:CRAb

[145] Upkar Varshney. Communications: Recent advances in wireless networking. Computer, 33(6):100–103, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r6100.pdf.

Macedonia:2000:ECE

[146] Michael Macedonia. Entertainment computing: The empire strikes back … with the X-box. Computer, 33(6):104–106, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r6104.pdf.

Waldo:2000:STA

[147] Jim Waldo. Software technologies: Alive and well: Jini technology today. Computer, 33(6):107–109, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r6107.pdf.

Lewis:2000:BCE

[148] Ted Lewis. Binary critic: The end of research as we know it? Computer, 33(6):112, 110–111, June 2000.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r6112.pdf.

Lyon:2000:LWS

[149] Douglas Lyon, Roger T. Alexander, James M. Bieman, John Viega, Bruce D. Wilner, Roger T. Alexander, James M. Bieman, John Viega, Derek B. Eckert, Ted Lewis, Howard C. Anderson, Thomas M. English, and Timothy C. Lethbridge. Letters: We shape our tools and our tools shape us; Java design flaws create programming stress; funny money’s role in the government’s favorite hobby; rigorous training teaches students how to think. Computer, 33(7):4–7, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7004.pdf.

Lawton:2000:ITV

[150] George Lawton. Industry trends: Video streams into the mainstream. Computer, 33(7):12–17, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7012.pdf.

Ortiz:2000:TNB

[151] Sixto Ortiz Jr. Technology news: Broadband fixed wireless travels the last mile. Computer, 33(7):18–21, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7018.pdf.

Lear:2000:NBD

[152] Anne C. Lear. News briefs: Do E-business patents help or hurt?; small wonder: MEMS will power optical Internet; rough seas ahead for US-centric Web sites; new protocol simplifies B2B communication; uncovering technology’s human side. Computer, 33(7):22–24, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7022.pdf.

Anonymous:2000:ASg

[153] Anonymous. Article summaries. Computer, 33(7):25–26, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7025.pdf.

Henning:2000:SCM

[154] John L. Henning. SPEC CPU2000: Measuring CPU performance in the new millennium. Computer, 33(7):28– 35, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2000/pdf/r7028.pdf;
http://www.computer.org/computer/
co2000/r7028abs.htm;   http://www.
spec.org/cpu2000/papers/COMPUTER_
200007.JLH.pdf;   http://www.specbench.
org/osg/cpu2000/.

Goodman:2000:WIP

[155] David J. Goodman. The wireless Internet: Promises and challenges. Computer, 33(7):36–41, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7036.pdf;
http://www.computer.org/computer/
co2000/r7036abs.htm.

Rabaey:2000:PSA

[156] Jan M. Rabaey, M. Josie Ammer, Julio L. da Silva, Jr., Danny Patel, and Shad Roundy. PicoRadio supports ad hoc ultra-low power wireless networking. Computer, 33(7):42–48, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7042.pdf;
http://www.computer.org/computer/
co2000/r7042abs.htm.

Snow:2000:RSW

[157] Andrew P. Snow, Upkar Varshney, and Alisha D. Malloy. Reliability and survivability of wireless and mobile networks. Computer, 33(7):49–55, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7049.pdf;
http://www.computer.org/computer/
co2000/r7049abs.htm.

Ball:2000:UPM

[158] Thomas Ball and James R. Larus. Using paths to measure, explain, and enhance program behavior. Computer, 33(7):57–65, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7057.pdf;
http://www.computer.org/computer/
co2000/r7057abs.htm.

Rajlich:2000:SMS

[159] Václav T. Rajlich and Keith H. Bennett. A staged model for the software life cycle. Computer, 33(7):66–71, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7066.pdf;
http://www.computer.org/computer/
co2000/r7066abs.htm.

Shull:2000:HPB

[160] Forrest Shull, Ioana Rus, and Victor Basili. How perspective-based reading can improve requirements inspections. Computer, 33(7):73–79, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7073.pdf;
http://www.computer.org/computer/
co2000/r7073abs.htm.

Piner:2000:CSCg

[161] Mary-Louise G. Piner. Computer Society connection: Ten teams to compete in CSIDC world finals: Editorin-Chief appointments made for 2001; candidates named for Society Board and Vice President offices; Board approves readings of two bylaws amendments; Computer Society members named to US National Academia of Engineering. Computer, 33(7):80–84, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7080.pdf.

Anonymous:2000:CCg

[162] Anonymous. Call and calendar. Computer, 33(7):85–88, July 2000.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7085.pdf.

Anonymous:2000:COf

[163] Anonymous. Career opportunities. Computer, 33(7):90–92, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7090.pdf.

Lutz:2000:NBMb

[164] Michael J. Lutz. New books: Modular design: Key to innovation and growth; 3G mobile communication; Bourne for Linux and Unix; electronic U; soft computing: Rolling with the punches; XML plans for managers. Computer, 33(7):93, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7093.pdf.

Kroeker:2000:PNS

[165] Kirk L. Kroeker. Products: L-3 Network Security’s proactive security measurement tool; US Software’s distributed operating environment; Microtest’s fiber-optic tester; Transim’s Internet-based IC simulation tool; CAD-UL’s toolkit for embedded processors; AMC’s embedded software visibility tools; International Test Technologies’s PowerPC testing tools. Computer, 33(7):94–96, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7094.pdf.

Holmes:2000:PFF

[166] Neville Holmes. The profession: Fashioning a foundation for the computing profession. Computer, 33(7):97–98, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2000/pdf/r7097.pdf.

Boehm:2000:SMR

[167] Barry Boehm. Software management: Requirements that handle IKIWISI, COTS, and rapid change. Computer, 33(7):99–102, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7099.pdf.

Milutinovic:2000:IEO

[168] Veljko Milutinovic, Predrag Knezevic, Bozidar Radunovic, Steve Casselman, and John Schewel. Integrated engineering: Obelix searches Internet using customer data. Computer, 33(7):104–107, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7104.pdf.

Michener:2000:IWM

[169] John R. Michener and Tolga Acar. Internet watch: Managing system and active-content integrity. Computer, 33(7):108–110, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7108.pdf.

Lewis:2000:BCWc

[170] Ted Lewis. Binary critic: The Web fosters a rust-belt renaissance. Computer, 33(7):112–112, 111–111, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r7112.pdf.

Thompson:2000:LMF

[171] J. Barrie Thompson, Gordon Knox, Thomas M. English, Gilda Pour, Martin L. Griss, Michael Lutz, Tom Rauscher, Ted Lewis, Richard A. Martin, Ted Lewis, and Gary Leavens. Letters: A maturing but fragmented profession; making the changes necessary to survive; creating incentives for academic research; academic interest in component technology. Computer, 33(8):4–4, 6–8, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8004.pdf.

Piner:2000:RMM

[172] Mary-Louise G. Piner. Report to members: McMaster students win top prize at CSIDC 2000: Heart Mate takes top honors; behind the secnes for CSIDC 2000; Computer Society Executive Director retires after 18 years. Computer, 33(8):9–11, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8009.pdf.

Leavitt:2000:ITW

[173] Neal Leavitt. Industry trends: Whatever happened to object-oriented databases? Computer, 33(8):16–19, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2000/pdf/r8016.pdf.

Clark:2000:TNE

[174] David Clark. Technology news: Encryption advances to meet Internet challenges. Computer, 33(8):20–24, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2000/pdf/r8020.pdf.

Garber:2000:NBC

[175] Lee Garber. News briefs: Companies power up networking; magnetic RAM research reaching critical stage; Micro Java is a platform under construction; Lotus offers on-the-fly translation. Computer, 33(8):26–28, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8026.pdf.

Anonymous:2000:ASh

[176] Anonymous. Article summaries. Computer, 33(8):29–30, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8029.pdf.

Cover:2000:DUB

[177] Jeffrey Voas. Developing a usagebased software certification process. Computer, 33(8):32–37, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8032.pdf;

http://www.computer.org/computer/
co2000/r8032abs.htm.

Jones:2000:CBS

[178] Anita Jones. The challenge of building survivable information-intensive systems. Computer, 33(8):39–43, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8039.pdf.

Amin:2000:TSH

[179] Massoud Amin. Toward selfhealing infrastructure systems. Computer, 33(8):44–53, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8044.pdf;
http://www.computer.org/computer/
co2000/r8044abs.htm.

Birman:2000:NGI

[180] Kenneth P. Birman. The nextgeneration Internet: Unsafe at any speed? Computer, 33(8):54–60, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8054.pdf;
http://www.computer.org/computer/
co2000/r8054abs.htm.

Wylie:2000:SIS

[181] Jay J. Wylie, Michael W. Bigrigg, John D. Strunk, Gregory R. Ganger, Han Kiliççöte, and Pradeep K. Khosla. Survivable information storage systems. Computer, 33(8):61–68, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8061.pdf;

http://www.computer.org/computer/
co2000/r8061abs.htm.

Brown:2000:IAC

[182] Donald E. Brown, Louise E. Gunderson, and Marc H. Evans. Interactive analysis of computer crimes. Computer, 33(8):69–77, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8069.pdf;
http://www.computer.org/computer/
co2000/r8069abs.htm.

Williams:2000:FRF

[183] T. W. Williams. Financial report: 1999 financial results. Computer, 33(8):78–83, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8078.pdf.

Anonymous:2000:CCh

[184] Anonymous. Call and calendar. Computer, 33(8):84–86, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8084.pdf.

Anonymous:2000:COg

[185] Anonymous. Career opportunities. Computer, 33(8):88–94, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8088.pdf.

Lutz:2000:NBCb

[186] Michael J. Lutz. New books: The comeback kid returns to Apple helm; behavioral patterns: Shaping business software; customizing software development; working with concurrent programs; improving algorithms; gathering and defining software requirements. Computer, 33(8):95, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8095.pdf.

Kroeker:2000:PSX

[187] Kirk L. Kroeker. Products: SoftQuad’s XML content creation tool; Premia’a application source code browser for MFC; Rogue Wave Software’s Linux middleware framework; eHelp’s help authoring system for serverbased applications; Bluestone Software’s total-e-server; NetworkView Software’s network mapping utility; Poet Software’s object database; Vital’s FreeBSD editor for programmers. Computer, 33(8):96–98, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8096.pdf.

Macedonia:2000:ECD

[188] Michael Macedonia. Entertainment computing: Distributed file sharing: Barbarians at the gates? Computer, 33(8):99–101, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8099.pdf.

Hay:2000:CCP

[189] Richard Hay. Communications: Comparing POS and ATM interfaces. Computer, 33(8):102–103, August 2000.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8102.pdf.

Danielsen:2000:STP

[190] Peter J. Danielsen. Software technologies: The promise of a voice-enabled Web. Computer, 33(8):104–106, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2000/pdf/r8104.pdf.

Chadwick:2000:SWT

[191] David Chadwick. Standards: Windows 2000: A threat to Internet diversity and open standards? Computer, 33(8):107–109, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8107.pdf.

Lewis:2000:BCWd

[192] Ted Lewis. Binary critic: Why WAP may never get off the ground. Computer, 33(8):112–112, 110–111, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r8112.pdf.

Swaing:2000:LCC

[193] Mark Swaing, Neville Holmes, Mark Wallace, Marc Lupien, and Ted Lewis. Letters: Computing: Craft, trade, profession; what components revolution?; software patent clarification. Computer, 33(9):4, 6, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r9004.pdf.

Lawton:2000:ITW

[194] George Lawton. Industry trends: The wild world of 3D graphics chips. Computer, 33(9):12–16, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r9012.pdf.

Anonymous:2000:ASi

[195] Anonymous. Article summaries. Computer, 33(9):17, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r9017.pdf.

Lee:2000:WAE

[196] Edward A. Lee. What’s ahead for embedded software? Computer, 33(9):18–26, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r9018.pdf;
                                        http://www.computer.org/computer/
                                        co2000/r9018abs.htm.

Olson:2000:SIE

[197] Michael A. Olson. Selecting and implementing an embedded database system. Computer, 33(9):27–34, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co2000/pdf/r9027.pdf;
                                         http://www.computer.org/computer/
                                        co2000/r9027abs.htm.

Lea:2000:NHE

[198] Rodger Lea, Simon Gibbs, Alec Dara-Abrams, and Edward Eytchison. Networking home entertainment devices with HAVi. Computer, 33(9):35–43, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r9035.pdf;
http://www.computer.org/computer/
co2000/r9035abs.htm.

Talia:2000:CPT

[199] Domenico Talia. Cellular processing tools for high-performance simulation. Computer, 33(9):44–52, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r9044.pdf;
http://www.computer.org/computer/
co2000/r9044abs.htm.

Anonymous:2000:CSE

[200] Anonymous. Computer Society election. Computer, 33(9):54–62, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2000/pdf/r9054.pdf.

Piner:2000:CSCh

[201] Mary-Louise G. Piner. Computer Society connection: IEEE President-Elect candidates share their views on key society issues. Computer, 33(9):64–69, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r9064.pdf.

Anonymous:2000:CCi

[202] Anonymous. Call and calendar. Computer, 33(9):70–72, 74, 76, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r9070.pdf.

Anonymous:2000:COh

[203] Anonymous. Career opportunities. Computer, 33(9):78–86, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r9078.pdf.

Lutz:2000:NBG

[204] Michael J. Lutz. New books: Getting the kinks out of Linux; words, words, words; understanding system on a chip; high-level digital receiver techniques; acing your technical presentation; intelligent behavior-based robotic systems. Computer, 33(9):87, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/r9087.pdf.

Kroeker:2000:PMD

[205] Kirk L. Kroeker. Products: Metrowerks’ development tools for novice programmers; EmmaSoft’s software installation tool; Eurosoft’s disk cloning tool; OnCore Systems’ tools for embedded processors; Paessier Tools’ server stress test tool; Logic Programming Associates’ AI development tools; FileStream’s archive tool now supports JAR format; SRC’s tool for converting 4GL code to Cobol. Computer, 33(9):88–90, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r9088.pdf.

Holmes:2000:POG

[206] Neville Holmes. The profession: Olympic Games reform: A study in system engineering. Computer, 33(9):91–93, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r9091.pdf.

Boehm:2000:SMP

[207] Barry Boehm. Software management project: Termination doesn’t equal project failure. Computer, 33(9):94– 96, September 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2000/pdf/
r9094.pdf.

Ronald:2000:IWC

[208] Edmund M. A. Ronald and Moshe Sipper. Internet watch: The challenge of tamperproof Internet computing. Computer, 33(9):98–99, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r9098.pdf.

Lewis:2000:BCX

[209] Ted Lewis. Binary critic: XM and Sirius get serious about broadcast satellite radio. Computer, 33 (9):104, 102–103, September 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r9104.pdf.

Bollinger:2000:LWW

[210] Terry Bollinger, Ted Lewis, Amit Sarkar, Ted Lewis, Vir V. Phoba, and Rekha Bangalore. Letters: To WAP or not to WAP?; global response infrastructure crucial for Web security. Computer, 33(10):4, 6, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rx004.pdf.

Paulson:2000:ITE

[211] Linda Dailey Paulson. Industry trends: Exploring the wireless: LANscape. Computer, 33(10):12–16, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rx012.pdf.

Garber:2000:NBL

[212] Lee Garber. News briefs: Linux advances on both desktop and palmtop?; does your computer know where you are?; changes to UCITA may not end controversy?; Cisco’s training program hits the streets. Computer, 33(10):17–20, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rx017.pdf;
http://www.gnome.org.

Anonymous:2000:ASj

[213] Anonymous. Article summaries. Computer, 33(10):21–22, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rx021.pdf.

Prechelt:2000:ECS

[214] Lutz Prechelt. An empirical comparison of seven programming languages. Computer, 33(10):23–29, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rx023.pdf;
http://www.computer.org/computer/
co2000/rx023abs.htm.

Weaver:2000:GEI

[215] Alfred C. Weaver, Ronald J. Vetter, Andrew B. Whinston, and Kathleen Swigger. Guest Editors’ introduction: The future of E-commerce. Computer, 33(10):30–31, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rx030.pdf.

Varshney:2000:MCN

[216] Upkar Varshney, Ronald J. Vetter, and Ravi Kalakota. Mobile commerce: A new frontier. Computer, 33(10):32–38, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rx032.pdf;
http://www.computer.org/computer/
co2000/rx032abs.htm.

Shim:2000:BBC

[217] Simon S. Y. Shim, Vishnu S. Pendyala, Meera Sundaram, and Jerry Z. Gao. Business-to-business E-commerce frameworks. Computer, 33(10):40–47, October 2000.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rx040.pdf;
http://www.computer.org/computer/
co2000/rx040abs.htm.

Wang:2000:PDA

[218] Wenli Wang, Zoltán Hidvégi, Andrew D. Bailey, Jr., and Andrew B. Whinston. E-process design and assurance using model checking. Computer, 33(10):48–53, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rx048.pdf;
http://www.computer.org/computer/
co2000/rx048abs.htm.

Ebringer:2000:PAP

[219] Tim Ebringer, Peter Thorne, and Yuliang Zheng. Parasitic authentication to protect your E-wallet. Computer, 33(10):54–60, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rx054.pdf;
http://www.computer.org/computer/
co2000/rx054abs.htm.

Piner:2000:CSCi

[220] Mary-Louise G. Piner. Computer Society connection: Awards recognize innovation, service; Computer Society prizes high school innovation; what’s new in Computer Society periodicals. Computer, 33(10):62–66, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rx062.pdf.

Anonymous:2000:CCj

[221] Anonymous. Call and calendar. Computer, 33(10):67–70, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rx067.pdf.

Anonymous:2000:COi

[222] Anonymous. Career opportunities. Computer, 33(10):72–92, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rx067.pdf.

Lutz:2000:NBU

[223] Michael J. Lutz. New books: Unified software history; speeding up the PC; Dell’s boy wonder; improving the machine; multiagent application and theory; working with the M-CORE embedded system; random methods. Computer, 33(10):93, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rx093.pdf.

Kroeker:2000:PNN

[224] Kirk L. Kroeker. Products: Novell’s newest remote management software; White Sands Technology’s highend database tool; Microsoft’s Visual Studio 6 for developing office apps; PassMark Software’s performance analysis tool for the masses; BSquare’s SDK for USB hardware control; Brain Patchwork’s Internet server developing toolkit; American Systems’ screen capture utility; Jn-Software’s SQL server dictionary reports tool; Microtest’s home network test tool. Computer, 33(10):94–97, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/rx094.pdf.

Lui-Kwan:2000:ECF

[225] Gerald Lui-Kwan. Entertainment computing: In-flight entertainment: The sky’s the limit. Computer, 33(10): 98–101, October 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2000/pdf/
                                        rx098.pdf.

Isaak:2000:SNL

[226] Jim Isaak. Standards: New laws and technologies morph Web page best practices. Computer, 33(10):102– 103, October 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2000/pdf/
                                        rx102.pdf.

Stork:2000:STU

[227] David G. Stork. Software technologies: Using open data collection for intelligent software. Computer, 33(10):104–106, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/rx104.pdf.

Samaha:2000:CGI

[228] Joseph Samaha, Ayman Kayssi, Ralph Achkar, and Mark Azar. Communications: G3 integrates three system technologies. Computer, 33(10):107–110, October 2000.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rx107.pdf.

Lewis:2000:BCI

[229] Ted Lewis. Binary critic: Innovation equals risk reduction; A farewell note. Computer, 33(10):112, 111–111, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2000/pdf/rx112.pdf.

StVille:2000:LBE

[230] Lyndall St. Ville, Neville Holmes, Wojciech Mikanik, Ted Lewis, Kevin J. Henderson, and George Lawton. Letters: Bird’s-eye view puts Olympic Games in focus; thanks to the critic; who’s the leader? Computer, 33(11):4, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/ry004.pdf.

Clark:2000:ITH

[231] David Clark. Industry trends: Handhelds drive mixed-signal chip development. Computer, 33(11):12– 15, November 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2000/pdf/
ry012.pdf.

Merton:2000:NBG

[232] Orren Merton and Linda Dailey Paulson. News briefs: Gamers jump into broadband technology; Intel has new chip design for handhelds; patent expiration begins new encryption era; privacy organization raises privacy concerns. Computer, 33(11):16–19, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/ry016.pdf.

Anonymous:2000:ASk

[233] Anonymous. Article summaries. Computer, 33(11):20–21, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/ry020.pdf.

Talagala:2000:AMS

[234] Nisha Talagala, Satoshi Asami, David Patterson, Bob Futernick, and Dakin Hart. The art of massive storage: A Web image archive. Computer, 33(11):22–28, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/ry022.pdf;
                                        http://www.computer.org/computer/
                                        co/ry022abs.htm.

Hahn:2000:CAS

[235] Udo Hahn and Inderjeet Mani. The challenges of automatic summarization. Computer, 33(11):29–36, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co2000/pdf/ry029.pdf;
                                         http://www.computer.org/computer/
                                        co/ry029abs.htm.

Ziviani:2000:CKN

[236] Nivio Ziviani, Edleno Silva de Moura, Gonzalo Navarro, and Ricardo Baeza-Yates. Compression: A key for next-generation text retrieval systems. Computer, 33(11):37–44, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/ry037.pdf;
http://www.computer.org/computer/
co/ry037abs.htm.

content:2000:NGW

[237] Michael S. Lew. Next-generation Web searches for visual content. Computer, 33(11):46–53, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/ry046.pdf;
http://www.computer.org/computer/
co/ry046abs.htm.

Brereton:2000:CBS

[238] Pearl Brereton and David Budgen. Component-based systems: A classification of issues. Computer, 33(11):54–62, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/ry054.pdf;
http://www.computer.org/computer/
co/ry054abs.htm.

Estrin:2000:NVN

[239] Deborah Estrin, Mark Handley, John Heidemann, Steven McCanne, Ya Xu, and Haobo Yu. Network visualization with Nam, the VINT network animator. Computer, 33(11):63–68, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/

                                        co/books/co2000/pdf/ry063.pdf;
                                        http://www.computer.org/computer/
                                        co/ry063abs.htm.

Piner:2000:CSCj

[240] Mary-Louise G. Piner. Computer Society connection: Second annual design competition seeks student participants; Society presents new award for undergraduate teachers: [Joseph Zachary recognized for innovative educational materials; Timothy Long and Ohio State colleague Bruce Weide honored]; IEEE Fellow Jack Kilby receives Nobel Prize in Physics. Computer, 33(11):70–73, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/ry070.pdf.

Anonymous:2000:CCk

[241] Anonymous. Call and calendar. Computer, 33(11):74–78, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/ry074.pdf.

Anonymous:2000:COj

[242] Anonymous. Career opportunities. Computer, 33(11):80–114, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lutz:2000:NBI

[243] Michael J. Lutz. New books: Improving your delivery; principles of system architecture; working with domain-oriented reuse; creating the perfect gaming engine; ARM design; rational agents’ architecture; building a better database. Computer, 33(11):115, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/ry115.pdf.

Anonymous:2000:PBA

[244] Anonymous. Products: Broadcom adds VoIP and home networking to cable modem chip; CodeWarrior 6.0 for the Mac; Inprise/Borland JBuilder 4; WinDK extension for Bluetooth; System Mechanic Mobile Toolkit; ActiveState Perl Dev Kit 2.0 for Perl 5.6; Pentium III with SpeedStep; Progress Software Apptivity version 3. Computer, 33(11):116–117, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/ry116.pdf.

Milutinovic:2000:IWG

[245] Veljko Milutinovic, Dragana Cvetkovic, and Jelena Mirkovic. Internet watch: Genetic search based on multiple mutations. Computer, 33(11):118– 119, November 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2000/pdf/
ry118.pdf.

Boehm:2000:SMAb

[246] Barry Boehm, Dan Port, and Mohammed Al-Said. Software management: Avoiding the software model-clash spiderweb. Computer, 33(11):120–122, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/ry120.pdf.

Borriello:2000:IEC

[247] Gaetano Borriello. Integrated engineering: The challenges to invisible computing. Computer, 33(11):123– 127, November 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2000/pdf/
                                        ry123.pdf.

Holmes:2000:PSC

[248] Neville Holmes. The profession: Some comments on the coding of programs. Computer, 33(11):128, 126– 127, November 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2000/pdf/
                                        ry128.pdf.

Budgen:2000:LTM

[249] David Budgen, Ted Lewis, Daniel Alex Finkelstein, and Lutz Prechelt. Letters: Are telephones made for ears or eyes?; when it comes to languages, people matter most. Computer, 33(12):4, 6, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/rz004.pdf.

Piner:2000:RMS

[250] Mary-Louise G. Piner. Report to members: Society elects new leadership for 2001. Computer, 33(12):8–9, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/rz008.pdf.

Anonymous:2000:ICS

[251] Anonymous. IEEE Computer Society membership application. Computer, 33(12):10–12, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rz010.pdf.

Ortiz:2000:ITW

[252] Sixto Ortiz Jr. Industry trends: Will PKI become a key to online security? Computer, 33(12):13–15, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rz013.pdf.

Leavitt:2000:TNM

[253] Neal Leavitt. Technology news: Malicious code moves to mobile devices. Computer, 33(12):16–19, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rz016.pdf.

Paulson:2000:NBU

[254] Linda Dailey Paulson and Orren Merton. News briefs: U.S. picks new encryption standard; better software with open source; taking a SIP of Internet telephony; schools may hold valuable spectrum; getting a feel for the Web. Computer, 33(12):20–23, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rz020.pdf.

Anonymous:2000:ASl

[255] Anonymous. Article summaries. Computer, 33(12):24–25, December 2000.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rz024.pdf.

Elliott:2000:RYC

[256] T. Michael Elliott. Reflections on 18 years of change at the Computer Society. Computer, 33(12):26–29, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2000/pdf/rz026.pdf.

Bjorkander:2000:GPU

[257] Morgan Björkander. Graphical programming using UML and SDL. Computer, 33(12):30–35, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rz030.pdf;
http://www.computer.org/computer/
co2000/rz030abs.htm.

Whittaker:2000:TMR

[258] James A. Whittaker and Jeffrey Voas. Toward a more reliable theory of software reliability. Computer, 33(12):36–42, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rz036.pdf;
http://www.computer.org/computer/
co2000/rz036abs.htm.

Longstaff:2000:WFR

[259] Thomas A. Longstaff, Clyde Chittister, Rich Pethia, and Yacov Y. Haimes. Are we forgetting the risks of information technology? Computer, 33(12):43–51, December 2000.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rz043.pdf;
http://www.computer.org/computer/
co2000/rz043abs.htm.

Arbaugh:2000:WVC

[260] William A. Arbaugh, William L. Fithen, and John McHugh. Windows of vulnerability: A case study analysis. Computer, 33(12):52–59, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rz052.pdf;
http://www.computer.org/computer/
co2000/rz052abs.htm.

Butler:2000:NSA

[261] Randy Butler, Von Welch, Douglas Engert, Ian Foster, Steven Tuecke, John Volmer, and Carl Kesselman. A national-scale authentication infrastructure. Computer, 33(12):60–66, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rz060.pdf;
http://www.computer.org/computer/
co2000/rz060abs.htm.

Chilimbi:2000:MPB

[262] Trishul M. Chilimbi, Mark D. Hill, and James R. Larus. Making pointerbased data structures cache conscious. Computer, 33(12):67–75, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rz067.pdf;
http://www.computer.org/computer/
co2000/rz067abs.htm.

Anonymous:2000:AIC

[263] Anonymous. Annual index: Computer, volume 33, 2000. Computer, 33(12):76–91, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/rz076.pdf.

Anonymous:2000:CSC

[264] Anonymous. Computer Society connection: St. Petersburg State professor wins CiSE educational software contest; Computer recognizes reviewers’ expertise; Merwin scholarship awarded to four students. Computer, 33(12):92–95, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/rz092.pdf.

Anonymous:2000:CCl

[265] Anonymous. Call and calendar. Computer, 33(12):96–99, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/rz096.pdf.

Anonymous:2000:COk

[266] Anonymous. Career opportunities. Computer, 33(12):100–143, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/rz100.pdf.

Anonymous:2000:POR

[267] Anonymous. Products: Oracle releases XDK update; Starbase’s code editing system; Arc Second’s palm PC CAD viewer; Minolta’s network document server for Windows 2000; Borland’s Java development tools for Palm OS; Rational’s code management tools; Blaxxun Interactive’s Web communications platform tools; Informix Software’s Linux database engine; ActiveState updates free Python distribution; KDE 2.0 released. Computer, 33(12):144–146, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rz144.pdf.

Lutz:2000:NBA

[268] Michael J. Lutz. New books: Atoms for tomorrow’s child?; satellite communications; keep on tracking!; A paradigm shift for speech processing; parallel programming portability. Computer, 33(12):147, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rz147.pdf.

Senn:2000:CEM

[269] James A. Senn. Communications: The emergence of M-commerce. Computer, 33(12):148–151, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/rz148.pdf.

Mason:2000:SSI

[270] Harry Mason. Standards: SCSI, the industry workhorse, is still working hard. Computer, 33(12):152–153, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/rz152.pdf.

Rhyne:2000:ECC

[271] Theresa-Marie Rhyne. Entertainment computing: Computer games’ influence on scientific and information visualization. Computer, 33(12):154–159, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co2000/pdf/rz154.pdf.

Holmes:2000:PWJ

[272] Neville Holmes. The profession: Why Johnny can’t program. Computer, 33(12):160, 158–159, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2000/pdf/rz160.pdf.

Aylor:2001:EMC

[273] James H. Aylor. EIC’s message: Can we work together? Computer, 34(1):4, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r1004.pdf.

Wah:2001:PMP

[274] Benjamin W. Wah. President’s message: Pushing forward into the third millennium. Computer, 34(1):6–8, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co2001/pdf/r1006.pdf.

Anonymous:2001:SME

[275] Anonymous. A salute to T. Michael Elliott. Computer, 34(1):7, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r1006.pdf.

Kropf:2001:LTD

[276] Markus Kropf, Michael Grunberg, Neville Holmes, Daniel Mall, Neville Holmes, and Ana Asuaga. Letters: Thoughts on distributed file sharing; is Holmes out to get us? Computer, 34(1):10–12, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r1010.pdf.

Clark:2001:ITF

[277] David Clark. Industry trends: Faceto-face with peer-to-peer networking. Computer, 34(1):18–21, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r1018.pdf.

Garber:2001:TNC

[278] Lee Garber. Technology news: Computer forensics: High-tech law enforcement. Computer, 34(1):22–27, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r1022.pdf.

Merton:2001:NBI

[279] Orren Merton and Linda Dailey Paulson. News briefs: Initiative takes long view of tele-immersion; will wireless be IPv6’s killer app?; Asian domain names to be accepted; chips hot off the printer; copyright ruling generates concern. Computer, 34(1):28–30, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r1028.pdf.

Anonymous:2001:ASa

[280] Anonymous. Article summaries. Computer, 34(1):31–32, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r1031.pdf.

Zhirnov:2001:NFS

[281] Victor V. Zhirnov and Daniel J. C. Herr. New frontiers: Self-assembly and nanoelectronics. Computer, 34(1): 34–43, January 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2001/pdf/
                                        r1034.pdf;   http://www.computer.
                                        org/computer/co2001/r1034abs.htm.

Warneke:2001:SDC

[282] Brett Warneke, Matt Last, Brian Liebowitz, and Kristofer S. J. Pister. Smart dust: Communicating with a cubic-millimeter computer. Computer, 34(1):44–51, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r1044.pdf;
                                        http://www.computer.org/computer/
                                        co2001/r1044abs.htm.

modeling:2001:PSC

[283] David Harel. From play-in scenarios to code: An achievable dream. Computer, 34(1):53–60, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r1053.pdf;
http://www.computer.org/computer/
co2001/r1053abs.htm.

Human:2001:TOB

[284] Scott Hamilton. Thinking outside the box at IHMC. Computer, 34(1):61–71, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r1061.pdf;
http://www.computer.org/computer/
co2001/r1061abs.htm.

Lindheim:2001:FNS

[285] Richard Lindheim and William Swartout. Forging a new simulation technology at the ICT. Computer, 34(1):72–79, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2001/pdf/r1072.pdf;
 http://www.computer.org/computer/
co2001/r1072abs.htm.

Back:2001:DIR

[286] Maribeth Back, Rich Gold, Anne Balsamo, Mark Chow, Matt Gorbet, Steve Harrison, Dale MacDonald, and Scott Minneman. Designing innovative reading experiences for a museum exhibition. Computer, 34(1):80–87, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r1080.pdf;
http://www.computer.org/computer/
co2001/r1080abs.htm.

Piner:2001:CSCa

[287] Mary-Louise G. Piner. Computer Society connection: Make the most of your membership in 2001: Editor-in-Chief positions open for 2002. Computer, 34(1):88–94, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r1088.pdf.

Anonymous:2001:CCa

[288] Anonymous. Call and calendar. Computer, 34(1):95–98, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r1095.pdf.

Anonymous:2001:COa

[289] Anonymous. Career opportunities. Computer, 34(1):100–128, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r1100.pdf.

Lutz:2001:NBIa

[290] Michael J. Lutz. New books: Impetus to componentization; computing in a data-centric world; digital certificates security; quantum computing/enigma — “jabberwocky” argument?; interfacing with biometry. Computer, 34(1):129, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r1129.pdf.

Anonymous:2001:PCP

[291] Anonymous. Products: Cross-platform toolkit for Bristol Technology; InstallShield updates Windows installer; Droplet offers unique Web application SDK; ObjectFX Corporation’s Web-based visualization software; Basis Technology updates C++ library; MathWorks unveils embedded control design suite; Intuitive Systems offers Java profiling tool; Computer Associates’ E-business development environment; Programming Research Ltd.’s static analysis tool. Computer, 34(1):130–131, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r1130.pdf.

Berghel:2001:IWC

[292] Hal Berghel. Internet watch: Cyberprivacy in the new millennium. Computer, 34(1):132–134, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r1132.pdf.

Boehm:2001:SMS

[293] Barry Boehm and Victor R. Basili. Software management: Software defect reduction top 10 list. Computer, 34(1):135–137, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r1135.pdf.

Sveda:2001:IEE

[294] Miroslav Sveda and Radimir Vrba. Integrated engineering: Executable specifications for embedded distributed systems. Computer, 34(1):138–140, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2001/pdf/r1138.pdf.

Holmes:2001:PKD

[295] Neville Holmes. The profession: The KWIC and the dead: A lesson in computing history. Computer, 34(1):144, 142–143, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r1144.pdf.

Holten:2001:LJC

[296] Jim Holten, Neville Holmes, John Fitzpatrick, Neville Holmes, Kevin Kornelson, Neville Holmes, William Adams, Neville Holmes, and Eric Weiss. Letters: Johnny’s craft requires training; foxes and hedgehogs; programming and animation; systems and managers; keep on keeping on. Computer, 34(2):4–7, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r2004.pdf.

Lawton:2001:ITK

[297] George Lawton. Industry trends: Knowledge management: Ready for prime time? Computer, 34(2):12– 15, February 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2001/pdf/
                                        r2012.pdf.

Miller:2001:TNT

[298] Sandra Kay Miller. Technology news: The trusted OS makes a comeback. Computer, 34(2):16–19, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r2016.pdf.

Paulson:2001:NBI

[299] Linda Dailey Paulson and Orren Merton. News briefs: Intel talks of tiny transistors; DOM upgrade bolsters XML; the Cyber Corps wants you; building the crash-proof computer. Computer, 34(2):20–22, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r2020.pdf.

Anonymous:2001:ASb

[300] Anonymous. Article summaries. Computer, 34(2):24–25, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r2024.pdf.

Lawrence:2001:PWR

[301] Steve Lawrence, David M. Pennock, Gary William Flake, Robert Krovetz, Frans M. Coetzee, Eric Glover, Finn Årup Nielsen, Andries Kruger, and C. Lee Giles. Persistence of Web references in scientific research. Computer, 34(2):26–31, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r2026.pdf;
http://www.computer.org/computer/
co2001/r2026abs.htm.

Elmagarmid:2001:OMT

[302] Ahmed K. Elmagarmid and William J. McIver, Jr. The ongoing march toward digital government. Computer, 34(2):32–38, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r2032.pdf.

Mecella:2001:EIG

[303] Massimo Mecella and Carlo Batini. Enabling Italian E-government through a cooperative architecture. Computer, 34(2):40–45, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r2040.pdf;
http://www.computer.org/computer/
co2001/r2040abs.htm.

Ambite:2001:SDA

[304] José Luis Ambite, Yigal Arens, Eduard Hovy, Andrew Philpot, Luis Gravano, Vasileios Hatzivassiloglou, and Judith Klavans. Simplifying data access: The energy data collection project. Computer, 34(2):47–54, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r2047.pdf;
http://www.computer.org/computer/
co2001/r2047abs.htm.

Bouguettaya:2001:MGD

[305] Athman Bouguettaya, Mourad Ouzzani, Brahim Medjahed, and Jerry Cameron. Managing government databases. Computer, 34(2):56–64, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2001/pdf/
r2056.pdf;   http://www.computer.
org/computer/co2001/r2056abs.htm.

Joshi:2001:DGS

[306] James Joshi, Arif Ghafoor, Walid G. Aref, and Eugene H. Spafford.

Digital government security infrastructure design challenges. Computer, 34(2):66–72, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r2066.pdf;
http://www.computer.org/computer/
co2001/r2066abs.htm.

Adomavicius:2001:UDM

[307] Gediminas Adomavicius and Alexander Tuzhilin. Using data mining methods to build customer profiles. Computer, 34(2):74–82, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r2074.pdf;
http://www.computer.org/computer/
co2001/r2074abs.htm.

Piner:2001:CSCb

[308] Mary-Louise G. Piner. Computer Society connection: Visionary Glen Culler named Cray Award winner; new IEEE membership option focuses on women in engineering; William R. Hewlett, 1913–2001; Cray Award recognizes creativity in highperformance computing; 2001 Fellows named for distinguished achievements; design competition teams selected for CSIDC 2001; Transactions on Software Engineering seeks Editor in Chief; new election process begins. Computer, 34(2):84–90, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r2084.pdf.

Anonymous:2001:CCb

[309] Anonymous. Call and calendar. Computer, 34(2):91–94, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r2091.pdf.

Anonymous:2001:APIa

[310] Anonymous. Advertiser/product index. Computer, 34(2):95, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r2091.pdf.

Anonymous:2001:COb

[311] Anonymous. Career opportunities. Computer, 34(2):96–111, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r2096.pdf.

Lutz:2001:NBIb

[312] Michael J. Lutz. New books: Investigating evidential computing; steering the computing juggernaut; in the mood for Java, Smalltalk, and algorithms?; meeting the computational intelligence challenge; assessing the digital economy. Computer, 34(2):112, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r2112.pdf.

Anonymous:2001:PVR

[313] Anonymous. Products: VentureCom’s real-time extension developer’s tool; PowerSteering’s software development kit; MathWorks’ technical development platform; Lineo’s embedded device SDK; RadView’s application testing package; StereoGraphic’s stereoscopic wireless visualization system. Computer, 34(2):113–114, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r2113.pdf.

Macedonia:2001:ECI

[314] Michael Macedonia. Entertainment computing: Innovative computing powers theme park adventures. Computer, 34(2):115–117, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r2115.pdf.

Haungs:2001:STP

[315] Jim Haungs. Software technologies: Pair programming on the C3 Project. Computer, 34(2):118–119, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r2118.pdf.

Spring:2001:SCD

[316] Michael B. Spring. Standards: The courtship dance of bulldogs and turtles. Computer, 34(2):120–122, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r2120.pdf.

Djuknic:2001:CGA

[317] Goran M. Djuknic and Robert E. Richton. Communications: Geolocation and assisted GPS. Computer, 34(2):123–125, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r2123.pdf.

Holmes:2001:PUE

[318] Neville Holmes. The profession: US electoral reform: The obvious obligation. Computer, 34 (2):128, 125–126, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r2128.pdf.

Castro:2001:LRN

[319] Maurice Castro, James Whittaker, Jeff Voas, Laurence E. LaForge, Jeff Voas, Richard H. Veith, Neville Holmes, and Bruce Geer. Letters: Rethinking the notion of time — again and is the aerospace industry ready for software reliability theory? and right pioneer, wrong pioneering and explicit parallelism. Computer, 34(3):4–7, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3004.pdf.

Ortiz:2001:ITI

[320] Sixto Ortiz Jr. Industry trends: Instant messaging: No longer just chat. Computer, 34(3):12–15, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3012.pdf.

Clark:2001:TNS

[321] David Clark. Technology news: Speech recognition: The wireless interface revolution. Computer, 34(3):16–18, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3016.pdf.

Paulson:2001:NBSa

[322] Linday Dailey Paulson and Orren Merton. News briefs: Streaming media initiative may be stymied and using AI to manage network traffic and is the Internet an energy hog? and making every cybervote count and two groups to focus on network security. Computer, 34(3):19–21, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3019.pdf.

Anonymous:2001:ASc

[323] Anonymous. Article summaries. Computer, 34(3):22–23, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3022.pdf.

Chen:2001:PDP

[324] Su-Shing Chen. The paradox of digital preservation. Computer, 34(3):24–28, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3024.pdf;
http://www.computer.org/computer/
co2001/r3024abs.htm.

Hull:2001:TZE

[325] Jonathan J. Hull and Peter E. Hart. Toward zero-effort personal document management. Computer, 34(3):30–35, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3030.pdf;
http://www.computer.org/computer/
co2001/r3030abs.htm.

Balakrishnan:2001:UIV

[326] Ravin Balakrishnan, George W. Fitzmaurice, and Gordon Kurtenbach. User interfaces for volumetric displays. Computer, 34(3):37–45, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3037.pdf;
http://www.computer.org/computer/
co2001/r3037abs.htm.

Hecht:2001:PED

[327] David L. Hecht. Printed embedded data graphical user interfaces. Computer, 34(3):47–55, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3047.pdf;
http://www.computer.org/computer/
co2001/r3047abs.htm.

Landay:2001:SIT

[328] James A. Landay and Brad A. Myers. Sketching interfaces: Toward more human interface design. Computer, 34(3):56–64, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3056.pdf;
http://www.computer.org/computer/
co2001/r3056abs.htm.

Chen:2001:VKD

[329] Chaomei Chen and Ray J. Paul. Visualizing a knowledge domain’s intellectual structure. Computer, 34(3):65–71, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2001/pdf/r3065.pdf;
 http://www.computer.org/computer/
co2001/r3065abs.htm.

Bellavista:2001:MAM

[330] Paolo Bellavista, Antonio Corradi, and Cesare Stefanelli. Mobile agent middleware for mobile computing. Computer, 34(3):73–81, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3073.pdf;
http://www.computer.org/computer/
co2001/r3073abs.htm.

Piner:2001:CSCc

[331] Mary-Louise G. Piner. Computer Society connection: Society to offer software engineering professional certification; computing leaders receive Society awards in 2000 [architecture prize award to [Edward S.] Davidson; Mc-Dowell winner [Raymond] Ozzie: Key Lotus Notes developer; [John] Iliffe receives Goode Award for a lifetime of work; Sandia’s [Stephen W.] Attaway named Fernbach winner]; Merwin Scholarship seeks students applicants by 31 May; portal provides access to online resources for the scientific computing field; online tutorial describes state-of-the-art testing method; CRA’s Taulbee Survey provides faculty salary data. Computer, 34(3):82–85, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3082.pdf.

Anonymous:2001:CCc

[332] Anonymous. Call and calendar. Computer, 34(3):86–89, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r3086.pdf.

Anonymous:2001:COc

[333] Anonymous. Career opportunities. Computer, 34(3):90–105, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r3090.pdf.

Lutz:2001:NBT

[334] Michael J. Lutz. New books: A theoretically rich and practical reasoning tool. Computer, 34(3):107, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r3107.pdf.

Anonymous:2001:PPS

[335] Anonymous. Products: ProxySource’s software design and collaboration application; YesSoftware’s code generation application; Persistence Software’s transactional application server; Instantiation’s Java productivity tools; JCanvas visual rapid application IDE; theKompany.com’s Python development environment; NeuVis updates Ebusiness visual modeling tools; LegacyJ’s Java-compliant Cobol compiler. Computer, 34(3):108–109, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r3108.pdf.

Shotsberger:2001:IWT

[336] Paul G. Shotsberger and Ron Vetter. Internet watch: Teaching and learning in the wireless classroom. Computer, 34(3):110–111, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3110.pdf.

Vahid:2001:IE

[337] Frank Vahid and Tony Givargis. Integrated engineering: Platform tuning for embedded systems design. Computer, 34(3):112–114, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3112.pdf.

Nuseibeh:2001:SMW

[338] Bashar Nuseibeh. Software management: Weaving together requirements and architectures. Computer, 34(3):115–117, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3115.pdf.

Nissenbaum:2001:PHC

[339] Helen Nissenbaum. The profession: How computer systems embody values. Computer, 34(3):120, 118–119, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r3120.pdf.

Anonymous:2001:CP

[340] Anonymous. Call for papers. Computer, 34(3):121, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

www.computer.org/computer/co2001/
r3029.pdf.

Whittle:2001:LAD

[341] Jon Whittle, David Harel, Jesse Ring, Neville Holmes, Shashank Garg, Jim Haynes, and George Sai-Halasz. Letters: An achievable dream encounters barriers and bottlenecks; reform or chaos?; using technology to help bridge the digital divide; chads, chadless, or chips?; processor performance nears the end of the road. Computer, 34(4):4, 6–8, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r4004.pdf.

Lawton:2001:ITN

[342] George Lawton. Industry trends: New technologies place video in your hand. Computer, 34(4):14–17, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r4014.pdf.

Miller:2001:TNA

[343] Sandra Kay Miller. Technology news: Aspect-oriented programming takes aim at software complexity. Computer, 34(4):18–21, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r4018.pdf.

Paulson:2001:NBH

[344] Linda Dailey Paulson and Orren Merton. News briefs: Hacker launches cyberattack from security site; new bus technology boosts performance; have your agent talk to my agent; peer-to-peer goes hand to hand; computing that takes a lot of nerve. Computer, 34(4):23–25, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r4023.pdf.

Anonymous:2001:ASd

[345] Anonymous. Article summaries. Computer, 34(4):26–27, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r4026.pdf.

Hamburgen:2001:ISB

[346] William R. Hamburgen, Deborah A. Wallach, Marc A. Viredaz, Lawrence S. Brakmo, Carl A. Waldspurger, Joel F. Bartlett, Timothy Mann, and Keith I. Farkas. Itsy: Stretching the bounds of mobile computing. Computer, 34 (4):28–37, April 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2001/pdf/
r4028.pdf;   http://www.computer.
org/computer/co2001/r4028abs.htm.

Makimoto:2001:CBN

[347] Tsugio Makimoto, Kazuhiko Eguchi, and Mitsugu Yoneyama. The cooler the better: New directions in the nomadic age. Computer, 34(4):38–42, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r4038.pdf;

                                        http://www.computer.org/computer/
                                        co2001/r4038abs.htm.

Chiang:2001:FDO

[348] Shang yi Chiang. Foundries and the dawn of an Open IP era. Computer, 34(4):43–46, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r4043.pdf;
                                        http://www.computer.org/computer/
                                        co2001/r4043abs.htm.

Vajapeyam:2001:ECP

[349] Sriram Vajapeyam and Mateo Valero. Early 21st century processors. Computer, 34(4):47–50, April 2001. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/co/books/
                                        co2001/pdf/r4047.pdf.

Mudge:2001:PFC

[350] Trevor Mudge. Power: A firstclass architectural design constraint. Computer, 34(4):52–58, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r4052.pdf;
                                        http://www.computer.org/computer/
                                        co2001/r4052abs.htm.

Morrow:2001:MUL

[351] Mike Morrow. Microarchitecture uses a low-power core. Computer, 34 (4):55, April 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:2001:ILD

[352] James E. Smith. Instruction-level distributed processing. Computer, 34 (4):59–65, April 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2001/pdf/
r4059.pdf;   http://www.computer.
org/computer/co2001/r4059abs.htm.

Sohi:2001:SMP

[353] Gurindar S. Sohi and Amir Roth. Speculative multithreaded processors. Computer, 34(4):66–73, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r4066.pdf;
http://www.computer.org/computer/
co2001/r4066abs.htm. See errata

[367].

Barrose:2001:PES

[354] Luiz André Barrose, Kourosh Charachorloo, Tom Heynemann, Dan Joyce, David Lowerll, Harland Maxwell, Joel McCormack, Ravishankar Mosur, Jeff Sprouse, Robert Stets, and Scott Smith. Piranha: Exploiting single-chip multiprocessing. Computer, 34(4):68, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r4059.pdf;
http://www.computer.org/computer/
co2001/r4059abs.htm.

Smith:2001:CMM

[355] Burton Smith. Cray MTA: Multithreading for latency response. Computer, 34(4):69, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r4059.pdf;

                                        http://www.computer.org/computer/
                                        co2001/r4059abs.htm.

Rau:2001:ECA

[356] B. Ramakrishna Rau and Michael S. Schlansker. Embedded computer architecture and automation. Computer, 34(4):75–82, April 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2001/pdf/
                                        r4075.pdf;   http://www.computer.
                                        org/computer/co2001/r4075abs.htm.

Piner:2001:CSCd

[357] Mary-Louise G. Piner. Computer Society connection: Strategic plan maps Society’s goals; IEEE Division VIII Director-Elect candidates named; student winners recognized for technical writing, academics, and involvement; membership reminder; EMC Corporation invests in CSIDC; Computer Society increases high school International Science Fair awards funding; Claude Shannon, 1916–2001. Computer, 34(4):84–87, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r4084.pdf.

Anonymous:2001:CCd

[358] Anonymous. Call and calendar. Computer, 34(4):88–90, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r4088.pdf.

Anonymous:2001:COd

[359] Anonymous. Career opportunities. Computer, 34(4):92–95, April 2001.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r4092.pdf.

Lutz:2001:NB

[360] Michael J. Lutz. New books: Who’s tapping into your computer? Computer, 34(4):96, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r4096.pdf.

Anonymous:2001:PBL

[361] Anonymous. Products: Borland’s Linux rapid application development environment; Platform Computing releases Web resource management suite; Empirix’s EJB load-testing software; Sanctum updates Web security model; Speedware’s mobile Internet software development tool; Magic Software unveils distributed application development environment; software development platform from Sun Microsystems and iPlanet; Aestiva’s Web software development system; I-Logix embeddedsystem software environment; Data Translation upgrades image analysis software. Computer, 34(4):97–99, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r4097.pdf.

Dunnigan:2001:ECC

[362] James F. Dunnigan. Entertainment computing: The curse of fast iron. Computer, 34(4):100–103, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r4100.pdf.

Elhanany:2001:CPS

[363] Itamar Elhanany, Michael Kahane, and Dan Sadot. Communications: Packet scheduling in nextgeneration multiterabit networks. Computer, 34(4):104–106, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r4104.pdf.

Maier:2001:SSA

[364] Mark W. Maier, David Emery, and Rich Hilliard. Standard: Software architecture: Introducing IEEE Standard 1471. Computer, 34(4):107–109, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r4107.pdf.

Mowshowitz:2001:PVO

[365] Abbe Mowshowitz. The profession virtual organization: The new feudalism. Computer, 34(4):112, 110–111, April 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r4112.pdf.

Cook:2001:LDD

[366] Michael Cook, Neville Holmes, and Wolf Nagl. Letters: The devil is in the details; Web reference persistence. Computer, 34(5):4–6, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r5004.pdf.

Anonymous:2001:ESM

[367] Anonymous. Errata: “Speculative Multithreaded Processors”. Computer, 34(5):7, May 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2001/pdf/

r5004.pdf. See [353].

Kiely:2001:ITW

[368] Don Kiely. Industry trends: Wanted: Programmers for handheld devices. Computer, 34(5):12–14, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r5012.pdf.

Edwards:2001:TNN

[369] John Edwards. Technology news: Next-generation viruses present new challenges. Computer, 34(5):16–18, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2001/pdf/r5016.pdf.

Paulson:2001:NBRa

[370] Linda Dailey Paulson. News briefs: Researchers develop new search techniques: Fake networks decoy hackers: Vendors begin laser networking. Computer, 34(5):19–21, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r5019.pdf.

Anonymous:2001:ASe

[371] Anonymous. Article summaries. Computer, 34(5):22–23, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r5022.pdf.

Kara:2001:PPR

[372] Atsushi Kara. Protecting privacy in remote-patient monitoring. Computer, 34(5):24–27, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r5024.pdf;
http://www.computer.org/computer/
co2001/r5024abs.htm.

Meyer:2001:SEA

[373] Bertrand Meyer. Software engineering in the Academy. Computer, 34 (5):28–35, May 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2001/pdf/
r5028.pdf;   http://www.computer.
org/computer/co2001/r5028abs.htm.

See letter by the author in [411].

Lutz:2001:GEI

[374] Michael J. Lutz. Guest Editor’s introduction: Software engineering on Internet time. Computer, 34(5):36, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r5036.pdf.

Griss:2001:ADA

[375] Martin L. Griss and Gilda Pour. Accelerating development with agent components. Computer, 34(5):37–43, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r5037.pdf;

http://www.computer.org/computer/
co2001/r5037abs.htm.

Sedigh-Ali:2001:SEM

[376] Sahra Sedigh-Ali, Arif Ghafoor, and Raymond A. Paul. Software engineering metrics for COTS-based systems. Computer, 34(5):44–50, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r5044.pdf;
http://www.computer.org/computer/
co2001/r5044abs.htm.

Baskerville:2001:HIS

[377] Richard Baskerville, Linda Levine, Jan Pries-Heje, Balasubramaniam Ramesh, and Sandra Slaughter. How Internet software companies negotiate quality. Computer, 34(5):51–57, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r5051.pdf;
http://www.computer.org/computer/
co2001/r5051abs.htm.

Lawlis:2001:FPE

[378] Patricia K. Lawlis, Kathryn E. Mark, Deborah A. Thomas, and Terry Courtheyn. A formal process for evaluating COTS software products. Computer, 34(5):58–63, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r5058.pdf;
http://www.computer.org/computer/
co2001/r5058abs.htm.

Wright:2001:EVD

[379] William E. Wright. An efficient videoon-demand model. Computer, 34 (5):64–70, May 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2001/pdf/
                                        r5064.pdf;   http://www.computer.
                                        org/computer/co2001/r5064abs.htm.

Anonymous:2001:CCe

[380] Anonymous. Call and calendar. Computer, 34(5):71–73, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r5071.pdf.

Piner:2001:CSCe

[381] Mary-Louise G. Piner. Computer Society connection: Society sponsors top computing conferences; two Editor-in-Chief appointments made for 2002; don’t miss these technical committee-sponsored events. Computer, 34(5):74–78, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r5074.pdf.

Anonymous:2001:COe

[382] Anonymous. Career opportunities. Computer, 34(5):80–83, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r5080.pdf.

Lutz:2001:NBBa

[383] Michael J. Lutz. New books: Backbone of computer applications. Computer, 34(5):84, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r5084.pdf.

Anonymous:2001:PPT

[384] Anonymous. Products: Planet 7 Technologies’ new XML development software; Apple Computer launches UNIX-based operating system; Codemesh releases Java/C++ integration software for Solaris; Telelogic AB’s C++ quality assessment environment; Green Hills ships embedded integrated development platform; Concurrent Computer introduces Linux application development environment. Computer, 34(5):85–86, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r5085.pdf.

Sobel:2001:IWW

[385] David L. Sobel. Internet watch: Will Carnivore devour online privacy? Computer, 34(5):87–88, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r5087.pdf.

Deniziak:2001:IED

[386] Stanislaw Deniziak and Krzysztof Sapiecha. Integrated engineering: Developing a high-level fault simulation standard. Computer, 34(5):89–90, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r5089.pdf.

Basili:2001:SMC

[387] Victor R. Basili and Barry Boehm. Software management: COTS-based systems top 10 list. Computer, 34 (5):91–95, May 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2001/pdf/
                                        r5091.pdf.

Holmes:2001:PGT

[388] Neville Holmes. The profession: The great term robbery. Computer, 34(5):96, 94–95, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r5096.pdf.

Gittins:2001:LCE

[389] Robert Gittins, Lynn H. Maxson, Robert W. Shirey, Neville Holmes, Robert W. Schwanke, and Neville Holmes. Letters: A call for engineering activism; shifting the economic equation; good thought, but wrong approach; distinguishing between humans and machines. Computer, 34(6):4, 6–7, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r6004.pdf.

Leavitt:2001:ITL

[390] Neal Leavitt. Industry trends: Linux: At a turning point? Computer, 34 (6):12–15, June 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2001/pdf/
                                        r6012.pdf.

Lawton:2001:TNN

[391] George Lawton. Technology news: New I/O technologies seek to end bottlenecks. Computer, 34(6):16–18, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6016.pdf.

Paulson:2001:NBC

[392] Linda Dailey Paulson. News briefs: Chip makers turn to new materials: Researchers push for optical Ethernet: Vendors hardwire piracy protection; technology used for liver research; US visa requests drop with tech economy. Computer, 34(6):19–21, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6019.pdf.

Anonymous:2001:ASf

[393] Anonymous. Article summaries. Computer, 34(6):22–23, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6022.pdf.

Geihs:2001:MCA

[394] Kurt Geihs. Middleware challenges ahead. Computer, 34(6):24–31, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6024.pdf;
http://www.computer.org/computer/
co2001/r6024abs.htm.

Wu:2001:OSS

[395] Ming-Wei Wu and Ying-Dar Lin COLUMNS. Open source software development: An overview. Computer, 34(6):33–38, June 2001.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6033.pdf;
http://www.computer.org/computer/
co2001/r6033abs.htm.

Leeper:2001:LTV

[396] David G. Leeper. A long-term view of short-range wireless. Computer, 34 (6):39–44, June 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2001/pdf/
r6039.pdf;   http://www.computer.
org/computer/co2001/r6039abs.htm.

Tseng:2001:LAA

[397] Yu-Chee Tseng, Shih-Lin Wu, Wen-Hwa Liao, and Chih-Min Chao. Location awareness in ad hoc wireless mobile networks. Computer, 34(6):46–52, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2001/pdf/r6046.pdf;
 http://www.computer.org/computer/
co2001/r6046abs.htm.

Saha:2001:BWI

[398] Subhasis Saha, Mark Jamtgaard, and John Villasenor RESEARCH. Bringing the wireless Internet to mobile devices. Computer, 34(6):54–58, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6054.pdf;
http://www.computer.org/computer/
co2001/r6054abs.htm.

Seligman:2001:XID

[399] Len Seligman and Arnon Rosenthal. XML’s impact on databases and data sharing. Computer, 34(6):59–67, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6059.pdf;
http://www.computer.org/computer/
co2001/r6059abs.htm.

Riedel:2001:ADL

[400] Erik Riedel, Christos Faloutsos, Garth A. Gibson, and David Nagle. Active disks for large-scale data processing. Computer, 34(6):68–74, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6068.pdf;
http://www.computer.org/computer/
co2001/r6068abs.htm.

Piner:2001:CSCf

[401] Mary-Louise G. Piner. Computer Society connection: Defining computing curricula for the modern age: Pointer variable inventor [Harold W. (Bud) Lawson] is pioneer winner. Computer, 34(6):75–78, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6075.pdf.

Anonymous:2001:CCf

[402] Anonymous. Call and calendar. Computer, 34(6):79–81, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6079.pdf.

Anonymous:2001:COf

[403] Anonymous. Career opportunities. Computer, 34(6):82–87, June 2001.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6082.pdf.

Anonymous:2001:APIb

[404] Anonymous. Advertiser/product index. Computer, 34(6):88, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6089.pdf.

Lutz:2001:NBWa

[405] Michael J. Lutz. New books: What computers can’t do. Computer, 34 (6):89, June 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2001/pdf/
r6089.pdf.

Anonymous:2001:PWB

[406] Anonymous. Products: Web-based remote administration tools; SGDL System’s 3D model development language kit; MigraTEC’s Solaris-to-Linux migration software; Visual Numerics updates C numerical library; Stardock’s Windows skin development software; InterNetwork’s new load capacity testing software; SuSE Linux for PowerPC; Raytheon updates network security tools; Tasking updates embedded development tools; ExoLab Group offers open-source data-binding software; Omnicore Software’s Java development environment; Basis International releases Java-based business basic; Zondigo’s wireless software development kit; MDD introduces password administration software; StatSoft revises data visualization tool; Abaco updates mobile application development framework. Computer, 34(6):90–93, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6090.pdf.

Varshney:2001:CIE

[407] Upkar Varshney and Radhika Jain. Communications: Issues in emerging 4G wireless networks. Computer, 34 (6):94–96, June 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2001/pdf/
r6094.pdf.

Ollis:2001:ECF

[408] Mark Ollis and Todd Williamson. Entertainment computing: The future of 3D video. Computer, 34(6):97–99, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6097.pdf.

Krechmer:2001:SNO

[409] Ken Krechmer. Standards: The need for openness in standards. Computer, 34(6):100–101, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6100.pdf.

Asuaga:2001:PEN

[410] Ana Asuaga. The profession: Engineering a new society. Computer, 34(6):104, 102–103, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r6104.pdf.

Phillips:2001:LMP

[411] Barbara Cavanaugh Phillips, Susan M. Polen, Louis C. Rose, Patricia K. Lawlis, Kathryn E. Mark, Deborah A. Thomas, Terry Courtheyn, Bertrand Meyer, David Gray, Neville Holmes, Robert Love, Neville Holmes, Stephen Gryphon, Michael Wallace, Ravin Balakrishnan, George W. Fitzmaurice, and Gordon Kurtenbach. Letters: Merging points of view in the COTS software evaluation process; the invention of monitors; A troubling phenomenon; volumetric interface options. Computer, 34(7):4, 6–7, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7004.pdf. See

[373].

Lawton:2001:ITI

[412] George Lawton. Industry trends: Internet appliances struggle for acceptance. Computer, 34(7):12–15, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7012.pdf.

Miller:2001:TNF

[413] Sandra Kay Miller. Technology news: Facing the challenge of wireless security. Computer, 34(7):16–18, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7016.pdf.

Paulson:2001:NBRb

[414] Linda Dailey Paulson. News briefs: Researchers work on improving memory: Tracking your E-mail: Handheldto-handheld: Fighting over Java; future searches; location, location, location; Michigan goes to cybercourt. Computer, 34(7):19–21, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7019.pdf.

Anonymous:2001:ASg

[415] Anonymous. Article summaries. Computer, 34(7):22–23, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7022.pdf.

Buttazzo:2001:ACU

[416] Giorgio Buttazzo. Artificial consciousness: Utopia or real possibility? Computer, 34(7):24–30, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7024.pdf;
http://www.computer.org/computer/
co2001/r7024abs.htm.

Parameswaran:2001:PNI

[417] Manoj Parameswaran, Anjana Susarla, and Andrew B. Whinston COVER. P2P networking: An informationsharing alternative. Computer, 34 (7):31–38, July 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2001/pdf/
r7031.pdf;   http://www.computer.
org/computer/co2001/r7031abs.htm.

Durfee:2001:SAC

[418] Edmund H. Durfee. Scaling up agent coordination strategies. Computer, 34 (7):39–46, July 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2001/pdf/
                                        r7039.pdf;   http://www.computer.
                                        org/computer/co2001/r7039abs.htm.

Bradshaw:2001:TC

[419] Jeffrey M. Bradshaw, Niranjan Suri, Alberto J. Cañas, Robert Davis, Kenneth Ford, Robert Hoffman, Renia Jeffers, and Thomas Reichherzer. Terraforming cyberspace. Computer, 34(7):48–56, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r7048.pdf;
                                        http://www.computer.org/computer/
                                        co2001/r7048abs.htm.

Lieberman:2001:AAA

[420] Henry Lieberman, Elizabeth Rosenzweig, and Push Singh. Aria: An agent for annotating and retrieving images. Computer, 34(7):57–62, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r7057.pdf;
                                        http://www.computer.org/computer/
                                        co2001/r7057abs.htm.

Case:2001:ECA

[421] Simon Case, Nader Azarmi, Marcus Thint, and Takeshi Ohtani. Enhancing E-communities with agent-based systems. Computer, 34(7):64–69, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7064.pdf;
http://www.computer.org/computer/
co2001/r7064abs.htm.

Laird:2001:UCG

[422] John E. Laird. Using a computer game to develop advanced AI. Computer, 34(7):70–75, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7070.pdf;
http://www.computer.org/computer/
co2001/r7070abs.htm.

Piner:2001:CSCg

[423] Mary-Louise G. Piner. Computer Society connection: Ten student teams advance to CSIDC world finals: IEEE Computer Society rewards high school student achievements. Computer, 34(7):76–79, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7076.pdf.

Anonymous:2001:CCg

[424] Anonymous. Call and calendar. Computer, 34(7):80–82, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7080.pdf.

Anonymous:2001:COg

[425] Anonymous. Career opportunities. Computer, 34(7):84–89, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7080.pdf.

Anonymous:2001:PSX

[426] Anonymous. Products: SoftQuad’s XML content creation software; Origin-Lab updates graphing tool; NuSphere’s enterprise Web development platform; MetaWare’s XScale programming tools; Aether Systems’ wireless development environment; Visual Numerics upgrades Java application deployment tools; C Level Design introduces C/C++ hardware design environment; ActiveState’s Perl development and deployment software; Advanced Software ships UML design tool; Borland’s Java 2 rapid application development environment; Web services application development platform; RidgeRun’s embedded Linux development kit; IONA modeling and development environment. Computer, 34(7):90–92, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7090.pdf.

Lutz:2001:NBBb

[427] Michael J. Lutz. New books: Building a framework for a sustainable world. Computer, 34(7):93, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7093.pdf.

Montgomery:2001:IWI

[428] Alan L. Montgomery and Christos Faloutsos. Internet watch: Identifying Web browsing trends and patterns. Computer, 34(7):94–95, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7094.pdf.

Sheard:2001:SME

[429] Sarah A. Sheard. Software management: Evolution of the frameworks quagmire. Computer, 34(7):96–98, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7096.pdf.

Pavan:2001:IE

[430] Allalaghatta Pavan, Rakesh Jha, Lee Graba, Saul Cooper, Ionut Cardei, Mihaela Cardei, Vipin Gopal, Sanjay Parthasarathy, and Saad Bedros. Integrated engineering. Computer, 34(7):99–101, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7099.pdf.

Holmes:2001:PVG

[431] Neville Holmes. The profession: Vanity and guilt, humility and pride. Computer, 34(7):104, 102–103, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r7104.pdf.

Pavlidis:2001:LTM

[432] Theo Pavlidis, Franco Favento, Ed Mechler, and Sarah A. Sheard. Letters: Tracing middleware’s historic origins; buzzword traps; additions to the frameworks quagmire. Computer, 34(8):4–5, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r8004.pdf;
                                        http://www.computer.org/computer/
                                        co2001/r8004abs.htm.

Aylor:2001:KK

[433] James H. Aylor. Keep on keepin’ on. Computer, 34(8):6–7, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r8006.pdf;
                                        http://www.computer.org/computer/
                                        co2001/r8006abs.htm.

Lawton:2001:IFG

[434] George Lawton. Is IPv6 finally gaining ground? Computer, 34(8):11–15, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co2001/pdf/r8011.pdf;
                                         http://www.computer.org/computer/
                                        co2001/r8011abs.htm.

Clark:2001:VBF

[435] David Clark. Visual Basic and the future of Web-application development. Computer, 34(8):16–18, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r8016.pdf;
                                        http://www.computer.org/computer/
                                        co2001/r8016abs.htm.

Paulson:2001:NBN

[436] Linda Dailey Paulson and Orren Merton. News briefs: NASA satellites will use AI; survey reveals PC abuse; tech slump cools once-hot job market; “porn-seeking” virus strikes Internet. Computer, 34(8):19–20, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r8019.pdf;
http://www.computer.org/computer/
co2001/r8019abs.htm.

Anonymous:2001:ASh

[437] Anonymous. Article summaries. Computer, 34(8):22–23, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r8022.pdf;
http://www.computer.org/computer/
co2001/r8022abs.htm.

Budd:2001:PME

[438] Timothy A. Budd. Protecting and managing electronic content with a digital battery. Computer, 34(8):24–30, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r8024.pdf;
http://www.computer.org/computer/
co2001/r8024abs.htm.

Want:2001:GEI

[439] Roy Want and Bill Schilit. Guest Editors’ introduction: Expanding the horizons of location-aware computing. Computer, 34(8):31–34, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r8031.pdf;
http://www.computer.org/computer/
co2001/r8031abs.htm.

Davies:2001:UDL

[440] Nigel Davies, Keith Cheverst, Keith Mitchell, and Alon Efrat. Using and determining location in a context-sensitive tour guide. Computer, 34(8):35–41, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r8035.pdf;
http://www.computer.org/computer/
co2001/r8035abs.htm.

Pradhan:2001:WHP

[441] Salil Pradhan, Cyril Brignone, Jun-Hong Cui, Alan McReynolds, and Mark T. Smith. Websigns: Hyperlinking physical locations to the Web. Computer, 34(8):42–48, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r8042.pdf;
http://www.computer.org/computer/
co2001/r8042abs.htm.

Addlesee:2001:ISC

[442] Mike Addlesee, Rupert Curwen, Steve Hodges, Joe Newman, Pete Steggles, Andy Ward, and Andy Hopper. Implementing a sentient computing system. Computer, 34(8):50–56, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r8050.pdf;
http://www.computer.org/computer/
co2001/r8050abs.htm.

Garber:2001:PSW

[443] Lee Garber. Polish students win Computer Society design competition. Computer, 34(8):67–71, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r8067.pdf;

http://www.computer.org/computer/
co2001/r8067abs.htm.

Piner:2001:CSCh

[444] Mary-Louise G. Piner. Computer Society connection: Society plans two new periodicals [IEEE Pervasive Computing and IEEE Transactions on Mobile Computing]; task forces address new technical disciplines. Computer, 34(8):72–75, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r8072.pdf;
http://www.computer.org/computer/
co2001/r8072abs.htm.

Anonymous:2001:CCh

[445] Anonymous. Call and calendar. Computer, 34(8):76–79, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r8076.pdf;
http://www.computer.org/computer/
co2001/r8076abs.htm.

Lutz:2001:NBWb

[446] Michael J. Lutz. New books: Winning the project management war. Computer, 34(8):83, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r8083.pdf;
http://www.computer.org/computer/
co2001/r8083abs.htm.

Anonymous:2001:PVL

[447] Anonymous. Products: Viosoft’s Linux embedded development environment; Popkin Software releases development modeling suite; Iopsis Software’s Forte for Java IDE; NQL releases scripting language components; Ascend Software updates delivery management system; Excel Software ships UML design tool; Hyperformix’s discrete-event simulation modeler; InCert’s application fault management software; BioconX releases biometric security software; Waterloo Maple’s mathematical computation software [Maple 7]; Equator Technologies’ application development toolkit. Computer, 34(8):84–86, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r8084.pdf;
http://www.computer.org/computer/
co2001/r8084abs.htm.

Geng:2001:PVA

[448] Xianjun Geng and Andrew B. Whinston. Profiting from value-added wireless services. Computer, 34(8):87–89, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2001/pdf/r8087.pdf;
 http://www.computer.org/computer/
co2001/r8087abs.htm.

Macedonia:2001:EBU

[449] Michael Macedonia. E3 2001: The birth of ubiquitous gaming? Computer, 34(8):90–91, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r8090.pdf;
http://www.computer.org/computer/
co2001/r8090abs.htm.

Meyer:2001:NC

[450] Bertrand Meyer. .NET is coming. Computer, 34(8):92–97, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r8092.pdf;
http://www.computer.org/computer/
co2001/r8092abs.htm.

Spinellis:2001:FCH

[451] Diomidis Spinellis. Fear of coding, and how to reduce it. Computer, 34(8):100, 98–99, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r8100.pdf;
http://www.computer.org/computer/
co2001/r8100abs.htm.

Cirillo:2001:LTW

[452] William Cirillo, Ming-Wei Wu, Ying-Dar Lin, Wayne Citrin, Michael J. Lewchuk, Jeff Redding, Neville Holmes, and Horst P. Richter. Letters: The trendy way needs a balanced critique; Internet appliance struggle; coding headaches and myths; confrontational communications; Utopia or reality. Computer, 34(9):4, 6–7, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r9004.pdf;
http://www.computer.org/computer/
co2001/r9004abs.htm.

Paulson:2001:TTT

[453] Linda Dailey Paulson. Translation technology tries to hurdle the language barrier. Computer, 34(9):12– 15, September 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2001/pdf/
                                        r9012.pdf;   http://www.computer.
                                        org/computer/co2001/r9012abs.htm.

Lawton:2001:TMP

[454] George Lawton. Is technology meeting the privacy challenge? Computer, 34(9):16–18, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r9016.pdf;
                                        http://www.computer.org/computer/
                                        co2001/r9016abs.htm.

Paulson:2001:NBSb

[455] Linda Dailey Paulson. News briefs: Storing data in wavelengths of light; Internet-traffic discovery could improve networks; microprocessor menagerie; Microsoft, Sun announce P2P technologies. Computer, 34(9):19–21, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co2001/pdf/r9019.pdf;
                                         http://www.computer.org/computer/
                                        co2001/r9019abs.htm; http://www.
                                        jxta.org;   http://www.microsoft.
                                        com/net/hailstorm.asp.

Anonymous:2001:ASi

[456] Anonymous. Article summaries. Computer, 34(9):22–23, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/r9022.pdf;
                                        http://www.computer.org/computer/
                                        co2001/r9022abs.htm.

Rajlich:2001:SCE

[457] Václav Rajlich, Norman Wilde, Michelle Buckellew, and Henry Page. Software cultures and evolution. Computer, 34(9):24–28, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r9024.pdf;
http://www.computer.org/computer/
co2001/r9024abs.htm.

Morris:2001:SCC

[458] John Morris, Gareth Lee, Kris Parker, Gary A. Bundell, and Chiou Peng Lam. Software component certification. Computer, 34(9):30–36, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r9030.pdf;
http://www.computer.org/computer/
co2001/r9030abs.htm.

Liu:2001:GEI

[459] Jonathan C. L. Liu and David H. C. Du. Guest Editors’ introduction: Continuous media on demand. Computer, 34(9):37–39, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r9037.pdf;
http://www.computer.org/computer/
co2001/r9037abs.htm.

Horn:2001:SRP

[460] Gavin B. Horn, Per Knudsgaard, Soren B. Lassen, Michael Luby, and Jens Eilstrup Rasmussen. A scalable and reliable paradigm for media on demand. Computer, 34(9):40– 45, September 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2001/pdf/
r9040.pdf;   http://www.computer.
org/computer/co2001/r9040abs.htm.

Elsen:2001:STM

[461] Ingo Elsen, Frank Hartung, Uwe Horn, Markus Kampmann, and Liliane Peters. Streaming technology in 3G mobile communication systems. Computer, 34(9):46–52, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r9046.pdf;
http://www.computer.org/computer/
co2001/r9046abs.htm.

Girgensohn:2001:KBU

[462] Andreas Girgensohn, John Boreczky, and Lynn Wilcox. Keyframe-based user interfaces for digital video. Computer, 34(9):61–67, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r9061.pdf;
http://www.computer.org/computer/
co2001/r9061abs.htm.

Pieper:2001:SMK

[463] Jan Pieper, Savitha Srinivasan, and Byron Dom. Streamingmedia knowledge discovery. Computer, 34(9):68–74, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r9068.pdf;
http://www.computer.org/computer/
co2001/r9068abs.htm.

Anonymous:2001:NCS

[464] Anonymous. Nominees for Computer Society office and Board of Governors. Computer, 34(9):76–83, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r9076.pdf;
http://www.computer.org/computer/
co2001/r9076abs.htm.

King:2001:IPE

[465] Willis K. King, Michael S. Adler, Paul J. Kostek, and Arthur Winston. IEEE President-Elect candidates address Computer Society concerns. Computer, 34(9):86–90, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r9086.pdf;
http://www.computer.org/computer/
co2001/r9086abs.htm.

Anonymous:2001:CCi

[466] Anonymous. Call and calendar. Computer, 34(9):91–95, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r9091.pdf;
http://www.computer.org/computer/
co2001/r9091abs.htm.

Anonymous:2001:PFS

[467] Anonymous. Products: Free Software Foundation updates compiler toolset; IT Factory’s Lotus Tools Suite; Cardiff Software’s real-time document verification application; Great Bridge updates open source database; Open-Path Products’ wireless application authoring tool; Curl’s Web application development environment; ThinAirApp ships mobile Visual Basic IDE; Princeton Softech updates server archiving software; Codemesh releases Java communications environment. Computer, 34(9):112–114, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r9112.pdf;
http://www.computer.org/computer/
co2001/r9112abs.htm.

Lutz:2001:NBC

[468] Michael J. Lutz. New books: Cloaking and uncloaking data. Computer, 34(9):115, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r9115.pdf;
http://www.computer.org/computer/
co2001/r9115abs.htm.

Michener:2001:CE

[469] John R. Michener and Steven D. Mohan. Clothing the E-emperor. Computer, 34(9):116–118, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r9116.pdf;
http://www.computer.org/computer/
co2001/r9116abs.htm.

Highsmith:2001:ASD

[470] Jim Highsmith and Alistair Cockburn. Agile software development: The business of innovation. Computer, 34(9):120–122, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r9120.pdf;
http://www.computer.org/computer/
co2001/r9120abs.htm.

Holmes:2001:CEH

[471] Neville Holmes. Crouching error, hidden markup. Computer, 34 (9):128, 126–127, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/r9128.pdf;
http://www.computer.org/computer/
co2001/r9128abs.htm.

Woodward:2001:LEN

[472] Stuart Woodward, Ivan Ryant, Stephen H. Unger, Ana Asuaga, Samuel Wong, Timothy A. Budd, and Vojislav Misic. Letters: Engineering a new society; digital battery application; metalanguage clarification. Computer, 34(10):4, 6–7, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rx004.pdf;
http://www.computer.org/computer/
co2001/rx004abs.htm.

Vaughan-Nichols:2001:ITT

[473] Steven J. Vaughan-Nichols. Industry trends: Tempest over Web-authoring tools. Computer, 34(10):12–14, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2001/pdf/rx012.pdf;
http://www.computer.org/computer/
co2001/rx012abs.htm.

Grado-Caffaro:2001:TN

[474] Maria-Angeles Grado-Caffaro and Martin Grado-Caffaro. Technology news: The challenges that XML faces. Computer, 34(10):15–18, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rx015.pdf;
http://www.computer.org/computer/
co2001/rx015abs.htm.

Garber:2001:NBT

[475] Lee Garber. News briefs: Is tech downturn changing education and employment trends; HTMT promises highperformance computing; controversial software law [UCITA] hist resistance. Computer, 34(10):19–21, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rx019.pdf;
http://www.computer.org/computer/
co2001/rx019abs.htm.

Anonymous:2001:ASj

[476] Anonymous. Article summaries. Computer, 34(10):22–23, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rx022.pdf;
http://www.computer.org/computer/
co2001/rx022abs.htm.

Hernandez:2001:SES

[477] Julio César Hernández, José María Sierra, Arturo Ribagorda, and Benjamín Ramos. Search engines as a security threat. Computer, 34(10): 25–30, October 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2001/pdf/
rx025.pdf;   http://www.computer.
org/computer/co2001/rx025abs.htm.

Tao:2001:SPA

[478] Lixin Tao. Shifting paradigms with the application service provider model. Computer, 34(10):32–39, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rx032.pdf;
http://www.computer.org/computer/
co2001/rx032abs.htm.

Arbaugh:2001:GEI

[479] William A. Arbaugh and Leendert van Doorn. Guest Editors’ introduction: Embedded security: Challenges and concerns. Computer, 34(10):40–47, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rx040.pdf;
http://www.computer.org/computer/
co2001/rx040abs.htm.

Traw:2001:PDC

[480] C. Brendan S. Traw. Protecting digital content within the home. Computer, 34(10):42–47, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rx042.pdf;
http://www.computer.org/computer/
co2001/rx042abs.htm.

Wallach:2001:CPT

[481] Dan S. Wallach. Copy protection technology is doomed. Computer, 34(10):48–49, October 2001.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rx048.pdf;
http://www.computer.org/computer/
co2001/rx048abs.htm.

Bergstrom:2001:MHA

[482] Peter Bergstrom, Kevin Driscoll, and John Kimball. Making home automation communications secure. Computer, 34(10):50–56, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rx050.pdf;
http://www.computer.org/computer/
co2001/rx050abs.htm.

Dyer:2001:BIS

[483] Joan G. Dyer, Mark Lindemann, Ronald Perez, Reiner Sailer, Leendert van Doorn, Sean W. Smith, and Steve Weingart. Building the IBM 4758 secure coprocessor. Computer, 34(10):57–66, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rx057.pdf;
http://www.computer.org/computer/
co2001/rx057abs.htm.

Bond:2001:ALA

[484] Mike Bond and Ross Anderson. APIlevel attacks on embedded systems. Computer, 34(10):67–75, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rx067.pdf;
http://www.computer.org/computer/
co2001/rx067abs.htm.

Piner:2001:CSCi

[485] Mary-Louise G. Piner. Computer Society connection: Society awards recognize variety of talents: 2000 Kanai Award to C. V. Ramamoorthy, 2000 Taylor L. Booth Education Award to Prithviraj Banerjee, 2001 Computer Pioneer Award to William H. Bridge and Vernon L. Schatz; 2001 Harry M. Goode Award to Oscar H. Ibarra; obituary: Michael L. Dertouzos, 1937– 2001; 2000 Richard E. Merwin Distinguished Service Award to Gerald L. Engel. Computer, 34(10):76–80, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rx076.pdf;
http://www.computer.org/computer/
co2001/rx076abs.htm.

Anonymous:2001:CCj

[486] Anonymous. Call and calendar. Computer, 34(10):81–83, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rx081.pdf;
http://www.computer.org/computer/
co2001/rx081abs.htm.

Anonymous:2001:PNI

[487] Anonymous. Products: National Instruments upgrades Measurement Studio; OnCore Systems’ tool for embedded processors; Excel Software’s modeling cards; EmmaSoft upgrades software installation tool; Sitraka’s JProbe for IBM AIX; InstallShield upgrades installation developer; EuroSoft ships disk-cloning tool. Computer, 34(10):104–106, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/rx104.pdf;
                                        http://www.computer.org/computer/
                                        co2001/rx104abs.htm.

Lutz:2001:NBD

[488] Michael J. Lutz. New books: Data modeling and mining. Computer, 34(10):107, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/rx107.pdf;
                                        http://www.computer.org/computer/
                                        co2001/rx107abs.htm.

Gerasenko:2001:BSW

[489] Sergei Gerasenko, Abhijit Joshi, Srinivas Rayaprolu, Kovendhan Ponnavaikko, and Dharma P. Agrawal. Beacon signals: What, why, how, and where? Computer, 34(10):108–110, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co2001/pdf/rx108.pdf;
                                         http://www.computer.org/computer/
                                        co2001/rx108abs.htm.

Day:2001:OGC

[490] Grantley Day. Online games: Crafting persistent-state worlds. Computer, 34(10):111–112, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/rx111.pdf;
                                        http://www.computer.org/computer/
                                        co2001/rx111abs.htm.

Bass:2001:SUT

[491] Len Bass and Bonnie E. John. Supporting usability through software architecture. Computer, 34(10):113–115, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2001/pdf/rx113.pdf;
http://www.computer.org/computer/
co2001/rx113abs.htm.

Cargill:2001:WWD

[492] Carl F. Cargill. Why are we doing this? Computer, 34(10):116–117, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rx116.pdf;
http://www.computer.org/computer/
co2001/rx116abs.htm.

Holmes:2001:PFL

[493] Neville Holmes. The profession’s future lies in its past. Computer, 34(10):120, 118–119, October 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rx120.pdf;
http://www.computer.org/computer/
co2001/rx120abs.htm.

Wah:2001:PMF

[494] Benjamin W. Wah. President’s message: Facing the challenges of the Computer Society’s future. Computer, 34(11):4–5, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry004.pdf;
http://www.computer.org/computer/
co2001/ry004abs.htm.

Leu:2001:LDB

[495] Joshua C. Leu, Timothy Budd, Frank Kearns, Neville Holmes, John Michael Williams, and Warren Montgomery. Letters: Digital battery piracy protection. Computer, 34(11):7–9, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry007.pdf;
http://www.computer.org/computer/
co2001/ry007abs.htm.

Ortiz:2001:TRP

[496] Sixto Ortiz Jr. 3D technology: Ready for the PC? Computer, 34(11):14–16, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry014.pdf;
http://www.computer.org/computer/
co2001/ry014abs.htm.

Leavitt:2001:EOG

[497] Neal Leavitt. Embedded OSs gain the inside track. Computer, 34(11):17–20, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry017.pdf;
http://www.computer.org/computer/
co2001/ry017abs.htm.

Paulson:2001:NBB

[498] Linda Dailey Paulson. News briefs: Breakthrough promises cheaper, faster chips; new viruses target P2P systems; AI program creates museum art; optical switches turn to holography. Computer, 34(11):21–23, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry021.pdf;

http://www.computer.org/computer/
co2001/ry021abs.htm.

Anonymous:2001:ASk

[499] Anonymous. Article summaries. Computer, 34(11):24–25, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry024.pdf;
http://www.computer.org/computer/
co2001/ry024abs.htm.

Zeller:2001:ADW

[500] Andreas Zeller. Automated debugging: Are we close? Computer, 34(11):26–31, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry026.pdf;
http://www.computer.org/computer/
co2001/ry026abs.htm.

Martin:2001:MVN

[501] Robert A. Martin. Managing vulnerabilities in networked systems. Computer, 34(11):32–38, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry032.pdf;
http://www.computer.org/computer/
co2001/ry032abs.htm.

Ledeczi:2001:CDS

[502] Ákos Lédeczi, Árpád Bakay, Miklós Maróti, Péter Völgyesi, Greg Nordstrom, Jonathan Sprinkle, and Gábor Karsai. Composing domain-specific design environments. Computer, 34(11): 44–51, November 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2001/pdf/
                                        ry044.pdf;   http://www.computer.
                                        org/computer/co2001/ry044abs.htm.

Mikkonen:2001:FDD

[503] Tommi Mikkonen and Peeter Pruuden. Flexibility as a design driver. Computer, 34(11):52–56, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/ry052.pdf;
                                        http://www.computer.org/computer/
                                        co2001/ry052abs.htm.

Pimentel:2001:EES

[504] Andy D. Pimentel, Louis O. Hertzberger, Paul Lieverse, Pieter van der Wolf, and Ed F. Deprettere. Exploring embedded-systems architectures with Artemis. Computer, 34(11):57– 63, November 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2001/pdf/
                                        ry057.pdf;   http://www.computer.
                                        org/computer/co2001/ry057abs.htm.

Alexander:2001:RSS

[505] Perry Alexander and Cindy Kong. Rosetta: Semantic support for modelcentered systems-level design. Computer, 34(11):64–70, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/ry064.pdf;
                                        http://www.computer.org/computer/
                                        co2001/ry064abs.htm.

Kerttula:2001:VDM

[506] Mikko Kerttula and Timo Tokkonen. Virtual design of multiengineering electronics systems. Computer, 34(11):71–79, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry071.pdf;
http://www.computer.org/computer/
co2001/ry071abs.htm.

Piner:2001:CSCj

[507] Mary-Louise G. Piner. Computer Society connection: What’s ahead for 2002 in Computer Society magazines; design competition underway: Entry rules revised; students win 2001 Merwin Scholarship Award; CSIDC changes seek to widen participation and focus on entry merit. Computer, 34(11):80–83, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry080.pdf;
http://www.computer.org/computer/
co2001/ry080abs.htm.

Anonymous:2001:CCk

[508] Anonymous. Call and calendar. Computer, 34(11):84–88, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry084.pdf;
http://www.computer.org/computer/
co2001/ry084abs.htm.

Anonymous:2001:COh

[509] Anonymous. Career opportunities. Computer, 34(11):90–123, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry090.pdf;
http://www.computer.org/computer/
co2001/ry090abs.htm.

Anonymous:2001:PGH

[510] Anonymous. Products: Green Hills ships StarCore development probe; Zayante’s FireWire support for Windows CE; Embedded Performance develops SoC debugger; Extended Systems updates Bluetooth development kit; First Embedded IPv6 router reference platform; Integrated Wireless Java technology; Java performance on wireless devices; Syclo’s mobile software development platform. Computer, 34(11):124–126, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry124.pdf;
http://www.computer.org/computer/
co2001/ry124abs.htm.

Lutz:2001:NBIc

[511] Michael J. Lutz. New books: Internetdriver telecommunications. Computer, 34(11):127, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry127.pdf;
http://www.computer.org/computer/
co2001/ry127abs.htm.

Kenyon:2001:EWC

[512] Chris Kenyon. The evolution of Web-caching markets. Computer, 34(11):128–130, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry128.pdf;
http://www.computer.org/computer/
co2001/ry128abs.htm.

Cockburn:2001:ASD

[513] Alistair Cockburn and Jim Highsmith. Agile software development: The people factor. Computer, 34(11):131–133, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry131.pdf;
http://www.computer.org/computer/
co2001/ry131abs.htm.

Holmes:2001:TTP

[514] Nevil Holmes. Terrorism, technology, and the profession. Computer, 34(11):136, 134–135, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/ry136.pdf;
http://www.computer.org/computer/
co2001/ry136abs.htm.

Rakitin:2001:LME

[515] Steven R. Rakitin, Werner Donné, Neville Holmes, and Berend de Boer. Letters: Manifesto elicits cynicism; more markup remarks. Computer, 34(12):4, 6–7, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz004.pdf;
http://www.computer.org/computer/
co2001/rz004abs.htm.

Wah:2001:PMR

[516] Benjamin W. Wah. President’s message: Reviewing our 2001 strategic accomplishments. Computer, 34(12): 8–9, December 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2001/pdf/
                                        rz008.pdf;   http://www.computer.
                                        org/computer/co2001/rz008abs.htm.

Anonymous:2001:RMS

[517] Anonymous. Report to members: Society elects officers and board members for 2002. Computer, 34(12):11–12, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/rz011.pdf;
                                        http://www.computer.org/computer/
                                        co2001/rz011abs.htm.

Lawton:2001:ITB

[518] George Lawton. Industry trends: Browsing the mobile Internet. Computer, 34(12):18–21, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/rz018.pdf;
                                        http://www.computer.org/computer/
                                        co2001/rz018abs.htm.

Vaughan-Nichols:2001:TNW

[519] Steven J. Vaughan-Nichols. Technology news: Will vector graphics finally make it on the Web? Computer, 34(12):22–24, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/rz022.pdf;

http://www.computer.org/computer/
co2001/rz022abs.htm.

Paulson:2001:NBW

[520] Linda Dailey Paulson and Lee Garber. News briefs: W3C proposal ignites standards controversy; Intel technology promises 20-GHz chips; Computer Society Editor experiences anthrax scare; Internet2 upgrades backbone. Computer, 34(12):25–27, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz025.pdf;
http://www.computer.org/computer/
co2001/rz025abs.htm.

Anonymous:2001:ASl

[521] Anonymous. Article summaries. Computer, 34(12):28–29, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz028.pdf;
http://www.computer.org/computer/
co2001/rz028abs.htm.

Zari:2001:URW

[522] Mazen Zari, Hossein Saiedian, and Muhammad Naeem. Understanding and reducing Web delays. Computer, 34(12):30–37, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz030.pdf;
http://www.computer.org/computer/
co2001/rz030abs.htm. See erratum

[566].

Blaha:2001:DWD

[523] Michael Blaha. Data warehouses and decision support systems. Computer, 34(12):38–39, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz038.pdf;
http://www.computer.org/computer/
co2001/rz038abs.htm.

Pedersen:2001:MDT

[524] Torben Bach Pedersen and Christian S. Jensen. Multidimensional database technology. Computer, 34(12):40– 46, December 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2001/pdf/
rz040.pdf;   http://www.computer.
org/computer/co2001/rz040abs.htm.

Chaudhuri:2001:DTD

[525] Surajit Chaudhuri, Umeshwar Dayal, and Venkatesh Ganti. Database technology for decision support systems. Computer, 34(12):48–55, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz048.pdf;
http://www.computer.org/computer/
co2001/rz048abs.htm.

Berndt:2001:HDW

[526] Donald J. Berndt, John W. Fisher, Alan R. Hevner, and James Studnicki. Healthcare data warehousing and quality assurance. Computer, 34(12):56– 65, December 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2001/pdf/
rz056.pdf;   http://www.computer.
org/computer/co2001/rz056abs.htm.

Trujillo:2001:DDW

[527] Juan Trujillo, Manuel Palomar, Jaime Gomez, and Il-Yeol Song. Designing data warehouses with OO conceptual models. Computer, 34(12):66–75, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz066.pdf;
http://www.computer.org/computer/
co2001/rz066abs.htm.

Armstrong:2001:SSO

[528] Rob Armstrong. Seven steps to optimizing data warehouse performance. Computer, 34(12):76–79, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz076.pdf;
http://www.computer.org/computer/
co2001/rz076abs.htm.

Pedone:2001:BSP

[529] Fernando Pedone. Boosting system performance with optimistic distributed protocols. Computer, 34(12): 80–86, December 2001. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2001/pdf/
rz080.pdf;   http://www.computer.
org/computer/co2001/rz080abs.htm.

Chang:2001:FCD

[530] Carl K. Chang, Jane Cleland-Haung, Shiyan Hua, and Annie Kuntzmann-Combelles. Functionclass decomposition: A hybrid software engineering method. Computer, 34(12):87–93, December 2001.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz087.pdf;
http://www.computer.org/computer/
co2001/rz087abs.htm.

Anonymous:2001:AIC

[531] Anonymous. Annual index: Computer, volume 34, 2001. Computer, 34(12):94–107, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz094.pdf;
http://www.computer.org/computer/
co2001/rz094abs.htm.

Piner:2001:CSCk

[532] Mary-Louise G. Piner. Computer Society connection: IEEE Computer Society welcomes new Executive Director [David W. Hennage]; Sakamura receives Takada Award; 8th International Software Metrics Symposium; ICDE — International Conference on Data Engineering; women in computing research: Distributed mentor approach; Computer recognizes reviewer’s expertise. Computer, 34(12):108–111, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz108.pdf;
http://www.computer.org/computer/
co2001/rz108abs.htm.

Anonymous:2001:COi

[533] Anonymous. Career opportunities. Computer, 34(12):116–149, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz116.pdf;

http://www.computer.org/computer/
co2001/rz116abs.htm.

Anonymous:2001:PEP

[534] Anonymous. Products: Empirix’ portable next-generation test platform; Novadigm updates adaptive IT management software; Zero G Software’s deployment tool; AireWeb ships wireless development kit; Ulead upgrades professional nonlinear video editor; Excel Software’s bug-tracking tool; MKS releases Impact and code analysis software. Computer, 34(12):150–152, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz150.pdf;
http://www.computer.org/computer/
co2001/rz150abs.htm.

Lutz:2001:NBF

[535] Michael J. Lutz. New books: Fighting cyber attacks. Computer, 34(12):153, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz153.pdf;
http://www.computer.org/computer/
co2001/rz153abs.htm.

Kagal:2001:CTB

[536] Lalana Kagal, Tim Finin, and Anupam Joshi. Communications: Trust-based security in pervasive computing environments. Computer, 34(12):154–157, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz154.pdf;

                                        http://www.computer.org/computer/
                                        co2001/rz154abs.htm.

Wall:2001:ST

[537] David A. E. Wall. Software technologies: Using Open Source for a profitable startup. Computer, 34(12):158–160, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/rz158.pdf;
                                        http://www.computer.org/computer/
                                        co2001/rz158abs.htm.

Macedonia:2001:ECW

[538] Michael Macedonia. Entertainment computing: Will Linux be computer games’ dark horse OS? Computer, 34(12):161–162, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/rz161.pdf;
                                        http://www.computer.org/computer/
                                        co2001/rz161abs.htm.

Rozenbit:2001:IEL

[539] Jerzy W. Rozenbit and Sanjaya Kumar. Integrated engineering: Looking back, moving forward. Computer, 34(12):164–165, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2001/pdf/rz164.pdf;
                                        http://www.computer.org/computer/
                                        co2001/rz164abs.htm.

Cifuentes:2001:PRE

[540] Cristina Cifuentes. The profession: Reverse engineering and the computing profession. Computer, 34 (12):168, 166–167, December 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2001/pdf/rz168.pdf;
http://www.computer.org/computer/
co2001/rz168abs.htm.

Aylor:2002:EME

[541] James H. Aylor. EIC’s message: The end of computing disciplines as we know them? Computer, 35(1):4–5, 8, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1004.pdf;
http://www.computer.org/computer/
co2002/r1004abs.htm.

King:2002:PMI

[542] Willis K. King. President’s message: Implementing our future vision in 2002. Computer, 35(1):6–8, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1006.pdf;
http://www.computer.org/computer/
co2002/r1006abs.htm.

Skowronski:2002:LTP

[543] Victor Skowronski, Norman Hines, Bob Giansiracusa, Neville Holmes, James S. Brown, Klaus Johannes Rusch, and Pearl Brereton. Letters: Terrorism and the profession: Asking the right questions; educational Luddism, privacy with security; search engines and security; software as a service. Computer, 35(1):9–11, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r1009.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r1009abs.htm.

Colwell:2002:RPA

[544] Bob Colwell. At random: Phrenology, astrology, risk, and product quality. Computer, 35(1):12–14, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r1012.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r1012abs.htm.

Anonymous:2002:ASa

[545] Anonymous. Article summaries. Computer, 35(1):15–16, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r1015.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r1015abs.htm.

Paulson:2002:NBA

[546] Linda Dailey Paulson. News briefs: Alliance seeks streaming-media standard; what do we do with millions of old PCs? new Nanotech Institute will sweat the small stuff; DSPs face new competition from CPUs; users create underground wireless networks. Computer, 35(1):22–25, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r1022.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r1022abs.htm.

Garber:2002:TNW

[547] Lee Garber. Technology news: Will 3G really be the next big wireless technology? Computer, 35(1):26–32, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1026.pdf;
http://www.computer.org/computer/
co2002/r1026abs.htm.

Narayanaswami:2002:ILW

[548] Chandra Narayanaswami, Noboru Kamijoh, Mandayam Raghunath, Tadanobu Inoue, Thomas Cipolla, Jim Sanford, Eugene Schlig, Sreekrishnan Venkiteswaran, Dinakar Guniguntala, Vishal Kulkarni, and Kazuhiko Yamazaki. IBM’s Linux watch: The challenge of miniaturization. Computer, 35(1):33–41, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1033.pdf;
http://www.computer.org/computer/
co2002/r1033abs.htm.

Allan:2002:TRS

[549] Alan Allan, Don Edenfeld, William H. Joyner, Jr., Andrew B. Kahng, Mike Rodgers, and Yervant Zorian. 2001 technology roadmap for semiconductors. Computer, 35(1):42–53, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1042.pdf;
http://www.computer.org/computer/
co2002/r1042abs.htm.

Schatz:2002:ICN

[550] Bruce R. Schatz. The interspace: Concept navigation across distributed communities. Computer, 35(1):54–62, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1054.pdf;
http://www.computer.org/computer/
co2002/r1054abs.htm.

Boehm:2002:GRA

[551] Barry Boehm. Get ready for agile methods, with care. Computer, 35(1):64–69, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1064.pdf;
http://www.computer.org/computer/
co2002/r1064abs.htm.

Benini:2002:NCN

[552] Luca Benini and Giovanni De Micheli. Networks on chips: A new SoC paradigm. Computer, 35(1):70–78, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2002/pdf/
r1070.pdf;   http://www.computer.
org/computer/co2002/r1070abs.htm.

Oskin:2002:PAR

[553] Mark Oskin, Frederic T. Chong, and Isaac L. Chuang. A practical architecture for reliable quantum computers. Computer, 35(1):79–87, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1079.pdf;
http://www.computer.org/computer/
co2002/r1079abs.htm.

Leen:2002:EAE

[554] Gabriel Leen and Donal Heffernan. Expanding automotive electronic systems. Computer, 35(1):88–93, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1088.pdf;
http://www.computer.org/computer/
co2002/r1088abs.htm.

Piner:2002:CSCa

[555] Mary-Louise G. Piner. Computer Society connection: Building on success: CS programs for 2002; conferences; board approves reading of bylaws amendment. Computer, 35(1):94–99, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1094.pdf;
http://www.computer.org/computer/
co2002/r1094abs.htm.

Anonymous:2002:CCa

[556] Anonymous. Call and calendar. Computer, 35(1):100–102, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1100.pdf;
http://www.computer.org/computer/
co2002/r1100abs.htm.

Anonymous:2002:COa

[557] Anonymous. Career opportunities. Computer, 35(1):104–129, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1104.pdf;

http://www.computer.org/computer/
co2002/r1104abs.htm.

Anonymous:2002:PRSa

[558] Anonymous. Products: Rational Software upgrades development suite; UniPress Software’s Web-based development and help desk tool; SGI’s OpenGL upgrade; ILOG’s Component Suite for building enterprise portals; Microsoft’s Windows XP Embedded; Borland application servers. Computer, 35(1):130–131, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1130.pdf;
http://www.computer.org/computer/
co2002/r1130abs.htm.

Lutz:2002:BIM

[559] Michael J. Lutz. Bookshelf: Intelligent machines and systems. Computer, 35(1):132, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1132.pdf;
http://www.computer.org/computer/
co2002/r1132abs.htm.

Starner:2002:WTT

[560] Thad Starner. Web technologies: Thick clients for personal wireless devices. Computer, 35(1):133–135, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1133.pdf;
http://www.computer.org/computer/
co2002/r1133abs.htm.

Wolf:2002:EC

[561] Wayne Wolf. Embedded computing: What is embedded computing? Computer, 35(1):136–137, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1136.pdf;
http://www.computer.org/computer/
co2002/r1136abs.htm.

Mathieu:2002:ISP

[562] Richard G. Mathieu. IT systems perspective: Top-down approach to computing. Computer, 35(1):138–139, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2002/pdf/r1138.pdf;
http://www.computer.org/computer/
co2002/r1138abs.htm.

Fuller:2002:SAF

[563] Sam Fuller and Tom Cox. Standards: Anatomy of a forwardlooking open standard. Computer, 35(1):140–141, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1140.pdf;
http://www.computer.org/computer/
co2002/r1140abs.htm.

Holmes:2002:PSO

[564] Neville Holmes. The profession: To see ourselves as others see us. Computer, 35(1):144, 142–143, January 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r1144.pdf;

http://www.computer.org/computer/
co2002/r1144abs.htm.

Schaefer:2002:LPT

[565] Robert Schaefer, Cristina Cifuentes, Irah H. Donner, David Parry, Neville Holmes, and Mark Wallace. Letters: Protecting technologies with copyrights and patents; the computing profession responds to global concerns. Computer, 35(2):6–8, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r2006.pdf;
http://www.computer.org/computer/
co2002/r2006abs.htm.

Anonymous:2002:EUR

[566] Anonymous. Erratum: “Understanding and Reducing Web Delays”. Computer, 35(2):8, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See [522].

Colwell:2002:CRR

[567] Bob Colwell. On computers and rock ’n’ roll. Computer, 35(2):9–11, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2002/pdf/r2009.pdf;
 http://www.computer.org/computer/
co2002/r2009abs.htm.

Anonymous:2002:ASb

[568] Anonymous. Article summaries. Computer, 35(2):12–13, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r2012.pdf;
http://www.computer.org/computer/
co2002/r2012abs.htm.

Vaughan-Nichols:2002:WSB

[569] Steven J. Vaughan-Nichols. Web services: Beyond the hype. Computer, 35(2):18–21, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r2018.pdf;
http://www.computer.org/computer/
co2002/r2018abs.htm.

Paulson:2002:WMN

[570] Linda Dailey Paulson. Wanted: More network-security graduates and research. Computer, 35(2):22–24, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2002/pdf/r2022.pdf;
http://www.computer.org/computer/
co2002/r2022abs.htm.

Paulson:2002:NBIa

[571] Linda Dailey Paulson. News briefs: IBM begins autonomic-computing project; malicious code hits instant messaging; handheld technology’s got your back; faster processors with optical connections. Computer, 35(2):25–27, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r2025.pdf;
http://www.computer.org/computer/
co2002/r2025abs.htm.

Kraemer:2002:EDC

[572] Kenneth L. Kraemer and Jason Dedrick. Enter the dragon: China’s computer industry. Computer, 35(2): 28–36, February 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2002/pdf/
r2028.pdf;   http://www.computer.
org/computer/co2002/r2028abs.htm.

Mukherjee:2002:PST

[573] Shubhendu S. Mukherjee, Sarita V. Adve, Todd Austin, Joel Emer, and Peter S. Magnusson. Performance simulation tools. Computer, 35(2):38–39, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r2038.pdf;
http://www.computer.org/computer/
co2002/r2038abs.htm.

Hughes:2002:RSS

[574] Christopher J. Hughes, Vijay S. Pai, Parthasarathy Ranganathan, and Sarita V. Adve. RSIM: Simulating shared-memory multiprocessors with ILP processors. Computer, 35(2):40–49, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r2040.pdf;
http://www.computer.org/computer/
co2002/r2040abs.htm.

Magnusson:2002:SFS

[575] Peter S. Magnusson, Magnus Christensson, Jesper Eskilson, Daniel Forsgren, Gustav Hållberg, Johan Högberg, Fredrik Larsson, Andreas Moestedt, and Bengt Werner. Simics: A full system simulation platform. Computer, 35(2):50–58, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r2050.pdf;

http://www.computer.org/computer/
co2002/r2050abs.htm.

Austin:2002:SIC

[576] Todd Austin, Eric Larson, and Dan Ernst. SimpleScalar: An infrastructure for computer system modeling. Computer, 35(2):59–67, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r2059.pdf;
http://www.computer.org/computer/
co2002/r2059abs.htm.

Emer:2002:APM

[577] Joel Emer, Pritpal Ahuja, Eric Borch, Artur Klauser, Chi-Keung Luk, Srilatha Manne, Shubhendu S. Mukherjee, Harish Patil, Steven Wallace, Nathan Binkert, Roger Espasa, and Toni Juan. Asim: A performance model framework. Computer, 35(2):68–76, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r2068.pdf;
http://www.computer.org/computer/
co2002/r2068abs.htm.

Schneidewind:2002:BKS

[578] Norman F. Schneidewind. Body of knowledge for software quality measurement. Computer, 35(2):77–83, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2002/pdf/
r2077.pdf;   http://www.computer.
org/computer/co2002/r2077abs.htm.

Piner:2002:CSCb

[579] Mary-Louise G. Piner. Computer Society connection: IEEE Fellows lead the engineering profession; Fellow nominations due 15 March; Society seeks Editor-in-Chief applicants; Society meeting and election schedule now available. Computer, 35(2):84–88, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r2084.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r2084abs.htm.

Anonymous:2002:CCb

[580] Anonymous. Call and calendar. Computer, 35(2):89–91, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r2089.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r2089abs.htm.

Anonymous:2002:COb

[581] Anonymous. Career opportunities. Computer, 35(2):92–103, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r2092.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r2092abs.htm.

Lutz:2002:BMK

[582] Michael J. Lutz. Bookshelf: Mining knowledge. Computer, 35(2): 105, February 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2002/pdf/

r2105.pdf;   http://www.computer.
org/computer/co2002/r2105abs.htm.

Anonymous:2002:PIR

[583] Anonymous. Products: Instantiations releases RAD tools for Java; Pencel updates XML mobile application interface; LynuxWorks’s new Linux distribution; Applied Microsystems’ embedded systems code tester; AppForge upgrades cross-platform design tool for handhelds; Computer Associates’ Java, XML, and wireless extensions; TurboPower;s bug detection and performance analysis tools. Computer, 35(2):106–108, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r2106.pdf;
http://www.computer.org/computer/
co2002/r2106abs.htm.

Arbaugh:2002:STS

[584] Bill Arbaugh. Security: Technical, social, and legal challenges. Computer, 35(2):109–111, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r2109.pdf;
http://www.computer.org/computer/
co2002/r2109abs.htm.

Macedonia:2002:LHS

[585] Michael Macedonia. Linux in Hollywood: A star is born. Computer, 35(2):112–114, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r2112.pdf;
http://www.computer.org/computer/
co2002/r2112abs.htm.

Varshney:2002:MSM

[586] Upkar Varshney. Multicast support in mobile commerce applications. Computer, 35(2):115–117, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r2115.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r2115abs.htm.

Holmes:2002:RDP

[587] Neville Holmes. Representative democracy and the profession. Computer, 35(2):120, 118–119, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r2120.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r2120abs.htm.

Leigh:2002:LTF

[588] Dave Leigh, Keith Southwell, Thomas F. Bitonti, Barry Boehm, Eric Matthews, Neville Holmes, and Benn Hemmen. Letters: Touting FCD’s benefits; professional approach to software development; unavoidable statistics; keeping us humble, honest, and human; A URL and a story. Computer, 35(3):6–7, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r3006.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r3006abs.htm.

Colwell:2002:ESQ

[589] Bob Colwell. Engineering, science, and quantum mechanics. Computer, 35 (3):8–10, March 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2002/pdf/
r3008.pdf;   http://www.computer.
org/computer/co2002/r3008abs.htm.

Anonymous:2002:ASc

[590] Anonymous. Article summaries. Computer, 35(3):11–12, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r3011.pdf;
http://www.computer.org/computer/
co2002/r3011abs.htm.

Lawton:2002:OSS

[591] George Lawton. Open source security: Opportunity or oxymoron? Computer, 35(3):18–21, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r3018.pdf;
http://www.computer.org/computer/
co2002/r3018abs.htm.

Clark:2002:MPB

[592] David Clark. Mobile processors begin to grow up. Computer, 35(3):22–25, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2002/pdf/r3022.pdf;
http://www.computer.org/computer/
co2002/r3022abs.htm.

Paulson:2002:NBK

[593] Linda Dailey Paulson. News briefs: Key snooping technology causes controversy; developments could boost optical networks; adding a third dimension to chips. Computer, 35(3):27–29, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2002/pdf/r3027.pdf;
 http://www.computer.org/computer/
co2002/r3027abs.htm.

Hauck:2002:UCA

[594] Roslin V. Hauck, Homa Atabakhsh, Pichai Ongvasith, Harsh Gupta, and Hsinchun Chen. Using Coplink to analyze criminal-justice data. Computer, 35(3):30–37, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r3030.pdf;
http://www.computer.org/computer/
co2002/r3030abs.htm.

Gueziec:2002:TPB

[595] André Guéziec. Tracking pitches for broadcast television. Computer, 35(3):38–43, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r3038.pdf;
http://www.computer.org/computer/
co2002/r3038abs.htm.

Poupyrev:2002:DGA

[596] Ivan Poupyrev, Desney S. Tan, Mark Billinghurst, Hirokazu Kato, Holger Regenbrecht, and Nobuji Tetsutani. Developing a generic augmented-reality interface. Computer, 35(3):44–50, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r3044.pdf;
http://www.computer.org/computer/
co2002/r3044abs.htm.

Fenstermacher:2002:LFC

[597] Kurt D. Fenstermacher and Mark Ginsburg. A lightweight framework for cross-application user monitoring. Computer, 35(3):51–59, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r3051.pdf;
http://www.computer.org/computer/
co2002/r3051abs.htm.

Allanson:2002:EIC

[598] Jennifer Allanson. Electrophysiologically interactive computer systems. Computer, 35(3):60–65, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r3060.pdf;
http://www.computer.org/computer/
co2002/r3060abs.htm.

Flake:2002:SOI

[599] Gary William Flake, Steve Lawrence, C. Lee Giles, and Frans M. Coetzee. Self-organization and identification of Web communities. Computer, 35(3):66–71, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r3066.pdf;
http://www.computer.org/computer/
co2002/r3066abs.htm.

Miranda:2002:PET

[600] Eduardo Miranda. Planning and executing time-bound projects. Computer, 35(3):73–79, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r3073.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r3073abs.htm.

Piner:2002:CSCc

[601] Mary-Louise G. Piner and Alan Clements. Computer Society connection: Online courses offer members expert instruction; revised student competition offers opportunity for increased participation. Computer, 35(3):80–84, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r3080.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r3080abs.htm.

Anonymous:2002:CCc

[602] Anonymous. Call and calendar. Computer, 35(3):85–88, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r3085.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r3085abs.htm.

Anonymous:2002:COc

[603] Anonymous. Career opportunities. Computer, 35(3):90–95, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r3090.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r3090abs.htm.

Lutz:2002:BBH

[604] Michael J. Lutz. Bookshelf: Beating hackers at their own game.

Computer, 35(3):96, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r3096.pdf;
http://www.computer.org/computer/
co2002/r3096abs.htm.

Anonymous:2002:PPJ

[605] Anonymous. Products: PrismTech’s JDO spec for transparent persistence; Altia’s graphics code generator for embedded applications; Design Science upgrades MathType for windows; PolarLake launches Enterprise XML platform for java; Syware’s database development software for PDAs; code generator for Web application development from YesSoftware; Embarcadero Technologies upgrades cross-platform job scheduler; Performance Technologies introduces telecom adapter; Rational Software’s latest IDE enhancement; Aprisa’s online research and design environment. Computer, 35(3):97–99, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r3097.pdf;
http://www.computer.org/computer/
co2002/r3097abs.htm.

Martin:2002:DCE

[606] Grant Martin and Frank Schirrmeister. A design chain for embedded systems. Computer, 35(3):100–103, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r3100.pdf;
http://www.computer.org/computer/
co2002/r3100abs.htm.

Steinfield:2002:RBV

[607] Charles Steinfield. Realizing the benefits of virtual teams. Computer, 35(3):104–106, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r3104.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r3104abs.htm.

Spink:2002:SCW

[608] Amanda Spink, Bernard J. Jansen, Dietmar Wolfram, and Tefko Saracevic. From E-sex to E-commerce: Web search changes. Computer, 35(3): 107–109, March 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2002/pdf/
                                        r3107.pdf;   http://www.computer.
                                        org/computer/co2002/r3107abs.htm.

Holmes:2002:CPP

[609] Neville Holmes. Computers, programming, and people. Computer, 35(3):112, 110–111, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r3112.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r3112abs.htm.

Lane:2002:LSD

[610] Paul D. Lane, Tom Pittman, John Marquet, Neville Holmes, Ben Stanley, and Jonah Z. Lavi. Letters: Software development risk metrics; on computers and a different drummer; democracy in Internet time; Linux game support; A call for ECBS educational resources. Computer, 35(4):6–8, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r4006.pdf;
http://www.computer.org/computer/
co2002/r4006abs.htm.

Colwell:2002:NMM

[611] Bob Colwell. Near misses: Murphy’s law is wrong. Computer, 35(4):9–12, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2002/pdf/r4009.pdf;
http://www.computer.org/computer/
co2002/r4009abs.htm.

Anonymous:2002:ASd

[612] Anonymous. Article summaries. Computer, 35(4):13, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r4013.pdf;
http://www.computer.org/computer/
co2002/r4013abs.htm.

Paulson:2002:FRT

[613] Linda Dailey Paulson. Faster RAM tackles data and marketplace bottlenecks. Computer, 35(4):17–19, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r4017.pdf;
http://www.computer.org/computer/
co2002/r4017abs.htm.

Paulson:2002:NBPa

[614] Linda Dailey Paulson. News briefs: Protocol-related problems threaten Internet security; faster storage connectivity technologies emerge; got to go googlewhacking; new wireless technology is in sight. Computer, 35(4):20–22, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r4020.pdf;
http://www.computer.org/computer/
co2002/r4020abs.htm.

Potlapally:2002:OFP

[615] Nachiketh R. Potlapally. Optical fingerprinting to protect data: A proposal. Computer, 35(4):23–28, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r4023.pdf;
http://www.computer.org/computer/
co2002/r4023abs.htm.

Welch:2002:TIA

[616] Donald Welch, Daniel Ragsdale, and Wayne Schepens. Training for information assurance. Computer, 35(4):30–37, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2002/pdf/r4030.pdf;
 http://www.computer.org/computer/
co2002/r4030abs.htm.

Srinivasan:2002:SRB

[617] Savitha Srinivasan and Eric Brown. Is speech recognition becoming mainstream? Computer, 35(4):38–41, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r4038.pdf;
http://www.computer.org/computer/
co2002/r4038abs.htm.

Padmanabhan:2002:LVS

[618] Mukund Padmanabhan and Michael Picheny. Large-vocabulary speech recognition algorithms. Computer, 35 (4):42–50, April 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2002/pdf/
r4042.pdf;   http://www.computer.
org/computer/co2002/r4042abs.htm.

Gorin:2002:ANS

[619] Allen L. Gorin, Alicia Abella, Tirso Alonso, Giuseppe Riccardi, and Jeremy H. Wright. Automated natural spoken dialog. Computer, 35(4):51–56, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2002/pdf/r4051.pdf;
http://www.computer.org/computer/
co2002/r4051abs.htm.

Moreno:2002:MRK

[620] Pedro J. Moreno, J. M. Van Thong, Beth Logan, and Gareth J. F. Jones. From multimedia retrieval to knowledge management. Computer, 35(4): 58–66, April 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2002/pdf/
r4058.pdf;   http://www.computer.
org/computer/co2002/r4058abs.htm.

Thayer:2002:SSE

[621] Richard H. Thayer. Software system engineering: A tutorial. Computer, 35(4):68–73, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r4068.pdf;

http://www.computer.org/computer/
co2002/r4068abs.htm.

Piner:2002:CSCd

[622] Mary-Louise G. Piner. Computer Society connection: Certification program provides mark of professionalism; CD-ROM supplement provides demos, interactive tutorials for computer graphics community; Academic Excellence Prize and Best Paper Scholarship awarded. Computer, 35(4):74–76, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r4074.pdf;
http://www.computer.org/computer/
co2002/r4074abs.htm.

Anonymous:2002:CCd

[623] Anonymous. Call and calendar. Computer, 35(4):77–79, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r4077.pdf;
http://www.computer.org/computer/
co2002/r4077abs.htm.

Anonymous:2002:COd

[624] Anonymous. Career opportunities. Computer, 35(4):80–84, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r4080.pdf;
http://www.computer.org/computer/
co2002/r4080abs.htm.

Meyer:2002:ACT

[625] Bertrand Meyer. Assessing a C# text. Computer, 35(4):86–88, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r4086.pdf;
http://www.computer.org/computer/
co2002/r4086abs.htm.

Anonymous:2002:B

[626] Michael J. Lutz. Bookshelf: Balancing speed and complexity with compilers; advanced cryptology text; computer architecture basics; mutation testing compendium. Computer, 35(4):89, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2002/pdf/r4089.pdf;
 http://www.computer.org/computer/
co2002/r4089abs.htm.

Anonymous:2002:PCC

[627] Anonymous. Products: Cape Clear Software’s Web services IDE; Metrowerk’s development tools for Symbian OS; Symantec’s new firewall and VPN software; PACT Technologies unveils parallel C compiler; Tonic Software updates business continuity software; Ampro’s processor module for embedded system designs. Computer, 35(4):90–91, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r4090.pdf;
http://www.computer.org/computer/
co2002/r4090abs.htm.

Bajaj:2002:GLT

[628] Rashmi Bajaj, Samantha Lalinda Ranaweera, and Dharma P. Agrawal. GPS: Location-tracking technology. Computer, 35(4):92–94, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r4092.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r4092abs.htm.

Korris:2002:ECD

[629] James Korris and Michael Macedonia. The end of celluloid: Digital cinema emerges. Computer, 35(4):96–98, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r4096.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r4096abs.htm.

McGraw:2002:MSS

[630] Gary McGraw. Managing software security risks. Computer, 35(4):99–101, April 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co2002/pdf/r4099.pdf;
                                         http://www.computer.org/computer/
                                        co2002/r4099abs.htm.

Holmes:2002:SED

[631] Neville Holmes. Side effects of digital technology. Computer, 35(4):104, 102–103, April 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2002/pdf/
                                        r4104.pdf;   http://www.computer.
                                        org/computer/co2002/r4104abs.htm.

Stom:2002:LQM

[632] Stephen Stom, Bob Colwell, Nicolas Gold, Steven J. Vaughan-Nicols, Werner Donné, Larry L. Gadeken, Neville Holmes, Tom Murray, Jason Che han Yip, V. Chitradevi, Mori Anvari, and Ray Findlay. Letters:

Quantum mystery mongering; beyond the hype; computers: Necessity or convenience? sanctions raise fairness issues. Computer, 35(5):6–10, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r5006.pdf;
http://www.computer.org/computer/
co2002/r5006abs.htm.

Colwell:2002:IYD

[633] Bob Colwell. If you didn’t test it, it doesn’t work. Computer, 35(5):11–13, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2002/pdf/r5011.pdf;
 http://www.computer.org/computer/
co2002/r5011abs.htm.

Anonymous:2002:ASe

[634] Anonymous. Article summaries. Computer, 35(5):17, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r5017.pdf;
http://www.computer.org/computer/
co2002/r5017abs.htm.

Edwards:2002:SGN

[635] John Edwards and Linda Dailey Paulson. Smart graphics: A new approach to meeting user needs. Computer, 35(5):18–21, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r5018.pdf;
http://www.computer.org/computer/
co2002/r5018abs.htm.

Leavitt:2002:DMC

[636] Neal Leavitt. Data mining for the corporate masses? Computer, 35 (5):22–24, May 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2002/pdf/
                                        r5022.pdf;   http://www.computer.
                                        org/computer/co2002/r5022abs.htm.

Paulson:2002:NBM

[637] Linda Dailey Paulson. News briefs: MPEG licensing dispute hurts Apple; will new chip revolutionize digital photography?; much of the Internet is out of reach; US approves new uses of wireless technology. Computer, 35(5):25–27, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r5025.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r5025abs.htm.

Blaha:2002:PSV

[638] Michael Blaha and Cheryl Smith. A pattern for softcoded values. Computer, 35(5):28–34, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r5028.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r5028abs.htm.

Eischen:2002:SDO

[639] Kyle Eischen. Software development: An outsider’s view. Computer, 35 (5):36–44, May 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2002/pdf/

r5036.pdf;   http://www.computer.
org/computer/co2002/r5036abs.htm.

Kim:2002:CCB

[640] Won Kim, Ki-Joon Chae, Dong-Sub Cho, Byoungju Choi, Anmo Jeong, Myung Kim, KiHo Lee, Meejeong Lee, Sang-Ho Lee, Seung-Soo Park, and Hwan-Seung Yong. The Chamois component-based knowledge engineering framework. Computer, 35(5):46–54, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co2002/pdf/r5046.pdf;
 http://www.computer.org/computer/
co2002/r5046abs.htm.

Aissi:2002:BPM

[641] Selim Aissi, Pallavi Malu, and Krishnamurthy Srinivasan. E-business process modeling: The next big step. Computer, 35(5):55–62, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r5055.pdf;
http://www.computer.org/computer/
co2002/r5055abs.htm.

Raghu:2002:NSM

[642] T. S. Raghu, R. Ramesh, and Andrew B. Whinston. Next steps for mobile entertainment portals. Computer, 35(5):63–70, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r5063.pdf;
http://www.computer.org/computer/
co2002/r5063abs.htm.

Jackson:2002:DPS

[643] Laura E. Jackson and George N. Rouskas. Deterministic preemptive scheduling of real-time tasks. Computer, 35(5):72–79, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r5072.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r5072abs.htm.

Shaikh:2002:EET

[644] Fayaz A. Shaikh, Stan McClellan, Manpreet Singh, and Sannedhi K. Chakravarthy. End-to-end testing of IP QoS mechanisms. Computer, 35 (5):80–87, May 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co2002/pdf/
                                        r5080.pdf;   http://www.computer.
                                        org/computer/co2002/r5080abs.htm.

Anonymous:2002:CCe

[645] Anonymous. Call and calendar. Computer, 35(5):88–91, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r5088.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r5088abs.htm.

Piner:2002:CSCe

[646] Mary-Louise G. Piner. Computer Society connection: Conferences offer findings in both new and traditional computing research fields. Computer, 35(5):92–94, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r5092.pdf;
http://www.computer.org/computer/
co2002/r5092abs.htm.

Anonymous:2002:COe

[647] Anonymous. Career opportunities. Computer, 35(5):96–98, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r5096.pdf;
http://www.computer.org/computer/
co2002/r5096abs.htm.

Lutz:2002:BAE

[648] Michael J. Lutz. Bookshelf: The art of evolutionary computation; protection for all media; breaking the gender barrier; software reuse: The state of the art. Computer, 35(5):99–??, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r5099.pdf;
http://www.computer.org/computer/
co2002/r5099abs.htm.

Anonymous:2002:PAU

[649] Anonymous. Products: Atinav upgrades Bluetooth software suite; new 3D terrain modeler from MultiGen-Paradigm; RIM introduces BlackBerry development environment for J2ME; Mercury Interactive adds Forte for Java to load-testing system; Wild-Packets’ real-time wireless analyzer; Novell updates mobile file management software; Layton Graphics unveils AutoCAD-to-PDF translator; Iomegas network attached storage servers. Computer, 35(5):100–101, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r5100.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r5100abs.htm.

Lyon:2002:TMT

[650] Douglas A. Lyon and Christopher L. Huntley. There’s more than one way to build a bridge. Computer, 35(5):102–103, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r5102.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r5102abs.htm.

McMillan:2002:TSP

[651] Kate McMillan. Technology standards pros aid homeland security. Computer, 35(5):104–105, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r5104.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r5104abs.htm.

Wolf:2002:HHE

[652] Wayne Wolf. Household hints for embedded systems designers. Computer, 35(5):106–108, May 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co2002/pdf/r5106.pdf;
                                        http://www.computer.org/computer/
                                        co2002/r5106abs.htm.

Holmes:2002:WDB

[653] Neville Holmes. Would a digital brain have a mind? Computer, 35(5):112, 110–111, May 2002.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2002/pdf/r5112.pdf;
http://www.computer.org/computer/
co2002/r5112abs.htm.

Scime:2002:LIS

[654] Anthony Scime, Bob Pedigo, Curtis Sjolander, Bob Colwell, Dick Brodine, Ted Lewis, Bruce D. Wilner, Gary McGraw, Jay Dearien, Neville Holmes, and Peter Stern. Letters: Information system engineering; quantum questions; is Murphy’s Law wrong?; Java questions for Ted; software security concerns; insightful observations on computers and society. Computer, 35(6):6–9, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/06/r6006.htm;
http://csdl.computer.org/dl/mags/
co/2002/06/r6006.pdf.

Colwell:2002:REF

[655] Bob Colwell. At random: Engineering follies. Computer, 35(6):10–12, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/06/r6010.htm;
http://csdl.computer.org/dl/mags/
co/2002/06/r6010.pdf;   http://
www.computer.org/computer/co2002/
r6010abs.htm.

Anonymous:2002:ASf

[656] Anonymous. Article summaries. Computer, 35(6):13–??, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/06/r6013.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/06/r6013.pdf.

Lawton:2002:MJM

[657] George Lawton. Moving Java into mobile phones. Computer, 35(6): 17–20, June 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2002/06/
                                        r6017.htm;   http://csdl.computer.
                                        org/dl/mags/co/2002/06/r6017.pdf.

Vaughan-Nichols:2002:WGE

[658] Steven J. Vaughan-Nichols. Will 10gigabit Ethernet have a bright future? Computer, 35(6):22–24, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/06/r6022.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/06/r6022.pdf.

Paulson:2002:NBL

[659] Linda Dailey Paulson. News briefs: Little chips promise big power. Computer, 35(6):25, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/06/r6025.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/06/r6025.pdf.

Paulson:2002:NBB

[660] Linda Dailey Paulson. News briefs: Bioinformatics experiences important breakthroughs. Computer, 35(6):26– 27, June 2002. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2002:NBC

[661] Linda Dailey Paulson. News briefs: Company unveils holographic DVD for 3D storage. Computer, 35(6):27, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anderson:2002:FSO

[662] Ross Anderson. Free speech online and offline. Computer, 35(6):28–30, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/06/r6028.htm;
http://csdl.computer.org/dl/mags/
co/2002/06/r6028.pdf;   http://
www.computer.org/computer/co2002/
r6028abs.htm.

Tripp:2002:SCD

[663] Leonard L. Tripp. Software certification debate: Benefits of certification. Computer, 35(6):31–33, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/06/r6031.htm;
http://csdl.computer.org/dl/mags/
co/2002/06/r6031.pdf;   http://
www.computer.org/computer/co2002/
r6031abs.htm.

Kolawa:2002:SCD

[664] Adam Kolawa. Software certification debate: Certification will do more harm than good. Computer, 35(6): 34–35, June 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/06/

r6034.htm;   http://csdl.computer.
org/dl/mags/co/2002/06/r6034.pdf;
 http://www.computer.org/computer/
co2002/r6034abs.htm.

Foster:2002:GSD

[665] Ian Foster, Carl Kesselman, Jeffrey M. Nick, and Steven Tuecke. Grid services for distributed system integration. Computer, 35(6):37–46, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/06/r6037.htm;
http://csdl.computer.org/dl/mags/
co/2002/06/r6037.pdf;   http://
www.computer.org/computer/co2002/
r6037abs.htm.

Kramer:2002:STH

[666] William T. C. Kramer. SCinet: Testbed for high-performance networked applications. Computer, 35 (6):47–55, June 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/06/
r6047.htm;   http://csdl.computer.
org/dl/mags/co/2002/06/r6047.pdf;
 http://www.computer.org/computer/
co2002/r6047abs.htm.

Shahabi:2002:YSG

[667] Cyrus Shahabi, Roger Zimmermann, Kun Fu, and Shu-Yuen Didi Yao. Yima: A second-generation continuous media server. Computer, 35 (6):56–64, June 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/06/
r6056.htm;   http://csdl.computer.
org/dl/mags/co/2002/06/r6056.pdf;

 http://www.computer.org/computer/
co2002/r6056abs.htm.

Bahn:2002:ERN

[668] Hyokyung Bahn, Kern Koh, Sam H. Noh, and Sang Lyul Min. Efficient replacement of nonuniform objects in Web caches. Computer, 35 (6):65–73, June 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/06/
r6065.htm;   http://csdl.computer.
org/dl/mags/co/2002/06/r6065.pdf;
 http://www.computer.org/computer/
co2002/r6065abs.htm.

Piner:2002:CSCf

[669] Mary-Louise G. Piner. Computer Society connection: IEEE Computer Society recognizes achievements of leading computer engineers: Per Brinch Hansen recognized for pioneering works; Stanford President Hennessy receives architecture award; Ian F. Akyildiz named Goode winner; Society digital library collection is expanded; Mills Award honors software engineering achievements. Computer, 35(6):74–76, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/06/r6074.htm;
http://csdl.computer.org/dl/mags/
co/2002/06/r6074.pdf.

Anonymous:2002:CCf

[670] Anonymous. Call and calendar. Computer, 35(6):77–80, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/06/r6077.htm;

                                        http://csdl.computer.org/dl/mags/
                                        co/2002/06/r6077.pdf.

Lutz:2002:BRI

[671] Michael J. Lutz. Bookshelf: Recipes for image analysis. Computer, 35 (6):81, June 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2002/06/
                                        r6081.htm;   http://csdl.computer.
                                        org/dl/mags/co/2002/06/r6081.pdf.

Anonymous:2002:PXO

[672] Anonymous. Products: XML-oriented FrameMaker from Adobe Systems; Breeze Factor’s XML data-binding tool for Java; Microsoft’s speech-enabling tools for Web applications; Agere Systems’s wireless LAN management tool; Macromedia debuts ColdFusion MX; OpenOffice.org unveils open source productivity suite; Integrated Measurement Systems’ validation tester. Computer, 35(6):82–83, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/06/r6082.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/06/r6082.pdf.

DeMarco:2002:STA

[673] Tom DeMarco and Barry Boehm. Software technologies: The agile methods fray. Computer, 35(6):90–92, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/06/r6090.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/06/r6090.pdf;   http://

www.computer.org/computer/co2002/
r6090abs.htm.

Kenneally:2002:SWL

[674] Erin Kenneally. Security: Who’s liable for insecure networks? Computer, 35(6):93–95, June 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/06/
r6093.htm;   http://csdl.computer.
org/dl/mags/co/2002/06/r6093.pdf;
 http://www.computer.org/computer/
co2002/r6093abs.htm.

Avancha:2002:CES

[675] Sasikanth Avancha, Anupam Joshi, and Timothy Finin. Communications: Enhanced service discovery in Bluetooth. Computer, 35(6):96–99, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/06/r6096.htm;
http://csdl.computer.org/dl/mags/
co/2002/06/r6096.pdf;   http://
www.computer.org/computer/co2002/
r6096abs.htm.

Norman:2002:ECH

[676] Donald A. Norman. Entertainment computing: Home theater: Not ready for prime time. Computer, 35(6):100–102, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/06/r6100.htm;
http://csdl.computer.org/dl/mags/
co/2002/06/r6100.pdf;   http://
www.computer.org/computer/co2002/
r6100abs.htm.

Asuaga:2002:PUC

[677] Ana Asuaga. The profession: Using computers in our daily life. Computer, 35(6):104, 103, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/06/r6104.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/06/r6104.pdf;   http://
                                        www.computer.org/computer/co2002/
                                        r6104abs.htm.

Colwell:2002:RWT

[678] Bob Colwell. At random: What technology is doing to music. Computer, 35(7):6–8, July 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2002/07/
                                        r7006.htm;   http://csdl.computer.
                                        org/dl/mags/co/2002/07/r7006.pdf;
                                         http://www.computer.org/computer/
                                        co2002/r7006abs.htm.

Anonymous:2002:ASg

[679] Anonymous. Article summaries. Computer, 35(7):9–10, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/07/r7009.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/07/r7009.pdf.

Ortiz:2002:ITB

[680] Sixto Ortiz Jr. Industry trends: Is business intelligence a smart move? Computer, 35(7):11–14, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/07/r7011.htm;

http://csdl.computer.org/dl/mags/
co/2002/07/r7011.pdf;   http://
www.computer.org/computer/co2002/
r7011abs.htm.

Lawton:2002:TNI

[681] George Lawton. Technology news: Invasive software: Who’s inside your computer? Computer, 35(7):15– 18, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/07/
r7015.htm;   http://csdl.computer.
org/dl/mags/co/2002/07/r7015.pdf;
 http://www.computer.org/computer/
co2002/r7015abs.htm.

Paulson:2002:NBTa

[682] Linda Dailey Paulson. News briefs: Taking a graphical approach to the password; improvements stream into multimedia technology; data at the speed of light. Computer, 35(7): 19–21, July 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/07/
r7019.htm;   http://csdl.computer.
org/dl/mags/co/2002/07/r7019.pdf.

Kim:2002:PCM

[683] Junhyong Kim. Perspectives: Computers are from Mars, organisms are from Venus. Computer, 35(7):25–32, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/07/r7025.htm;
http://csdl.computer.org/dl/mags/
co/2002/07/r7025.pdf;   http://
www.computer.org/computer/co2002/
r7025abs.htm.

Feitelson:2002:PBL

[684] Dror G. Feitelson and Millet Treinin. Perspectives: The blueprint for life? Computer, 35(7):34–40, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/07/r7034.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/07/r7034.pdf;   http://
                                        www.computer.org/computer/co2002/
                                        r7034abs.htm.

Heath:2002:GEI

[685] Lenwood S. Heath and Naren Ramakrishnan. Guest Editors’ introduction: The emerging landscape of bioinformatics software systems. Computer, 35(7):41–45, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/07/r7041.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/07/r7041.pdf;   http://
                                        www.computer.org/computer/co2002/
                                        r7041abs.htm.

Pop:2002:CFG

[686] Mihai Pop, Steven L. Salzberg, and Martin Shumway. Cover feature: Genome sequence assembly: Algorithms and issues. Computer, 35 (7):47–54, July 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2002/07/
                                        r7047.htm;   http://csdl.computer.
                                        org/dl/mags/co/2002/07/r7047.pdf;
                                         http://www.computer.org/computer/
                                        co2002/r7047abs.htm.

Moret:2002:CFT

[687] Bernard M. E. Moret, Li-San Wang, and Tandy Warnow. Cover feature: Toward new software for computational phylogenetics. Computer, 35 (7):55–64, July 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/07/
r7055.htm;   http://csdl.computer.
org/dl/mags/co/2002/07/r7055.pdf;
 http://www.computer.org/computer/
co2002/r7055abs.htm.

Parvin:2002:CFB

[688] Bahram Parvin, Qing Yang, Gerald Fontenay, and Mary Helen Barcellos-Hoff. Cover feature: BioSig: An imaging bioinformatic system for studying phenomics. Computer, 35(7): 65–71, July 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/07/
r7065.htm;   http://csdl.computer.
org/dl/mags/co/2002/07/r7065.pdf;
 http://www.computer.org/computer/
co2002/r7065abs.htm.

Paxia:2002:CFR

[689] Salvatore Paxia, Archisman Rudra, Yi Zhou, and Bud Mishra. Cover feature: A random walk down the genomes: DNA evolution in Valis. Computer, 35(7):73–79, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/07/r7073.htm;
http://csdl.computer.org/dl/mags/
co/2002/07/r7073.pdf;   http://
www.computer.org/computer/co2002/
r7073abs.htm.

Seo:2002:CFI

[690] Jinwook Seo and Ben Shneiderman. Cover feature: Interactively exploring hierarchical clustering results. Computer, 35(7):80–86, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/07/r7080.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/07/r7080.pdf;   http://
                                        www.computer.org/computer/co2002/
                                        r7080abs.htm.

Piner:2002:CSCg

[691] Mary-Louise G. Piner. Computer Society connection: Ten student teams advance to CSIDC world finals; IEEE Computer Society election coming soon; Board of Governors approves reading of bylaws amendment; new Editors in Chief selected for Computer, six other publications. Computer, 35(7):87–91, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/07/r7087.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/07/r7087.pdf.

Anonymous:2002:CCg

[692] Anonymous. Call and calender. Computer, 35(7):92–94, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/07/r7092.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/07/r7092.pdf.

Lutz:2002:BPD

[693] Michael J. Lutz. Bookshelf: Putting discrepancy theory to work. Computer, 35(7):95–??, July 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/07/
r7095.htm;   http://csdl.computer.
org/dl/mags/co/2002/07/r7095.pdf.

Anonymous:2002:PNR

[694] Anonymous. Products: New rackmount server from Apple; OSDL’s Linux database test tool; DigiTerra ships back-office integration toolset; Oracle’s Developer Suite for business applications and Web services; Macromedia release J2EE-compatible application server and IDE; MotionBuilder 4.0 from Kaydara; Zero G Software’s InstallAnywhere 5. Computer, 35(7):96–97, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/07/r7096.htm;
http://csdl.computer.org/dl/mags/
co/2002/07/r7096.pdf.

Anonymous:2002:COf

[695] Anonymous. Career opportunities. Computer, 35(7):98–100, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/07/r7098.pdf.

Palmer:2002:ISP

[696] Jonathan Palmer. IT systems perspectives: Designing for Web site usability. Computer, 35(7):102–103, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/07/r7102.htm;

                                        http://csdl.computer.org/dl/mags/
                                        co/2002/07/r7102.pdf;   http://
                                        www.computer.org/computer/co2002/
                                        r7102abs.htm.

Rigden:2002:SNY

[697] Christine Rigden. Standards: Now you see it, now you don’t. Computer, 35(7):104–105, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/07/r7104.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/07/r7104.pdf;   http://
                                        www.computer.org/computer/co2002/
                                        r7104abs.htm.

Conte:2002:ECC

[698] Thomas M. Conte. Embedded computing: Choosing the brain(s) of an embedded system. Computer, 35(7):106–107, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/07/r7106.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/07/r7106.pdf;   http://
                                        www.computer.org/computer/co2002/
                                        r7106abs.htm.

Shim:2002:WTI

[699] Simon S. Y. Shim and Yen-Jen Lee. Web technologies: Interactive TV: VoD meets the Internet. Computer, 35(7):108–109, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/07/r7108.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/07/r7108.pdf;   http://

www.computer.org/computer/co2002/
r7108abs.htm.

Holmes:2002:PSG

[700] Neville Holmes. The profession: Seven great blunders of the computing world. Computer, 35(7):112, 110–111, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/07/r7112.htm;
http://csdl.computer.org/dl/mags/
co/2002/07/r7112.pdf;   http://
www.computer.org/computer/co2002/
r7112abs.htm.

Boyes:2002:LPF

[701] David Boyes, Bob Pedigo, Andrew Borland, Elliott Levinson, Adam Kolawa, Zoran Obradovic, Adam Kolawa, Enrique C. Madrona, Gene Cyprych, Brian Berenbach, Tom Dafis, Philip Mason, Kenneth Blake, Kristen Schroeder, Tom DeMarco, Garth Klatt, and Neville Holmes. Letters: Privacy and freedom of speech; not ready for prime time; distinguishing between programming and software engineering; misspent training days; the agile methods fray; dodging thunderbolts. Computer, 35(8):6–9, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/08/r8006.htm;
http://csdl.computer.org/dl/mags/
co/2002/08/r8006.pdf.

Colwell:2002:RWB

[702] B. Colwell. At random: Weird books: Dancing naked; the O. J. Simpson trial; what is E.T.’s home phone number?; traveling the astral plane; physics of Star Trek; great mambo chicken.

Computer, 35(8):10–12, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/08/r8010.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/08/r8010.pdf.

Garber:2002:CCR

[703] Lee Garber. CSIDC competition: Romanian disaster-response project wins design competition. Computer, 35(8):13–15, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/08/r8013.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/08/r8013.pdf.

Anonymous:2002:ASh

[704] Anonymous. Article summaries. Computer, 35(8):19, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/08/r8019.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/08/r8019.pdf.

Paulson:2002:ITC

[705] Linda Dailey Paulson. Industry trends: Computer system, heal thyself. Computer, 35(8):20–22, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/08/r8020.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/08/r8020.pdf.

Paulson:2002:NBD

[706] Linda Dailey Paulson. News briefs: Distributed storage challenges data glut; professors use technology to fight plagiarism; startup wants to automate chip design. Computer, 35(8): 23–25, August 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/08/
r8023.htm;   http://csdl.computer.
org/dl/mags/co/2002/08/r8023.pdf.

Lacity:2002:CPL

[707] Mary Lacity. Computing practices: Lessons in global information technology sourcing. Computer, 35(8):26–33, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/08/r8026.htm;
http://csdl.computer.org/dl/mags/
co/2002/08/r8026.pdf;   http://
www.computer.org/computer/co2002/
r8026abs.htm.

Gupta:2002:CPJ

[708] Rahul Gupta, Sumeet Talwar, and Dharma P. Agrawal. Computing practices: Jini home networking: a step toward pervasive computing. Computer, 35(8):34–40, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/08/r8034.htm;
http://csdl.computer.org/dl/mags/
co/2002/08/r8034.pdf;   http://
www.computer.org/computer/co2002/
r8034abs.htm.

Gutman:2002:CFP

[709] Peter Gutman. Cover feature: PKI: It’s not dead, just resting. Computer, 35(8):41–49, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/08/r8041.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/08/r8041.pdf;   http://
                                        www.computer.org/computer/co2002/
                                        r8041abs.htm.

Cybenko:2002:CFC

[710] George Cybenko, Annarita Giani, and Paul Thompson. Cover feature: Cognitive hacking: A battle for the mind. Computer, 35(8):50–56, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/08/r8050.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/08/r8050.pdf;   http://
                                        www.computer.org/computer/co2002/
                                        r8050abs.htm.

Lotspiech:2002:CFB

[711] Jeffrey Lotspiech, Stefan Nusser, and Florian Pestoni. Cover feature: Broadcast encryption’s bright future. Computer, 35(8):57–63, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/08/r8057.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/08/r8057.pdf;   http://
                                        www.computer.org/computer/co2002/
                                        r8057abs.htm.

Piner:2002:CSCh

[712] Mary-Louise G. Piner. Computer Society connection: High school student from China receives IEEE Computer Society prize; distance learning campus offers new courses; John Cocke, 1925– 2002. Computer, 35(8):64–66, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/08/r8064.htm;
http://csdl.computer.org/dl/mags/
co/2002/08/r8064.pdf.

Anonymous:2002:CCh

[713] Anonymous. Call and calendar. Computer, 35(8):67–69, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/08/r8067.htm;
http://csdl.computer.org/dl/mags/
co/2002/08/r8067.pdf.

Anonymous:2002:COg

[714] Anonymous. Career opportunities. Computer, 35(8):70–72, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/08/r8070.pdf.

Anonymous:2002:PPD

[715] Anonymous. Products: Palm Digital Media’s eBook publishing tools; Veridicom’s fingerprint authentication module; high-speed debugging tool for Macraigor Systems; QNX launches IDE for embedded programming; Extreme Networks’ network management tools; Parascript’s handwriting recognition SDK; system debugger tool from Compuware. Computer, 35(8):74–??, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/08/r8074.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/08/r8074.pdf.

Lutz:2002:BLE

[716] Michael J. Lutz. Bookshelf: Learning by example. Computer, 35(8): 76, August 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2002/08/
                                        r8076.htm;   http://csdl.computer.
                                        org/dl/mags/co/2002/08/r8076.pdf.

Arbaugh:2002:SIT

[717] Bill Arbaugh. Security: Improving the TCPA specification. Computer, 35(8):77–79, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/08/r8077.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/08/r8077.pdf.

Demestichas:2002:CWB

[718] Panagiotis Demestichas, Louisa Papadopoulou, Vera Stavroulaki, Michael Theologou, Guillaume Vivier, Georges Martinez, and Fabio Galliano. Communications: Wireless beyond 3G: Managing services and network resources. Computer, 35(8):80–82, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/08/r8080.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/08/r8080.pdf.

Chang:2002:ECH

[719] Dean Chang. Entertainment computing: Haptics: Gaming’s new sensation. Computer, 35(8):84–86, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/08/r8084.htm;
http://csdl.computer.org/dl/mags/
co/2002/08/r8084.pdf.

Memon:2002:STG

[720] Atif M. Memon. Software technologies: GUI testing: Pitfalls and process. Computer, 35(8):87–91, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/08/r8087.htm;
http://csdl.computer.org/dl/mags/
co/2002/08/r8087.pdf.

Orrick:2002:PTA

[721] Diana Mayer Orrick. The profession: Toward adequate online privacy safeguards. Computer, 35(8):92, August 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/08/r8092.htm;
http://csdl.computer.org/dl/mags/
co/2002/08/r8092.pdf.

Anonymous:2002:ASi

[722] Anonymous. Article summaries. Computer, 35(9):6, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/09/r9006.htm;
http://csdl.computer.org/dl/mags/
co/2002/09/r9006.pdf.

Kolbe:2002:LHT

[723] Ray Kolbe, Mark S. Hutchenreuther, William D. Jackson, Barry M. Baker, Todd Kolb, Franco Favento, Bert Craytor, and Louis H. R. Muller. Letters: Home theater strategies; professional follies; uncovering invasive software; certification debate. Computer, 35(9):8–10, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/09/r9008.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/09/r9008.pdf.

Colwell:2002:REP

[724] Bob Colwell. At random: Employee performance reviews. Computer, 35(9):12–15, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/09/r9012.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/09/r9012.pdf.

Vaughan-Nichols:2002:ITD

[725] Steven J. Vaughan-Nichols. Industry trends: Developing the distributedcomputing OS. Computer, 35(9):19– 21, September 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2002/09/
                                        r9019.htm;   http://csdl.computer.
                                        org/dl/mags/co/2002/09/r9019.pdf;
                                         http://www.computer.org/computer/
                                        co2002/r9019abs.htm.

Paulson:2002:NBTb

[726] Linda Dailey Paulson. News briefs: Tiny “punch cards” boost storage capacity; researchers bring wireless communications to the chip; war chalking for free wireless networking; new chips will help computers visualize in three dimensions. Computer, 35(9):22, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/09/r9022.htm;
http://csdl.computer.org/dl/mags/
co/2002/09/r9022.pdf.

Bimber:2002:PMF

[727] Oliver Bimber, Stephen M. Gatesy, Lawrence M. Witmer, Ramesh Raskar, and L. Miguel Encarnação. Perspectives: Merging fossil specimens with computer-generated information. Computer, 35(9):25–30, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/09/r9025.htm;
http://csdl.computer.org/dl/mags/
co/2002/09/r9025.pdf;   http://
www.computer.org/computer/co2002/
r9025abs.htm.

Walrad:2002:CPI

[728] Chuck Walrad and Darrel Strom. Computing practices: The importance of branching models in SCM. Computer, 35(9):31–38, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/09/r9031.htm;
http://csdl.computer.org/dl/mags/
co/2002/09/r9031.pdf;   http://
www.computer.org/computer/co2002/
r9031abs.htm.

Kathail:2002:CFP

[729] Vinod Kathail, Shail Aditya, Robert Schreiber, B. Ramakrishna Rau, Darren C. Cronquist, and Mukund Sivaraman. Cover feature: PICO: Automatically designing custom computers. Computer, 35(9):39–47, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/09/r9039.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/09/r9039.pdf;   http://
                                        www.computer.org/computer/co2002/
                                        r9039abs.htm.

Wolf:2002:CFS

[730] Wayne Wolf, Burak Ozer, and Tiehan Lv. Cover feature: Smart cameras as embedded systems. Computer, 35(9):48–53, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/09/r9048.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/09/r9048.pdf;   http://
                                        www.computer.org/computer/co2002/
                                        r9048abs.htm.

Cao:2002:RFM

[731] Jiannong Cao, Xinyu Feng, Jian Lu, and Sajal K. Das. Research feature: Mailbox-based scheme for designing mobile agent communication protocols. Computer, 35(9):54–60, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/09/r9054.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/09/r9054.pdf;   http://

www.computer.org/computer/co2002/
r9054abs.htm.

Anonymous:2002:ICS

[732] Anonymous. IEEE Computer Society election. Computer, 35(9):62– 68, September 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/09/
r9062.htm;   http://csdl.computer.
org/dl/mags/co/2002/09/r9062.pdf.

Diamond:2002:CSC

[733] S. L. Diamond. Computer Society connection: IEEE President-Elect candidates address Computer Society concerns; draft computer engineering curriculum volume available for review. Computer, 35(9):70–76, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/09/r9070.htm;
http://csdl.computer.org/dl/mags/
co/2002/09/r9070.pdf.

Anonymous:2002:CCi

[734] Anonymous. Call and calendar. Computer, 35(9):77–80, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/09/r9077.htm;
http://csdl.computer.org/dl/mags/
co/2002/09/r9077.pdf.

Lutz:2002:BRB

[735] Michael J. Lutz. Bookshelf: Riskbased approach to mission-critical software testing; the TLA+ specification language primer [Specifying Systems: The TLA+ Language and Tools for Hardware and Software Engineers]; interpolation tutorial [Nonuniform Sampling: Theory and Practice]; pattern-matching toolkit [Flexible Pattern Matching in Strings: Practical On-Line Search Algorithms for Texts and Biological Sequences]. Computer, 35(9):81, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/09/r9081.htm;
http://csdl.computer.org/dl/mags/
co/2002/09/r9081.pdf.

Anonymous:2002:PEB

[736] Anonymous. Products: Enhanced Bluetooth test tool from Tektronix; NEXIQ Technologies’s intelligent display software; Actel’s FGPA development IDE; Parasoft’s automated Java classes testing unit; Packeteer upgrades central reporting application; Visi-Comp releases Java debugger; Compuware’s driver development suite for Windows; Silas Technologies upgrades application-monitoring software. Computer, 35(9):82–83, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/09/r9082.htm;
http://csdl.computer.org/dl/mags/
co/2002/09/r9082.pdf.

Anonymous:2002:COh

[737] Anonymous. Career opportunities. Computer, 35(9):84–91, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/computer/co2002/r9084.pdf.

Abdelsalam:2002:WTE

[738] Abdelsalam “Solom” Heddaya. Web technologies: An economically scalable Internet. Computer, 35(9):93– 95, September 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/09/
r9093.htm;   http://csdl.computer.
org/dl/mags/co/2002/09/r9093.pdf.

Wolf:2002:ECW

[739] Wayne Wolf. Embedded computing: Whither Warhol’s Law? Computer, 35(9):96–97, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/09/r9096.htm;
http://csdl.computer.org/dl/mags/
co/2002/09/r9096.pdf.

Shirky:2002:ISP

[740] Clay Shirky. IT systems perspective: Web services and context horizons. Computer, 35(9):98–100, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/09/r9098.htm;
http://csdl.computer.org/dl/mags/
co/2002/09/r9098.pdf.

Holmes:2002:JTS

[741] Neville Holmes. Jobs, trades, skills, and the profession. Computer, 35 (9):104, 102–103, September 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/09/r9104.htm;
http://csdl.computer.org/dl/mags/
co/2002/09/r9104.pdf.

Anonymous:2002:ASj

[742] Anonymous. Article summaries. Computer, 35(10):6, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/10/rx006.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/10/rx006.pdf.

Schneck:2002:LCP

[743] Paul Schneck, Jeffrey Lotspiech, Christopher Bird, Jeffrey R. Carter, Eric A. Weiss, Neville Holmes, and Mario L. Epp. Letters: Content protection alternatives; software development viewpoints; Java questions; seven great blunders; technological musings. Computer, 35(10):8–??, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/10/rx008.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/10/rx008.pdf.

Colwell:2002:RPM

[744] Bob Colwell. At random: The power of the most likely. Computer, 35(10):12–14, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/10/rx012.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/10/rx012.pdf.

Clark:2002:ITD

[745] David Clark. Industry trends: Do Web standards and patents mix? Computer, 35(10):19–22, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/10/rx019.htm;
http://csdl.computer.org/dl/mags/
co/2002/10/rx019.pdf.

Leavitt:2002:TNL

[746] Neal Leavitt. Technology news: Let’s hear it for audio mining. Computer, 35(10):23–25, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/10/rx023.htm;
http://csdl.computer.org/dl/mags/
co/2002/10/rx023.pdf.

Paulson:2002:NBPb

[747] Linda Dailey Paulson. News briefs: Project promises accessible technology for the disabled; two efforts aim to upgrade mobile memory; new haptics approach lets parents-to-be “touch” their unborn children; prime breakthrough may improve encryption. Computer, 35(10):26, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/10/rx026.htm;
http://csdl.computer.org/dl/mags/
co/2002/10/rx026.pdf;   http://
www.cse.iitk.ac.in/news/primality.
html.

Kalakota:2002:PED

[748] Ravi Kalakota, Pavan Gundepudi, Jonathan Wareham, Arun Rai, and Richard Welke. Perspectives: The economics of DSL regulation. Computer, 35(10):29–36, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/10/rx029.htm;

                                        http://csdl.computer.org/dl/mags/
                                        co/2002/10/rx029.pdf;   http://
                                        www.computer.org/computer/co2002/
                                        rx029abs.htm.

Schilit:2002:CFW

[749] Bill N. Schilit, Jonathan Trevor, David M. Hilbert, and Tzu Khiau Koh. Cover feature: Web interaction using very small Internet devices. Computer, 35(10):37–45, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/10/rx037.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/10/rx037.pdf;   http://
                                        www.computer.org/computer/co2002/
                                        rx037abs.htm.

Banerjee:2002:CFR

[750] Suman Banerjee, Sulabh Agarwal, Kevin Kamel, Andrzej Kochut, Christopher Kommareddy, Tamer Nadeem, Pankaj Thakkar, Bao Trinh, Adel Youssef, Moustafa Youssef, Ronald L. Larsen, A. Udaya Shankar, and Ashok Agrawala. Cover feature: Rover: Scalable location-aware computing. Computer, 35(10):46–53, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/10/rx046.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/10/rx046.pdf;   http://
                                        www.computer.org/computer/co2002/
                                        rx046abs.htm.

Wood:2002:CFD

[751] Anthony D. Wood and John A. Stankovic. Cover feature: Denial of service in sensor networks. Computer, 35(10):54–62, October 2002.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/10/rx054.htm;
http://csdl.computer.org/dl/mags/
co/2002/10/rx054.pdf;   http://
www.computer.org/computer/co2002/
rx054abs.htm.

Samaranayake:2002:RFD

[752] Samitha Samaranayake, Nodari Sitchinava, Rohit Kapur, Minesh B. Amin, and Thomas W. Williams. Research feature: Dynamic scan: Driving down the cost of test. Computer, 35(10):63–68, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/10/rx063.htm;
http://csdl.computer.org/dl/mags/
co/2002/10/rx063.pdf;   http://
www.computer.org/computer/co2002/
rx063abs.htm.

Anonymous:2002:PRSb

[753] Anonymous. Products: Rational Software’s .NET development software; SGI’s development APIs for visualization; Java development tool from Visual Numerics; SUSSMicroTec’s MEMS testing vacuum prober; Face-Time’s IM information security system; MX development resource kit from Macromedia; .NET data integration tool from XAWare; Voice-Genie’s VoiceXML gateway. Computer, 35(10):70–??, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/10/rx070.htm;
http://csdl.computer.org/dl/mags/
co/2002/10/rx070.pdf.

Lutz:2002:BXC

[754] Michael J. Lutz. Bookshelf: The XML culture clash [System Architecture with XML]; broadband’s core technologies [Integrated Broadband Networks: TCP/IP, ATM, SDH/ SONET, and WDM/Optics]; efficient IT management [Building Operational Excellence]; making CORBA secure [Developing Secure Distributed Systems with CORBA]. Computer, 35(10):72, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/10/rx072.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/10/rx072.pdf.

Anonymous:2002:CSC

[755] Anonymous. Computer Society connection: Partnerships improve society publications; lecture programs seeks presenters; standards news; Ole-Johan Dahl, 1931–2002, and Kristen Nygaard, 1926–2002. Computer, 35(10):74–77, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/10/rx074.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/10/rx074.pdf.

Anonymous:2002:CCj

[756] Anonymous. Call and calendar. Computer, 35(10):78–79, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/10/rx078.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/10/rx078.pdf.

Anonymous:2002:COi

[757] Anonymous. Career opportunities. Computer, 35(10):82–95, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/computer/co2002/rx082.pdf.

Macedonia:2002:ECC

[758] Michael Macedonia. Entertainment computing: The computer graphics wars heat up. Computer, 35(10):97–99, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/10/rx097.htm;
http://csdl.computer.org/dl/mags/
co/2002/10/rx097.pdf.

Huber:2002:CTM

[759] Josef F. Huber. Communications: Toward the mobile Internet. Computer, 35(10):100–102, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/10/rx100.htm;
http://csdl.computer.org/dl/mags/
co/2002/10/rx100.pdf.

Huang:2002:STP

[760] Andrew ‘bunnie” Huang. Security: The trusted PC: Skin-deep security. Computer, 35(10):103–105, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/10/rx103.htm;
http://csdl.computer.org/dl/mags/
co/2002/10/rx103.pdf.

Briggs:2002:PRD

[761] Laura Briggs and Kari Boyd McBride. The profession: Rethinking the digital divide. Computer, 35 (10):108, 106–107, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/10/rx108.htm;
http://csdl.computer.org/dl/mags/
co/2002/10/rx108.pdf.

Anonymous:2002:ASk

[762] Anonymous. Article summaries. Computer, 35(11):6–7, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry006.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry006.pdf.

Johnson:2002:LCH

[763] Jeff Johnson, Don Dwiggins, Michael Kaebling, and Neville Holmes. Letters: Cognitive hacking of a different sort; computing blunders. Computer, 35(11):8–9, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry008.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry008.pdf.

Colwell:2002:RCT

[764] Bob Colwell. At random: The coolest thing on Earth. Computer, 35(11):10–12, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry010.htm;

http://csdl.computer.org/dl/mags/
co/2002/11/ry010.pdf.

Vaughan-Nichols:2002:ITB

[765] Steven J. Vaughan-Nichols. Industry trends: Bull market for IEEE 802.11 WLAN chipsets. Computer, 35(11):17–19, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry017.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry017.pdf.

Paulson:2002:TNS

[766] Linda D. Paulson. Technology news: Stopping intruders outside the gates. Computer, 35(11):20–22, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry020.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry020.pdf.

Paulson:2002:NBIb

[767] Linda Dailey Paulson. News briefs: IBM works on a universal virtual computer; battling intrusions with a “network telescope”; choreographing Web services. Computer, 35(11):23–25, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry023.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry023.pdf;   http://
www.computer.org/computer/co2002/
ry023abs.htm.

Zhong:2002:GEI

[768] Ning Zhong, Jiming Liu, and Yiyu Yao. Guest Editors’ introduction: In search of the Wisdom Web. Computer, 35(11):27–31, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry027.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry027.pdf;   http://
www.computer.org/computer/co2002/
ry027abs.htm.

Kumar:2002:CFW

[769] Ravi Kumar, Prabhakar Raghavan, Sridhar Rajagopalan, and Andrew Tomkins. Cover feature: The Web and social networks. Computer, 35(11):32–36, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry032.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry032.pdf;   http://
www.computer.org/computer/co2002/
ry032abs.htm.

Nishida:2002:CFS

[770] Toyoaki Nishida. Cover feature: Social intelligence design for the Web. Computer, 35(11):37–41, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry037.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry037.pdf;   http://
www.computer.org/computer/co2002/
ry037abs.htm.

Ishida:2002:CFQ

[771] Toru Ishida. Cover feature: Q: A scenario description language for interactive agents. Computer, 35(11): 42–47, November 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/11/
ry042.htm;   http://csdl.computer.
org/dl/mags/co/2002/11/ry042.pdf;
http://www.computer.org/computer/
co2002/ry042abs.htm.

Bryson:2002:CFT

[772] Joanna J. Bryson, David L. Martin, Sheila A. McIlraith, and Lynn A. Stein. Cover feature: Toward behavioral intelligence in the Semantic Web. Computer, 35(11):48–54, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry048.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry048.pdf;   http://
www.computer.org/computer/co2002/
ry048abs.htm.

Fensel:2002:CFO

[773] Dieter Fensel. Cover feature: Ontologybased knowledge management. Computer, 35(11):56–59, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry056.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry056.pdf;   http://
www.computer.org/computer/co2002/
ry056abs.htm.

Missikof:2002:CFI

[774] Michele Missikof, Roberto Navigli, and Paola Velardi. Cover feature: Integrated approach to Web ontology learning and engineering. Computer, 35(11):60–63, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry060.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry060.pdf;   http://
www.computer.org/computer/co2002/
ry060abs.htm.

Han:2002:CFD

[775] Jiawei Han and Kevin C.-C. Chang. Cover feature: Data mining for Web intelligence. Computer, 35(11):64– 70, November 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/11/
ry064.htm;   http://csdl.computer.
org/dl/mags/co/2002/11/ry064.pdf;
 http://www.computer.org/computer/
co2002/ry064abs.htm.

Cercone:2002:CFC

[776] Nick Cercone, Lijun Hou, Vlado Keselj, Aijun An, Kanlaya Naruedomkul, and Xiaohua Hu. Cover feature: From computational intelligence to Web intelligence. Computer, 35(11):72–76, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry072.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry072.pdf;   http://
www.computer.org/computer/co2002/
ry072abs.htm.

Lutz:2002:BAN

[777] Michael J. Lutz. Bookshelf: Architecture by the numbers [Computer Architecture: A Quantitative Approach]; half a Java tone [Algorithms in Java]; embedded systems simulation [Embedded Systems and Computer Architecture]; brandname multicasting [Interdomain Multicast Routing: Practical Juniper Networks and Cisco Systems Solutions]. Computer, 35(11):77, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry077.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry077.pdf.

Anonymous:2002:POU

[778] Anonymous. Products: Omnicore upgrades Java IDE CodeGuide emWare’s SDE for intelligent device management; Metrowerks’ CodeWarrior for Embedded Linux; integrated software environment form Xilinx; new version of InstallShield Professional; Motorola’s 32-bit CAN reference design; Utopia-LVDS bridge reference design kit from National Semiconductor; First Silicon Solutions’ analysis tool for flash-based FPGAs. Computer, 35(11):78–79, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry078.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry078.pdf.

Piner:2002:CSCi

[779] Mary-Louise G. Piner. Computer Society connection: Partnership with Wiley will showcase expert texts; online collection offers information technology resource for libraries; recent awards reinforce Society’s commitment to service, teaching. Computer, 35(11):80–83, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/11/ry080.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/11/ry080.pdf.

Anonymous:2002:CCk

[780] Anonymous. Call and calendar. Computer, 35(11):84–86, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/11/ry084.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/11/ry084.pdf.

Anonymous:2002:COj

[781] Anonymous. Career opportunities. Computer, 35(11):88–105, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/11/ry088.pdf.

Norton:2002:ISP

[782] Roger Norton. IT systems perspective: Using virtual Linux servers. Computer, 35(11):106–107, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/11/ry106.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/11/ry106.pdf.

Madsen:2002:ECS

[783] Jan Madsen, Jørgen Steensgaard-Madsen, and Lars Christensen. Embedded computing: A sophomore course in codesign. Computer, 35(11):108– 110, November 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/11/
ry108.htm;   http://csdl.computer.
org/dl/mags/co/2002/11/ry108.pdf.

Ransom:2002:WTU

[784] Stephen Ransom and Xindong Wu. Web technologies: Using knowledge anchors to reduce cognitive overheads. Computer, 35(11):111–112, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry111.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry111.pdf.

Ricketts:2002:SIR

[785] Anita C. Ricketts. Standards: IEEE registration authority. Computer, 35(11):113, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry113.htm;
http://csdl.computer.org/dl/mags/
co/2002/11/ry113.pdf.

Holmes:2002:PPW

[786] Neville Holmes. The profession: The profession and the world. Computer, 35(11):116, 114–115, November 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/11/ry116.htm;

                                        http://csdl.computer.org/dl/mags/
                                        co/2002/11/ry116.pdf.

Anonymous:2002:ASl

[787] Anonymous. Article summaries. Computer, 35(12):4–5, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/12/rz004.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/12/rz004.pdf.

Swarr:2002:LOW

[788] Robert Swarr, Kevin J. Crandall, Andrew “bunnie” Huang, John Michael Williams, Peter Grogono, Bob Colwell, Michael Gile, Neville Holmes, Rick Schrenker, Neville Holmes, Norton Hines, and Neville Holmes. Letters: Open Web standards stimulate innovation; redefining acronyms; debug strategies; jobs, trades, skills, and the profession. Computer, 35(12): 6–9, December 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2002/12/
                                        rz006.htm;   http://csdl.computer.
                                        org/dl/mags/co/2002/12/rz006.pdf.

King:2002:PMY

[789] Willis K. King. President’s message: Year in review. Computer, 35(12):10–11, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2002/12/rz010.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2002/12/rz010.pdf.

Aylor:2002:EMI

[790] James H. Aylor. EIC’s message: It has been a great ride! Computer, 35(12):12, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz012.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz012.pdf.

Colwell:2002:RSP

[791] Bob Colwell. At random: Sciencephilosophy celebrity death match. Computer, 35(12):13–15, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz013.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz013.pdf.

Vaughan-Nichols:2002:ITO

[792] Steven J. Vaughan-Nichols. Industry trends: OFDM: Back to the wireless future. Computer, 35(12):19– 21, December 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/12/
rz019.htm;   http://csdl.computer.
org/dl/mags/co/2002/12/rz019.pdf.

Lawton:2002:VWF

[793] George Lawton. Virus wars: Fewer attacks, new threats. Computer, 35(12): 22–24, December 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/12/
rz022.htm;   http://csdl.computer.
org/dl/mags/co/2002/12/rz022.pdf.

Paulson:2002:NBR

[794] Linda Dailey Paulson. News briefs: Researchers upgrade Smart Card technology; new optical clock could synchronize chips; blind, deaf engineer develops computerized Braille machine. Computer, 35(12):25–27, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz025.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz025.pdf.

Rowen:2002:CPR

[795] Chris Rowen. Computing practices: Reducing SoC simulation and development time. Computer, 35(12):29–34, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz029.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz029.pdf;   http://
www.computer.org/computer/co2002/
rz029abs.htm.

Robinson:2002:CFT

[796] Gary S. Robinson. Cover feature: Toward the age of smarter storage. Computer, 35(12):35–41, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz035.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz035.pdf;   http://
www.computer.org/computer/co2002/
rz035abs.htm.

Scheible:2002:CFS

[797] John P. Scheible. Cover feature:

A survey of storage options. Computer, 35(12):42–46, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz042.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz042.pdf;   http://
www.computer.org/computer/co2002/
rz042abs.htm.

LoBue:2002:CFS

[798] Michael T. LoBue, Harry Mason, Thomas Hammond-Doel, Eric Anderson, Mike Alexenko, and Tom Clark. Cover feature: Surveying today’s most popular storage interfaces. Computer, 35(12):48–55, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz048.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz048.pdf;   http://
www.computer.org/computer/co2002/
rz048abs.htm.

Griswold:2002:CFS

[799] Robert Griswold. Cover feature: Storage topologies. Computer, 35(12):56– 63, December 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/12/
rz056.htm;   http://csdl.computer.
org/dl/mags/co/2002/12/rz056.pdf;
http://www.computer.org/computer/
co2002/rz056abs.htm.

Cummings:2002:CFE

[800] Roger Cummings. Cover feature: The evolution of information assurance. Computer, 35(12):65–72, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz065.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz065.pdf;   http://
www.computer.org/computer/co2002/
rz065abs.htm.

Piner:2002:RMC

[801] Mary-Louise G. Piner. Report to members: Chang chosen IEEE Computer Society President-Elect; Arthur W. Winston chosen IEEE President-Elect for 2003. Computer, 35(12):73–74, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz073.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz073.pdf.

Aylor:2002:CSC

[802] James H. Aylor. Computer Society connection: Reviewers contribute to Computer’s success. Computer, 35(12):75–76, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz075.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz075.pdf.

Anonymous:2002:CCl

[803] Anonymous. Call and calendar. Computer, 35(12):77–79, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz077.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz077.pdf.

Anonymous:2002:AIC

[804] Anonymous. 2002 annual index: Computer volume 35, 2002. Computer, 35 (12):80–91, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2002/12/rz080.htm;   http://www.
computer.org/computer/co2002/rz080.
pdf.

Anonymous:2002:COk

[805] Anonymous. Career opportunities. Computer, 35(12):92–116, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz092.pdf.

Lutz:2002:BGB

[806] Michael J. Lutz. Bookshelf: Getting the bugs out [Debugging: The 9 Indispensable Rules for Finding Even the Most Elusive Software and Hardware Problems]; saving time with arithmetic and logic [Hacker’s Delight]; probability and computing [Probability and Statistics with Reliability, Queuing and Computer Science Applications, 2nd edition]; graphics toolbox [Guide to Graphics Software Tools]. Computer, 35(12):117, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz117.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz117.pdf.

Anonymous:2002:PSS

[807] Anonymous. Products: SOISIC ships design kit for SOI structures; systems and software development tools from Telelogic; RSA Security’s Web access management system; Altera’s free embedded processor portfolio; signal integrity measurement tools from tektronix; Oracle upgrades Java development tool; Xilinx delivers EDK for FPGA processor; Westbridge’s tool to sniff unauthorized XML; Speech-Studio’s telephony development tools. Computer, 35(12):118–119, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz118.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz118.pdf.

Varshney:2002:CMP

[808] Upkar Varshney. Communications: Mobile payments. Computer, 35(12):120–121, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz120.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz120.pdf.

Stapleton:2002:ECA

[809] Christopher Stapleton, Charles Hughes, Michael Moshell, Paulius Micikevicius, and Marty Altman. Entertainment computing: Applying mixed reality to entertainment. Computer, 35(12):122–124, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz122.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz122.pdf.

Saydjari:2002:SDC

[810] O. Sami Saydjari. Security: Defending cyberspace. Computer, 35(12): 125, December 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2002/12/
rz125.htm;   http://csdl.computer.
org/dl/mags/co/2002/12/rz125.pdf.

Santini:2002:PUL

[811] Simone Santini. The profession: Using language more responsibly. Computer, 35(12):128, 126–127, December 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2002/12/rz128.htm;
http://csdl.computer.org/dl/mags/
co/2002/12/rz128.pdf.

Anonymous:2003:ASa

[812] Anonymous. Article summaries. Computer, 36(1):4, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/01/r1004.htm;   http://www.
computer.org/computer/co2003/r1004.
pdf.

Carver:2003:EMB

[813] Doris L. Carver. EIC’s message: Building on the past, looking to the future. Computer, 36(1):6–7, 10, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1006.htm;
http://csdl.computer.org/dl/mags/
co/2003/01/r1006.pdf.

Diamond:2003:PMB

[814] Stephen L. Diamond. President’s message: Building a better Computer Society. Computer, 36(1):8–10, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1008.htm;
http://csdl.computer.org/dl/mags/
co/2003/01/r1008.pdf.

Appleton:2003:LSB

[815] Brad Appleton, Chuck Walrd, Darrel Strom, Jeffrey R. Carter, Neville Holmes, Franco Favento, Neville Holmes, Julian Weiss, and Neville Holmes. Letters: SCM branching strategies; software development division of labor; balancing conflicting goals. Computer, 36(1):11– 13, January 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/01/r1011.htm;   http://www.
computer.org/computer/co2003/r1011.
pdf.

Colwell:2003:RMD

[816] Bob Colwell. At random: A man, a dog, and a machine. Computer, 36(1):14–16, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1014.htm;
http://csdl.computer.org/dl/mags/
co/2003/01/r1014.pdf.

Paulson:2003:ITL

[817] Linda Dailey Paulson. Industry trends: Low-power chips for high-powered handhelds. Computer, 36(1):21–23, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1021.htm;
http://csdl.computer.org/dl/mags/
co/2003/01/r1021.pdf.

Leavitt:2003:TNW

[818] Neal Leavitt. Technology news: Will wireless gaming be a winner? Computer, 36(1):24–27, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1024.htm;
http://csdl.computer.org/dl/mags/
co/2003/01/r1024.pdf.

Paulson:2003:NBV

[819] Linda Dailey Paulson. News briefs: Vendors push wireless LAN security; picturing a new encryption technique; think tank targets better software. Computer, 36(1):28–30, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1028.htm;
http://csdl.computer.org/dl/mags/
co/2003/01/r1028.pdf.

Hamilton:2003:CFI

[820] Scott Hamilton. Cover feature: Intel research extends Moore’s Law. Computer, 36(1):31–40, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1031.htm;
http://csdl.computer.org/dl/mags/
co/2003/01/r1031.pdf;   http://

                                        www.computer.org/computer/co2003/
                                        r1031abs.htm.

Kephart:2003:CFV

[821] Jeffrey O. Kephart and David M. Chess. Cover feature: The vision of autonomic computing. Computer, 36(1):41–50, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/01/r1041.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/01/r1041.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r1041abs.htm.

Katz:2003:CFN

[822] Daniel S. Katz and Raphael R. Some. Cover feature: NASA advances robotic space exploration. Computer, 36(1):52–61, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/01/r1052.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/01/r1052.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r1052abs.htm.

Reed:2003:CFG

[823] Daniel A. Reed. Cover feature: Grids, the TeraGrid, and beyond. Computer, 36(1):62–68, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/01/r1062.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/01/r1062.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r1062abs.htm.

Diorio:2003:CFC

[824] Chris Diorio and Jaideep Mavoori. Cover feature: Computer electronics meet animal brains. Computer, 36(1):69–75, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1069.htm;
http://csdl.computer.org/dl/mags/
co/2003/01/r1069.pdf;   http://
www.computer.org/computer/co2003/
r1069abs.htm.

Janowiak:2003:CFC

[825] Robert M. Janowiak. Cover feature: Computers and communications: a symbiotic relationship. Computer, 36(1):76–79, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1076.htm;
http://csdl.computer.org/dl/mags/
co/2003/01/r1076.pdf;   http://
www.computer.org/computer/co2003/
r1076abs.htm.

Piner:2003:CSCa

[826] Mary-Louise G. Piner. Computer Society connection: Look ahead to 2003 programs; publications; conferences. Computer, 36(1):81–87, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1081.htm;
http://csdl.computer.org/dl/mags/
co/2003/01/r1081.pdf.

Anonymous:2003:CCa

[827] Anonymous. Call and calendar. Computer, 36(1):89–91, January 2003.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1089.htm;
http://csdl.computer.org/dl/mags/
co/2003/01/r1089.pdf.

Anonymous:2003:COa

[828] Anonymous. Career opportunities. Computer, 36(1):92–111, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/computer/co2003/r1092.pdf.

Anonymous:2003:PSR

[829] Anonymous. Products: Starbase releases decision-support software; OC Systems extends analysis tool to J2EE; InstallShield streamlines software installation app; Silicon Defense counters stealth scans; compuware upgrades Java profiling tool;’ Pervasive Software releases V8 database engine; Xilinx ships DSP design tool; MKS adds wizards to monitoring system. Computer, 36(1):112–113, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1112.htm;
http://csdl.computer.org/dl/mags/
co/2003/01/r1112.pdf.

Lutz:2003:BFP

[830] Michael J. Lutz. Bookshelf: False positives and true negatives [Facts and Fallacies of Software Engineering]. Computer, 36(1):114, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1114.htm;

http://csdl.computer.org/dl/mags/
co/2003/01/r1114.pdf.

Tirri:2003:WTS

[831] Henry Tirri. Web technologies: Search in vain: Challenges for Internet search. Computer, 36(1):115–116, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1115.htm;
http://csdl.computer.org/dl/mags/
co/2003/01/r1115.pdf.

Loiacono:2003:ISP

[832] Eleanor T. Loiacono. IT systems perspective: Improving Web accessibility. Computer, 36(1):117–119, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1117.htm;
http://csdl.computer.org/dl/mags/
co/2003/01/r1117.pdf.

Dutt:2003:ECC

[833] Nikil Dutt and Kiyoung Choi. Embedded computing: Configurable processors for embedded computing. Computer, 36(1):120–123, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/01/r1120.htm;
http://csdl.computer.org/dl/mags/
co/2003/01/r1120.pdf.

Hughes:2003:SSS

[834] Jim Hughes and Jack Cole. Standards: Security in storage. Computer, 36(1):124–125, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/01/r1124.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/01/r1124.pdf.

Holmes:2003:PRP

[835] Neville Holmes. The profession: Revising the principles of technorealism. Computer, 36(1):128, 126– 127, January 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2003/01/
                                        r1128.htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/01/r1128.pdf.

Anonymous:2003:ASb

[836] Anonymous. Article summaries. Computer, 36(2):4, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //csdl.computer.org/dl/mags/co/
                                        2003/02/r2004.htm;   http://www.
                                        computer.org/computer/co2003/r2004.
                                        pdf.

Doucette:2003:LRV

[837] Doug Doucette, Howard Marsh, Daniel Leong, Roger Coombs, Rick Siple, and Neville Holmes. Letters: Risk versus threat; science and philosophy; inflation, greed, and chattels. Computer, 36(2):5–7, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //csdl.computer.org/dl/mags/co/
                                        2003/02/r2005.htm;   http://www.
                                        computer.org/computer/co2003/r2005.
                                        pdf.

Colwell:2003:RMM

[838] Bob Colwell. At random: Me and my theremin. Computer, 36(2):8– 9, February 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/02/
r2008.htm;   http://csdl.computer.
org/dl/mags/co/2003/02/r2008.pdf.

Ortiz:2003:ITN

[839] Sixto Ortiz Jr. Industry trends: New monitor technologies are on display. Computer, 36(2):13–16, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/02/r2013.htm;
http://csdl.computer.org/dl/mags/
co/2003/02/r2013.pdf.

Vaughan-Nichols:2003:TNM

[840] Steven J. Vaughan-Nichols. Technology news: Mobile IPv6 and the future of wireless Internet access. Computer, 36(2):18–20, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/02/r2018.htm;
http://csdl.computer.org/dl/mags/
co/2003/02/r2018.pdf.

Paulson:2003:NBD

[841] Linda Dailey Paulson. News briefs: Databases that heal themselves; chip makers turn to multigate transistors; going mobile with Web services. Computer, 36(2):22–24, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/02/r2022.htm;
http://csdl.computer.org/dl/mags/
co/2003/02/r2022.pdf.

Lin:2003:PDT

[842] Daniel Lin. Perspectives: Dual tragedies: IP rights in industry standards. Computer, 36(2):25–27, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2003/02/
                                        r2025.htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/02/r2025.pdf;
                                         http://www.computer.org/computer/
                                        co2003/r2025abs.htm.

Keeton:2003:GEI

[843] Kimberly Keeton, Russell Clapp, and Ashwini Nanda. Guest Editors’ introduction: Evaluating servers with commercial workloads. Computer, 36(2):29–32, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/02/r2029.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/02/r2029.pdf.

Luo:2003:CFB

[844] Yue Luo, Juan Rubio, Lizy Kurian John, Pattabi Seshadri, and Alex Mericas. Cover feature: Benchmarking Internet servers on superscalar machines. Computer, 36(2):34–40, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/02/r2034.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/02/r2034.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r2034abs.htm.

Garcia:2003:CFT

[845] Daniel F. García and Javier García.

Cover feature: TPC-W E-commerce benchmark evaluation. Computer, 36 (2):42–48, February 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/02/
r2042.htm;   http://csdl.computer.
org/dl/mags/co/2003/02/r2042.pdf;
 http://www.computer.org/computer/
co2003/r2042abs.htm.

Alameldeen:2003:CFS

[846] Alaa R. Alameldeen, Milo M. K. Martin, Carl J. Mauer, Kevin E. Moore, Min Xu, Mark D. Hill, David A. Wood, and Daniel J. Sorin. Cover feature: Simulating a US$2M commercial server on a US$2K PC. Computer, 36(2):50–57, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/02/r2050.htm;
http://csdl.computer.org/dl/mags/
co/2003/02/r2050.pdf;   http://
www.computer.org/computer/co2003/
r2050abs.htm.

Tsuei:2003:CFQ

[847] Thin-Fong Tsuei and Wayne Yamamoto. Cover feature: Queuing simulation model for multiprocessor systems. Computer, 36(2):58–64, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/02/r2058.htm;
http://csdl.computer.org/dl/mags/
co/2003/02/r2058.pdf;   http://
www.computer.org/computer/co2003/
r2058abs.htm.

Eeckhout:2003:CFD

[848] Lieven Eeckhout, Hans Vandierendonck, and Koen De Bosschere. Cover feature: Designing computer architecture research workloads. Computer, 36(2):65–71, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/02/r2065.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/02/r2065.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r2065abs.htm.

Anonymous:2003:CSCa

[849] Anonymous. Computer Society connection: IEEE names Fellows for 2003; Fellow nominations due 15 March; IEEE Foundation donates over US$425,000 to projects. Computer, 36(2):72–76, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/02/r2072.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/02/r2072.pdf.

Anonymous:2003:CCb

[850] Anonymous. Call and calendar. Computer, 36(2):77–79, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/02/r2077.htm;
                                        http://www.computer.org/computer/
                                        co2003/r2077.pdf.

Anonymous:2003:PSA

[851] Anonymous. Products: SGI announces record-performing Linux servers; Pacific Nanotechnogy releases atomic force microscope; Nvidia extends Cg programming environment; Empirix introduces VoIP analyzer; Codagen Technologies upgrades UML tool; Perforce Software enhances SCM system; Compuware upgrades Java development tool. Computer, 36(2):80–81, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/02/r2080.htm;
http://csdl.computer.org/dl/mags/
co/2003/02/r2080.pdf.

Lutz:2003:BBO

[852] Michael J. Lutz. Bookshelf: Bringing one champion up and another down [Behind Deep Blue: Building the Computer that Defeated the World Chess Champion]; putting tests in the driver’s seat [Test-Driven Development: By Example]; designing mixed-technology systems [The System Designer’s Guide to VHDL-AMS: Analog, Mixed-Signal, and Mixed-Technology Modeling]; practical interface primer [Display Interfaces: Fundamentals and Standards]. Computer, 36(2):82–83, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/02/r2082.htm;
http://csdl.computer.org/dl/mags/
co/2003/02/r2082.pdf.

Anonymous:2003:COb

[853] Anonymous. Career opportunities. Computer, 36(2):84–93, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/computer/co2003/r2084.pdf.

Macedonia:2003:ECR

[854] Michael Macedonia. Entertainment computing: Revitalizing museums with digital technology. Computer, 36(2):94–96, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/02/r2094.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/02/r2094.pdf.

Schilit:2003:ICM

[855] Bill N. Schilit. Invisible computing: Mega-utilities drive invisible technologies. Computer, 36(2):97–99, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/02/r2097.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/02/r2097.pdf.

Tung:2003:CMO

[856] Y. Alex Tung, Ram D. Gopal, and Andrew B. Whinston. Communications: Multiple online auctions. Computer, 36(2):100–102, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/02/r2100.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/02/r2100.pdf.

Shapiro:2003:SUW

[857] Jonathan S. Shapiro. Security: Understanding the Windows EAL4 evaluation. Computer, 36(2):103–105, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/02/
r2103.htm;   http://csdl.computer.
org/dl/mags/co/2003/02/r2103.pdf.

Holmes:2003:PTC

[858] Neville Holmes. The profession: Truth and clarity in arithmetic. Computer, 36(2):108, 106–107, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/02/r2108.htm;
http://csdl.computer.org/dl/mags/
co/2003/02/r2108.pdf.

Anonymous:2003:ASc

[859] Anonymous. Article summaries. Computer, 36(3):4, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/03/r3004.htm;   http://www.
computer.org/computer/co2003/r3004.
pdf.

Isaak:2003:LWA

[860] Jim Isaak, James N. Porter, Gary S. Robinson, Andreas Siebert, Simone Santini, George Brooke, Simone Santini, Fabrizio Luccio, and George Brooke. Letters: Web accessibility guidelines; storage data point updates; fighting decadent jargon. Computer, 36(3):5–10, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/03/r3005.htm;   http://www.
computer.org/computer/co2003/r3005.
pdf.

Colwell:2003:RGB

[861] Bob Colwell. At random: Ground bounce. Computer, 36(3):11–13, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/03/r3011.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/03/r3011.pdf.

Lawton:2003:ITI

[862] Geroge Lawton. Industry trends: Instant Messaging puts on a business suit. Computer, 36(3):14–16, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/03/r3014.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/03/r3014.pdf.

Vaughan-Nichols:2003:TNHa

[863] Steven J. Vaughan-Nichols. Technology news: How trustworthy is trusted computing? Computer, 36(3):18–20, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/03/r3018.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/03/r3018.pdf.

Paulson:2003:NBI

[864] Linda Dailey Paulson. News briefs: IBM seeks to unify data-access approaches; researcher puts the network to music; DARPA creating self-aware computing; students help put a computer in every classroom. Computer, 36(3):22, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/03/r3022.htm;
http://csdl.computer.org/dl/mags/
co/2003/03/r3022.pdf.

Saha:2003:PPC

[865] Debashis Saha and Amitava Mukherjee. Perspectives: Pervasive computing: A paradigm for the 21st Century. Computer, 36(3):25–31, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/03/r3025.htm;
http://csdl.computer.org/dl/mags/
co/2003/03/r3025.pdf;   http://
www.computer.org/computer/co2003/
r3025abs.htm.

Boehm:2003:CFV

[866] Barry Boehm and Li Guo Huang. Cover feature: Value-based software engineering: A case study. Computer, 36(3):33–41, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/03/r3033.htm;
http://csdl.computer.org/dl/mags/
co/2003/03/r3033.pdf;   http://
www.computer.org/computer/co2003/
r3033abs.htm.

Davis:2003:CFA

[867] Alan M. Davis. Cover feature: The art of requirements triage. Computer, 36(3):42–49, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/03/r3042.htm;
http://csdl.computer.org/dl/mags/
co/2003/03/r3042.pdf;   http://

                                        www.computer.org/computer/co2003/
                                        r3042abs.htm.

Gorton:2003:CFR

[868] Ian Gorton, Anna Liu, and Paul Brebner. Cover feature: Rigorous evaluation of COTS middleware technology. Computer, 36(3):50–55, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/03/r3050.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/03/r3050.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r3050abs.htm.

Lopez:2003:RFA

[869] Pedro García López and Antonio F. Gómez Skarmeta. Research feature: ANTS framework for cooperative work environments. Computer, 36(3):56–62, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/03/r3056.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/03/r3056.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r3056abs.htm.

Chen:2003:RFP

[870] Xin Chen and Xiaodong Zhang. Research feature: A popularity-based prediction model for Web prefetching. Computer, 36(3):63–70, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/03/r3063.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/03/r3063.pdf;   http://

www.computer.org/computer/co2003/
r3063abs.htm.

Anonymous:2003:CSCb

[871] Anonymous. Computer Society connection: New distance learning catalog features wireless, Microsoft .NET courses; universities from around the world compete in CDIDC 2003; salaries rise slightly for US, Canadian computing faculty. Computer, 36(3):72–74, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/03/r3072.htm;
http://csdl.computer.org/dl/mags/
co/2003/03/r3072.pdf.

Anonymous:2003:CCc

[872] Anonymous. Call and calendar. Computer, 36(3):75–78, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/03/r3075.htm;
http://www.computer.org/computer/
co2003/r3075.pdf.

Lutz:2003:BMD

[873] Michael J. Lutz. Bookshelf: Making digital telephones networks smarter [Intelligent Networks: Principles and Applications]; developing business components [Component-Based Development: Principles and Planning for Business Systems]; A pyramid scheme for computer graphics [Pyramid Algorithms: A Dynamic Programming Approach to Curves and Surfaces for Geometric Modeling]; drafting requirements that work [Writing Better Requirements]. Computer, 36(3):79, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/03/r3079.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/03/r3079.pdf.

Anonymous:2003:PJU

[874] Anonymous. Products: JetBrains upgrades IntelliJ Java IDE; Catalyst’s USB analyzer supports device emulation; VMware releases Enterprise Server VM software; Motorola offers free soft modem reference design; RealNetworks releases source for Helix DNA Server; Packeteer accelerates intranet and Internet applications. Computer, 36(3):80–81, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/03/r3080.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/03/r3080.pdf.

Anonymous:2003:COc

[875] Anonymous. Career opportunities. Computer, 36(3):84–88, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://www.computer.
                                        org/computer/co2003/r3084.pdf.

Walczak:2003:WTX

[876] Krzysztof Walczak and Wojciech Cellary. Web technologies: X-VRML for advanced virtual reality applications. Computer, 36(3):89–92, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/03/r3089.htm;

http://csdl.computer.org/dl/mags/
co/2003/03/r3089.pdf.

Wolf:2003:ECH

[877] Wayne Wolf. Embedded computing: How many system architectures? Computer, 36(3):93–95, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/03/r3093.htm;
http://csdl.computer.org/dl/mags/
co/2003/03/r3093.pdf.

Dutta:2003:ISP

[878] Amitava Dutta. IT system perspective: Managing IT with systems thinking. Computer, 36(3):96–99, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/03/r3096.htm;
http://csdl.computer.org/dl/mags/
co/2003/03/r3096.pdf.

Milutinovic:2003:PWD

[879] Veljko Milutinovic and Nikola Skundric. The profession: Will distance learning create a global university? Computer, 36(3):100, 98–99, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/03/r3100.htm;
http://csdl.computer.org/dl/mags/
co/2003/03/r3100.pdf.

Anonymous:2003:ASd

[880] Anonymous. Article summaries. Computer, 36(4):4, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/

2003/04/r4004.htm;   http://www.
computer.org/computer/co2003/r4004.
pdf.

Vaughan:2003:LIR

[881] Tom Vaughan, Mark Whitt, Christophe Alviset, Neville Holmes, Mike Barnett, Neville Holmes, Thomas R. Leith, and Neville Holmes. Letters: IP rights: Innovation versus standardization; research model options; revised technorealism principles. Computer, 36(4):7–9, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/04/r4007.htm;   http://www.
computer.org/computer/co2003/r4007.
pdf.

Colwell:2003:RSF

[882] Bob Colwell. At random: Science fairs. Computer, 36(4):13–16, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/04/r4013.htm;
http://csdl.computer.org/dl/mags/
co/2003/04/r4013.pdf.

Paulson:2003:ITO

[883] Linda Dailey Paulson. Industry trends: Optical networking chips get faster in a slow market. Computer, 36(4):18–20, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/04/r4018.htm;
http://csdl.computer.org/dl/mags/
co/2003/04/r4018.pdf.

Vaughan-Nichols:2003:TNBa

[884] Steven J. Vaughan-Nichols. Technology news: The battle over the universal Java IDE. Computer, 36(4): 21–23, April 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/04/
r4021.htm;   http://csdl.computer.
org/dl/mags/co/2003/04/r4021.pdf.

Paulson:2003:NBRa

[885] Linda Dailey Paulson. News briefs: Researchers automate the digital animation process; new compression technology would improve online video quality; Sun tries to guide Mobile Java; new search engines keep humming along. Computer, 36(4):24–26, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/04/r4024.htm;
http://csdl.computer.org/dl/mags/
co/2003/04/r4024.pdf.

Vahid:2003:SH

[886] Frank Vahid. Perspectives: The softening of hardware. Computer, 36 (4):27–34, April 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/04/
r4027.htm;   http://csdl.computer.
org/dl/mags/co/2003/04/r4027.pdf;
http://www.computer.org/computer/
co2003/r4027abs.htm.

Henkel:2003:GEI

[887] Jörg Henkel, Xiaobo Sharon Hu, and Shuvra S. Bhattacharyya. Guest Editors’ introduction: Taking on the embedded system design challenge. Computer, 36(4):35–37, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/04/r4035.htm;
http://csdl.computer.org/dl/mags/
co/2003/04/r4035.pdf;   http://
www.computer.org/computer/co2003/
r4035abs.htm.

Wolf:2003:CFD

[888] Wayne Wolf. Cover feature: A decade of hardware/software codesign. Computer, 36(4):38–43, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/04/r4038.htm;
http://csdl.computer.org/dl/mags/
co/2003/04/r4038.pdf;   http://
www.computer.org/computer/co2003/
r4038abs.htm.

Balarin:2003:CFM

[889] Felice Balarin, Yosinori Watanabe, Harry Hsieh, Luciano Lavagno, Claudio Passerone, and Alberto Sangiovanni-Vincentelli. Cover feature: Metropolis: an integrated electronic system design environment. Computer, 36(4):45–52, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/04/r4045.htm;
http://csdl.computer.org/dl/mags/
co/2003/04/r4045.pdf;   http://
www.computer.org/computer/co2003/
r4045abs.htm.

Benini:2003:CFS

[890] Luca Benini, Davide Bertozzi, Davide Bruni, Nicola Drago, Franco Fummi, and Massimo Poncino. Cover feature: SystemC cosimulation and emulation of multiprocessor SoC designs. Computer, 36(4):53–59, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/04/r4053.htm;
http://csdl.computer.org/dl/mags/
co/2003/04/r4053.pdf;   http://
www.computer.org/computer/co2003/
r4053abs.htm.

Richter:2003:CFF

[891] Kai Richter, Marek Jersak, and Rolf Ernst. Cover feature: A formal approach to MpSoC performance verification. Computer, 36(4):60–67, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/04/r4060.htm;
http://csdl.computer.org/dl/mags/
co/2003/04/r4060.pdf;   http://
www.computer.org/computer/co2003/
r4060abs.htm.

Schaumont:2003:DSC

[892] Patrick Schaumont and Ingrid Verbauwhede. Cover feature: Domainspecific codesign for embedded security. Computer, 36(4):68–74, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/04/r4068.htm;
http://csdl.computer.org/dl/mags/
co/2003/04/r4068.pdf;   http://
www.computer.org/computer/co2003/
r4068abs.htm.

Piner:2003:CSCb

[893] Mary-Louise G. Piner. Computer Society connection: Students from around the world receive scholarship funds; IFIP TC-2 Chair Armando Haeberer, 1947–2003; new Education Board mission statement highlights increased vitality; Board of Governors approves reading of bylaws amendment; IEEE Division VIII Directorelect candidates named; research funding available for undergraduate women. Computer, 36(4):75–78, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/04/r4075.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/04/r4075.pdf.

Anonymous:2003:CCd

[894] Anonymous. Call and calendar. Computer, 36(4):79–81, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/04/r4079.htm;
                                        http://www.computer.org/computer/
                                        co2003/r4079.pdf.

Anonymous:2003:COd

[895] Anonymous. Career opportunities. Computer, 36(4):82–87, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://www.computer.
                                        org/computer/co2003/r4082.pdf.

Lutz:2003:BDA

[896] Michael J. Lutz. Bookshelf: In-depth approach to configuration management [Configuration Management Principles and Practice]; fuzzy control, clearly explained [A First Course in Fuzzy and Neural Control]; crime fighting through data mining [Investigative Data Mining for Security and Criminal Detection]; multilevel database tweaking [Database Performance Tuning and Optimization]. Computer, 36(4):88, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/04/r4088.htm;
http://csdl.computer.org/dl/mags/
co/2003/04/r4088.pdf.

Anonymous:2003:PSU

[897] Anonymous. Products: Sybase upgrades RAD tool; haptic workstation uses two hands; TimeSys launches Linux 4.0. Computer, 36(4):89, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/04/r4089.htm;
http://csdl.computer.org/dl/mags/
co/2003/04/r4089.pdf.

Ranganathan:2003:CCA

[898] Anand Ranganathan and Hui Lei. Communications: Context-aware communication. Computer, 36(4):90– 92, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/04/
r4090.htm;   http://csdl.computer.
org/dl/mags/co/2003/04/r4090.pdf.

Gehorsam:2003:ECC

[899] Robert Gehorsam. Entertainment computing: The coming revolution in massively multiuser persistent worlds. Computer, 36(4):93–95, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/04/r4093.htm;

                                        http://csdl.computer.org/dl/mags/
                                        co/2003/04/r4093.pdf.

Naedele:2003:SSX

[900] Martin Naedele. Security: Standards for XML and Web services security. Computer, 36(4):96–98, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/04/r4096.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/04/r4096.pdf.

Johanson:2003:ICI

[901] Brad Johanson, Terry Winograd, and Armando Fox. Invisible computing: Interactive workspaces. Computer, 36(4):99–103, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/04/r4099.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/04/r4099.pdf.

Holmes:2003:PCP

[902] Neville Holmes. The profession: The case for perspicuous programming. Computer, 36(4):104, 102–103, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/04/r4104.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/04/r4104.pdf.

Anonymous:2003:ASe

[903] Anonymous. Article summaries. Computer, 36(5):4, May 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2003/05/

r5004.htm;   http://csdl.computer.
org/dl/mags/co/2003/05/r5004.pdf.

DeMarco:2003:LCC

[904] Tom DeMarco, Neville Holmes, William Adams, Neville Holmes, Neal Laurance, Neville Holmes, Simone Santini, Bob Colwell, Bill Dietrich, Al Davis, and David Resnic. Letters: Calculator column kudos and complaints; bouncing between metaphor and abstraction; requirements triage; Instant Messaging. Computer, 36(5):5–9, May 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/05/r5005.htm;
http://csdl.computer.org/dl/mags/
co/2003/05/r5005.pdf.

Colwell:2003:RLB

[905] Bob Colwell. At random: Leave Bernoulli out of this. Computer, 36 (5):10–12, May 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/05/
r5010.htm;   http://csdl.computer.
org/dl/mags/co/2003/05/r5010.pdf.

Vaughan-Nichols:2003:ITX

[906] Steven J. Vaughan-Nichols. Industry trends: XML raises concerns as it gains prominence. Computer, 36 (5):14–16, May 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/05/
r5014.htm;   http://csdl.computer.
org/dl/mags/co/2003/05/r5014.pdf.

Geer:2003:TNF

[907] David Geer. Technology news: Federated approach expands databaseaccess technology. Computer, 36(5): 18–20, May 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2003/05/
                                        r5018.htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/05/r5018.pdf.

Paulson:2003:NBNa

[908] Linda Dailey Paulson. News briefs: New techniques for speeding wireless streaming video; vendors support new fixed wireless broadband technology; researchers improve parallel processing. Computer, 36(5):21, May 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/05/r5021.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/05/r5021.pdf.

Hopgood:2003:PAI

[909] Adrian A. Hopgood. Perspectives: Artificial intelligence: Hype or reality? Computer, 36(5):24–28, May 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/05/r5024.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/05/r5024.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r5024abs.htm.

Yunos:2003:CPW

[910] Hassim Mohamed Yunos, Jerry Zeyu Gao, and Simon Shim. Computing practices: Wireless advertising’s challenges and opportunities.

Computer, 36(5):30–37, May 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/05/r5030.htm;
http://csdl.computer.org/dl/mags/
co/2003/05/r5030.pdf;   http://
www.computer.org/computer/co2003/
r5030abs.htm.

Milenkovic:2003:CFT

[911] Milan Milenkovic, Scott H. Robinson, Rob C. Knauerhase, David Barkai, Sharad Garg, Vijay Tewari, Todd A. Anderson, and Mic Bowman. Cover feature: Toward Internet distributed computing. Computer, 36 (5):38–46, May 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/05/
r5038.htm;   http://csdl.computer.
org/dl/mags/co/2003/05/r5038.pdf;
 http://www.computer.org/computer/
co2003/r5038abs.htm.

Tseng:2003:CFI

[912] Yu-Chee Tseng, Chia-Ching Shen, and Wen-Tsuen Chen. Cover feature: Integrating mobile IP with ad hoc networks. Computer, 36(5):48–55, May 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/05/r5048.htm;
http://csdl.computer.org/dl/mags/
co/2003/05/r5048.pdf;   http://
www.computer.org/computer/co2003/
r5048abs.htm.

Chau:2003:RFC

[913] Michael Chau and Hsinchun Chen. Research feature: Comparison of three vertical search spiders. Computer, 36 (5):56–62, May 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2003/05/
                                        r5056.htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/05/r5056.pdf;
                                         http://www.computer.org/computer/
                                        co2003/r5056abs.htm.

Robertazzi:2003:RFT

[914] Thomas G. Robertazzi. Research feature: Ten reasons to use divisible load theory. Computer, 36(5):63–68, May 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/05/r5063.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/05/r5063.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r5063abs.htm.

Anonymous:2003:POU

[915] Anonymous. Products: Omnicore upgrades Java development tools; TimeSys announces first RTSJcompliant JVM; Xilinx adds hardwarein-the-loop cosimulation to DSP design tool; gInstall’s upgrades setup system platform; Motorola launches motocoder developer platform; ARM upgrades compilation tools for embedded apps; SGI server leads HPC benchmarks. Computer, 36(5):69–70, May 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/05/r5069.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/05/r5069.pdf.

Lutz:2003:BIC

[916] Michael J. Lutz. Bookshelf: Integrating capability maturity models [CMMI: Guidelines for Process Integration and Product Improvement]; UML with style [The Elements of UML Style]; interviewing the new media [Uncanny Networks: Dialogues with the Virtual Intelligentsia]; evolutionary buffer [Advances in Evolutionary Computing]. Computer, 36(5):71, May 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/05/r5071.htm;
http://csdl.computer.org/dl/mags/
co/2003/05/r5071.pdf.

Anonymous:2003:CCe

[917] Anonymous. Call and calendar. Computer, 36(5):75–77, May 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/05/r5075.htm;   http://www.
computer.org/computer/co2003/r5075.
pdf.

Anonymous:2003:CSCd

[918] Anonymous. Computer Society connection: Conferences explore key computing topics. Computer, 36(5): 78–81, May 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/05/
r5078.htm;   http://csdl.computer.
org/dl/mags/co/2003/05/r5078.pdf.

Anonymous:2003:COe

[919] Anonymous. Career opportunities. Computer, 36(5):82–84, May 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/computer/co2003/r5082.pdf.

Ross:2003:ISP

[920] Ronald G. Ross. IT systems perspective: The business rule approach. Computer, 36(5):85–87, May 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/05/r5085.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/05/r5085.pdf.

Kunkel:2003:ECT

[921] Joachim Kunkel. Embedded computing: Toward IP-based system-level SoC design. Computer, 36(5):88–89, May 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/05/r5088.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/05/r5088.pdf.

Gudivada:2003:PCP

[922] Venkat N. Gudivada. The profession: The computing profession at a crossroads. Computer, 36(5):92, 90–91, May 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/05/r5092.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/05/r5092.pdf.

Anonymous:2003:ASf

[923] Anonymous. Article summaries. Computer, 36(6):4–5, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //csdl.computer.org/dl/mags/co/
                                        2003/06/r6004.htm;   http://www.
                                        computer.org/computer/co2003/r6004.
                                        pdf.

Colwell:2003:RTC

[924] Bob Colwell. At random: That’s not computing—This is computing. Computer, 36(6):6–8, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/06/r6006.htm;
http://csdl.computer.org/dl/mags/
co/2003/06/r6006.pdf.

Vaughan-Nichols:2003:ITO

[925] Steven J. Vaughan-Nichols. Industry trends: OSs battle in the Smart-Phone market. Computer, 36(6):10–12, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/06/r6010.htm;
http://csdl.computer.org/dl/mags/
co/2003/06/r6010.pdf.

Leavitt:2003:TNT

[926] Neal Leavitt. Technology news: Two technologies vie for recognition in speech market. Computer, 36(6): 13–16, June 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/06/
r6013.htm;   http://csdl.computer.
org/dl/mags/co/2003/06/r6013.pdf.

Paulson:2003:NBX

[927] Linda Dailey Paulson. News briefs: Xerox hopes chip will cut costs of fiber services; Sharp develops a 3D display for cell phones and computers; IETF will continue using patented technologies in standards; computer-assisted tuna fishing. Computer, 36(6):18–19, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/06/r6018.htm;
http://csdl.computer.org/dl/mags/
co/2003/06/r6018.pdf.

Parrish:2003:CPC

[928] Allen S. Parrish, Brandon Dixon, David Cordes, Susan Vrbsky, and David Brown. Computing practices: CARE: An automobile crash data analysis tool. Computer, 36(6):22–30, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/06/r6022.htm;
http://csdl.computer.org/dl/mags/
co/2003/06/r6022.pdf;   http://
www.computer.org/computer/co2003/
r6022abs.htm.

Eischen:2003:PAP

[929] Kyle Eischen. Perspectives: Andhra Pradesh: lessons for global software development. Computer, 36 (6):31–37, June 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/06/
r6031.htm;   http://csdl.computer.
org/dl/mags/co/2003/06/r6031.pdf;
 http://www.computer.org/computer/
co2003/r6031abs.htm.

Williams:2003:GEI

[930] Laurie Williams and Alistair Cockburn. Guest Editors’ introduction: Agile software development: It’s about feedback and change. Computer, 36(6):39–43, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/06/r6039.htm;   http://www.

computer.org/computer/co2003/r6039.
pdf.

Beck:2003:CFA

[931] Kent Beck and Barry Boehm. Cover feature: Agility through discipline: A debate. Computer, 36(6):44–46, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/06/r6044.htm;
http://csdl.computer.org/dl/mags/
co/2003/06/r6044.pdf;   http://
www.computer.org/computer/co2003/
r6044abs.htm.

Larman:2003:CFI

[932] Craig Larman and Victor R. Basili. Cover feature: Iterative and incremental development: A brief history. Computer, 36(6):47–56, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/06/r6047.htm;
http://csdl.computer.org/dl/mags/
co/2003/06/r6047.pdf;   http://
www.computer.org/computer/co2003/
r6047abs.htm.

Boehm:2003:CFU

[933] Barry Boehm and Richard Turner. Cover feature: Using risk to balance agile and plan-driven methods. Computer, 36(6):57–66, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/06/r6057.htm;
http://csdl.computer.org/dl/mags/
co/2003/06/r6057.pdf;   http://
www.computer.org/computer/co2003/
r6057abs.htm.

Lippert:2003:CFD

[934] Martin Lippert, Petra Becker-Pechau, Holger Breitling, Jörn Koch, Andreas Kornstädt, Stefan Roock, Axel Schmolitzky, Henning Wolf, and Heinz Züllighoven. Cover feature: Developing complex projects using XP with extensions. Computer, 36(6):67–73, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/06/r6067.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/06/r6067.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r6067abs.htm.

Cohn:2003:CFI

[935] Mike Cohn and Doris Ford. Cover feature: Introducing an agile process to an organization. Computer, 36 (6):74–78, June 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2003/06/
                                        r6074.htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/06/r6074.pdf;
                                         http://www.computer.org/computer/
                                        co2003/r6074abs.htm.

Lycett:2003:CFM

[936] Mark Lycett, Robert D. Macredie, Chaitali Patel, and Ray J. Paul. Cover feature: Migrating agile methods to standardized development practice. Computer, 36(6):79–85, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/06/r6079.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/06/r6079.pdf;   http://

www.computer.org/computer/co2003/
r6079abs.htm.

Ward:2003:CKA

[937] Bob Ward. Cray, Kanai awards recognize achievement in high-performance computing and distributed systems: [Monty Denneau, IBM Blue Gene architect; Steven S. Yao; Alfred Z. Spector]. Computer, 36(6):89–91, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/06/r6089.htm;
http://csdl.computer.org/dl/mags/
co/2003/06/r6089.pdf.

Anonymous:2003:CCf

[938] Anonymous. Call and calendar. Computer, 36(6):92–94, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/06/r6092.htm;   http://www.
computer.org/computer/co2003/r6092.
pdf.

Anonymous:2003:COf

[939] Anonymous. Career opportunities. Computer, 36(6):96–97, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/computer/co2003/r6096.pdf.

Anonymous:2003:PVF

[940] Anonymous. Products: VMware’s fourth-generation desktop virtualization software; automated design reviews with Reviewer for Rose; CodeWeavers debues CrossOver Office; Corel Smart Graphics Studio now available; IronGrid’s Java performance tool; Infragistics releases NetAdvantage component collection. Computer, 36(6):98–99, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/06/r6098.htm;
http://csdl.computer.org/dl/mags/
co/2003/06/r6098.pdf.

Lutz:2003:BOS

[941] Michael J. Lutz. Bookshelf: Offbeat software history lesson [From Airline Reservations to Sonice the Hedgehog: A History of the Software Industry]; real-world approach to user research [Observing the User Experience: A Practitioner’s Guide to User Research]; database design roadmap [Database Design and Development: An Essential Guide for IT Professionals]; in-depth look at compiler design [ The Compiler Design Handbook: Optimizations and Machine Code Generation]. Computer, 36(6):100, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/06/r6100.htm;
http://csdl.computer.org/dl/mags/
co/2003/06/r6100.pdf.

Varshney:2003:CSF

[942] Upkar Varshney. Communications: The status and future of 802.11based WLANs. Computer, 36(6): 102–105, June 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/06/
r6102.htm;   http://csdl.computer.
org/dl/mags/co/2003/06/r6102.pdf.

Fogel:2003:ECE

[943] David B. Fogel. Entertainment computing: Evolutionary entertainment with intelligent agents. Computer, 36(6):106–108, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/06/r6106.htm;
http://csdl.computer.org/dl/mags/
co/2003/06/r6106.pdf.

Lamming:2003:ICS

[944] Mik Lamming and Denis Bohm. Invisible computing: SPECs: Personal pervasive systems. Computer, 36(6):109–111, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/06/r6109.htm;
http://csdl.computer.org/dl/mags/
co/2003/06/r6109.pdf.

Arbaugh:2003:SFO

[945] William A. Arbaugh. Security: Firewalls: An outdated defense. Computer, 36(6):112–113, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/06/r6112.htm;
http://csdl.computer.org/dl/mags/
co/2003/06/r6112.pdf.

Holmes:2003:PPU

[946] Neville Holmes. The profession: The problem with Unicode. Computer, 36(6):116, 114–115, June 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/06/r6116.htm;

                                        http://csdl.computer.org/dl/mags/
                                        co/2003/06/r6116.pdf.

Anonymous:2003:ASg

[947] Anonymous. Article summaries. Computer, 36(7):4, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //csdl.computer.org/dl/mags/co/
                                        2003/07/r7004.htm;   http://www.
                                        computer.org/computer/co2003/r7004.
                                        pdf.

Raskin:2003:LWH

[948] Jef Raskin, William R. Dunn, Peter Bissmire, and Yibing Wang. Letters: What holds up a plane? putting the profession in perspective. Computer, 36(7):5–6, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //csdl.computer.org/dl/mags/co/
                                        2003/07/r7005.htm;   http://www.
                                        computer.org/computer/co2003/r7005.
                                        pdf.

Colwell:2003:RTT

[949] Bob Colwell. At random: Tools of the trade. Computer, 36(7):7–9, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/07/r7007.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/07/r7007.pdf.

Paulson:2003:ITI

[950] Linda Dailey Paulson. Industry trends: The ins and outs of new local I/O trends. Computer, 36(7):13–16, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/07/r7013.htm;
http://csdl.computer.org/dl/mags/
co/2003/07/r7013.pdf.

Vaughan-Nichols:2003:TNC

[951] Steven J. Vaughan-Nichols. Technology news: The challenge of Wi-Fi roaming. Computer, 36(7):17–19, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/07/r7017.htm;
http://csdl.computer.org/dl/mags/
co/2003/07/r7017.pdf.

Paulson:2003:NBCa

[952] Linda Dailey Paulson. News briefs: Chip making to dye for; using AI in games that adapt to users; group considers drastic measures to stop spam; technology enables online voice auditions. Computer, 36(7):20–22, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/07/r7020.htm;
http://csdl.computer.org/dl/mags/
co/2003/07/r7020.pdf.

Blanchette:2003:PSU

[953] Stephen Blanchette Jr. Perspectives: Surviving unemployment in the high-tech downturn. Computer, 36 (7):24–28, July 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/07/
r7024.htm;   http://csdl.computer.
org/dl/mags/co/2003/07/r7024.pdf;
 http://www.computer.org/computer/
co2003/r7024abs.htm.

Oppliger:2003:CPM

[954] Rolf Oppliger. Computing practices: Microsoft .NET passport: A security analysis. Computer, 36(7):29–35, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/07/r7029.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/07/r7029.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r7029abs.htm.

Srinivasan:2003:GEI

[955] Savitha Srinivasan. Guest Editor’s introduction: On piracy and privacy. Computer, 36(7):36–38, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/07/r7036.htm;
                                        http://www.computer.org/computer/
                                        co2003/r7036.pdf.

Eskicioglu:2003:CFP

[956] Ahmet M. Eskicioglu. Cover feature: Protecting intellectual property in digital multimedia networks. Computer, 36(7):39–45, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/07/r7039.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/07/r7039.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r7039abs.htm.

Naor:2003:CFP

[957] Dalit Naor and Moni Naor. Cover feature: Protecting cryptographic keys: the trace-and-revoke approach. Computer, 36(7):47–53, July 2003.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/07/r7047.htm;
http://csdl.computer.org/dl/mags/
co/2003/07/r7047.pdf;   http://
www.computer.org/computer/co2003/
r7047abs.htm.

England:2003:CFT

[958] Paul England, Butler Lampson, John Manferdelli, Marcus Peinado, and Bryan Willman. Cover feature: A trusted open platform. Computer, 36(7):55–62, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/07/r7055.htm;
http://csdl.computer.org/dl/mags/
co/2003/07/r7055.pdf;   http://
www.computer.org/computer/co2003/
r7055abs.htm.

Naumovich:2003:CFP

[959] Gleb Naumovich and Nasir Memon. Cover feature: Preventing piracy, reverse engineering, and tampering. Computer, 36(7):64–71, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/07/r7064.htm;
http://csdl.computer.org/dl/mags/
co/2003/07/r7064.pdf;   http://
www.computer.org/computer/co2003/
r7064abs.htm.

Brendan:2003:CFT

[960] C. Brendan and S. Traw. Cover feature: Technical challenges of protecting digital entertainment content. Computer, 36(7):72–78, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/07/r7072.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/07/r7072.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r7072abs.htm.

Anonymous:2003:CSCe

[961] Anonymous. Computer Society connection: Service, education, and technical awards honor outstanding achievements. Computer, 36(7):79–80, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/07/r7079.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/07/r7079.pdf.

Anonymous:2003:CCg

[962] Anonymous. Call and calendar. Computer, 36(7):81–83, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //csdl.computer.org/dl/mags/co/
                                        2003/07/r7081.htm;   http://www.
                                        computer.org/computer/co2003/r7081.
                                        pdf.

Anonymous:2003:PAS

[963] Anonymous. Products: Artifact Software offers personal code manager; Whirling Dervishes releases namespace extension library; MathWorks announces imaging application toolboxes; Compuware enhances driver development tools; Netscout Systems extends storage network performance monitors; PTC upgrades routing schematics software; VoiceGenie releases graphical IVR application environment. Computer, 36(7):84–85, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/07/r7084.htm;
http://csdl.computer.org/dl/mags/
co/2003/07/r7084.pdf.

Lutz:2003:BSW

[964] Michael J. Lutz. Bookshelf: Surveying the Web’s best and brightest [Deep Sites: Intelligent Innovation in Contemporary Web Design]; moving voice communications to the Web [Implementing Voice over IP]; developing custom .NET management tools [.NET Systems Management Services]; evolving Java [Applied Evolutionary Algorithms in Java]. Computer, 36(7):86, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/07/r7086.htm;
http://csdl.computer.org/dl/mags/
co/2003/07/r7086.pdf.

Bradshaw:2003:WTT

[965] Jeffrey M. Bradshaw, Giacomo Cabri, and Rebecca Montanari. Web technologies: Taking back cyberspace. Computer, 36(7):89–92, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/07/r7089.htm;
http://csdl.computer.org/dl/mags/
co/2003/07/r7089.pdf.

Langdon:2003:ISP

[966] Christoph Schlueter Langdon. IT systems perspective: The state of Web services. Computer, 36(7):93–94, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/07/r7093.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/07/r7093.pdf.

Gupta:2003:ECD

[967] Rajesh Gupta. Embedded computing: Driving research in system-chip design technology. Computer, 36 (7):95–97, July 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2003/07/
                                        r7095.htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/07/r7095.pdf.

Holmes:2003:PSP

[968] Neville Holmes. The profession: Should professionals be political? Computer, 36(7):100, 98–99, July 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/07/r7100.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/07/r7100.pdf.

Anonymous:2003:ASh

[969] Anonymous. Article summaries. Computer, 36(8):4, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //csdl.computer.org/dl/mags/co/
                                        2003/08/r8004.htm;   http://www.
                                        computer.org/computer/co2003/r8004.
                                        pdf.

Heires:2003:LEV

[970] James T. Heires, Barry Boehm, Li Guo Huang, Paul C. Jorgensen, Laurie Williams, Alistair Cockburn, Rita Creel, Rob Schaaf, Steven Hong, William A. Arbaugh, and Mikhail S. Brikman. Letters: Earned-value management terminology; putting agile methods in perspective; standards clarification; balancing feedback and foresight; defending firewalls; analyzing XML development trends. Computer, 36(8):5–7, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/08/r8005.htm;   http://www.
computer.org/computer/co2003/r8005.
pdf.

Colwell:2003:RED

[971] Bob Colwell. At random: Engineering decisions. Computer, 36(8): 9–11, August 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/08/
r8009.htm;   http://csdl.computer.
org/dl/mags/co/2003/08/r8009.pdf.

Ward:2003:CWC

[972] Bob Ward. CSIDC winners: CSIDC WINNERS: Tablet-PC classroom system wins design competition. Computer, 36(8):15–18, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/08/r8015.htm;
http://csdl.computer.org/dl/mags/
co/2003/08/r8015.pdf.

Vaughan-Nichols:2003:ITS

[973] Steven J. Vaughan-Nichols. Industry trends: Seeking better Web search technologies. Computer, 36(8):19– 21, August 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2003/08/
                                        r8019.htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/08/r8019.pdf.

Ortiz:2003:TNS

[974] Sixto Ortiz Jr. Technology news: Software radios add flexibility to wireless technology. Computer, 36(8): 23–25, August 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2003/08/
                                        r8023.htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/08/r8023.pdf.

Paulson:2003:NBS

[975] Linda Dailey Paulson. News briefs: Spammers begin sending viruses with their junk mail; “ballistic” technique boosts storage capacity; FAST approach could accelerate Internet downloads; software cuts the “bull” out of documents. Computer, 36(8):26–28, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/08/r8026.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/08/r8026.pdf.

Skadron:2003:PCC

[976] Kevin Skadron, Margaret Martonosi, David I. August, Mark D. Hill, David J. Lilja, and Vijay S. Pai. Perspectives: Challenges in computer architecture evaluation. Computer, 36(8):30–36, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/08/r8030.htm;

http://csdl.computer.org/dl/mags/
co/2003/08/r8030.pdf;   http://
www.computer.org/computer/co2003/
r8030abs.htm.

Wood:2003:CPS

[977] Alan P. Wood. Computing practices: Software reliability from the customer view. Computer, 36(8):37–42, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/08/r8037.htm;
http://csdl.computer.org/dl/mags/
co/2003/08/r8037.pdf;   http://
www.computer.org/computer/co2003/
r8037abs.htm.

Bourianoff:2003:CFF

[978] George Bourianoff. Cover feature: The future of nanocomputing. Computer, 36(8):44–53, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/08/r8044.htm;
http://csdl.computer.org/dl/mags/
co/2003/08/r8044.pdf;   http://
www.computer.org/computer/co2003/
r8044abs.htm.

Kapasi:2003:CFP

[979] Ujval J. Kapasi, Scott Rixner, William J. Dally, Brucek Khailany, Jung Ho Ahn, Peter Mattson, and John D. Owens. Cover feature: Programmable stream processors. Computer, 36(8):54–62, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/08/r8054.htm;
http://csdl.computer.org/dl/mags/
co/2003/08/r8054.pdf;   http://

                                        www.computer.org/computer/co2003/
                                        r8054abs.htm.

Najjar:2003:RFH

[980] Walid A. Najjar, Wim Böhm, Bruce A. Draper, Jeff Hammes, Robert Rinker, J. Ross Beveridge, Monica Chawathe, and Charles Ross. Research feature: High-level language abstraction for reconfigurable computing. Computer, 36(8):63–69, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/08/r8063.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/08/r8063.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r8063abs.htm.

Thomas:2003:RFP

[981] Fred Thomas. Research feature: Penny tag technologies for removable data storage. Computer, 36(8): 70–76, August 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2003/08/
                                        r8070.htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/08/r8070.pdf;
                                         http://www.computer.org/computer/
                                        co2003/r8070abs.htm.

Ward:2003:CSCa

[982] Bob Ward. Computer Society connection: Russian teen wins top IEEE Computer Society prize at Intel Science Fair; Computer Society Board of Governors approves reading of proposed bylaw changes. Computer, 36(8):77–80, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/08/r8077.htm;
http://csdl.computer.org/dl/mags/
co/2003/08/r8077.pdf.

Anonymous:2003:CCh

[983] Anonymous. Call and calendar. Computer, 36(8):81–83, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/08/r8081.htm;
http://www.computer.org/computer/
co2003/r8081.pdf.

Anonymous:2003:COg

[984] Anonymous. Career opportunities. Computer, 36(8):84–87, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/computer/co2003/r8084.pdf.

Anonymous:2003:PAI

[985] Anonymous. Products: Apple introduces 64-bit PC with 1-GHz bus; Altera upgrades PLD design tool; StrikeIron’s initial product analyzes Web services; Metrowerks teams up with AMD on Linux platform. Computer, 36(8):88, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/08/r8088.htm;
http://csdl.computer.org/dl/mags/
co/2003/08/r8088.pdf.

Lutz:2003:BMB

[986] Michael J. Lutz. Bookshelf: Mathbased introduction to 3D graphics [3-D Computer Graphics: A Mathematical Introduction with OpenGL]; thinking tools for agile development [Lean software development: an agile toolkit]; copyright versus creativity [Copyrights and Copywrongs: The Rise of Intellectual Property and How It Threatens Creativity]; Web design for Scalable Vector Graphics [SVG for Web Developers]. Computer, 36(8):89, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/08/r8089.htm;
http://csdl.computer.org/dl/mags/
co/2003/08/r8089.pdf.

Jain:2003:CWC

[987] Saurabh Jain and Dharma P. Agrawal. Communications: Wireless community networks. Computer, 36(8): 90–92, August 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/08/
r8090.htm;   http://csdl.computer.
org/dl/mags/co/2003/08/r8090.pdf.

Landay:2003:ICD

[988] James A. Landay and Gaetano Borriello. Invisible computing: Design patterns for ubiquitous computing. Computer, 36(8):93–95, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/08/r8093.htm;
http://csdl.computer.org/dl/mags/
co/2003/08/r8093.pdf.

Schell:2003:ECS

[989] Jesse Schell. Entertainment computing: Shaping an entertaining future at Carnegie Mellon. Computer, 36(8):96–98, August 2003.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/08/r8096.htm;
http://csdl.computer.org/dl/mags/
co/2003/08/r8096.pdf.

Arbaugh:2003:SWS

[990] William A. Arbaugh. Security: Wireless security is different. Computer, 36(8):99–101, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/08/r8099.htm;
http://csdl.computer.org/dl/mags/
co/2003/08/r8099.pdf.

Santini:2003:PBC

[991] Simone Santini. The profession: Bringing copyright into the information age. Computer, 36(8):104, 102–103, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/08/r8104.htm;
http://csdl.computer.org/dl/mags/
co/2003/08/r8104.pdf.

Anonymous:2003:ASi

[992] Anonymous. Article summaries. Computer, 36(9):4, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/09/r9004.htm;
http://www.computer.org/computer/
co2003/r9004.pdf.

Colwell:2003:REM

[993] Bob Colwell. At random: Engineering, marketing teams, and management. Computer, 36(9):5–7, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/09/r9005.htm;
http://csdl.computer.org/dl/mags/
co/2003/09/r9005.pdf.

Paulson:2003:ITR

[994] Linda Dailey Paulson. Industry trends: Reconfiguring wireless phones with adaptive chips. Computer, 36(9):9–11, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/09/r9009.htm;
http://csdl.computer.org/dl/mags/
co/2003/09/r9009.pdf.

Vaughan-Nichols:2003:TNBb

[995] Steven J. Vaughan-Nichols. Technology news: Building better software with better tools. Computer, 36(9):12–14, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/09/r9012.htm;
http://csdl.computer.org/dl/mags/
co/2003/09/r9012.pdf.

Paulson:2003:NBNb

[996] Linda Dailey Paulson. News briefs: Nanotech RAM holds promise for universal memory; US court establishes protection against reverse engineering; using Web crawlers to fight piracy; networked piano-to-piano communication to help with music education. Computer, 36(9):15–17, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/

dl/mags/co/2003/09/r9015.htm;
http://csdl.computer.org/dl/mags/
co/2003/09/r9015.pdf.

Feitelson:2003:CPM

[997] Dror G. Feitelson. Computing practices: Metric and workload effects on computer systems evaluation. Computer, 36(9):18–25, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/09/r9018.htm;
http://csdl.computer.org/dl/mags/
co/2003/09/r9018.pdf;   http://
www.computer.org/computer/co2003/
r9018abs.htm.

Myers:2003:GEI

[998] Brad A. Myers and Michael Beigl. Guest Editors’ introduction: Handheld computing. Computer, 36(9): 27–29, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/09/r9027.htm;   http://www.
computer.org/computer/co2003/r9027.
pdf.

Tatar:2003:CFH

[999] Deborah Tatar, Jeremy Roschelle, Phil Vahey, and William R. Penuel. Cover feature: Handhelds go to school: Lessons learned. Computer, 36(9):30–37, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/09/r9030.htm;
http://csdl.computer.org/dl/mags/
co/2003/09/r9030.pdf;   http://
www.computer.org/computer/co2003/
r9030abs.htm.

Munoz:2003:CFC

[1000] Miguel A. Muñoz, Marcela Rodríguez, Jesus Favela, Ana I. Martinez-Garcia, and Victor M. González. Cover feature: Context-aware mobile communication in hospitals. Computer, 36(9):38–46, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/09/r9038.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/09/r9038.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r9038abs.htm.

Bellotti:2003:CFV

[1001] Francesco Bellotti, Riccardo Berta, Alessandro De Gloria, Edmondo Ferretti, and Massimiliano Margarone. Cover feature: VeGame: Exploring art and history in Venice. Computer, 36(9):48–55, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/09/r9048.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/09/r9048.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        r9048abs.htm.

Raghunath:2003:CFF

[1002] Mandayam Raghunath, Chandra Narayanaswami, and Claudio Pinhanez. Cover feature: Fostering a symbiotic handheld environment. Computer, 36(9):56–65, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/09/r9056.htm;
                                        http://csdl.computer.org/dl/mags/

co/2003/09/r9056.pdf;   http://
www.computer.org/computer/co2003/
r9056abs.htm.

Medvidovic:2003:CFS

[1003] Nenad Medvidovic, Marija Mikic-Rakic, Nikunj R. Mehta, and Sam Malek. Cover feature: Software architectural support for handheld computing. Computer, 36(9):66–73, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/09/r9066.htm;
http://csdl.computer.org/dl/mags/
co/2003/09/r9066.pdf;   http://
www.computer.org/computer/co2003/
r9066abs.htm.

Anonymous:2003:ICS

[1004] Anonymous. IEEE Computer Society election. Computer, 36(9):75– 82, September 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/09/
r9075.htm;   http://csdl.computer.
org/dl/mags/co/2003/09/r9075.pdf.

Chang:2003:CSC

[1005] Carl K. Chang. Computer Society connection: IEEE President-Elect candidates answer questions on critical governance issues. Computer, 36(9):83–89, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/09/r9083.htm;
http://csdl.computer.org/dl/mags/
co/2003/09/r9083.pdf.

Anonymous:2003:CCi

[1006] Anonymous. Call and calendar. Computer, 36(9):90–93, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/09/r9090.htm;
http://www.computer.org/computer/
co2003/r9090.pdf.

Anonymous:2003:PCU

[1007] Anonymous. Products: Compuware upgrades J2EE development environment; Ektron releases browser-based image tool; IronGrid offers JDBC performance tool; Microsoft enhances Java conversion assistant; Broadcom announces single-chip 10-Gigabit Ethernet switch; SGI finalizes OpenGL 1.5 specification; Adaptec extends Serial ATA RAID product family. Computer, 36(9):94–95, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/09/r9094.htm;
http://csdl.computer.org/dl/mags/
co/2003/09/r9094.pdf.

Lutz:2003:BBC

[1008] Michael J. Lutz. Bookshelf: Building C# programs with patterns [C# Design Patterns]; using Star Schema in Oracle [Oracle DBA Guide to Data Warehousing and Star Schemas]; immersive Java text [Java Outside In]; applying AI to the Web [Web Intelligence]; moronic marketing [In Search of Stupidity: Over 20 Years of High-Tech Marketing Disasters]. Computer, 36(9):96–97, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/09/r9096.htm;
http://csdl.computer.org/dl/mags/
co/2003/09/r9096.pdf.

Anonymous:2003:COh

[1009] Anonymous. Career opportunities.

Computer, 36(9):98–105, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/computer/co2003/r9098.pdf.

Elhanany:2003:SSF

[1010] Itamar Elhanany, Kurt Busch, and Derek Chiou. Standards: Switch fabric interfaces. Computer, 36(9):106– 108, September 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/09/
r9106.htm;   http://csdl.computer.
org/dl/mags/co/2003/09/r9106.pdf.

Ghahramani:2003:ISP

[1011] Bahador Ghahramani and Mark A.

Pauley. IT systems perspective: Java in high-performance environments. Computer, 36(9):109–111, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/09/r9109.htm;
http://csdl.computer.org/dl/mags/
co/2003/09/r9109.pdf.

Bayardo:2003:WTT

[1012] Roberto J. Bayardo and Ramakrishnan Srikant. Web technologies: Technological solutions for protecting privacy. Computer, 36(9):115–118, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/09/r9115.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/09/r9115.pdf.

Henkel:2003:ECC

[1013] Jörg Henkel. Embedded computing: Closing the SoC design gap. Computer, 36(9):119–121, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/09/r9119.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/09/r9119.pdf.

Cellary:2003:PPR

[1014] Wojciech Cellary. The profession: The profession’s role in the global information society. Computer, 36 (9):124, 122–123, September 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/09/r9124.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/09/r9124.pdf.

Anonymous:2003:AIP

[1015] Anonymous. About this issue: The promise of Web services computing and point out gaps in the technologies. Computer, 36(10):1, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2003:ASj

[1016] Anonymous. Article summaries. Computer, 36(10):5–6, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/10/rx005.htm;

http://www.computer.org/computer/
co2003/rx005.pdf.

Kolb:2003:LAM

[1017] Todd Kolb, Timothy C. Lommasson, and Doug Domeny. Letters: Academia meets industry; stopping spam; A breath of fresh air. Computer, 36(10):7, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/10/rx007.htm;   http://www.
computer.org/computer/co2003/rx007.
pdf.

Colwell:2003:RDR

[1018] Bob Colwell. At random: Design reviews. Computer, 36(10):8–10, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx008.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx008.pdf;   http://
www.computer.org/computer/co2003/
rx008abs.htm.

Vaughan-Nichols:2003:ITP

[1019] Steven J. Vaughan-Nichols. Industry trends: Presence technology: More than just instant messaging. Computer, 36(10):11–13, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx011.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx011.pdf;   http://
www.computer.org/computer/co2003/
rx011abs.htm.

Geer:2003:TNT

[1020] David Geer. Technology news: Taking steps to secure Web services. Computer, 36(10):14–16, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx014.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx014.pdf;   http://
www.computer.org/computer/co2003/
rx014abs.htm.

Paulson:2003:NBCb

[1021] Linda Dailey Paulson. News briefs: Common fabric dyes promise to help optical communications; proponents drop support for controversial software proposal; Intel says: Read my lips; companies look for unit of computing consumption. Computer, 36(10):17, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx017.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx017.pdf.

Halpern:2003:PVT

[1022] Mark Halpern. Perspectives: VIM: Taming software with hardware. Computer, 36(10):21–25, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx021.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx021.pdf;   http://
www.computer.org/computer/co2003/
rx021abs.htm.

Romein:2003:CPS

[1023] John W. Romein and Henri E. Bal.

Computing practices: Solving awari with parallel retrograde analysis. Computer, 36(10):26–33, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx026.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx026.pdf;   http://
www.computer.org/computer/co2003/
rx026abs.htm.

Chung:2003:GEI

[1024] Jen-Yao Chung, Kwei-Jay Lin, and Richard G. Mathieu. Guest Editors’ introduction: Web services computing: Advancing software interoperability. Computer, 36(10):35– 37, October 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/10/rx035.htm;   http://www.
computer.org/computer/co2003/rx035.
pdf.

Turner:2003:CFT

[1025] Mark Turner, David Budgen, and Pearl Brereton. Cover feature: Turning software into a service. Computer, 36(10):38–44, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx038.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx038.pdf;   http://
www.computer.org/computer/co2003/
rx038abs.htm.

Peltz:2003:CFW

[1026] Chris Peltz. Cover feature: Web services orchestration and choreography. Computer, 36(10):46–52, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx046.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx046.pdf;   http://
www.computer.org/computer/co2003/
rx046abs.htm.

Thompson:2003:CFC

[1027] Timothy Thompson, Rick Weil, and Mark D. Wood. Cover feature: CPXe: Web services for Internet imaging. Computer, 36(10):54–62, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx054.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx054.pdf;   http://
www.computer.org/computer/co2003/
rx054abs.htm.

Tsai:2003:CFO

[1028] Tse-Ming Tsai, Han-Kuan Yu, Hsin-Te Shih, Ping-Yao Liao, Ren-Dar Yang, and Seng cho T. Chou. Cover feature: Ontology-mediated integration of intranet Web services. Computer, 36(10):63–71, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx063.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx063.pdf;   http://
www.computer.org/computer/co2003/
rx063abs.htm.

Dubnov:2003:RFU

[1029] Shlomo Dubnov, Gerard Assayag, Olivier Lartillot, and Gill Bejerano. Research feature: Using machine-learning methods for musical style modeling. Computer, 36(10):73–80, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx073.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx073.pdf;   http://
www.computer.org/computer/co2003/
rx073abs.htm.

Ward:2003:CSCb

[1030] Bob Ward. Computer Society connection: New journals address biology, security; IEEE Computer Society recognizes certified software developers professionals. Computer, 36(10):81–83, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx081.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx081.pdf.

Anonymous:2003:CCj

[1031] Anonymous. Call and calendar. Computer, 36(10):87–89, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx087.htm;
http://www.computer.org/computer/
co2003/rx087.pdf.

Anonymous:2003:COi

[1032] Anonymous. Career opportunities.

Computer, 36(10):90–99, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/computer/co2003/rx090.pdf.

Anonymous:2003:PIU

[1033] Anonymous. Products: Intel updates multimedia performance primitives library; Lindows.com launches low-cost Linux PC; Metrowerks adds Linux kernel-level debugging; Rogue Wave announces C/C++ tool for Web services; ScanSoft announces toolkits to support Windows Mobile 2003; CoCreate launches collaboration toolset; Sybase ships IDE for Windows CE-based devices. Computer, 36(10):100–101, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx100.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx100.pdf.

Lutz:2003:BLF

[1034] Michael J. Lutz. Bookshelf: Laying the foundation for pervasive computing [Pervasive Computing]; integrated approach to data handling [Exploratory Data Mining and Data Cleaning]; making theory practical [Modern Cryptography: Theory and Practice]; detailed look at systems testing [Testing of Digital Systems]. Computer, 36(10):102, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/10/rx102.htm;   http://www.
computer.org/computer/co2003/rx102.
pdf.

Chan:2003:SSP

[1035] Haowen Chan and Adrian Perrig. Security: Security and privacy in sensor networks. Computer, 36(10):103–105, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx103.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx103.pdf;   http://
www.computer.org/computer/co2003/
rx103abs.htm.

Macedonia:2003:ECG

[1036] Michael Macedonia. Entertainment computing: The GPU enters computing’s mainstream. Computer, 36(10):106–108, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx106.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx106.pdf;   http://
www.computer.org/computer/co2003/
rx106abs.htm.

Elhanany:2003:CBN

[1037] Itamar Elhanany and Vahid Tabatabaee.

Communications: Benchmarking nextgeneration switch fabrics. Computer, 36(10):109–110, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/10/rx109.htm;
http://csdl.computer.org/dl/mags/
co/2003/10/rx109.pdf;   http://
www.computer.org/computer/co2003/
rx109abs.htm.

McDonald:2003:ICU

[1038] David W. McDonald. Invisible computing: Ubiquitous recommendation systems. Computer, 36(10):111–112, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2003/10/
                                        rx111.htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/10/rx111.pdf;
                                         http://www.computer.org/computer/
                                        co2003/rx111abs.htm.

Zhuge:2003:PFE

[1039] Hai Zhuge and Xiaoqing Shi. The profession: Fighting epidemics in the information and knowledge age. Computer, 36(10):116, 114–115, October 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/10/rx116.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/10/rx116.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        rx116abs.htm.

Anonymous:2003:ASk

[1040] Anonymous. Article summaries. Computer, 36(11):4, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/11/ry004.htm;
                                        http://www.computer.org/computer/
                                        co2003/ry004.pdf.

Colwell:2003:RHT

[1041] Bob Colwell. At random: How not to teach science. Computer, 36(11): 7–9, November 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/11/
ry007.htm;   http://csdl.computer.
org/dl/mags/co/2003/11/ry007.pdf.

Paulson:2003:ITWa

[1042] Linda Dailey Paulson. Industry trends: Will fuel cells replace batteries in mobile devices? Computer, 36(11):10–12, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/11/ry010.htm;
http://csdl.computer.org/dl/mags/
co/2003/11/ry010.pdf.

Vaughan-Nichols:2003:TNS

[1043] Steven J. Vaughan-Nichols. Technology news: Statistical language approach translates into success. Computer, 36(11):14–16, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/11/ry014.htm;
http://csdl.computer.org/dl/mags/
co/2003/11/ry014.pdf.

Paulson:2003:NBG

[1044] Linda Dailey Paulson. New briefs:

Getting in touch with interactive haptic technology; finding optical fiber in the deep blue sea; researcher has new approach to making music with computers; W3C works on Semantic Web proposal. Computer, 36(11):18–20, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/11/ry018.htm;
http://csdl.computer.org/dl/mags/
co/2003/11/ry018.pdf.

Menzies:2003:CPD

[1045] Tim Menzies and Ying Hu. Computing practices: Data mining for very busy people. Computer, 36(11):22–29, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/11/ry022.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/11/ry022.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        ry022abs.htm.

Ashok:2003:CPN

[1046] Roy L. Ashok and Dharma P. Agrawal. Computing practices: Nextgeneration wearable networks. Computer, 36(11):31–39, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/11/ry031.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/11/ry031.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        ry031abs.htm.

Dunn:2003:CFD

[1047] William R. Dunn. Cover feature: Designing safety-critical computer systems. Computer, 36(11):40– 46, November 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2003/11/
                                        ry040.htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/11/ry040.pdf;
                                         http://www.computer.org/computer/
                                        co2003/ry040abs.htm.

Arnout:2003:CFU

[1048] Karine Arnout and Bertrand Meyer.

Cover feature: Uncovering hidden contracts: The .NET example. Computer, 36(11):48–55, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/11/ry048.htm;
http://csdl.computer.org/dl/mags/
co/2003/11/ry048.pdf;   http://
www.computer.org/computer/co2003/
ry048abs.htm.

Caro:2003:RFS

[1049] Armando L. Caro Jr., Janardhan R.

Iyengar, Paul D. Amer, Sourabh Ladha, Gerard J. Heinz II, and Keyur C. Shah. Research feature: SCTP: A proposed standard for robust Internet data transport. Computer, 36(11):56–63, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/11/ry056.htm;
http://csdl.computer.org/dl/mags/
co/2003/11/ry056.pdf;   http://
www.computer.org/computer/co2003/
ry056abs.htm.

Geng:2003:RFS

[1050] Xianjun Geng, Ram D. Gopal, R. Ramesh, and Andrew B. Whinston. Research feature: Scaling Web services with capacity provision networks. Computer, 36(11):64–72, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/11/ry064.htm;
http://csdl.computer.org/dl/mags/
co/2003/11/ry064.pdf;   http://

                                        www.computer.org/computer/co2003/
                                        ry064abs.htm.

Ward:2003:CSCc

[1051] Bob Ward. Computer Society connection: Computer Society magazines look ahead to 2004. Computer, 36(11):74–77, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/11/ry074.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/11/ry074.pdf.

Anonymous:2003:CCk

[1052] Anonymous. Call and calendar. Computer, 36(11):79–81, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/11/ry079.htm;
                                        http://www.computer.org/computer/
                                        co2003/ry079.pdf.

Lutz:2003:BFE

[1053] Michael J. Lutz. Bookshelf: The five essential metrics [Five Core Metrics: The Intelligence Behind Successful Software Management]; hunting bugs in concurrent systems [The SPIN Model Checker: Primer and Reference Model]; revised digital testing tome [Digital Logic Testing and Simulation]; Java for science and engineering [Technical Java: Developing Scientific and Engineering Applications]. Computer, 36(11):85, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/11/ry085.htm;

http://csdl.computer.org/dl/mags/
co/2003/11/ry085.pdf.

Anonymous:2003:PCN

[1054] Anonymous. Products: ClearSight Networks releases application-layer analyzer; Intervoice announces first SALT-based components; VoiceGenie Technologies upgrades VoiceXML platform; AppForce enhances mobileplatform design software; Metrowerks upgrades tools for embedded Linux products; OpenOffice.org updates Linux office tool suite; Quest Software releases Java analysis tools. Computer, 36(11):86–87, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/11/ry086.htm;
http://csdl.computer.org/dl/mags/
co/2003/11/ry086.pdf.

Anonymous:2003:COj

[1055] Anonymous. Career opportunities.

Computer, 36(11):88–102, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/computer/co2003/ry088.pdf.

Johnson:2003:ECF

[1056] Steven D. Johnson. Embedded computing: Formal methods in embedded design. Computer, 36(11):104– 106, November 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/11/
ry104.htm;   http://csdl.computer.
org/dl/mags/co/2003/11/ry104.pdf;
 http://www.computer.org/computer/
co2003/ry104abs.htm.

McClelland:2003:SMS

[1057] Marilyn McClelland. Standards: Metadata standards for educational resources. Computer, 36(11):107–109, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2003/11/
                                        ry107.htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/11/ry107.pdf;
                                         http://www.computer.org/computer/
                                        co2003/ry107abs.htm.

Hevner:2003:ISP

[1058] Alan R. Hevner and Salvatore T. March. IT systems perspectives: The information systems research cycle. Computer, 36(11):111–113, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/11/ry111.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/11/ry111.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        ry111abs.htm.

Lindahl:2003:WTW

[1059] Charlie Lindahl and Elise Blount. Web technologies: Weblogs: Simplifying Web publishing. Computer, 36(11):114–116, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2003/11/ry114.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2003/11/ry114.pdf;   http://
                                        www.computer.org/computer/co2003/
                                        ry114abs.htm.

Holmes:2003:PAI

[1060] Neville Holmes. The profession:

Artificial intelligence: Arrogance or ignorance? Computer, 36 (11):120, 118–119, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2003/11/ry120.htm;
http://csdl.computer.org/dl/mags/
co/2003/11/ry120.pdf;   http://
www.computer.org/computer/co2003/
ry120abs.htm.

Anonymous:2003:ASl

[1061] Anonymous. Article summaries. Computer, 36(12):4, December 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://www.computer.org/computer/
co2003/rz004.pdf.

Wallace:2003:LRD

[1062] Mark Wallace, Raghavendra Rao Loka, Steven J. Vaughan-Nichols, Jesse N. Alexander, Mark Neidengard, Mark Halpern, Eric Haines, Behrooz Parhami, Hai Zhuge, and Xiaoqing Shi. Letters: Rearranging the deck chairs; successful design reviews; better software versus more software; GPU clarification; challenging a theoretic discussion. Computer, 36(12):6–8, December 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/computer/co2003/rz006.pdf.

Colwell:2003:RBC

[1063] Bob Colwell. At random: Benchmarketing competition. Computer, 36 (12):9–11, December 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //csdl.computer.org/dl/mags/co/
                                        2003/12/rz009.pdf;   http://www.
                                        computer.org/computer/co2003/rz009abs.
                                        htm.

Diamond:2003:PML

[1064] Stephen L. Diamond. President’s message: Looking forward. Computer, 36 (12):12–14, December 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //csdl.computer.org/dl/mags/co/
                                        2003/12/rz012.pdf.

Paulson:2003:ITWb

[1065] Linda Dailey Paulson. Industry trends: Will ultrawideband technology connect in the marketplace? Computer, 36(12): 15–17, December 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2003/12/
                                        rz015.pdf.

Vaughan-Nichols:2003:TNHb

[1066] Steven J. Vaughan-Nichols. Technology news: Hard drive technology reaches a turning point. Computer, 36(12): 21–23, December 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2003/12/
                                        rz021.pdf.

Paulson:2003:NBRb

[1067] Linda Dailey Paulson. News briefs: Researchers plan experimental optical network; using magnets for wireless device connectivity; two projects offer high-tech help to the needy; new electronic paper can display moving images. Computer, 36(12):24– 26, December 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/12/
rz024.pdf.

Thomson:2003:CPM

[1068] Judi Thomson, Dan Adams, Paula J.

Cowley, and Kevin Walker. Computing practices: Metadata’s role in a scientific archive. Computer, 36 (12):27–34, December 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/12/rz027.pdf;   http://www.
computer.org/computer/co2003/rz027abs.
htm.

Stan:2003:GEI

[1069] Mircea R. Stan and Kevin Skadron.

Guest Editors’ introduction: Poweraware computing. Computer, 36(12): 35–38, December 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/computer/co2003/rz035.
pdf;   http://www.computer.org/
computer/co2003/rz035abs.htm.

Lefurgy:2003:CFE

[1070] Charles Lefurgy, Karthick Rajamani, Freeman Rawson, Wes Felter, Michael Kistler, and Tom W. Keller. Cover feature: Energy management for commercial servers. Computer, 36(12): 39–48, December 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/12/rz039.pdf;   http://www.
computer.org/computer/co2003/rz039abs.
htm.

Albonesi:2003:CFD

[1071] David H. Albonesi, Rajeev Balasubramonian, Steven G. Dropsho, Sandhya Dwarkadas, Eby G. Friedman, Michael C. Huang, Volkan Kursun, Grigorios Magklis, Michael L. Scott, Greg Semeraro, Pradip Bose, Alper Buyuktosunoglu, Peter W. Cook, and Stanley E. Schuster. Cover feature: Dynamically tuning processor resources with adaptive processing. Computer, 36(12):49–58, December 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //csdl.computer.org/dl/mags/co/
                                        2003/12/rz049.pdf;   http://www.
                                        computer.org/computer/co2003/rz049abs.
                                        htm.

Gurumurthi:2003:CFR

[1072] Sudhanva Gurumurthi, Anand Sivasubramaniam, Mahmut Kandemir, and Hubertus Franke. Cover feature: Reducing disk power consumption in servers with DRPM. Computer, 36 (12):59–66, December 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //csdl.computer.org/dl/mags/co/
                                        2003/12/rz059.pdf;   http://www.
                                        computer.org/computer/co2003/rz059abs.
                                        htm.

Kim:2003:CFL

[1073] Nam Sung Kim, Todd Austin, David Blaauw, Trevor Mudge, Krisztián Flautner, Jie S. Hu, Mary Jane Irwin, Mahmut Kandemir, and Vijaykrishnan Narayanan. Cover feature: Leakage current: Moore’s Law meets static power. Computer, 36 (12):68–75, December 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/12/rz068.pdf;   http://www.
computer.org/computer/co2003/rz068abs.
htm.

Rao:2003:CFB

[1074] Ravishankar Rao, Sarma Vrudhula, and Daler N. Rakhmatov. Cover feature: Battery modeling for energyaware system design. Computer, 36 (12):77–87, December 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/12/rz077.pdf;   http://www.
computer.org/computer/co2003/rz077abs.
htm.

Anonymous:2003:AIC

[1075] Anonymous. Annual index, Computer, volume 36, 2003. Computer, 36(12): 88–99, December 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

computer.org/computer/co2003/rz088.
pdf.

Ward:2003:RME

[1076] Bob Ward. Report to members: Engel voted IEEE Computer Society President-Elect. Computer, 36(12): 100–101, December 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/12/
rz100.pdf.

Carver:2003:CSC

[1077] Doris Carver. Computer Society connection: Reviewers play key role in publishing Computer; Carver Mead receives presidential honors [US National Medal of Technology]; Robert Allen Short (1927–2003). Computer, 36(12): 102–104, December 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/12/
rz102.pdf.

Anonymous:2003:CCl

[1078] Anonymous. Call and calendar. Computer, 36(12):105–107, December 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/computer/co2003/rz105.pdf.

Anonymous:2003:PBS

[1079] Anonymous. Products: BEA Systems releases preview of Java API for XML; Panda appliance defends network perimeters; Parasoft enhances Java test and analysis tool. Computer, 36(12):108, December 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/12/rz108.pdf.

Lutz:2003:BAA

[1080] Michael J. Lutz. Bookshelf: Agile approach to database management [Agile Database Techniques: Effective Strategies for the Agile Software Developer]; appealing to women gamers [Gender Inclusive Game Design: Expanding the Market]; computing organically [Molecular Computing]; ins and outs of outsourcing [Successful IT Outsourcing]. Computer, 36(12): 109, December 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2003/12/
rz109.pdf.

Anonymous:2003:COk

[1081] Anonymous. Career opportunities.

Computer, 36(12):110–128, December 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.computer.
org/computer/co2003/rz110.pdf.

Geer:2003:SRM

[1082] Dan Geer. Security: Risk management is still where the money is. Computer, 36(12):129–131, December 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/12/rz129.pdf;   http://www.
computer.org/computer/co2003/rz129abs.
htm.

Fishwick:2003:ECN

[1083] Paul Fishwick. Entertainment computing: Nurturing next-generation computer scientists. Computer, 36(12): 132–134, December 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2003/12/rz132.pdf;   http://www.
computer.org/computer/co2003/rz132abs.
htm.

Schilit:2003:ICW

[1084] Bill Schilit, Jason Hong, and Marco Gruteser. Invisible computing: Wireless location privacy protection. Computer, 36(12):135–137, December 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //csdl.computer.org/dl/mags/co/
                                        2003/12/rz135.pdf;   http://www.
                                        computer.org/computer/co2003/rz135abs.
                                        htm.

Varshney:2003:CPH

[1085] Upkar Varshney. Communications: Pervasive healthcare. Computer, 36 (12):138–140, December 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //csdl.computer.org/dl/mags/co/
                                        2003/12/rz138.pdf;   http://www.
                                        computer.org/computer/co2003/rz138abs.
                                        htm.

Holmes:2003:PDD

[1086] Neville Holmes. The profession: The digital divide, the UN, and the computing profession. Computer, 36(12): 144, 142–143, December 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //csdl.computer.org/dl/mags/co/
                                        2003/12/rz144.pdf;   http://www.
                                        computer.org/computer/co2003/rz144abs.
                                        htm.

Anonymous:2004:ASa

[1087] Anonymous. Article summaries. Computer, 37(1):4, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/01/r1004.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/01/r1004.htm.

Welch:2004:LPA

[1088] Gerad Welch, Neville Holmes, Jan Giezen, Neville Holmes, and Thomas Robertazzi. Letters: The problem with AI; defining memory; divisible load-scheduling discovery. Computer, 37(1):6–7, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/01/r1006.pdf;
http://csdl.computer.org/dl/mags/
co/2004/01/r1006.htm.

Carver:2004:EMC

[1089] Doris L. Carver. EIC’s message: Computer in 2004. Computer, 37(1):8–9, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/01/r1008.pdf;
http://csdl.computer.org/dl/mags/
co/2004/01/r1008.htm.

Chang:2004:PME

[1090] Carl K. Chang. President’s message: Evolving the world with the world’s Computer Society. Computer, 37(1):10–12, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/01/r1010.pdf;
http://csdl.computer.org/dl/mags/
co/2004/01/r1010.htm.

Colwell:2004:RDF

[1091] Bob Colwell. At random: Design fragility. Computer, 37(1):13– 16, January 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
01/r1013abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/01/
r1013.htm;   http://csdl.computer.
org/dl/mags/co/2004/01/r1013.pdf.

Holmes:2004:YA

[1092] Neville Holmes. 32 and 16 years ago. Computer, 37(1):17, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/01/r1017.htm;
http://csdl.computer.org/dl/mags/
co/2004/01/r1017.pdf.

Vaughan-Nichols:2004:ITV

[1093] Steven J. Vaughan-Nichols. Industry trends: Vendors go to extreme lengths for new chips. Computer, 37(1):18–20, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/01/r1018.htm;
http://csdl.computer.org/dl/mags/
co/2004/01/r1018.pdf.

Lawton:2004:TNP

[1094] George Lawton. Technology news: Is peer-to-peer secure enough for corporate use? Computer, 37(1):22– 25, January 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/01/
r1022.htm;   http://csdl.computer.
org/dl/mags/co/2004/01/r1022.pdf.

Paulson:2004:NBM

[1095] Linda Dailey Paulson. News briefs: E-mail security gets personal; ceiling for controversial US H-1B visas drops; technique speeds data transmission in computers; viewing the world through interactive panoramic images; Web sites that offer panoramic images. Computer, 37(1):26–28, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
01/r1026abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/01/
r1026.htm;   http://csdl.computer.
org/dl/mags/co/2004/01/r1026.pdf.

Poore:2004:PTT

[1096] Jesse H. Poore. Perspectives: A tale of three disciplines… and a revolution. Computer, 37(1):30–36, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
01/r1030abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/01/
r1030.htm;   http://csdl.computer.
org/dl/mags/co/2004/01/r1030.pdf.

Horgan:2004:PES

[1097] John Horgan. Perspectives: The end of science revisited. Computer, 37(1): 37–43, January 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
01/r1037abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/01/
r1037.htm;   http://csdl.computer.
org/dl/mags/co/2004/01/r1037.pdf.

Edenfeld:2004:TRS

[1098] Don Edenfeld, Andrew B. Kahng, Mike Rodgers, and Yervant Zorian. 2003 technology roadmap for semiconductors. Computer, 37(1):47– 56, January 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/

                                        01/r1047abs.htm;   http://csdl.
                                        computer.org/dl/mags/co/2004/01/
                                        r1047.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/01/r1047.pdf.

Regan:2004:NMR

[1099] Patrick Regan and Scott Hamilton. NASA’s mission reliable. Computer, 37 (1):59–68, January 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        01/r1059abs.htm;   http://csdl.
                                        computer.org/dl/mags/co/2004/01/
                                        r1059.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/01/r1059.pdf.

Niederman:2004:IEP

[1100] Fred Niederman. IT employment prospects in 2004: A mixed bag. Computer, 37(1):69–77, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        01/r1069abs.htm;   http://csdl.
                                        computer.org/dl/mags/co/2004/01/
                                        r1069.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/01/r1069.pdf.

Pentland:2004:DRC

[1101] Alex (Sandy) Pentland, Richard Fletcher, and Amir Hasson. DakNet: Rethinking connectivity in developing nations. Computer, 37(1):78– 83, January 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        01/r1078abs.htm;   http://csdl.
                                        computer.org/dl/mags/co/2004/01/
                                        r1078.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/01/r1078.pdf.

Bimber:2004:COH

[1102] Oliver Bimber. Combining optical holograms with interactive computer graphics. Computer, 37(1): 85–91, January 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
01/r1085abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/01/
r1085.htm;   http://csdl.computer.
org/dl/mags/co/2004/01/r1085.pdf.

Anonymous:2004:CCa

[1103] Anonymous. Call and calendar. Computer, 37(1):92–94, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/01/r1092.pdf;
http://csdl.computer.org/dl/mags/
co/2004/01/r1092.htm.

Ward:2004:CSC

[1104] Bob Ward. Computer Society connection: Paving the wave for the future: New programs for 2004. Computer, 37(1):95–99, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/01/r1095.htm;
http://csdl.computer.org/dl/mags/
co/2004/01/r1095.pdf.

Anonymous:2004:COa

[1105] Anonymous. Career opportunities.

Computer, 37(1):100–112, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/01/r1100.
pdf.

Anonymous:2004:PIU

[1106] Anonymous. Products: IMSI updates popular CAD tool; immersive data visualization for PC users; Zero G releases InstallAnywhere 6; Code-Futures updates database persistence tool; new streaming-media system from Xiran; Parasoft offers software for Web services security; ClearMail 2.0 now available; 10x Software’s open source IDE. Computer, 37(1):114, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/01/r1114.htm;
http://csdl.computer.org/dl/mags/
co/2004/01/r1114.pdf.

Lutz:2004:BIV

[1107] Michael J. Lutz. Bookshelf: Implementing verification biometrics [Practical Biometrics: From Aspiration to Implementation]; from adventure to zork [Twisty Little Passages]; architecting enterprises [A Practical Guide to Enterprise Architecture]; compiler writer’s handbook [Engineering a Compiler]. Computer, 37(1): 115, January 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
01/r1115abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/01/
r1115.htm;   http://csdl.computer.
org/dl/mags/co/2004/01/r1115.pdf.

Ghanem:2004:WTD

[1108] Thanaa M. Ghanem and Walid G. Aref. Web technologies: Databases deepen the Web. Computer, 37(1): 116–117, January 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
01/r1116abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/01/
r1116.htm;   http://csdl.computer.
org/dl/mags/co/2004/01/r1116.pdf.

Andrews:2004:ECP

[1109] David Andrews, Douglas Niehaus, and Peter Ashenden. Embedded computing: Programming models for hybrid CPU/FPGA chips. Computer, 37(1): 118–120, January 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
01/r1118abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/01/
r1118.htm;   http://csdl.computer.
org/dl/mags/co/2004/01/r1118.pdf.

Alavi:2004:ISP

[1110] Maryam Alavi. IT systems perspectives: Distributed learning environments. Computer, 37(1):121– 122, January 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
01/r1121abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/01/
r1121.htm;   http://csdl.computer.
org/dl/mags/co/2004/01/r1121.pdf.

LoBue:2004:SBB

[1111] Michael LoBue. Standards: Best business practices for standards groups. Computer, 37(1):124–127, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
01/r1124abs.htm;   http://csdl.

                                        computer.org/dl/mags/co/2004/01/
                                        r1124.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/01/r1124.pdf.

Babu:2004:POM

[1112] Mohan Babu. The profession: From organization man to free agent. Computer, 37(1):128, 126–127, January 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        01/r1128abs.htm;   http://csdl.
                                        computer.org/dl/mags/co/2004/01/
                                        r1128.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/01/r1128.pdf.

Anonymous:2004:ASb

[1113] Anonymous. Article summaries. Computer, 37(2):4–??, February 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/02/r2004.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/02/r2004.htm.

Anonymous:2004:La

[1114] Anonymous.  Letters.  Com-

puter, 37(2):6–??, February 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/02/r2006.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/02/r2006.htm.

Colwell:2004:CLP

[1115] Bob Colwell. Computer lessons from pop songs. Computer, 37(2):7–??, February 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
02/r2007abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/02/
r2007.htm;   http://csdl.computer.
org/dl/mags/co/2004/02/r2007.pdf.

Vaughan-Nichols:2004:NTR

[1116] Steven J. Vaughan-Nichols. New trends revive supercomputing industry. Computer, 37(2):10–??, February 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/02/r2010.htm;
http://csdl.computer.org/dl/mags/
co/2004/02/r2010.pdf.

Geer:2004:WNS

[1117] David Geer. Will new standards help curb spam? Computer, 37(2):14–??, February 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/02/r2014.htm;
http://csdl.computer.org/dl/mags/
co/2004/02/r2014.pdf.

Paulson:2004:NBa

[1118] Linda Dailey Paulson. News briefs.

Computer, 37(2):17–??, February 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/02/r2017.htm;
http://csdl.computer.org/dl/mags/
co/2004/02/r2017.pdf.

Blaha:2004:CBS

[1119] Michael Blaha. A copper bullet for software quality improvement. Computer, 37(2):21–??, February 2004. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
02/r2021abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/02/
r2021.htm;   http://csdl.computer.
org/dl/mags/co/2004/02/r2021.pdf.

Wu:2004:GEI

[1120] Jie Wu and Ivan Stojmenovic. Guest Editors’ introduction: Ad hoc networks. Computer, 37(2):29–??, February 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/02/r2029.
pdf;   http://csdl.computer.org/
comp/mags/co/2004/02/r2029abs.htm;
 http://csdl.computer.org/dl/mags/
co/2004/02/r2029.htm.

Cao:2004:CCB

[1121] Guohong Cao, Liangzhong Yin, and Chita R. Das. Cooperative cachebased data access in ad hoc networks. Computer, 37(2):32–??, February 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
02/r2032abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/02/
r2032.htm;   http://csdl.computer.
org/dl/mags/co/2004/02/r2032.pdf.

Carle:2004:EEA

[1122] Jean Carle and David Simplot-Ryl. Energy-efficient area monitoring for sensor networks. Computer, 37(2): 40–??, February 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/

02/r2040abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/02/
r2040.htm;   http://csdl.computer.
org/dl/mags/co/2004/02/r2040.pdf.

Conti:2004:CLM

[1123] Marco Conti, Gaia Maselli, Giovanni Turi, and Silvia Giordano. Cross-layering in mobile ad hoc network design. Computer, 37(2):48– ??, February 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
02/r2048abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/02/
r2048.htm;   http://csdl.computer.
org/dl/mags/co/2004/02/r2048.pdf.

Mohapatra:2004:GCM

[1124] Prasant Mohapatra, Chao Gui, and Jian Li. Group communications in mobile ad hoc networks. Computer, 37(2): 52–??, February 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
02/r2052abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/02/
r2052.htm;   http://csdl.computer.
org/dl/mags/co/2004/02/r2052.pdf.

Milanovic:2004:RSM

[1125] Nikola Milanovic, Miroslaw Malek, Anthony Davidson, and Veljko Milutinovic. Routing and security in mobile ad hoc networks. Computer, 37(2): 61–??, February 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
02/r2061abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/02/

                                        r2061.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/02/r2061.pdf.

Xiao:2004:POM

[1126] Li Xiao, Abhishek Patil, Yunhao Liu, Lionel M. Ni, and Abdol-Hossein Esfahanian. Prioritized overlay multicast in mobile ad hoc environments. Computer, 37(2):67– ??, February 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        02/r2067abs.htm;   http://csdl.
                                        computer.org/dl/mags/co/2004/02/
                                        r2067.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/02/r2067.pdf.

Anonymous:2004:CSCa

[1127] Anonymous. Computer Society connection. Computer, 37(2):75–??, February 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2004/02/r2075.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2004/02/r2075.pdf.

Anonymous:2004:CCb

[1128] Anonymous. Call and calendar. Computer, 37(2):80–??, February 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2004/02/r2080.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2004/02/r2080.htm.

Anonymous:2004:C

[1129] Anonymous. Careers. Computer, 37 (2):84–??, February 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2004/02/r2084.pdf.

Hazas:2004:LAC

[1130] Mike Hazas, James Scott, and John Krumm. Location-aware computing comes of age. Computer, 37(2):95– ??, February 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
02/r2095abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/02/
r2095.htm;   http://csdl.computer.
org/dl/mags/co/2004/02/r2095.pdf.

Plowman:2004:IIS

[1131] Lydia Plowman and Rosemary Luckin.

Interactivity, interfaces, and smart toys. Computer, 37(2):98–??, February 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
02/r2098abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/02/
r2098.htm;   http://csdl.computer.
org/dl/mags/co/2004/02/r2098.pdf.

Gutmann:2004:SPK

[1132] Peter Gutmann. Simplifying public key management. Computer, 37(2): 101–??, February 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
02/r2101abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/02/
r2101.htm;   http://csdl.computer.
org/dl/mags/co/2004/02/r2101.pdf.

Vasiu:2004:MAW

[1133] Luminita Vasiu and Qusay H. Mahmoud. Mobile agents in wireless devices. Computer, 37(2):104–??, February 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        02/r2104abs.htm;   http://csdl.
                                        computer.org/dl/mags/co/2004/02/
                                        r2104.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/02/r2104.pdf.

Mahmoud:2004:MCP

[1134] Qusay H. Mahmoud, Wlodek Dobosiewicz, and David Swayne. Making computer programming fun and accessible. Computer, 37(2):108– ??, February 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        02/r2108abs.htm;   http://csdl.
                                        computer.org/dl/mags/co/2004/02/
                                        r2108.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/02/r2108.pdf.

Anonymous:2004:ASc

[1135] Anonymous. Article summaries. Computer, 37(3):4–??, March 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/03/r3004.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/03/r3004.htm.

Anonymous:2004:Lb

[1136] Anonymous. Letters. Computer, 37 (3):6–??, March 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/

03/r3006.pdf;   http://csdl.computer.
org/dl/mags/co/2004/03/r3006.htm.

Anonymous:2004:YAa

[1137] Anonymous. 32 and 16 years ago. Computer, 37(3):8–??, March 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/03/r3008.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/03/r3008.htm.

Colwell:2004:ZO

[1138] Bob Colwell. The zen of overclocking. Computer, 37(3):9–??, March 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
03/r3009abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/03/
r3009.htm;   http://csdl.computer.
org/dl/mags/co/2004/03/r3009.pdf.

Vaughan-Nichols:2004:VAS

[1139] Steven J. Vaughan-Nichols. Voice authentication speaks to the marketplace. Computer, 37(3):13–??, March 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/03/r3013.htm;
http://csdl.computer.org/dl/mags/
co/2004/03/r3013.pdf.

Leavitt:2004:WIP

[1140] Neal Leavitt. Will interoperability problems give IP telephony a busy signal? Computer, 37(3):16–??, March 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2004/03/r3016.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2004/03/r3016.pdf.

Paulson:2004:NBb

[1141] Linda Dailey Paulson. News briefs. Computer, 37(3):19–??, March 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2004/03/r3019.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2004/03/r3019.pdf.

Burger:2004:BTA

[1142] Doug Burger and James R. Goodman. Billion-transistor architectures: There and back again. Computer, 37 (3):22–??, March 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        03/r3022abs.htm;   http://csdl.
                                        computer.org/dl/mags/co/2004/03/
                                        r3022.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/03/r3022.pdf.

Maurer:2004:MPU

[1143] Peter M. Maurer. Metamorphic programming: Unconventional high. Computer, 37(3):30–??, March 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/03/r3030abs.htm;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/03/r3030.htm;   http://
                                        csdl.computer.org/dl/mags/co/2004/
                                        03/r3030.pdf.

Colwell:2004:WMN

[1144] Bob Colwell. We may need a new box. Computer, 37(3):40–??, March 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/03/r3040.pdf;
 http://csdl.computer.org/comp/
mags/co/2004/03/r3040abs.htm;
http://csdl.computer.org/dl/mags/
co/2004/03/r3040.htm.

Shanbhag:2004:RES

[1145] Naresh R. Shanbhag. Reliable and efficient system-on-chip design. Computer, 37(3):42–??, March 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
03/r3042abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/03/
r3042.htm;   http://csdl.computer.
org/dl/mags/co/2004/03/r3042.pdf.

Uht:2004:GBW

[1146] Augustus K. Uht. Going beyond worstcase specs with TEAtime. Computer, 37(3):51–??, March 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
03/r3051abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/03/
r3051.htm;   http://csdl.computer.
org/dl/mags/co/2004/03/r3051.pdf.

Austin:2004:MTS

[1147] Todd Austin, David Blaauw, Trevor Mudge, and Krisztián Flautner. Making typical silicon matter with razor. Computer, 37(3):57–??, March 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        03/r3057abs.htm;   http://csdl.
                                        computer.org/dl/mags/co/2004/03/
                                        r3057.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/03/r3057.pdf.

Anonymous:2004:SPA

[1148] Anonymous. Speeding up processing with approximation circuits. Computer, 37(3):67–??, March 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/03/r3067abs.htm;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/03/r3067.htm;   http://
                                        csdl.computer.org/dl/mags/co/2004/
                                        03/r3067.pdf.

Anonymous:2004:CSCb

[1149] Anonymous. Computer Society connection. Computer, 37(3):77–??, March 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2004/03/r3077.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2004/03/r3077.htm.

Anonymous:2004:CCc

[1150] Anonymous. Call and calendar. Computer, 37(3):81–??, March 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/03/r3081.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/03/r3081.htm.

Anonymous:2004:COb

[1151] Anonymous. Career opportunities. Computer, 37(3):84–??, March 2004.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/03/r3084.pdf.

Anonymous:2004:Pa

[1152] Anonymous.  Products.  Com-

puter, 37(3):91–??, March 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/03/r3091.pdf;
http://csdl.computer.org/dl/mags/
co/2004/03/r3091.htm.

Anonymous:2004:Ba

[1153] Anonymous.  Bookshelf.  Com-

puter, 37(3):92–??, March 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/03/r3092.pdf;
http://csdl.computer.org/dl/mags/
co/2004/03/r3092.htm.

Richardson:2004:MSD

[1154] Gary Richardson and Blake Ives. Managing systems development. Computer, 37(3):93–??, March 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
03/r3093abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/03/
r3093.htm;   http://csdl.computer.
org/dl/mags/co/2004/03/r3093.pdf.

Zambonelli:2004:AHS

[1155] Franco Zambonelli and Michael Luck.

Agent hell: A scenario of worst practices. Computer, 37(3):96–??, March 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/03/r3096.pdf;
 http://csdl.computer.org/comp/
mags/co/2004/03/r3096abs.htm;
http://csdl.computer.org/dl/mags/
co/2004/03/r3096.htm.

Wolf:2004:ENP

[1156] Wayne Wolf. Embedded is the new paradigm(s). Computer, 37(3): 99–??, March 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
03/r3099abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/03/
r3099.htm;   http://csdl.computer.
org/dl/mags/co/2004/03/r3099.pdf.

Holmes:2004:LCP

[1157] Neville Holmes. Languages and the computing profession. Computer, 37 (3):104–??, March 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
03/r3104abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/03/
r3104.htm;   http://csdl.computer.
org/dl/mags/co/2004/03/r3104.pdf.

Anonymous:2004:ASd

[1158] Anonymous. Article summaries. Computer, 37(4):4–??, April 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/04/r4004.pdf;
 http://csdl.computer.org/dl/mags/
co/2004/04/r4004.htm.

Anonymous:2004:Lc

[1159] Anonymous. Letters. Computer, 37 (4):6–??, April 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
04/r4006.pdf;   http://csdl.computer.
org/dl/mags/co/2004/04/r4006.htm.

Colwell:2004:BII

[1160] Bob Colwell. Brainstorming, influence, and icebergs. Computer, 37 (4):9–??, April 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
04/r4009abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/04/
r4009.htm;   http://csdl.computer.
org/dl/mags/co/2004/04/r4009.pdf.

Lawton:2004:WNP

[1161] George Lawton. Will network processor units live up to their promise? Computer, 37(4):13–??, April 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/04/r4013.htm;
http://csdl.computer.org/dl/mags/
co/2004/04/r4013.pdf.

Paulson:2004:NBc

[1162] Linda Dailey Paulson. News briefs.

Computer, 37(4):16–??, April 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/04/r4016.htm;
http://csdl.computer.org/dl/mags/
co/2004/04/r4016.pdf.

Sanden:2004:CJT

[1163] Bo Sandén. Coping with Java threads. Computer, 37(4):20–??, April 2004. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
04/r4020abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/04/
r4020.htm;   http://csdl.computer.
org/dl/mags/co/2004/04/r4020.pdf.

Blundo:2004:SIA

[1164] Carlo Blundo and Stelvio Cimato. A software infrastructure for authenticated Web metering. Computer, 37 (4):28–??, April 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
04/r4028abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/04/
r4028.htm;   http://csdl.computer.
org/dl/mags/co/2004/04/r4028.pdf.

Rangarajan:2004:ANN

[1165] Santosh K. Rangarajan, Vir V. Phoha, Kiran S. Balagani, Rastko R. Selmic, and S. S. Iyengar. Adaptive neural network clustering of Web users. Computer, 37(4):34– ??, April 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
04/r4034abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/04/
r4034.htm;   http://csdl.computer.
org/dl/mags/co/2004/04/r4034.pdf.

Bhatti:2004:XBS

[1166] Rafae Bhatti, Elisa Bertino, Arif Ghafoor, and James B. D. Joshi. XMLbased specification for Web services document security. Computer, 37 (4):41–??, April 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
04/r4041abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/04/
r4041.htm;   http://csdl.computer.
org/dl/mags/co/2004/04/r4041.pdf.

Chen:2004:CDM

[1167] Hsinchun Chen, Wingyan Chung, Jennifer Jie Xu, Gang Wang, Yi Qin, and Michael Chau. Crime data mining: A general framework and some examples. Computer, 37(4): 50–??, April 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
04/r4050abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/04/
r4050.htm;   http://csdl.computer.
org/dl/mags/co/2004/04/r4050.pdf.

Megiddo:2004:OLA

[1168] Nimrod Megiddo and Dharmendra S.

Modha. Outperforming LRU with an adaptive replacement cache algorithm. Computer, 37(4):58–??, April 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
04/r4058abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/04/
r4058.htm;   http://csdl.computer.
org/dl/mags/co/2004/04/r4058.pdf.

Anonymous:2004:CSCc

[1169] Anonymous. Computer Society connection. Computer, 37(4):69–??, April 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/04/r4069.htm;

                                        http://csdl.computer.org/dl/mags/
                                        co/2004/04/r4069.pdf.

Anonymous:2004:CCd

[1170] Anonymous. Call and calendar. Computer, 37(4):71–??, April 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/04/r4071.pdf.

Anonymous:2004:Bb

[1171] Anonymous. Bookshelf. Computer, 37 (4):73–??, April 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        04/r4073.pdf;   http://csdl.computer.
                                        org/dl/mags/co/2004/04/r4073.htm.

Anonymous:2004:COc

[1172] Anonymous. Career opportunities. Computer, 37(4):74–??, April 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/04/r4074.pdf.

Anonymous:2004:Pb

[1173] Anonymous. Products. Computer, 37 (4):77–??, April 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        04/r4077.pdf;   http://csdl.computer.
                                        org/dl/mags/co/2004/04/r4077.htm.

Chakraborty:2004:ERB

[1174] Dipanjan Chakraborty and Hui Lei. Extending the reach of business processes. Computer, 37(4):78–??, April 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
04/r4078abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/04/
r4078.htm;   http://csdl.computer.
org/dl/mags/co/2004/04/r4078.pdf.

Overmars:2004:TCS

[1175] Mark Overmars. Teaching computer science through game design. Computer, 37(4):81–??, April 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/04/r4081abs.htm;
 http://csdl.computer.org/dl/mags/
co/2004/04/r4081.htm;   http://
csdl.computer.org/dl/mags/co/2004/
04/r4081.pdf.

Want:2004:EUS

[1176] Roy Want. Enabling ubiquitous sensing with RFID. Computer, 37(4): 84–??, April 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
04/r4084abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/04/
r4084.htm;   http://csdl.computer.
org/dl/mags/co/2004/04/r4084.pdf.

Schneier:2004:HBC

[1177] Bruce Schneier. Hacking the business climate for network security. Computer, 37(4):87–??, April 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
04/r4087abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/04/
r4087.htm;   http://csdl.computer.
org/dl/mags/co/2004/04/r4087.pdf.

Johnson:2004:BTC

[1178] Margaret L. Johnson. Biometrics and the threat to civil liberties. Computer, 37(4):92–??, April 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        04/r4092abs.htm;   http://csdl.
                                        computer.org/dl/mags/co/2004/04/
                                        r4092.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/04/r4092.pdf.

Anonymous:2004:ASe

[1179] Anonymous. Article summaries. Computer, 37(5):4–??, May 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/05/r5004.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/05/r5004.htm.

Anonymous:2004:Ld

[1180] Anonymous. Letters. Computer, 37 (5):6–??, May 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        05/r5006.pdf;   http://csdl.computer.
                                        org/dl/mags/co/2004/05/r5006.htm.

Colwell:2004:ESW

[1181] Bob Colwell. Engineering a sustainable world. Computer, 37(5):9– ??, May 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        05/r5009abs.htm;   http://csdl.
                                        computer.org/dl/mags/co/2004/05/
                                        r5009.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/05/r5009.pdf.

Anonymous:2004:YAb

[1182] Anonymous. 32 and 16 years ago. Computer, 37(5):13–??, May 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/05/r5013.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/05/r5013.htm.

Paulson:2004:UMU

[1183] Linda Dailey Paulson. Using MPLS to unify multiple network types. Computer, 37(5):15–??, May 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/05/r5015.htm;
http://csdl.computer.org/dl/mags/
co/2004/05/r5015.pdf.

Vaughan-Nichols:2004:WMG

[1184] Steven J. Vaughan-Nichols. Wireless middleware: Glue for the mobile infrastructure. Computer, 37(5):18–??, May 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/05/r5018.htm;
http://csdl.computer.org/dl/mags/
co/2004/05/r5018.pdf.

Anonymous:2004:NBa

[1185] Anonymous. News briefs. Computer, 37(5):22–??, May 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/05/
r5022.htm;   http://csdl.computer.
org/dl/mags/co/2004/05/r5022.pdf.

Finkelstein:2004:CCS

[1186] Anthony Finkelstein, James Hetherington, Linzhong Li, Ofer Margoninski, Peter Saffrey, Rob Seymour, and Anne Warner. Computational challenges of systems biology. Computer, 37(5):26–??, May 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
05/r5026abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/05/
r5026.htm;   http://csdl.computer.
org/dl/mags/co/2004/05/r5026.pdf.

Dishman:2004:IWS

[1187] Eric Dishman. Inventing wellness systems for aging in place. Computer, 37(5):34–??, May 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
05/r5034abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/05/
r5034.htm;   http://csdl.computer.
org/dl/mags/co/2004/05/r5034.pdf.

Pentland:2004:HMT

[1188] Alex (Sandy) Pentland. Healthwear: Medical technology becomes wearable. Computer, 37(5):42–??, May 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
05/r5042abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/05/
r5042.htm;   http://csdl.computer.
org/dl/mags/co/2004/05/r5042.pdf.

Kara:2004:PPC

[1189] Atsushi Kara. Private-to-private communications over the Internet. Computer, 37(5):53–??, May 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2004/05/r5053abs.htm;   http://
csdl.computer.org/dl/mags/co/2004/
05/r5053.htm;   http://csdl.computer.
org/dl/mags/co/2004/05/r5053.pdf.

Eugster:2004:EID

[1190] Patrick T. Eugster, Rachid Guerraoui, Anne-Marie Kermarrec, and Laurent Massoulié. Epidemic information dissemination in distributed systems. Computer, 37(5):60–??, May 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
05/r5060abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/05/
r5060.htm;   http://csdl.computer.
org/dl/mags/co/2004/05/r5060.pdf.

Anonymous:2004:CSCd

[1191] Anonymous. Computer Society connection. Computer, 37(5):68–??, May 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/05/r5068.htm;
http://csdl.computer.org/dl/mags/
co/2004/05/r5068.pdf.

Anonymous:2004:CCe

[1192] Anonymous. Call and calendar.

Computer, 37(5):72–??, May 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/05/r5072.pdf;
 http://csdl.computer.org/dl/mags/
co/2004/05/r5072.htm.

Anonymous:2004:COd

[1193] Anonymous. Career opportunities. Computer, 37(5):74–??, May 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/05/r5074.pdf.

Anonymous:2004:Bc

[1194] Anonymous. Bookshelf. Computer, 37(5):80–??, May 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        05/r5080.pdf;   http://csdl.computer.
                                        org/dl/mags/co/2004/05/r5080.htm.

Austin:2004:MS

[1195] Todd Austin, David Blaauw, Scott Mahlke, Trevor Mudge, Chaitali Chakrabarti, and Wayne Wolf. Mobile supercomputers. Computer, 37 (5):81–??, May 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        05/r5081abs.htm;   http://csdl.
                                        computer.org/dl/mags/co/2004/05/
                                        r5081.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/05/r5081.pdf.

Richardson:2004:SDP

[1196] Gary Richardson and Blake Ives. Systems development processes. Computer, 37(5):84–??, May 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2004/05/r5084abs.htm;   http://
                                        csdl.computer.org/dl/mags/co/2004/
                                        05/r5084.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/05/r5084.pdf.

Cellary:2004:VWS

[1197] Wojciech Cellary, Wojciech Wiza, and Krzysztof Walczak. Visualizing Web search results in 3D. Computer, 37 (5):87–??, May 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
05/r5087abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/05/
r5087.htm;   http://csdl.computer.
org/dl/mags/co/2004/05/r5087.pdf.

Holmes:2004:DIP

[1198] Neville Holmes. Data and information as property. Computer, 37 (5):92–??, May 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
05/r5092abs.htm;   http://csdl.
computer.org/dl/mags/co/2004/05/
r5092.htm;   http://csdl.computer.
org/dl/mags/co/2004/05/r5092.pdf.

Anonymous:2004:ASf

[1199] Anonymous. Article summaries.

Computer, 37(6):4–??, June 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/06/r6004.pdf;
 http://csdl.computer.org/dl/mags/
co/2004/06/r6004.htm.

Colwell:2004:LHB

[1200] Bob Colwell. Let’s hear from those in the back. Computer, 37(6):6–??, June 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/06/r6006.htm;

                                        http://csdl.computer.org/dl/mags/
                                        co/2004/06/r6006.pdf.

Vaughan-Nichols:2004:AWB

[1201] Steven J. Vaughan-Nichols. Achieving wireless broadband with WiMax. Computer, 37(6):10–??, June 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2004/06/r6010.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2004/06/r6010.pdf.

Geer:2004:JHS

[1202] David Geer. Just how secure are security products? Computer, 37 (6):14–??, June 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2004/06/
                                        r6014.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/06/r6014.pdf.

Paulson:2004:NBd

[1203] Linda Dailey Paulson. News briefs. Computer, 37(6):18–??, June 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2004/06/r6018.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2004/06/r6018.pdf.

Strong:2004:RES

[1204] Diane M. Strong and Olga Volkoff. A roadmap for enterprise system implementation. Computer, 37(6):22– ??, June 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2004/06/

r6022.htm;   http://csdl.computer.
org/dl/mags/co/2004/06/r6022.pdf.

Shim:2004:GEI

[1205] Simon S. Y. Shim, Li Gong, Aviel D.

Rubin, and Linley Gwennap. Guest Editors’ introduction: Securing the high-speed Internet. Computer, 37 (6):33–??, June 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
06/r6033.pdf;   http://csdl.computer.
org/dl/mags/co/2004/06/r6033.htm.

Lampson:2004:CSR

[1206] Butler W. Lampson. Computer security in the real world. Computer, 37 (6):37–??, June 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/06/
r6037.htm;   http://csdl.computer.
org/dl/mags/co/2004/06/r6037.pdf.

Chen:2004:WEH

[1207] Thomas M. Chen and Jean-Marc Robert. Worm epidemics in high-speed networks. Computer, 37(6):48–??, June 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/06/r6048.htm;
http://csdl.computer.org/dl/mags/
co/2004/06/r6048.pdf.

Friend:2004:MGI

[1208] Robert Friend. Making the gigabit IPsec VPN architecture secure. Computer, 37(6):54–??, June 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2004/06/r6054.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2004/06/r6054.pdf.

Wool:2004:QSF

[1209] Avishai Wool. A quantitative study of firewall configuration errors. Computer, 37(6):62–??, June 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2004/06/r6062.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2004/06/r6062.pdf.

Anonymous:2004:CSCe

[1210] Anonymous. Computer Society connection. Computer, 37(6):68–??, June 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2004/06/r6068.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2004/06/r6068.htm.

Anonymous:2004:CCf

[1211] Anonymous. Call and calendar. Computer, 37(6):71–??, June 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/06/r6071.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/06/r6071.htm.

Anonymous:2004:Pc

[1212] Anonymous. Products. Computer, 37 (6):73–??, June 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2004/06/

r6073.htm;   http://csdl.computer.
org/dl/mags/co/2004/06/r6073.pdf.

Anonymous:2004:Bd

[1213] Anonymous. Bookshelf. Computer, 37 (6):74–??, June 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
06/r6074.pdf;   http://csdl.computer.
org/dl/mags/co/2004/06/r6074.htm.

Anonymous:2004:COe

[1214] Anonymous. Career opportunities.

Computer, 37(6):76–??, June 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/06/r6076.pdf.

Yi:2004:EPC

[1215] Byung K. Yi and S. R. Subramanya.

Enhancing personal communications with multimedia. Computer, 37(6): 79–??, June 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/06/
r6079.htm;   http://csdl.computer.
org/dl/mags/co/2004/06/r6079.pdf.

Arbaugh:2004:PNS

[1216] William A. Arbaugh. A patch in nine saves time? Computer, 37 (6):82–??, June 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/06/
r6082.htm;   http://csdl.computer.
org/dl/mags/co/2004/06/r6082.pdf.

Smith:2004:ISC

[1217] Joshua R. Smith. Imperceptible sensory channels. Computer, 37(6): 84–??, June 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/06/
r6084.htm;   http://csdl.computer.
org/dl/mags/co/2004/06/r6084.pdf.

Rizzo:2004:DAD

[1218] Albert A. Rizzo, Todd Bowerly, Cyrus Shahabi, J. Galen Buckwalter, Dean Klimchuk, and Roman Mitura. Diagnosing attention disorders in a virtual classroom. Computer, 37(6):87–??, June 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/06/r6087.htm;
http://csdl.computer.org/dl/mags/
co/2004/06/r6087.pdf.

Olsen:2004:EIC

[1219] Kai A. Olsen. The economics of international conferences. Computer, 37 (6):92–??, June 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/06/
r6092.htm;   http://csdl.computer.
org/dl/mags/co/2004/06/r6092.pdf.

Anonymous:2004:ASg

[1220] Anonymous. Article summaries. Computer, 37(7):4–??, July 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/07/r7004.pdf;
 http://csdl.computer.org/dl/mags/
co/2004/07/r7004.htm.

Anonymous:2004:Le

[1221] Anonymous. Letters. Computer, 37 (7):6–??, July 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2004/
07/r7006.pdf;   http://csdl.computer.
org/dl/mags/co/2004/07/r7006.htm.

Colwell:2004:TS

[1222] Bob Colwell. Travel stories. Computer, 37(7):8–??, July 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/07/
r7008.htm;   http://csdl.computer.
org/dl/mags/co/2004/07/r7008.pdf.

Anonymous:2004:YAc

[1223] Anonymous. 32 and 16 years ago. Computer, 37(7):11–??, July 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/07/r7011.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/07/r7011.htm.

Paulson:2004:OSD

[1224] Linda Dailey Paulson. Open source databases move into the marketplace. Computer, 37(7):13–??, July 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/07/r7013.htm;
http://csdl.computer.org/dl/mags/
co/2004/07/r7013.pdf.

Lawton:2004:MFW

[1225] George Lawton. Is MIMO the future of wireless communications?

Computer, 37(7):20–??, July 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/07/r7020.htm;
http://csdl.computer.org/dl/mags/
co/2004/07/r7020.pdf.

Paulson:2004:NBe

[1226] Linda Dailey Paulson. News briefs. Computer, 37(7):24–??, July 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/07/r7024.htm;
http://csdl.computer.org/dl/mags/
co/2004/07/r7024.pdf.

Mead:2004:WLI

[1227] Nancy R. Mead. Who is liable for insecure systems? Computer, 37 (7):27–??, July 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/07/
r7027.htm;   http://csdl.computer.
org/dl/mags/co/2004/07/r7027.pdf.

Jungck:2004:IHS

[1228] Peder Jungck and Simon S. Y. Shim. Issues in high-speed Internet security. Computer, 37(7):36–??, July 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/07/r7036.htm;
http://csdl.computer.org/dl/mags/
co/2004/07/r7036.pdf.

Burger:2004:SES

[1229] Doug Burger, Stephen W. Keckler, Kathryn S. McKinley, Mike Dahlin, Lizy K. John, Calvin Lin, Charles R.

Moore, James Burrill, Robert G. McDonald, William Yoder, and the TRIPS Team. Scaling to the end of silicon with EDGE architectures. Computer, 37(7):44–??, July 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/07/r7044.htm;
http://csdl.computer.org/dl/mags/
co/2004/07/r7044.pdf.

McKinley:2004:CAS

[1230] Philip K. McKinley, Seyed Masoud Sadjadi, Eric P. Kasten, and Betty H. C. Cheng. Composing adaptive software. Computer, 37(7):56–??, July 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/07/r7056.htm;
http://csdl.computer.org/dl/mags/
co/2004/07/r7056.pdf.

Kozuch:2004:SMC

[1231] Michael Kozuch, Mahadev Satyanarayanan, Thomas Bressoud, Casey Helfrich, and Shafeeq Sinnamohideen. Seamless mobile computing on fixed infrastructure. Computer, 37(7):65– ??, July 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/07/
r7065.htm;   http://csdl.computer.
org/dl/mags/co/2004/07/r7065.pdf.

Montanari:2004:PBD

[1232] Rebecca Montanari, Emil Lupu, and Cesare Stefanelli. Policy-based dynamic reconfiguration of mobile-code applications. Computer, 37(7):73– ??, July 2004. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2004/07/
                                        r7073.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/07/r7073.pdf.

Anonymous:2004:CCg

[1233] Anonymous. Call and calendar. Computer, 37(7):81–??, July 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/07/r7081.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/07/r7081.htm.

Anonymous:2004:Be

[1234] Anonymous. Bookshelf. Computer, 37(7):82–??, July 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        07/r7082.pdf;   http://csdl.computer.
                                        org/dl/mags/co/2004/07/r7082.htm.

Anonymous:2004:Pd

[1235] Anonymous. Products. Computer, 37 (7):83–??, July 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2004/
                                        07/r7083.pdf;   http://csdl.computer.
                                        org/dl/mags/co/2004/07/r7083.htm.

Anonymous:2004:CSCf

[1236] Anonymous. Computer Society connection. Computer, 37(7):84–??, July 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2004/07/r7084.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2004/07/r7084.pdf.

Anonymous:2004:COf

[1237] Anonymous. Career opportunities.

Computer, 37(7):88–??, July 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/07/r7088.pdf.

Fitzgerald:2004:CLO

[1238] Brian Fitzgerald. A critical look at open source. Computer, 37(7): 92–??, July 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/07/
r7092.htm;   http://csdl.computer.
org/dl/mags/co/2004/07/r7092.pdf.

Koopman:2004:ESS

[1239] Philip Koopman. Embedded system security. Computer, 37(7):95–??, July 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/07/r7095.htm;
http://csdl.computer.org/dl/mags/
co/2004/07/r7095.pdf.

Holmes:2004:DP

[1240] Neville Holmes. In defense of Power-Point. Computer, 37(7):100–??, July 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/07/r7100.htm;
http://csdl.computer.org/dl/mags/
co/2004/07/r7100.pdf.

Anonymous:2004:ASh

[1241] Anonymous. Article summaries.

Computer, 37(8):4–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/08/r8004.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/08/r8004.htm.

Anonymous:2004:Lf

[1242] Anonymous.  Letters.  Com-

puter, 37(8):6–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/08/r8006.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/08/r8006.htm.

Colwell:2004:AP

[1243] Bob Colwell. The art of the possible. Computer, 37(8):8–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2004/08/r8008.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2004/08/r8008.pdf.

Anonymous:2004:PUT

[1244] Anonymous. Poznan University team wins US$18,000 at CSIDC 2004. Computer, 37(8):11–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2004/08/r8011.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2004/08/r8011.htm.

Geer:2004:MSS

[1245] David Geer. E-micropayments sweat the small stuff. Computer, 37(8):19–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2004/08/r8019.

pdf;   http://csdl.computer.org/
dl/mags/co/2004/08/r8019.htm.

Ortiz:2004:SST

[1246] Sixto Ortiz Jr. 3D searching starts to take shape. Computer, 37(8):24–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/08/r8024.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/08/r8024.htm.

Paulson:2004:NBf

[1247] Linda Dailey Paulson. News briefs.

Computer, 37(8):28–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/08/r8028.pdf;
http://csdl.computer.org/dl/mags/
co/2004/08/r8028.htm.

Dybkjaer:2004:MCS

[1248] Hans Dybkjær and Laila Dybkjær.

Modeling complex spoken dialog. Computer, 37(8):32–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/08/r8032.htm;
http://csdl.computer.org/dl/mags/
co/2004/08/r8032.pdf.

Culler:2004:GEI

[1249] David Culler, Deborah Estrin, and Mani Srivastava. Guest Editors’ introduction: Overview of sensor networks. Computer, 37(8):41–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/08/r8041.

pdf;   http://csdl.computer.org/
dl/mags/co/2004/08/r8041.htm.

Martinez:2004:ESN

[1250] Kirk Martinez, Jane K. Hart, and Royan Ong. Environmental sensor networks. Computer, 37(8):50–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/08/r8050.htm;
http://csdl.computer.org/dl/mags/
co/2004/08/r8050.pdf.

Brennan:2004:RDD

[1251] Sean M. Brennan, Angela M. Mielke, David C. Torney, and Arthur B. Maccabe. Radiation detection with distributed sensor networks. Computer, 37(8):57–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/08/r8057.htm;
http://csdl.computer.org/dl/mags/
co/2004/08/r8057.pdf.

Maroti:2004:SLU

[1252] Miklos Maroti, Gyula Simon, Akos Ledeczi, and Janos Sztipanovits. Shooter localization in urban terrain. Computer, 37(8):60–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/08/r8060.htm;
http://csdl.computer.org/dl/mags/
co/2004/08/r8060.pdf.

Enz:2004:WUP

[1253] Christian C. Enz, Amre El-Hoiydi, Jean-Dominique Decotignie, and Vincent Peiris. WiseNET: An ultralowpower wireless sensor network solution. Computer, 37(8):62–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/08/r8062.htm;
http://csdl.computer.org/dl/mags/
co/2004/08/r8062.pdf.

Hemingway:2004:FMS

[1254] Bruce Hemingway, Waylon Brunette, Tom Anderl, and Gaetano Borriello. The flock: Mote sensors sing in undergraduate curriculum. Computer, 37(8):72–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/08/r8072.htm;
http://csdl.computer.org/dl/mags/
co/2004/08/r8072.pdf.

Anonymous:2004:CCh

[1255] Anonymous. Call and calendar.

Computer, 37(8):84–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/08/r8084.pdf;
 http://csdl.computer.org/dl/mags/
co/2004/08/r8084.htm.

Anonymous:2004:COg

[1256] Anonymous. Career opportunities.

Computer, 37(8):86–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/08/r8086.pdf.

Anonymous:2004:Bf

[1257] Anonymous.  Bookshelf.  Com-

puter, 37(8):88–??, August 2004.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/08/r8088.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/08/r8088.htm.

Anonymous:2004:Pe

[1258] Anonymous.  Products.  Com-

puter, 37(8):89–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/08/r8089.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/08/r8089.htm.

Christensen:2004:PMN

[1259] Ken Christensen, Bruce Nordman, and Rich Brown. Power management in networked devices. Computer, 37(8):91–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2004/08/r8091.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2004/08/r8091.pdf.

Keromytis:2004:QPD

[1260] Angelos D. Keromytis. “Patch on Demand” saves even more time? Computer, 37(8):94–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2004/08/r8094.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2004/08/r8094.pdf.

Clarke:2004:CAT

[1261] Siobhán Clarke and Cormac Driver. Context-aware trails. Computer, 37 (8):97–??, August 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/08/
r8097.htm;   http://csdl.computer.
org/dl/mags/co/2004/08/r8097.pdf.

Macedonia:2004:DWM

[1262] Michael Macedonia. The digital world’s midlife crisis. Computer, 37(8):100–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/08/r8100.htm;
http://csdl.computer.org/dl/mags/
co/2004/08/r8100.pdf.

Swayne:2004:QOR

[1263] David A. Swayne, Qusay H. Mahmoud, and Włodek Dobosiewicz. An “offshore-resistant” degree program. Computer, 37(8):104–??, August 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/08/r8104.htm;
http://csdl.computer.org/dl/mags/
co/2004/08/r8104.pdf.

Anonymous:2004:ASi

[1264] Anonymous. Article summaries.

Computer, 37(9):4–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/09/r9004.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/09/r9004.htm.

Colwell:2004:ES

[1265] Bob Colwell. Engineers as soothsayers. Computer, 37(9):6–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2004/09/r9006.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2004/09/r9006.htm.

Anonymous:2004:YAd

[1266] Anonymous. 32 and 16 years ago. Computer, 37(9):10–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2004/09/r9010.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2004/09/r9010.htm.

Lawton:2004:MMT

[1267] George Lawton. Machine-to-machine technology gears up for growth. Computer, 37(9):12–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2004/09/r9012.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2004/09/r9012.htm.

Leavitt:2004:SAS

[1268] Neal Leavitt. Scob attack: A sign of bad things to come? Computer, 37(9):16–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/09/r9016.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/09/r9016.htm.

Anonymous:2004:NBb

[1269] Anonymous. News briefs. Computer, 37(9):20–??, September 2004.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/09/r9020.pdf;
 http://csdl.computer.org/dl/mags/
co/2004/09/r9020.htm.

Kobylarz:2004:BCW

[1270] Thaddeus J. A. Kobylarz. Beyond 3G:

Compound wireless services. Computer, 37(9):23–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/09/r9023.htm;
http://csdl.computer.org/dl/mags/
co/2004/09/r9023.pdf.

Wang:2004:VDC

[1271] Runsheng Wang and Xiaorui Hu.

VoIP development in China. Computer, 37(9):30–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/09/r9030.htm;
http://csdl.computer.org/dl/mags/
co/2004/09/r9030.pdf.

Thakoor:2004:BEM

[1272] Sarita Thakoor, John Michael Morookian, Javaan Chahl, Butler Hine, and Steve Zornetzer. BEES: Exploring Mars with bioinspired technologies. Computer, 37(9):38–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/09/r9038.htm;
http://csdl.computer.org/dl/mags/
co/2004/09/r9038.pdf.

Paulson:2004:BR

[1273] Linda Dailey Paulson. Biomimetic robots. Computer, 37(9):48–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2004/09/
                                        r9048.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/09/r9048.pdf.

Avanzini:2004:DUS

[1274] Federico Avanzini, Davide Rocchesso, Alberto Belussi, Alessandro Dal Palù, and Agostino Dovier. Designing an urban-scale auditory alert system. Computer, 37(9):55–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2004/09/r9055.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2004/09/r9055.pdf.

Obrenovic:2004:MMH

[1275] Zeljko Obrenovic and Dusan Starcevic. Modeling multimodal human-computer interaction. Computer, 37(9):65– ??, September 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2004/09/
                                        r9065.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/09/r9065.pdf.

Anonymous:2004:COh

[1276] Anonymous. Career opportunities. Computer, 37(9):74–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2004/09/r9074.
                                        pdf.

Anonymous:2004:CSCg

[1277] Anonymous. Computer Society connection. Computer, 37(9):82–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/09/r9082.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/09/r9082.htm.

Anonymous:2004:Bg

[1278] Anonymous.  Bookshelf.  Com-

puter, 37(9):88–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/09/r9088.pdf;
http://csdl.computer.org/dl/mags/
co/2004/09/r9088.htm.

Anonymous:2004:Pf

[1279] Anonymous.  Products.  Com-

puter, 37(9):89–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/09/r9089.pdf;
http://csdl.computer.org/dl/mags/
co/2004/09/r9089.htm.

Anonymous:2004:CCi

[1280] Anonymous. Call and calendar.

Computer, 37(9):91–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/09/r9091.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/09/r9091.htm.

Hofstader:2004:IAB

[1281] Chris Hofstader. Internet accessibility: Beyond disability. Computer, 37(9):103–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/09/r9103.htm;
http://csdl.computer.org/dl/mags/
co/2004/09/r9103.pdf.

Hayden:2004:ESF

[1282] Tim Hayden. Empowering sports fans with technology. Computer, 37(9):106– ??, September 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/09/
r9106.htm;   http://csdl.computer.
org/dl/mags/co/2004/09/r9106.pdf.

Serpanos:2004:SHS

[1283] Dimitrios N. Serpanos. Speeding up high-speed protocol processors. Computer, 37(9):108–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/09/r9108.htm;
http://csdl.computer.org/dl/mags/
co/2004/09/r9108.pdf.

Cole:2004:RB

[1284] Jack Cole. Raising the bar. Computer, 37(9):112–??, September 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/09/r9112.htm;
http://csdl.computer.org/dl/mags/
co/2004/09/r9112.pdf.

Holmes:2004:RDT

[1285] Neville Holmes. Rationality and digital technology. Computer, 37(9):116– ??, September 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/09/
r9116.htm;   http://csdl.computer.
org/dl/mags/co/2004/09/r9116.pdf.

Anonymous:2004:ASj

[1286] Anonymous. Article summaries. Computer, 37(10):4–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/10/rx004.pdf;
http://csdl.computer.org/dl/mags/
co/2004/10/rx004.htm.

Anonymous:2004:Lg

[1287] Anonymous.  Letters.  Com-

puter, 37(10):6–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/10/rx006.pdf;
http://csdl.computer.org/dl/mags/
co/2004/10/rx006.htm.

Colwell:2004:TCF

[1288] Bob Colwell. Trusting a chaotic future. Computer, 37(10):8–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/10/rx008.htm;
http://csdl.computer.org/dl/mags/
co/2004/10/rx008.pdf.

Anonymous:2004:RM

[1289] Anonymous. Report to members.

Computer, 37(10):15–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/10/rx015.

pdf;   http://csdl.computer.org/
dl/mags/co/2004/10/rx015.htm.

Geer:2004:WGR

[1290] David Geer. Will gesture-recognition technology point the way? Computer, 37(10):20–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/10/rx020.pdf;
 http://csdl.computer.org/dl/mags/
co/2004/10/rx020.htm.

Lawton:2004:LHR

[1291] George Lawton. The lowdown on highrise chips. Computer, 37(10):24–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/10/rx024.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/10/rx024.htm.

Paulson:2004:NBg

[1292] Linda Dailey Paulson. News briefs. Computer, 37(10):28–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/10/rx028.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/10/rx028.htm.

Chua:2004:FIA

[1293] Cecil Eng Huang Chua and Jonathan Wareham. Fighting Internet auction fraud: An assessment and proposal. Computer, 37(10):31–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/10/rx031.htm;

http://csdl.computer.org/dl/mags/
co/2004/10/rx031.pdf.

Whitworth:2004:SST

[1294] Brian Whitworth and Elizabeth Whitworth. Spam and the social-technical gap. Computer, 37(10):38–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/10/rx038.htm;
http://csdl.computer.org/dl/mags/
co/2004/10/rx038.pdf.

Garlan:2004:RAB

[1295] David Garlan, Shang-Wen Cheng, An-Cheng Huang, Bradley Schmerl, and Peter Steenkiste. Rainbow: Architecture-based self-adaptation with reusable infrastructure. Computer, 37 (10):46–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/mags/co/
2004/10/rx046.htm;   http://csdl.
computer.org/dl/mags/co/2004/10/
rx046.pdf.

Saiedian:2004:UUB

[1296] Hossein Saiedian and Srikrishnan Raguraman. Using UML-based rate monotonic analysis to predict schedulability. Computer, 37(10):56–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/10/
rx056.htm;   http://csdl.computer.
org/dl/mags/co/2004/10/rx056.pdf.

Harel:2004:MMW

[1297] David Harel and Bernhard Rumpe.

Meaningful modeling: What’s the semantics of “semantics”? Computer, 37(10):64–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/10/rx064.htm;
http://csdl.computer.org/dl/mags/
co/2004/10/rx064.pdf.

Griswold:2004:AEC

[1298] William G. Griswold, Patricia Shanahan, Steven W. Brown, Robert Boyer, Matt Ratto, R. Benjamin Shapiro, and Tan Minh Truong. Active-Campus: Experiments in communityoriented ubiquitous computing. Computer, 37(10):73–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/10/rx073.htm;
http://csdl.computer.org/dl/mags/
co/2004/10/rx073.pdf.

Anonymous:2004:CSCh

[1299] Anonymous. Computer Society connection. Computer, 37(10):82–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/10/rx082.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/10/rx082.htm.

Anonymous:2004:CCj

[1300] Anonymous. Call and calendar. Computer, 37(10):86–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/10/rx086.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/10/rx086.htm.

Anonymous:2004:COi

[1301] Anonymous. Career opportunities.

Computer, 37(10):92–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/10/rx092.pdf.

Anonymous:2004:Bh

[1302] Anonymous.  Bookshelf.  Com-

puter, 37(10):104–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/10/rx104.pdf;
http://csdl.computer.org/dl/mags/
co/2004/10/rx104.htm.

Anonymous:2004:Pg

[1303] Anonymous.  Products.  Com-

puter, 37(10):105–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/10/rx105.pdf;
http://csdl.computer.org/dl/mags/
co/2004/10/rx105.htm.

Smith:2004:MBB

[1304] Sean Smith. Magic boxes and boots: Security in hardware. Computer, 37(10):106–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/10/rx106.htm;
http://csdl.computer.org/dl/mags/
co/2004/10/rx106.pdf.

Schmandt:2004:UCL

[1305] Chris Schmandt and Natalia Marmasse. User-centered location awareness. Computer, 37(10):110–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/10/
rx110.htm;   http://csdl.computer.
org/dl/mags/co/2004/10/rx110.pdf.

Goslin:2004:PGE

[1306] Mike Goslin and Mark R. Mine. The Panda3D graphics engine. Computer, 37(10):112–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/10/rx112.htm;
http://csdl.computer.org/dl/mags/
co/2004/10/rx112.pdf.

Maheswaran:2004:TNC

[1307] Muthucumaru Maheswaran and Shoukat Ali. A taxonomy of network computing systems. Computer, 37(10): 115–??, October 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/10/
rx115.htm;   http://csdl.computer.
org/dl/mags/co/2004/10/rx115.pdf.

Skowronski:2004:DAM

[1308] Victor Skowronski. Do agile methods marginalize problem solvers? Computer, 37(10):120–??, October 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/10/rx120.htm;
http://csdl.computer.org/dl/mags/
co/2004/10/rx120.pdf.

Anonymous:2004:ASk

[1309] Anonymous. Article summaries.

Computer, 37(11):4–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/11/ry004.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/11/ry004.htm.

Anonymous:2004:Lh

[1310] Anonymous.  Letters.  Com-

puter, 37(11):6–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/11/ry006.pdf;
http://csdl.computer.org/dl/mags/
co/2004/11/ry006.htm.

Anonymous:2004:YAe

[1311] Anonymous. 32 and 16 years ago.

Computer, 37(11):8–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/11/ry008.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/11/ry008.htm.

Colwell:2004:ATC

[1312] Bob Colwell. All things are connected.

Computer, 37(11):10–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/11/ry010.htm;
http://csdl.computer.org/dl/mags/
co/2004/11/ry010.pdf.

Leavitt:2004:WSF

[1313] Neal Leavitt. Are Web services finally ready to deliver? Computer, 37(11):14–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/11/ry014.pdf;
 http://csdl.computer.org/dl/mags/
co/2004/11/ry014.htm.

Paulson:2004:NBh

[1314] Linda Dailey Paulson. News briefs. Computer, 37(11):20–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/11/ry020.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/11/ry020.htm.

Rosenbloom:2004:NFC

[1315] Paul S. Rosenbloom. A new framework for computer science and engineering. Computer, 37(11):23–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/11/ry023.htm;
http://csdl.computer.org/dl/mags/
co/2004/11/ry023.pdf.

ONeal:2004:RCP

[1316] Mike O’Neal. Restructuring computing programs to meet employment challenges. Computer, 37(11): 29–??, November 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/11/
ry029.htm;   http://csdl.computer.
org/dl/mags/co/2004/11/ry029.pdf.

Kant:2004:GEI

[1317] Krishna Kant and Prasant Mohapatra. Guest Editors’ introduction: Internet data centers. Computer, 37(11):35–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/11/ry035.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/11/ry035.htm.

Kallahalla:2004:SSB

[1318] Mahesh Kallahalla, Mustafa Uysal, Ram Swaminathan, David E. Lowell, Mike Wray, Tom Christian, Nigel Edwards, Chris I. Dalton, and Frederic Gittler. SoftUDC: A softwarebased data center for utility computing. Computer, 37(11):38–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/11/ry038.htm;
http://csdl.computer.org/dl/mags/
co/2004/11/ry038.pdf.

Regnier:2004:TOD

[1319] Greg Regnier, Srihari Makineni, Ramesh Illikkal, Ravi Iyer, Dave Minturn, Ram Huggahalli, Don Newell, Linda Cline, and Annie Foong. TCP onloading for data center servers. Computer, 37(11):48–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/11/ry048.htm;
http://csdl.computer.org/dl/mags/
co/2004/11/ry048.pdf.

Candea:2004:ROC

[1320] George Candea, Aaron B. Brown, Armando Fox, and David Patterson. Recovery-oriented computing: Building multitier dependability. Computer, 37(11):60–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2004/11/ry060.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2004/11/ry060.pdf.

Bianchini:2004:PEM

[1321] Ricardo Bianchini and Ram Rajamony. Power and energy management for server systems. Computer, 37(11):68–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2004/11/ry068.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2004/11/ry068.pdf.

Anonymous:2004:Ph

[1322] Anonymous.  Products.  Com-

puter, 37(11):77–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/11/ry077.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/11/ry077.htm.

Anonymous:2004:Bi

[1323] Anonymous.  Bookshelf.  Com-

puter, 37(11):78–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/11/ry078.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/11/ry078.htm.

Anonymous:2004:CSCi

[1324] Anonymous. Computer Society connection. Computer, 37(11):80–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/11/ry080.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/11/ry080.htm.

Anonymous:2004:CCk

[1325] Anonymous. Call and calendar.

Computer, 37(11):87–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/11/ry087.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/11/ry087.htm.

Anonymous:2004:COj

[1326] Anonymous. Career opportunities.

Computer, 37(11):90–??, November 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/11/ry090.
pdf.

Hassler:2004:OCP

[1327] Vesna Hassler. Online collaboration products. Computer, 37(11):106– ??, November 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/11/
ry106.htm;   http://csdl.computer.
org/dl/mags/co/2004/11/ry106.pdf.

Hughes:2004:ISE

[1328] Jim Hughes. IEEE standard for encrypted storage. Computer, 37(11): 110–??, November 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/11/

                                        ry110.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/11/ry110.pdf.

Wolf:2004:AA

[1329] Wayne Wolf. Applications and architectures. Computer, 37(11):114– ??, November 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2004/11/
                                        ry114.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/11/ry114.pdf.

Holmes:2004:UH

[1330] Neville Holmes. The usefulness of hindsight. Computer, 37(11):120– ??, November 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2004/11/
                                        ry120.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/11/ry120.pdf.

Anonymous:2004:ASl

[1331] Anonymous. Article summaries. Computer, 37(12):4–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2004/12/rz004.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2004/12/rz004.htm.

Anonymous:2004:Li

[1332] Anonymous.  Letters.  Com-

puter, 37(12):6–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/12/rz006.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/12/rz006.htm.

Colwell:2004:IQI

[1333] Bob Colwell. Interface quotas and Internet-derived value. Computer, 37(12):10–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/12/rz010.pdf;
http://csdl.computer.org/dl/mags/
co/2004/12/rz010.htm.

Chang:2004:TWC

[1334] Carl K. Chang. Transforming the world’s Computer Society: Together, we will get it done! Computer, 37(12):14–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/12/rz014.pdf;
http://csdl.computer.org/dl/mags/
co/2004/12/rz014.htm.

Ortiz:2004:ITJ

[1335] Sixto Ortiz Jr. Internet telephony jumps off the wires. Computer, 37(12):16–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/12/rz016.pdf;
http://csdl.computer.org/dl/mags/
co/2004/12/rz016.htm.

Geer:2004:DRT

[1336] David Geer. Digital rights technology sparks interoperability concerns. Computer, 37(12):20–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/12/rz020.

pdf;   http://csdl.computer.org/
dl/mags/co/2004/12/rz020.htm.

Paulson:2004:NBi

[1337] Linda Dailey Paulson. News briefs. Computer, 37(12):23–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/12/rz023.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/12/rz023.htm.

Lindvall:2004:ASD

[1338] Mikael Lindvall, Dirk Muthig, Aldo Dagnino, Christina Wallin, Michael Stupperich, David Kiefer, John May, and Tuomo Kähkönen. Agile software development in large organizations. Computer, 37(12):26–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/12/rz026.htm;
http://csdl.computer.org/dl/mags/
co/2004/12/rz026.pdf.

Myers:2004:THD

[1339] Brad A. Myers, Jeffrey Nichols, Jacob O. Wobbrock, and Robert C. Miller. Taking handheld devices to the next level. Computer, 37(12):36– ??, December 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/12/
rz036.htm;   http://csdl.computer.
org/dl/mags/co/2004/12/rz036.pdf.

Prante:2004:RCD

[1340] Thorsten Prante, Norbert Streitz, and Peter Tandler. Roomware: Computers disappear and interaction evolves.

Computer, 37(12):47–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/12/rz047.htm;
http://csdl.computer.org/dl/mags/
co/2004/12/rz047.pdf.

Schilit:2004:DE

[1341] Bill N. Schilit and Uttam Sengupta. Device ensembles. Computer, 37(12):56–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/12/rz056.htm;
http://csdl.computer.org/dl/mags/
co/2004/12/rz056.pdf.

Narayanaswami:2004:EDC

[1342] Chandra Narayanaswami and M. T.

Raghunath. Expanding the digital camera’s reach. Computer, 37(12): 65–??, December 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/12/
rz065.htm;   http://csdl.computer.
org/dl/mags/co/2004/12/rz065.pdf.

Anonymous:2004:AI

[1343] Anonymous. 2004 annual index.

Computer, 37(12):74–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/12/rz074.
pdf;   http://csdl.computer.org/
dl/mags/co/2004/12/rz074.htm.

Anonymous:2004:CCl

[1344] Anonymous. Call and calendar.

Computer, 37(12):87–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2004/12/rz087.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2004/12/rz087.htm.

Carver:2004:RMV

[1345] Doris Carver. Reviewers make vital contributions to computer. Computer, 37(12):89–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2004/12/rz089.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2004/12/rz089.htm.

Anonymous:2004:CNC

[1346] Anonymous. Cooper named 2005 Computer Society President-Elect. Computer, 37(12):92–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2004/12/rz092.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2004/12/rz092.htm.

Anonymous:2004:Bj

[1347] Anonymous.  Bookshelf.  Com-

puter, 37(12):94–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2004/12/rz094.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2004/12/rz094.htm.

Anonymous:2004:Pi

[1348] Anonymous.  Products.  Com-

puter, 37(12):95–??, December 2004.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2004/12/rz095.pdf;
 http://csdl.computer.org/dl/mags/
co/2004/12/rz095.htm.

Anonymous:2004:COk

[1349] Anonymous. Career opportunities.

Computer, 37(12):96–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2004/12/rz096.
pdf.

Varshney:2004:VMC

[1350] Upkar Varshney. Vehicular mobile commerce. Computer, 37(12):116– ??, December 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/12/
rz116.htm;   http://csdl.computer.
org/dl/mags/co/2004/12/rz116.pdf.

Audeh:2004:MSW

[1351] Malik Audeh. Metropolitan-scale wifi mesh networks. Computer, 37(12): 119–??, December 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2004/12/
rz119.htm;   http://csdl.computer.
org/dl/mags/co/2004/12/rz119.pdf.

Macedonia:2004:SB

[1352] Michael Macedonia. Small is beautiful.

Computer, 37(12):122–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2004/12/rz122.htm;

                                        http://csdl.computer.org/dl/mags/
                                        co/2004/12/rz122.pdf.

Arbaugh:2004:RRD

[1353] William A. Arbaugh. The real risk of digital voting? Computer, 37(12):124–??, December 2004. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2004/12/rz124.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2004/12/rz124.pdf.

Santini:2004:DCS

[1354] Simone Santini. Determining computing science’s role. Computer, 37(12): 128–??, December 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2004/12/
                                        rz128.htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/12/rz128.pdf.

Anonymous:2005:ASa

[1355] Anonymous. Article summaries. Computer, 38(1):4–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/01/r1004.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2005/01/r1004.htm.

Anonymous:2005:La

[1356] Anonymous.  Letters.  Com-

puter, 38(1):6–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/01/r1006.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2005/01/r1006.htm.

Engel:2005:BPL

[1357] Gerald L. Engel. Building on the past–looking to the future. Computer, 38(1):7–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/01/r1007.pdf;
 http://csdl.computer.org/dl/mags/
co/2005/01/r1007.htm.

Carver:2005:MTC

[1358] Doris L. Carver. Maintaining technical currency in today’s technology climate. Computer, 38(1):10–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/01/r1010.
pdf;   http://csdl.computer.org/
dl/mags/co/2005/01/r1010.htm.

Colwell:2005:MIM

[1359] Bob Colwell. Machine intelligence meets neuroscience. Computer, 38(1):12–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/01/r1012.htm;
http://csdl.computer.org/dl/mags/
co/2005/01/r1012.pdf.

Anonymous:2005:YAa

[1360] Anonymous. 32 and 16 years ago.

Computer, 38(1):16–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/01/r1016.htm;
http://csdl.computer.org/dl/mags/
co/2005/01/r1016.pdf.

Geer:2005:MBT

[1361] David Geer. Malicious bots threaten network security. Computer, 38(1): 18–??, January 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2005/
                                        01/r1018.pdf;   http://csdl.computer.
                                        org/dl/mags/co/2005/01/r1018.htm.

Paulson:2005:SSC

[1362] Linda Dailey Paulson. Squeezing supercomputers onto a chip. Computer, 38(1):21–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/01/r1021.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2005/01/r1021.htm.

Anonymous:2005:NBa

[1363] Anonymous. News briefs. Computer, 38(1):24–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/01/r1024.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2005/01/r1024.htm.

Rehnmark:2005:RSL

[1364] Fredrik Rehnmark, William Bluethmann, Joshua Mehling, Robert O. Ambrose, Myron Diftler, Mars Chu, and Ryan Necessary. Robonaut: The ’short list’ of technology hurdles. Computer, 38(1):28–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2005/01/r1028.htm;

http://csdl.computer.org/dl/mags/
co/2005/01/r1028.pdf.

Efroni:2005:RAR

[1365] Sol Efroni, David Harel, and Irun R.

Cohen. Reactive animation: Realistic modeling of complex dynamic systems. Computer, 38(1):38–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/01/r1038.htm;
http://csdl.computer.org/dl/mags/
co/2005/01/r1038.pdf.

Bimber:2005:EES

[1366] Oliver Bimber, Andreas Emmerling, and Thomas Klemmer. Embedded entertainment with smart projectors. Computer, 38(1):48–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/01/r1048.htm;
http://csdl.computer.org/dl/mags/
co/2005/01/r1048.pdf.

Dwyer:2005:SAA

[1367] Chris Dwyer, Alvin R. Lebeck, and Daniel J. Sorin. Self-assembled architectures and the temporal aspects of computing. Computer, 38(1):56– ??, January 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2005/01/
r1056.htm;   http://csdl.computer.
org/dl/mags/co/2005/01/r1056.pdf.

Keyes:2005:CQC

[1368] Robert W. Keyes. Challenges for quantum computing with solid-state devices. Computer, 38(1):65–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2005/01/r1065.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2005/01/r1065.pdf.

Bharati:2005:IIS

[1369] Pratyush Bharati. India’s IT services industry: A comparative analysis. Computer, 38(1):71–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2005/01/r1071.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2005/01/r1071.pdf.

Anonymous:2005:CSO

[1370] Anonymous. Computer Society offers a wide array of products and services in 2005. Computer, 38(1):79–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/01/r1079.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2005/01/r1079.htm.

Anonymous:2005:CCa

[1371] Anonymous. Call and calendar. Computer, 38(1):84–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/01/r1084.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2005/01/r1084.htm.

Anonymous:2005:COa

[1372] Anonymous. Career opportunities. Computer, 38(1):88–??, January 2005.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/01/r1088.pdf.

Anonymous:2005:Ba

[1373] Anonymous.  Bookshelf.  Com-

puter, 38(1):100–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/01/r1100.pdf;
 http://csdl.computer.org/dl/mags/
co/2005/01/r1100.htm.

Anonymous:2005:Pa

[1374] Anonymous.  Products.  Com-

puter, 38(1):101–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/01/r1101.pdf;
 http://csdl.computer.org/dl/mags/
co/2005/01/r1101.htm.

Hellestrand:2005:ES

[1375] Graham Hellestrand. The engineering of supersystems. Computer, 38(1):103–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/01/r1103.htm;
http://csdl.computer.org/dl/mags/
co/2005/01/r1103.pdf.

Johnson:2005:JDF

[1376] Rod Johnson. J2EE development frameworks. Computer, 38(1):107– ??, January 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2005/01/

                                        r1107.htm;   http://csdl.computer.
                                        org/dl/mags/co/2005/01/r1107.pdf.

Croll:2005:SSS

[1377] Paul R. Croll and Susan K. Land. S2ESC: Setting standards for three decades. Computer, 38(1):111–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2005/01/
                                        r1111.htm;   http://csdl.computer.
                                        org/dl/mags/co/2005/01/r1111.pdf.

Cellary:2005:PSK

[1378] Wojciech Cellary. People and software in a knowledge-based economy. Computer, 38(1):116–??, January 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2005/01/r1116.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2005/01/r1116.pdf.

Anonymous:2005:ASb

[1379] Anonymous. Article summaries. Computer, 38(2):4–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/02/r2004.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2005/02/r2004.htm.

Anonymous:2005:Lb

[1380] Anonymous.  Letters.  Com-

puter, 38(2):6–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/02/r2006.pdf;

 http://csdl.computer.org/dl/mags/
co/2005/02/r2006.htm.

Colwell:2005:BSM

[1381] Bob Colwell. Behind the scenes at Micro-37. Computer, 38(2):8– ??, February 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2005/02/
r2008.htm;   http://csdl.computer.
org/dl/mags/co/2005/02/r2008.pdf.

Anonymous:2005:YAb

[1382] Anonymous. 32 and 16 years ago.

Computer, 38(2):12–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/02/r2012.htm;
http://csdl.computer.org/dl/mags/
co/2005/02/r2012.pdf.

Geer:2005:DAW

[1383] David Geer. Developments advance Web conferencing. Computer, 38(2):14–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/02/r2014.pdf;
 http://csdl.computer.org/dl/mags/
co/2005/02/r2014.htm.

Lawton:2005:TCA

[1384] George Lawton. Telecom carriers actively pursue passive optical networks. Computer, 38(2):19–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/02/r2019.
pdf;   http://csdl.computer.org/
dl/mags/co/2005/02/r2019.htm.

Anonymous:2005:NBb

[1385] Anonymous. News briefs. Computer, 38(2):22–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/02/r2022.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2005/02/r2022.htm.

Himmelstein:2005:LSI

[1386] Marty Himmelstein. Local search: The Internet is the Yellow Pages. Computer, 38(2):26–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2005/02/r2026.htm;
                                        http://csdl.computer.org/dl/mags/
                                        co/2005/02/r2026.pdf.

Zorian:2005:NDT

[1387] Yervant Zorian. Nanoscale design and test challenges. Computer, 38(2):36–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/02/r2036.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2005/02/r2036.htm.

Mitra:2005:RSD

[1388] Subhasish Mitra, Norbert Seifert, Ming Zhang, Quan Shi, and Kee Sup Kim. Robust system design with built-in soft-error resilience. Computer, 38(2):43–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2005/02/r2043.htm;

http://csdl.computer.org/dl/mags/
co/2005/02/r2043.pdf.

Roy:2005:TLO

[1389] Rob Roy, Debashis Bhattacharya, and Vamsi Boppana. Transistor-level optimization of digital designs with flex cells. Computer, 38(2):53–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/02/r2053.htm;
http://csdl.computer.org/dl/mags/
co/2005/02/r2053.pdf.

Jerraya:2005:HSI

[1390] Ahmed A. Jerraya and Wayne Wolf.

Hardware/software interface codesign for embedded systems. Computer, 38(2):63–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/02/r2063.htm;
http://csdl.computer.org/dl/mags/
co/2005/02/r2063.pdf.

Talarico:2005:NFP

[1391] Claudio Talarico, Jerzy W. Rozenblit, Vinod Malhotra, and Albert Stritter. A new framework for power estimation of embedded systems. Computer, 38(2):71–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/02/r2071.htm;
http://csdl.computer.org/dl/mags/
co/2005/02/r2071.pdf.

Anonymous:2005:Pb

[1392] Anonymous.  Products.  Com-

puter, 38(2):79–??, February 2005.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/02/r2079.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2005/02/r2079.htm.

Anonymous:2005:Bb

[1393] Anonymous.  Bookshelf.  Com-

puter, 38(2):80–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/02/r2080.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2005/02/r2080.htm.

Anonymous:2005:CSCa

[1394] Anonymous. Computer Society connection. Computer, 38(2):81–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/02/r2081.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2005/02/r2081.htm.

Anonymous:2005:CCb

[1395] Anonymous. Call and calendar. Computer, 38(2):86–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/02/r2086.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2005/02/r2086.htm.

Anonymous:2005:COb

[1396] Anonymous. Career opportunities. Computer, 38(2):88–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/02/r2088.pdf.

Macedonia:2005:EGR

[1397] Michael Macedonia. Ender’s game redux. Computer, 38(2):95–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/02/r2095.htm;
http://csdl.computer.org/dl/mags/
co/2005/02/r2095.pdf.

Schilit:2005:CPD

[1398] Bill N. Schilit and Roy Want. Creating and protecting digital worlds. Computer, 38(2):99–??, February 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/02/r2099.htm;
http://csdl.computer.org/dl/mags/
co/2005/02/r2099.pdf.

Holmes:2005:PBP

[1399] Neville Holmes. The profession and the big picture. Computer, 38(2): 104–??, February 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2005/02/
r2104.htm;   http://csdl.computer.
org/dl/mags/co/2005/02/r2104.pdf.

Anonymous:2005:ASc

[1400] Anonymous. Article summaries.

Computer, 38(3):4–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/03/r3004.pdf;
 http://csdl.computer.org/dl/mags/
co/2005/03/r3004.htm.

Anonymous:2005:Lc

[1401] Anonymous. Letters. Computer, 38 (3):6–??, March 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2005/
                                        03/r3006.pdf;   http://csdl.computer.
                                        org/dl/mags/co/2005/03/r3006.htm.

Colwell:2005:EPB

[1402] Bob Colwell. Engineers, programmers, and black boxes. Computer, 38(3):8–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/03/r3008.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2005/03/r3008.htm.

Anonymous:2005:YAc

[1403] Anonymous. 32 and 16 years ago. Computer, 38(3):12–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/03/r3012.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2005/03/r3012.htm.

Cole:2005:SET

[1404] Bernard Cole. Search engines tackle the desktop. Computer, 38(3):14–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/03/r3014.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2005/03/r3014.htm.

Geer:2005:ITCa

[1405] David Geer. Is it time for clockless chips? Computer, 38(3):18–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/03/r3018.
pdf;   http://csdl.computer.org/
dl/mags/co/2005/03/r3018.htm.

Paulson:2005:NBa

[1406] Linda Dailey Paulson. News briefs.

Computer, 38(3):22–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/03/r3022.pdf;
http://csdl.computer.org/dl/mags/
co/2005/03/r3022.htm.

Gao:2005:IBR

[1407] Hong Tina Gao, Jane Huffman Hayes, and Henry Cai. Integrating biological research through Web services. Computer, 38(3):26–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/03/r3026.htm;
http://csdl.computer.org/dl/mags/
co/2005/03/r3026.pdf.

Pentland:2005:SAC

[1408] Alex (Sandy) Pentland. Socially aware computation and communication. Computer, 38(3):33–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/03/r3033.htm;
http://csdl.computer.org/dl/mags/
co/2005/03/r3033.pdf.

Streitz:2005:DSA

[1409] Norbert A. Streitz, Carsten Röcker, Thorsten Prante, Daniel van Alphen, Richard Stenzel, and Carsten Magerkurth. Designing smart artifacts for smart environments. Computer, 38(3):41– ??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2005/03/
r3041.htm;   http://csdl.computer.
org/dl/mags/co/2005/03/r3041.pdf.

Helal:2005:GTS

[1410] Sumi Helal, William Mann, Hicham El-Zabadani, Jeffrey King, Youssef Kaddoura, and Erwin Jansen. The Gator Tech smart house: A programmable pervasive space. Computer, 38(3):50–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/03/r3050.htm;
http://csdl.computer.org/dl/mags/
co/2005/03/r3050.pdf.

Srinivasan:2005:WLD

[1411] Savitha Srinivasan, Vikas Krishna, and Scott Holmes. Web-log-driven business activity monitoring. Computer, 38(3):61–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/03/r3061.htm;
http://csdl.computer.org/dl/mags/
co/2005/03/r3061.pdf.

Anonymous:2005:COc

[1412] Anonymous. Career opportunities. Computer, 38(3):70–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/03/r3070.pdf.

Anonymous:2005:CSCb

[1413] Anonymous. Computer Society connection. Computer, 38(3):75–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/03/r3075.
pdf;   http://csdl.computer.org/
dl/mags/co/2005/03/r3075.htm.

Anonymous:2005:CCc

[1414] Anonymous. Call and calendar.

Computer, 38(3):80–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/03/r3080.pdf;
http://csdl.computer.org/dl/mags/
co/2005/03/r3080.htm.

Anonymous:2005:Pc

[1415] Anonymous.  Products.  Com-

puter, 38(3):83–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/03/r3083.pdf;
http://csdl.computer.org/dl/mags/
co/2005/03/r3083.htm.

Anonymous:2005:Bc

[1416] Anonymous.  Bookshelf.  Com-

puter, 38(3):86–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/03/r3086.pdf;
http://csdl.computer.org/dl/mags/
co/2005/03/r3086.htm.

Wolf:2005:BSR

[1417] Wayne Wolf. Building the software radio. Computer, 38(3):87–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/03/r3087.htm;
http://csdl.computer.org/dl/mags/
co/2005/03/r3087.pdf.

Deutsch:2005:POI

[1418] Herb Deutsch. Public opinion’s influence on voting system technology. Computer, 38(3):93–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/03/r3093.htm;
http://csdl.computer.org/dl/mags/
co/2005/03/r3093.pdf.

Anandalingam:2005:WCH

[1419] G. Anandalingam and Henry C. Lucas Jr. The winner’s curse in high tech. Computer, 38(3):96–??, March 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/03/r3096.htm;
http://csdl.computer.org/dl/mags/
co/2005/03/r3096.pdf.

Johnson:2005:OSV

[1420] Thomas K. Johnson. An open-secret voting system. Computer, 38(3): 100–??, March 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2005/03/
r3100.htm;   http://csdl.computer.
org/dl/mags/co/2005/03/r3100.pdf.

Anonymous:2005:ASd

[1421] Anonymous. Article summaries. Computer, 38(4):7–??, April 2005.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/04/r4007.pdf;
 http://csdl.computer.org/dl/mags/
co/2005/04/r4007.htm.

Anonymous:2005:Ld

[1422] Anonymous. Letters. Computer, 38 (4):8–??, April 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
04/r4008.pdf;   http://csdl.computer.
org/dl/mags/co/2005/04/r4008.htm.

Anonymous:2005:YAd

[1423] Anonymous. 32 and 16 years ago. Computer, 38(4):10–??, April 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/04/r4010.
pdf;   http://csdl.computer.org/
dl/mags/co/2005/04/r4010.htm.

Colwell:2005:PHL

[1424] Bob Colwell. Point of highest leverage. Computer, 38(4):12–??, April 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/04/r4012.htm;
http://csdl.computer.org/dl/mags/
co/2005/04/r4012.pdf.

Geer:2005:WBX

[1425] David Geer. Will binary XML speed network traffic? Computer, 38(4): 16–??, April 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/

                                        04/r4016.pdf;   http://csdl.computer.
                                        org/dl/mags/co/2005/04/r4016.htm.

Leavitt:2005:MPN

[1426] Neal Leavitt. Mobile phones: The next frontier for hackers? Computer, 38 (4):20–??, April 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2005/
                                        04/r4020.pdf;   http://csdl.computer.
                                        org/dl/mags/co/2005/04/r4020.htm.

Paulson:2005:NBb

[1427] Linda Dailey Paulson. News briefs. Computer, 38(4):24–??, April 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/04/r4024.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2005/04/r4024.htm.

Zhuge:2005:FIE

[1428] Hai Zhuge. The future interconnection environment. Computer, 38(4): 27–??, April 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/dl/mags/co/2005/04/
                                        r4027.htm;   http://csdl.computer.
                                        org/dl/mags/co/2005/04/r4027.pdf.

Anderson:2005:OII

[1429] Thomas Anderson, Larry Peterson, Scott Shenker, and Jonathan Turner. Overcoming the Internet impasse through virtualization. Computer, 38(4):34–??, April 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/co/2005/04/r4034.htm;

http://csdl.computer.org/dl/mags/
co/2005/04/r4034.pdf.

Baker:2005:EGS

[1430] Mark Baker, Amy Apon, Clayton Ferner, and Jeff Brown. Emerging grid standards. Computer, 38 (4):43–??, April 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2005/04/
r4043.htm;   http://csdl.computer.
org/dl/mags/co/2005/04/r4043.pdf.

Schuba:2005:SNS

[1431] Christoph L. Schuba, Jason Goldschmidt, Michael F. Speer, and Mohamed Hefeeda. Scaling network services using programmable network devices. Computer, 38(4):52–??, April 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/04/r4052.htm;
http://csdl.computer.org/dl/mags/
co/2005/04/r4052.pdf.

Boykin:2005:LSN

[1432] P. Oscar Boykin and Vwani P. Roychowdhury. Leveraging social networks to fight spam. Computer, 38 (4):61–??, April 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/dl/mags/co/2005/04/
r4061.htm;   http://csdl.computer.
org/dl/mags/co/2005/04/r4061.pdf.

Anonymous:2005:COd

[1433] Anonymous. Career opportunities.

Computer, 38(4):70–??, April 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/04/r4070.pdf.

Anonymous:2005:Bd

[1434] Anonymous. Bookshelf. Computer, 38 (4):71–??, April 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2005/
                                        04/r4071.pdf;   http://csdl.computer.
                                        org/dl/mags/co/2005/04/r4071.htm.

Anonymous:2005:CCd

[1435] Anonymous. Call and calendar. Computer, 38(4):72–??, April 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/04/r4072.pdf;
                                         http://csdl.computer.org/dl/mags/
                                        co/2005/04/r4072.htm.

Anonymous:2005:Pd

[1436] Anonymous. Products. Computer, 38 (4):75–??, April 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2005/
                                        04/r4075.pdf;   http://csdl.computer.
                                        org/dl/mags/co/2005/04/r4075.htm.

Anonymous:2005:CSCc

[1437] Anonymous. Computer Society connection. Computer, 38(4):76–??, April 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/04/r4076.
                                        pdf;   http://csdl.computer.org/
                                        dl/mags/co/2005/04/r4076.htm.

Jacobson:2005:GEV

[1438] Jeffrey Jacobson and Michael Lewis.

Game engine virtual reality with CaveUT. Computer, 38(4):79–??, April 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/04/r4079.htm;
http://csdl.computer.org/dl/mags/
co/2005/04/r4079.pdf.

Smith:2005:SMA

[1439] Ian Smith. Social-mobile applications. Computer, 38(4):84–??, April 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/04/r4084.htm;
http://csdl.computer.org/dl/mags/
co/2005/04/r4084.pdf.

Holmes:2005:DS

[1440] Neville Holmes. In defense of spam.

Computer, 38(4):88–??, April 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
dl/mags/co/2005/04/r4088.htm;
http://csdl.computer.org/dl/mags/
co/2005/04/r4088.pdf.

Anonymous:2005:ASe

[1441] Anonymous. Article summaries. Computer, 38(5):4–??, May 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/05/r5004.pdf.

Anonymous:2005:YAe

[1442] Anonymous. 32 and 16 years ago. Computer, 38(5):6–??, May 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2005/05/r5006.pdf.

Colwell:2005:PO

[1443] Bob Colwell. Pecking orders. Computer, 38(5):8–??, May 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/05/r5008.pdf;
                                         http://csdl.computer.org/comp/
                                        mags/co/2005/05/r5008abs.htm.

Geer:2005:ITCb

[1444] David Geer. Industry trends: Chip makers turn to multicore processors. Computer, 38(5):11–??, May 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/05/r5011.pdf;
                                         http://csdl.computer.org/comp/
                                        mags/co/2005/05/r5011abs.htm.

Paulson:2005:TNW

[1445] Linda Dailey Paulson. Technology news: Will hard drives finally stop shrinking? Computer, 38(5):14–??, May 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/05/r5014.
                                        pdf;   http://csdl.computer.org/
                                        comp/mags/co/2005/05/r5014abs.htm.

Paulson:2005:NBc

[1446] Linda Dailey Paulson. News briefs. Computer, 38(5):17–??, May 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/05/r5017.pdf.

Kostelijk:2005:MAT

[1447] Ton Kostelijk. Misleading architecting tradeoffs. Computer, 38(5): 20–??, May 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
05/r5020abs.htm;   http://csdl.
computer.org/dl/mags/co/2005/05/
r5020.pdf.

Figueiredo:2005:GEI

[1448] Renato Figueiredo, Peter A. Dinda, and José Fortes. Guest Editors’ introduction: Resource virtualization renaissance. Computer, 38(5):28–??, May 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/05/r5028.pdf;
http://csdl.computer.org/comp/
mags/co/2005/05/r5028abs.htm.

Smith:2005:AVM

[1449] James E. Smith and Ravi Nair.

The architecture of virtual machines. Computer, 38(5):32–??, May 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
05/r5032abs.htm;   http://csdl.
computer.org/dl/mags/co/2005/05/
r5032.pdf.

Rosenblum:2005:VMM

[1450] Mendel Rosenblum and Tal Garfinkel.

Virtual machine monitors: Current technology and future trends. Computer, 38(5):39–??, May 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/05/r5039abs.htm;   http://
csdl.computer.org/dl/mags/co/2005/
05/r5039.pdf.

Uhlig:2005:IVT

[1451] Rich Uhlig, Gil Neiger, Dion Rodgers, Amy L. Santoni, Fernando C. M. Martins, Andrew V. Anderson, Steven M. Bennett, Alain Kägi, Felix H. Leung, and Larry Smith. Intel virtualization technology. Computer, 38 (5):48–??, May 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
05/r5048abs.htm;   http://csdl.
computer.org/dl/mags/co/2005/05/
r5048.pdf.

Whitaker:2005:RDV

[1452] Andrew Whitaker, Richard S. Cox, Marianne Shaw, and Steven D. Gribble. Rethinking the design of virtual machine monitors. Computer, 38 (5):57–??, May 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
05/r5057abs.htm;   http://csdl.
computer.org/dl/mags/co/2005/05/
r5057.pdf.

Ruth:2005:VDE

[1453] Paul Ruth, Xuxian Jiang, Dongyan Xu, and Sebastien Goasguen. Virtual distributed environments in a shared infrastructure. Computer, 38 (5):63–??, May 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
05/r5063abs.htm;   http://csdl.
computer.org/dl/mags/co/2005/05/
r5063.pdf.

Anonymous:2005:COe

[1454] Anonymous. Career opportunities.

Computer, 38(5):70–??, May 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/05/r5070.pdf.

Anonymous:2005:CSCd

[1455] Anonymous. Computer Society connection. Computer, 38(5):74–??, May 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/05/r5074.
pdf.

Anonymous:2005:CCe

[1456] Anonymous. Call and calendar. Computer, 38(5):79–??, May 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/05/r5079.pdf.

Anonymous:2005:Pe

[1457] Anonymous. Products. Computer, 38 (5):83–??, May 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
05/r5083.pdf.

Anonymous:2005:Be

[1458] Anonymous. Bookshelf. Computer, 38(5):85–??, May 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
05/r5085.pdf.

Kroenke:2005:BRD

[1459] David M. Kroenke. Beyond the relational database model. Computer, 38 (5):89–??, May 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
05/r5089abs.htm;   http://csdl.
computer.org/dl/mags/co/2005/05/
r5089.pdf.

Kahan:2005:OQD

[1460] William Kahan and Dan Zuras. An open question to developers of numerical software. Computer, 38(5): 91–94, May 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
05/r5091abs.htm;   http://csdl.
computer.org/comp/mags/co/2005/
05/r5toc.htm;   http://csdl.computer.
org/dl/mags/co/2005/05/r5091.pdf.

Magnusson:2005:VTL

[1461] Peter S. Magnusson. The virtual test lab. Computer, 38(5): 95–??, May 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
05/r5095abs.htm;   http://csdl.
computer.org/dl/mags/co/2005/05/
r5095.pdf.

Mahmoud:2005:RCS

[1462] Qusay H. Mahmoud. Revitalizing computing science education. Computer, 38(5):100–??, May 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
05/r5100abs.htm;   http://csdl.
computer.org/dl/mags/co/2005/05/
r5100.pdf.

Wirth:2006:GIT

[1463] Niklaus Wirth. Good ideas, through the looking glass. Computer, 39(1): 28–39, January 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
01/r1toc.xml;   http://csdl2.computer.
org/dl/mags/co/2006/01/r1028.pdf.

Anonymous:2005:ASf

[1464] Anonymous. Article summaries. Computer, 38(6):4, June 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

csdl2.computer.org/comp/mags/co/
2005/06/r6004.pdf.

Anonymous:2005:Le

[1465] Anonymous. Letters. Computer, 38 (6):6–8, June 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl2.

computer.org/comp/mags/co/2005/
06/r6006.pdf.

Colwell:2005:FR

[1466] Bob Colwell. Frames of reference. Computer, 38(6):9–11, June 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

csdl2.computer.org/comp/mags/co/
2005/06/r6009.pdf.

Anonymous:2005:YAf

[1467] Anonymous. 32 & 16 years ago. Computer, 38(6):12–13, June 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

                                        csdl2.computer.org/comp/mags/co/
                                        2005/06/r6012.pdf.

Lawton:2005:WLA

[1468] George Lawton. What lies ahead for cellular technology? Computer, 38 (6):14–17, June 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl2.

                                        computer.org/comp/mags/co/2005/
                                        06/r6014.pdf.

Geer:2005:STG

[1469] David Geer. Security technologies go phishing. Computer, 38(6):18–21, June 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl2.computer.org/
                                        comp/mags/co/2005/06/r6018.pdf.

Anonymous:2005:NBc

[1470] Anonymous. News briefs. Computer, 38(6):22–24, June 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

                                        csdl2.computer.org/comp/mags/co/
                                        2005/06/r6022.pdf.

Brewer:2005:CTD

[1471] Eric Brewer, Michael Demmer, Bowei Du, Melissa Ho, Matthew Kam, Sergiu Nedevschi, Joyojeet Pal, Rabin Patra, Sonesh Surana, and Kevin Fall. The case for technology in developing regions. Computer, 38(6):25–38, June 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hazzan:2005:RAL

[1472] Orit Hazzan and James E. Tomayko. Reflection and abstraction in learning software engineering’s human aspects. Computer, 38(6):39–45, June 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Heath:2005:PNE

[1473] Barbara P. Heath, Russell L. Herman, Gabriel G. Lugo, James H. Reeves, Ronald J. Vetter, and Charles R. Ward. Project Numina: Enhancing student learning with handheld computers. Computer, 38(6):46–53, June 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Friedland:2005:AME

[1474] Gerald Friedland and Karl Pauls. Architecting multimedia environments for teaching. Computer, 38(6):57–64, June 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mourad:2005:TEM

[1475] Magda Mourad, Gerard L. Hanley, Merlot Barbra Bied Sperling, and Jack Gunther. Toward an electronic marketplace for higher education. Computer, 38(6):66–75, June 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Conrad:2005:SRE

[1476] James M. Conrad. Stiquito for robotics and embedded systems education. Computer, 38(6):77–81, June 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2005:CSCe

[1477] Anonymous. Computer Society connection. Computer, 38(6):82–84, June 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl2.computer.org/
                                        comp/mags/co/2005/06/r6082.pdf.

Anonymous:2005:CCf

[1478] Anonymous. Call and calendar. Computer, 38(6):85–87, June 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

                                        csdl2.computer.org/comp/mags/co/
                                        2005/06/r6085.pdf.

Anonymous:2005:COf

[1479] Anonymous. Career opportunities. Computer, 38(6):88–92, June 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://csdl2.computer.org/comp/
                                        mags/co/2005/06/r6088.pdf.

Anonymous:2005:Pf

[1480] Anonymous. Products. Computer, 38(6):93, June 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl2.

                                        computer.org/comp/mags/co/2005/
                                        06/r6093.pdf.

Anonymous:2005:Bf

[1481] Anonymous. Bookshelf. Computer, 38(6):94, June 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl2.

                                        computer.org/comp/mags/co/2005/
                                        06/r6094.pdf.

Subramanya:2005:UMD

[1482] S. R. Subramanya and Byung K. Yi. Utility model for on-demand digital content. Computer, 38(6):95–98, June 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goldstein:2005:PM

[1483] Seth Copen Goldstein, Jason D. Campbell, and Todd C. Mowry. Programmable matter. Computer, 38(6): 99–101, June 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schaaf:2005:WAA

[1484] Rob Schaaf. What’s all this about systems? Computer, 38(6):104, 102– 103, June 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl2.

computer.org/persagen/DLPublication.
jsp?pubtype=m{\&}acronym=co;   http:
//search2.computer.org/advanced/
Advanced_Search_logged.jsp;   http:
//www.computer.org; http://www.
computer.org/contact.htm;   http:
//www.computer.org/cspress/; http:
//www.computer.org/help.htm;   http:
//www.computer.org/pr/; http://
www.computer.org/publications/dlib/
;   http://www.computer.org/toc.
htm.

Anonymous:2005:ASg

[1485] Anonymous. Article summaries. Computer, 38(7):4, July 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/07/r7004.pdf.

Anonymous:2005:Lf

[1486] Anonymous. Letters. Computer, 38 (7):6–9, July 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
07/r7006.pdf.

Anonymous:2005:YAg

[1487] Anonymous. 32 & 16 years ago. Computer, 38(7):10–11, July 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2005/07/r7010.pdf.

Colwell:2005:JSF

[1488] Bob Colwell. Judging science fairs. Computer, 38(7):12–15, July 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geer:2005:EBD

[1489] David Geer. Eclipse becomes the dominant Java IDE. Computer, 38 (7):16–18, July 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2005/
                                        07/r7016.pdf.

Leavitt:2005:IMN

[1490] Neal Leavitt. Instant messaging: A new target for hackers. Computer, 38 (7):20–23, July 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2005/
                                        07/r7020.pdf.

Paulson:2005:NBd

[1491] Linda Dailey Paulson. News briefs. Computer, 38(7):24–26, July 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/07/r7024.pdf.

Hole:2005:SWF

[1492] Kjell J. Hole, Erlend Dyrnes, and Per Thorsheim. Securing wi-fi networks. Computer, 38(7):28–34, July 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jerraya:2005:GEI

[1493] Ahmed Jerraya and Wayne Wolf. Guest Editors’ introduction: Multiprocessor systems-on-chips. Computer, 38 (7):36–40, July 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
07/r7036.pdf.

Goodacre:2005:PAI

[1494] John Goodacre and Andrew N. Sloss.

Parallelism and the ARM instruction set architecture. Computer, 38 (7):42–50, July 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leibson:2005:CPN

[1495] Steve Leibson and James Kim. Configurable processors: A new era in chip design. Computer, 38(7):51–59, July 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nava:2005:OPD

[1496] Mario Diaz Nava, Patrick Blouet, Philippe Teninge, Marcello Coppola, Tarek Ben-Ismail, Samuel Picchiottino, and Robin Wilson. An open platform for developing multiprocessor SoCs. Computer, 38(7):60–67, July 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Levy:2005:EDE

[1497] Markus Levy. Evaluating digital entertainment system performance. Computer, 38(7):68–72, July 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2005:CCg

[1498] Anonymous. Call and calendar. Computer, 38(7):73–75, July 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2005/07/r7073.pdf.

Anonymous:2005:CSCf

[1499] Anonymous. Computer Society connection. Computer, 38(7):76–78, July 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/07/r7076.
                                        pdf.

Anonymous:2005:COg

[1500] Anonymous. Career opportunities. Computer, 38(7):80–82, July 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/07/r7080.pdf.

Anonymous:2005:Pg

[1501] Anonymous. Products. Computer, 38(7):83, July 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2005/
                                        07/r7083.pdf.

Anonymous:2005:Bg

[1502] Anonymous. Bookshelf. Computer, 38(7):84, July 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2005/
                                        07/r7084.pdf.

Lee:2005:APT

[1503] Edward A. Lee. Absolutely positively on time: What would it take? Computer, 38(7):85–87, July 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stoneburner:2005:DFA

[1504] Gary Stoneburner. Developer-focused assurance requirements. Computer, 38 (7):91–93, July 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mangold:2005:TBG

[1505] Bernard Mangold, Mike Dooley, Gary W. Flake, Havi Hoffman, Tejaswi Kasturi, David M. Pennock, and Rael Dornfest. The tech buzz game. Computer, 38(7):94–97, July 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2005:TW

[1506] Neville Holmes. The turning of the wheel. Computer, 38(7):100, 98–99, July 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2005:ASh

[1507] Anonymous. Article summaries. Computer, 38(8):4, August 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/08/r8004.pdf.

Ward:2005:RMN

[1508] Bob Ward. Report to members: NC State team wins $20,000 first prize at CSIDC. Computer, 38(8):5–8, August 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/08/r8005.
pdf.

Anonymous:2005:YAh

[1509] Anonymous. 32 & 16 years ago. Computer, 38(8):10–11, August 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/08/r8010.pdf.

Colwell:2005:WWC

[1510] Bob Colwell. What’s the worst that can happen? Computer, 38(8): 12–15, August 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2005:STG

[1511] Linda Dailey Paulson. Search technology goes mobile. Computer, 38 (8):19–22, August 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/08/r8019.pdf.

Geer:2005:MTU

[1512] David Geer. Making 3D technology universal. Computer, 38(8):23–25, August 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/08/r8023.
pdf.

Paulson:2005:NBe

[1513] Linda Dailey Paulson. News briefs. Computer, 38(8):26–28, August 2005.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/08/r8026.pdf.

Bimber:2005:GEI

[1514] Oliver Bimber. Guest Editor’s introduction: The ultimate display— what will it be? Computer, 38(8): 29–30, August 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
08/r8029.pdf.

Dodgson:2005:AD

[1515] Neil A. Dodgson. Autostereoscopic 3D displays. Computer, 38(8):31–36, August 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Favalora:2005:VDA

[1516] Gregg E. Favalora and Actuality Systems. Volumetric 3D displays and application infrastructure. Computer, 38 (8):37–44, August 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Slinger:2005:CGH

[1517] Chris Slinger, Colin Cameron, and Maurice Stanley. Computer-generated holography as a generic display technology. Computer, 38(8):46–53, August 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Watson:2005:UDW

[1518] Benjamin Watson and David Luebke.

The ultimate display: Where will all the pixels come from? Computer, 38 (8):54–61, August 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jepsen:2005:SMM

[1519] Mary Lou Jepsen. Smoke, mirrors, and manufacturable displays. Computer, 38 (8):63–67, August 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2005:CSC

[1520] Bob Ward. Computer Society connection. Computer, 38(8):68–72, August 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/08/r8068.
pdf.

Anonymous:2005:CCh

[1521] Anonymous. Call and calendar. Computer, 38(8):73–76, August 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/08/r8073.pdf.

Anonymous:2005:COh

[1522] Anonymous. Career opportunities. Computer, 38(8):78–81, August 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/08/r8078.pdf.

Anonymous:2005:Bh

[1523] Anonymous. Bookshelf. Computer, 38(8):82, August 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
08/r8082.pdf.

Anonymous:2005:Ph

[1524] Anonymous. Products. Computer, 38(8):83–84, August 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/08/r8083.pdf.

Manocha:2005:GPC

[1525] Dinesh Manocha. General-purpose computations using graphics processors. Computer, 38(8):85–88, August 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Harper:2005:TNC

[1526] Richard Harper and Microsoft Research Cambridge. Toward a new communications genre. Computer, 38(8): 89–91, August 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chung:2005:BEC

[1527] Jen-Yao Chung, Kwei-Jay Lin, and Simon S. Y. Shim. Beyond electronic commerce. Computer, 38(8): 92–93, August 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Skowronski:2005:TSG

[1528] Victor Skowronski. Talking to the solitary genius. Computer, 38(8):96, 94–95, August 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2005:ASi

[1529] Anonymous. Article summaries.

Computer, 38(9):4, September 2005.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/09/r9004.pdf.

Anonymous:2005:Lg

[1530] Anonymous. Letters. Computer, 38 (9):6–7, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/09/r9006.pdf.

Anonymous:2005:YAi

[1531] Anonymous. 32 & 16 years ago. Computer, 38(9):8–9, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/09/r9008.pdf.

Colwell:2005:PLC

[1532] Bob Colwell. Presentation lessons from comedians. Computer, 38(9): 10–13, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/09/r9010.pdf.

Geer:2005:TGP

[1533] David Geer. Taking the graphics processor beyond graphics. Computer, 38 (9):14–16, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/09/r9014.pdf.

Lawton:2005:LLE

[1534] George Lawton. LAMP lights enterprise development efforts. Computer, 38(9):18–20, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/09/r9018.pdf.

Paulson:2005:NBf

[1535] Linda Dailey Paulson. News briefs.

Computer, 38(9):22–24, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/09/r9022.
pdf.

Zyda:2005:VSV

[1536] Michael Zyda and USC Information Sciences Institute. From visual simulation to virtual reality to games. Computer, 38(9):25–32, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fairley:2005:IRG

[1537] Richard E. Fairley and Mary Jane Willshire. Iterative rework: The good, the bad, and the ugly. Computer, 38(9): 34–41, September 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Janzen:2005:TDD

[1538] David Janzen and Hossein Saiedian.

Test-driven development: Concepts, taxonomy, and future direction. Computer, 38(9):43–50, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Meservy:2005:TSD

[1539] Thomas O. Meservy and Kurt D. Fenstermacher. Transforming software development: An MDA road map. Computer, 38(9):52–58, September 2005.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhatti:2005:DAM

[1540] Rafae Bhatti, Basit Shafiq, Mohamed Shehab, and Arif Ghafoor. Distributed access management in multimedia IDCs. Computer, 38(9):60– 69, September 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2005:IPE

[1541] Bob Ward. IEEE President-Elect candidates address Computer Society concerns. Computer, 38(9):70–75, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/09/r9070.
                                        pdf.

Anonymous:2005:COi

[1542] Anonymous. Career opportunities. Computer, 38(9):76–80, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/09/r9076.
                                        pdf.

Anonymous:2005:CCi

[1543] Anonymous. Call and calendar. Computer, 38(9):85–88, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/09/r9085.
                                        pdf.

Anonymous:2005:ICS

[1544] Anonymous. IEEE Computer Society election. Computer, 38(9):89– 96, September 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
09/r9089.pdf.

Anonymous:2005:Bi

[1545] Anonymous. Bookshelf. Computer, 38(9):98, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/09/r9098.pdf.

Anonymous:2005:Pi

[1546] Anonymous. Products. Computer, 38(9):99, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/09/r9099.pdf.

Hardaway:2005:OUC

[1547] Don Hardaway, Mark J. Hogan, and Richard G. Mathieu. Outsourcing the University Computer Lab. Computer, 38(9):100–102, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cole:2005:SSC

[1548] Jack Cole. Security in storage: A call for participation. Computer, 38 (9):103–105, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2005:IWC

[1549] Neville Holmes. The Internet, the Web, and the chaos. Computer, 38(9):108, 106–107, September 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2005:ASj

[1550] Anonymous. Article summaries. Computer, 38(10):4, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2005/10/rx004.pdf.

Anonymous:2005:YAj

[1551] Anonymous. 32 & 16 years ago. Computer, 38(10):6–7, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/10/rx006.pdf.

Anonymous:2005:Lh

[1552] Anonymous. Letters. Computer, 38 (10):8, October 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2005/
                                        10/rx008.pdf.

Colwell:2005:CD

[1553] Bob Colwell. Complexity in design. Computer, 38(10):10–12, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/10/rx010.
                                        pdf.

Paulson:2005:BRW

[1554] Linda Dailey Paulson. Building rich Web applications with Ajax. Computer, 38(10):14–17, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geer:2005:SMT

[1555] David Geer. Statistical machine translation gains respect. Computer, 38 (10):18–21, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/10/rx018.pdf.

Paulson:2005:NBg

[1556] Linda Dailey Paulson. News briefs.

Computer, 38(10):22–24, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/10/rx022.pdf.

Ramakrishnan:2005:AEF

[1557] Naren Ramakrishnan. From the Area Editor: Frontiers of search. Computer, 38(10):26–27, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/10/rx026.pdf.

Gyongyi:2005:SJI

[1558] Zoltán Gyöngyi and Hector Garcia-Molina. Spam: It’s not just for inboxes anymore. Computer, 38(10): 28–34, October 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Capra:2005:UWS

[1559] Robert G. Capra, III and Manuel A.

Pérez-Quiñones. Using Web search engines to find and refind information. Computer, 38(10):36–42, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ferreira:2005:ISA

[1560] Anita Ferreira and John Atkinson.

Intelligent search agents using Webdriven natural-language explanatory dialogs. Computer, 38(10):44–52, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mohan:2005:SAN

[1561] Bharath Kumar Mohan and Indian Institute of Science. Searching association networks for nurturers. Computer, 38(10):54–60, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ding:2005:SSW

[1562] Li Ding, Tim Finin, Anupam Joshi, Yun Peng Rong Pan, and Pavan Reddivari. Search on the Semantic Web. Computer, 38(10):62–69, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leong:2005:CPL

[1563] Monk-Ping Leong, Chi-Chiu Cheung, Chin-Wang Cheung, Polly P. M. Wan, Ivan K. H. Leung, Winnie M. M. Yeung, Wing-Seung Yuen, Kenneth S. K. Chow, Kwong-Sak Leung, and Philip H. W. Leong. CPE: A parallel library for financial engineering applications. Computer, 38(10):70–77, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2005:COj

[1564] Anonymous. Career opportunities. Computer, 38(10):78–90, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/10/rx078.pdf.

Ward:2005:CAT

[1565] Bob Ward. Computer awards top achievers. Computer, 38(10):91–94, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/10/rx091.
pdf.

Anonymous:2005:CCj

[1566] Anonymous. Call and calendar. Computer, 38(10):95–98, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2005/10/rx095.pdf.

Anonymous:2005:Pj

[1567] Anonymous. Products. Computer, 38(10):99, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/10/rx099.pdf.

Anonymous:2005:Bj

[1568] Anonymous. Bookshelf. Computer, 38(10):100, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/10/rx100.pdf.

Smith:2005:PFV

[1569] Brian K. Smith. Physical fitness in virtual worlds. Computer, 38(10):101– 103, October 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Birman:2005:CWS

[1570] Ken Birman. Can Web services scale up? Computer, 38(10):107–110, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Klemmer:2005:IPD

[1571] Scott R. Klemmer. Integrating physical and digital interactions. Computer, 38(10):111–113, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kalles:2005:IPC

[1572] Dimitris Kalles. Improving professional conduct in publishing. Computer, 38(10):116, 114–115, October 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2005:ASk

[1573] Anonymous. Article summaries. Computer, 38(11):4, November 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/11/ry004.pdf.

Anonymous:2005:Li

[1574] Anonymous. Letters. Computer, 38(11):6, November 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2005/11/ry006.pdf.

Colwell:2005:BES

[1575] Bob Colwell. Books engineers should read. Computer, 38(11):7–10, November 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2005:YAk

[1576] Anonymous. 32 & 16 years ago.

Computer, 38(11):11–12, November 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/11/ry011.
pdf.

Geer:2005:BCN

[1577] David Geer. Building converged networks with IMS technology. Computer, 38(11):14–16, November 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/11/ry014.pdf.

Lawton:2005:MAH

[1578] George Lawton. E-mail authentication is here, but has it arrived yet? Computer, 38(11):17–19, November 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/11/ry017.
pdf.

Paulson:2005:NBh

[1579] Linda Dailey Paulson. News briefs.

Computer, 38(11):20–22, November 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/11/ry020.
pdf.

Stankovic:2005:OOP

[1580] John A. Stankovic, Insup Lee, Aloysius Mok, and Raj Rajkumar. Opportunities and obligations for physical computing systems. Computer, 38(11):

23–31, November 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kumar:2005:HCM

[1581] Rakesh Kumar, Dean M. Tullsen, Norman P. Jouppi, and Parthasarathy Ranganathan. Heterogeneous chip multiprocessors. Computer, 38(11):32– 38, November 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cameron:2005:HPP

[1582] Kirk W. Cameron, Virginia Tech Rong Ge, Virginia Tech Xizhou Feng, and Virginia Tech. High-performance, power-aware distributed computing for scientific applications. Computer, 38 (11):40–47, November 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Casas:2005:BSE

[1583] Roberto Casas and Oscar Casas. Battery sensing for energy-aware system design. Computer, 38(11):48– 54, November 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ye:2005:SAC

[1584] Nong Ye and Toni Farley. A scientific approach to cyberattack detection. Computer, 38(11):55–61, November 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2005:COk

[1585] Anonymous. Career opportunities. Computer, 38(11):62–79, November 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/11/ry062.
pdf.

Ward:2005:CSM

[1586] Bob Ward. Computer Society magazines set coverage for 2006. Computer, 38(11):81–86, November 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/11/ry081.pdf.

Anonymous:2005:CCk

[1587] Anonymous. Call and calendar.

Computer, 38(11):87–90, November 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/11/ry087.
pdf.

Anonymous:2005:Pk

[1588] Anonymous. Products. Computer, 38(11):95, November 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/11/ry095.pdf.

Anonymous:2005:Bk

[1589] Anonymous. Bookshelf. Computer, 38(11):96, November 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/11/ry096.pdf.

Hoskinson:2005:CUR

[1590] Andy Hoskinson. Creating the ultimate research assistant. Computer, 38(11):

97–99, November 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Baggi:2005:ISS

[1591] Denis L. Baggi. An IEEE standard for symbolic music. Computer, 38(11): 100–102, November 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Arazi:2005:RPK

[1592] Benjamin Arazi, Itamar Elhanany, Ortal Arazi, and Hairong Qi. Revisiting public-key cryptography for wireless sensor networks. Computer, 38(11): 103–105, November 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2005:DTA

[1593] Neville Holmes. Digital technology, age, and gaming. Computer, 38(11): 108, 106–107, November 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2005:ASl

[1594] Anonymous. Article summaries. Computer, 38(12):4, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2005/12/rz004.pdf.

Anonymous:2005:Lj

[1595] Anonymous. Letters. Computer, 38 (12):5–6, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2005/12/rz005.pdf.

Engel:2005:EVI

[1596] Gerald L. Engel. The end of a very interesting year. Computer, 38(12): 7–8, December 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
12/rz007.pdf.

Colwell:2005:TM

[1597] Bob Colwell. It’s time to move on. Computer, 38(12):9–12, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2005:YAl

[1598] Anonymous. 32 & 16 years ago.

Computer, 38(12):13–14, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/12/rz013.
pdf.

Geer:2005:UMB

[1599] David Geer. Users make a beeline for ZigBee technology. Computer, 38 (12):16–19, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2005/12/rz016.pdf.

Leavitt:2005:WPS

[1600] Neal Leavitt. Will proposed standard make mobile phones more secure? Computer, 38(12):20–22, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/12/rz020.
pdf.

Paulson:2005:NBi

[1601] Linda Dailey Paulson. News briefs. Computer, 38(12):23–25, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/12/rz023.
                                        pdf.

Cushing:2005:GEI

[1602] Judy Cushing and Theresa Pardo. Guest Editors’ introduction: Research in the digital government realm. Computer, 38(12):26–32, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/12/rz026.
                                        pdf.

Delcambre:2005:DGR

[1603] Lois Delcambre and Genevieve Giuliano. Digital government research in academia. Computer, 38(12):33– 39, December 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pantel:2005:DAI

[1604] Patrick Pantel, Andrew Philpot, and Eduard Hovy. Data alignment and integration. Computer, 38(12):43– 50, December 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marchionini:2005:AGS

[1605] Gary Marchionini, Stephanie W. Haas, Junliang Zhang, and Jonathan Elsas. Accessing government statistical information. Computer, 38(12):

52–61, December 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Harrison:2005:BCI

[1606] Teresa M. Harrison, James P. Zappen, and Sibel Adali. Building community information systems: The connected kids case. Computer, 38(12): 62–69, December 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lau:2005:AGR

[1607] Gloria T. Lau, Kincho H. Law, and Gio Wiederhold. Analyzing government regulations using structural and domain information. Computer, 38(12): 70–76, December 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carver:2005:CRE

[1608] Doris Carver. Computer recognizes expert reviewers. Computer, 38(12): 77–79, December 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2005/
12/rz077.pdf.

Ward:2005:WVC

[1609] Bob Ward. Williams voted 2006 Computer Society President-Elect. Computer, 38(12):80–81, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2005/12/rz080.
pdf.

Anonymous:2005:CCl

[1610] Anonymous. Call and calendar. Computer, 38(12):82–83, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/12/rz082.
                                        pdf.

Anonymous:2005:Bl

[1611] Anonymous. Bookshelf. Computer, 38(12):84, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2005/12/rz084.pdf.

Anonymous:2005:Pl

[1612] Anonymous. Products. Computer, 38(12):85, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2005/12/rz085.pdf.

Anonymous:2005:AI

[1613] Anonymous. Annual index. Computer, 38(12):86–97, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2005/12/rz086.pdf.

Anonymous:2005:COl

[1614] Anonymous. Career opportunities. Computer, 38(12):98–116, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2005/12/rz098.
                                        pdf.

Hong:2005:MSR

[1615] Jason I. Hong. Minimizing security risks in Ubicomp systems. Computer, 38(12):118–119, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shim:2005:FIM

[1616] Simon S. Y. Shim, Geetanjali Bhalla, and Vishnu Pendyala. Federated identity management. Computer, 38(12): 120–122, December 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Macedonia:2005:PE

[1617] Michael Macedonia. Power from the edge. Computer, 38(12):123–125, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Santini:2005:WSI

[1618] Simone Santini. We are sorry to inform you …. Computer, 38(12):128, 126–127, December 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2006:ASa

[1619] Anonymous. Article summaries. Computer, 39(1):4, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/01/r1004.pdf.

Anonymous:2006:La

[1620] Anonymous. Letters. Computer, 39 (1):6–7, January 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carver:2006:CPL

[1621] Doris L. Carver. Celebrating the past, looking to the future. Computer, 39 (1):8–10, January 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2006:GSV

[1622] David Alan Grier. George Stibitz’s values and R.C. Archibald’s slide rule. Computer, 39(1):11–13, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2006:YAa

[1623] Anonymous. 32 & 16 years ago. Computer, 39(1):14–15, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2006/01/r1014.pdf.

Lawton:2006:IFM

[1624] George Lawton. Improved flash memory grows in popularity. Computer, 39 (1):16–18, January 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geer:2006:SCC

[1625] David Geer. Security of critical control systems sparks concern. Computer, 39(1):20–23, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2006/01/r1020.pdf.

Paulson:2006:NBa

[1626] Linda Dailey Paulson. News briefs. Computer, 39(1):24–26, January 2006.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/01/r1024.pdf.

Bowen:2006:TCF

[1627] Jonathan P. Bowen and Michael G.

Hinchey. Ten commandments of formal methods … ten years later. Computer, 39(1):40–48, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Colwell:2006:PCI

[1628] Robert P. Colwell. The Pentium Chronicles: Introduction. Computer, 39(1):49–54, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rottman:2006:SOE

[1629] Joseph W. Rottman. Successfully outsourcing embedded software development. Computer, 39(1):55–61, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cooke:2006:NEA

[1630] Daniel E. Cooke, Matt Barryl Michael Lowry, and Cordell Green. NASA’s exploration agenda and capability engineering. Computer, 39(1):63–73, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Svore:2006:LSA

[1631] Krysta M. Svore, Alfred V. Aho, Andrew W. Cross, Isaac Chuang, and Igor L. Markov. A layered software architecture for quantum computing design tools. Computer, 39(1):

74–83, January 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2006:CSO

[1632] Bob Ward. Computer Society offers full slate of programs in 2006. Computer, 39(1):87–91, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2006/01/r1087.pdf.

Anonymous:2006:CCa

[1633] Anonymous. Call and calendar. Computer, 39(1):92–95, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2006/01/r1092.pdf.

Anonymous:2006:Ba

[1634] Anonymous. Bookshelf. Computer, 39(1):96, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2006/01/r1096.pdf.

Anonymous:2006:Pa

[1635] Anonymous. Products. Computer, 39(1):97, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2006/01/r1097.pdf.

Bell:2006:PCS

[1636] Gordon Bell, Jim Gray, and Alex Szalay. Petascale computational systems. Computer, 39(1):110–112, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Huntley:2006:DVS

[1637] Christopher L. Huntley. A developmental view of system security. Computer, 39(1):113–114, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Morse:2006:ESI

[1638] Katherine L. Morse, Mike Lightner, Reed Little, Bob Lutz, and Roy Scrudder. Enabling simulation interoperability. Computer, 39(1):115–117, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Narayanan:2006:RCE

[1639] Vijaykrishnan Narayanan and Yuan Xie. Reliability concerns in embedded system designs. Computer, 39(1): 118–120, January 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Meyer:2006:URS

[1640] Bertrand Meyer. The unspoken revolution in software engineering. Computer, 39(1):124, 122–123, January 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2006:ASb

[1641] Anonymous. Article summaries. Computer, 39(2):4, February 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/02/r2004.pdf.

Cooper:2006:MOS

[1642] Deborah M. Cooper. Making our society even better. Computer, 39(2): 5–7, February 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
02/r2005.pdf.

Grier:2006:IC

[1643] David Alan Grier. The innovation curve. Computer, 39(2):8–10, February 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2006:YAb

[1644] Anonymous. 32 & 16 years ago. Computer, 39(2):11–12, February 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/02/r2011.pdf.

Emigh:2006:NFP

[1645] Jacqueline Emigh. New Flash player rises in the Web-video market. Computer, 39(2):14–16, February 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geer:2006:WSD

[1646] David Geer. Will software developers ride Ruby on Rails to success? Computer, 39(2):18–20, February 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/02/r2018.
pdf.

Paulson:2006:NBb

[1647] Linda Dailey Paulson. News briefs. Computer, 39(2):21–23, February 2006.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/02/r2021.pdf.

Schmidt:2006:GEI

[1648] Douglas C. Schmidt. Guest Editor’s introduction: Model-driven engineering. Computer, 39(2):25–31, February 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Balasubramanian:2006:DAU

[1649] Krishnakumar Balasubramanian, Aniruddha Gokhale, Gabor Karsai, Janos Sztipanovits, and Sandeep Neema. Developing applications using modeldriven design environments. Computer, 39(2):33–40, February 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Childs:2006:CCM

[1650] Adam Childs, Jesse Greenwald, Georg Jung, Matthew Hoosier, and John Hatcliff. CALM and Cadena: Metamodeling for component-based productline development. Computer, 39(2): 42–50, February 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gray:2006:ACE

[1651] Jeff Gray, Yuehua Lin, and Jing Zhang.

Automating change evolution in modeldriven engineering. Computer, 39(2): 51–58, February 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

France:2006:MDD

[1652] Robert B. France, Sudipto Ghosh, Trung Dinh-Trong, and Arnor Solberg.

Model-driven development using UML 2.0: Promises and pitfalls. Computer, 39(2):59–66, February 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bramberger:2006:DES

[1653] Michael Bramberger, Andreas Doblander, Arnold Maier, Bernhard Rinner, and Helmut Schwabach. Distributed embedded smart cameras for surveillance applications. Computer, 39(2): 68–75, February 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2006:IAF

[1654] Bob Ward. IEEE announces 2006 Fellows. Computer, 39(2):79–83, February 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2006/02/r2079.
                                        pdf.

Anonymous:2006:CCb

[1655] Anonymous. Call and calendar. Computer, 39(2):84–87, February 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2006/02/r2084.pdf.

Anonymous:2006:COa

[1656] Anonymous. Career opportunities. Computer, 39(2):88–94, February 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2006/02/r2088.pdf.

Anonymous:2006:Bb

[1657] Anonymous. Bookshelf. Computer, 39(2):95, February 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/02/r2095.pdf.

Weaver:2006:BA

[1658] Alfred C. Weaver. Biometric authentication. Computer, 39(2):96– 97, February 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Birman:2006:UWS

[1659] Ken Birman. The untrustworthy Web services revolution. Computer, 39(2): 98–100, February 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Macedonia:2006:GBE

[1660] Michael Macedonia. Google becomes an entertainment company. Computer, 39(2):101–103, February 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Newman:2006:RDL

[1661] Mark W. Newman, Trevor F. Smith, and Bill N. Schilit. Recipes for digital living. Computer, 39(2):104– 106, February 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hinchey:2006:SMS

[1662] Michael G. Hinchey and Roy Sterritt.

Self-managing software. Computer, 39 (2):107–109, February 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2006:FFC

[1663] Neville Holmes. The farmer, the fairies, and the computers. Computer, 39(2): 112, 110–111, February 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //bell.computer.org/dlcomments/;
                                         http://csdl.computer.org/comp/
                                        mags/co/2006/02/r2112.pdf.

Anonymous:2006:ASc

[1664] Anonymous. Article summaries. Computer, 39(3):4, March 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2006/03/r3004.pdf.

Anonymous:2006:Lb

[1665] Anonymous. Letters. Computer, 39 (3):6–7, March 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2006:LL

[1666] David Alan Grier. The lay of the land. Computer, 39(3):9–11, March 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2006:YAc

[1667] Anonymous. 32 & 16 years ago. Computer, 39(3):12–13, March 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2006/03/r3012.pdf.

Geer:2006:BBN

[1668] David Geer. Behavior-based network security goes mainstream. Computer, 39(3):14–17, March 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2006:NFC

[1669] Sixto Ortiz, Jr. Is near-field communication close to success? Computer, 39(3):18–20, March 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/03/r3018.pdf.

Paulson:2006:NBc

[1670] Linda Dailey Paulson. News briefs.

Computer, 39(3):21–24, March 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/03/r3021.pdf.

Roussos:2006:ERR

[1671] George Roussos. Enabling RFID in retail. Computer, 39(3):25–30, March 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ranganathan:2006:EAU

[1672] Parthasarathy Ranganathan, Erik Geelhoed, Meera Manahan, and Ken Nicholas. Energy-aware user interfaces and energy-adaptive displays. Computer, 39(3):31–38, March 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chittaro:2006:VIM

[1673] Luca Chittaro. Visualizing information on mobile devices. Computer, 39 (3):40–45, March 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tolia:2006:QIU

[1674] Niraj Tolia, David G. Andersen, and M. Satyanarayanan. Quantifying interactive user experience on thin clients. Computer, 39(3):46–52, March 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Italiano:2006:SOD

[1675] Isabel Cristina Italiano and João Eduardo Ferreira. Synchronization options for data warehouse designs. Computer, 39(3):53–57, March 2006. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Badra:2006:KEA

[1676] Mohamad Badra and Ibrahim Hajjeh. Key-exchange authentication using shared secrets. Computer, 39(3): 58–66, March 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2006:OAF

[1677] Bob Ward. Organizers announce field for CSIDC 2006. Computer, 39(3): 67–70, March 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
03/r3067.pdf.

Anonymous:2006:CCc

[1678] Anonymous. Call and calendar. Computer, 39(3):71–73, March 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/03/r3071.pdf.

Anonymous:2006:Pb

[1679] Anonymous. Products. Computer, 39(3):75, March 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
03/r3075.pdf.

Anonymous:2006:COb

[1680] Anonymous. Career opportunities.

Computer, 39(3):76–81, March 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/03/r3076.pdf.

Gilbert:2006:MCB

[1681] Juan E. Gilbert. Making a case for BPC. Computer, 39(3):83–86, March 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paul:2006:WN

[1682] JoAnn M. Paul. What’s in a name. Computer, 39(3):87–89, March 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/03/r3087.
pdf.

Walczak:2006:VME

[1683] Krzysztof Walczak, Wojciech Cellary, and Martin White. Virtual museum exhibitions. Computer, 39(3): 93–95, March 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Duval:2006:SUO

[1684] Erik Duval and Wayne Hodgins. Standardized uniqueness: Oxymoron or vision of the future? Computer, 39 (3):96–98, March 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bichler:2006:SOC

[1685] Martin Bichler and Kwei-Jay Lin. Service-oriented computing. Computer, 39(3):99–101, March 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kowalik:2006:AMC

[1686] Janusz Kowalik. The applied mathematics and computer science schism. Computer, 39(3):104, 102–103, March 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://bell.computer.
                                        org/dlcomments/;   http://csdl.
                                        computer.org/comp/mags/co/2006/
                                        03/r3104.pdf.

Anonymous:2006:ASd

[1687] Anonymous. Article summaries. Computer, 39(4):4, April 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2006/04/r4004.pdf.

Grier:2006:CA

[1688] David Alan Grier. Coming of age. Computer, 39(4):5–7, April 2006. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2006:YAd

[1689] Anonymous. 32 & 16 years ago. Computer, 39(4):8–9, April 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/04/r4008.pdf.

Greer:2006:DMF

[1690] David Greer. Digital music faces incompatible formats. Computer, 39 (4):10–12, April 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2006:TCM

[1691] Linda Dailey Paulson. TV comes to the mobile phone. Computer, 39 (4):13–16, April 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2006:NBd

[1692] Linda Daily Paulson. News briefs.

Computer, 39(4):17–19, April 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/04/r4017.pdf.

Meyer:2006:TRU

[1693] Bertrand Meyer. Testable, reusable units of cognition. Computer, 39 (4):20–24, April 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schrenker:2006:GEI

[1694] Richard A. Schrenker. Guest Editor’s introduction: Software engineering for future healthcare and clinical systems. Computer, 39(4):26–32, April 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:2006:HCM

[1695] Insup Lee, George J. Pappas, Rance Cleaveland, John Hatcliff, Bruce H. Krogh, Peter Lee, Harvey Rubin, and Lui Sha. High-confidence medical device software and systems. Computer, 39(4):33–38, April 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rakitin:2006:CDS

[1696] Steven R. Rakitin. Coping with defective software in medical devices. Computer, 39(4):40–45, April 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gary:2006:IOS

[1697] Kevin Gary, Luis Ibáñez, Stephen Aylward, David Gobbi, M. Brian Blake, and Keven Cleary. IGSTK: An open source software toolkit for image-guided surgery. Computer, 39 (4):46–53, April 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gehlot:2006:EPS

[1698] Vijay Gehlot and Elliot B. Sloane. Ensuring patient safety in wireless medical device networks. Computer, 39 (4):54–60, April 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jetley:2006:FMA

[1699] Raoul Jetley, S. Purushothaman Iyer, and Paul L. Jones. A formal methods approach to medical device review. Computer, 39(4):61–67, April 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2006:COc

[1700] Anonymous. Career opportunities.

Computer, 39(4):68–73, April 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/04/r4068.pdf.

Ward:2006:CSS

[1701] Bob Ward. Computer Society seeks nominations for Technology of the Decade Award. Computer, 39(4): 75–77, April 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
04/r4075.pdf.

Anonymous:2006:CCd

[1702] Anonymous. Call and calendar.

Computer, 39(4):78–79, April 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/04/r4078.pdf.

Anonymous:2006:Pc

[1703] Anonymous. Products. Computer, 39(4):83, April 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
04/r4083.pdf.

Anonymous:2006:Bc

[1704] Anonymous. Bookshelf. Computer, 39(4):84, April 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
04/r4084.pdf.

Subramanya:2006:UIM

[1705] S. R. Subramanya and Byung K. Yi. User interfaces for mobile content. Computer, 39(4):85–87, April 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weaver:2006:SSL

[1706] Alfred C. Weaver. Secure sockets layer. Computer, 39(4):88–90, April 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hasselbring:2006:TTS

[1707] Wilhelm Hasselbring and Ralf Reussner. Toward trustworthy software systems. Computer, 39(4):91–92, April 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:2006:VSG

[1708] Cliff Jones, Peter O’Hearn, and Jim Woodcock. Verified software: A grand challenge. Computer, 39(4): 93–95, April 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Baudisch:2006:ILD

[1709] Patrick Baudisch. Interacting with large displays. Computer, 39(4):96–97, April 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2006:PPP

[1710] Neville Holmes. In praise of professional precision. Computer, 39(4): 100, 98–99, April 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //bell.computer.org/dlcomments/;

 http://csdl.computer.org/comp/
mags/co/2006/04/r4100.pdf.

Anonymous:2006:ASe

[1711] Anonymous. Article summaries. Computer, 39(5):4, May 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/05/r5004.pdf.

Anonymous:2006:Lc

[1712] Anonymous. Letters. Computer, 39 (5):6–7, May 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
05/r5006.pdf.

Grier:2006:WHS

[1713] David Alan Grier. A winter of hope and a spring of despair. Computer, 39(5):8–10, May 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2006:YAe

[1714] Anonymous. 32 & 16 years ago. Computer, 39(5):11–12, May 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/05/r5011.pdf.

Leavitt:2006:RTW

[1715] Neal Leavitt. Recommendation technology: Will it boost E-commerce? Computer, 39(5):13–16, May 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geer:2006:HGR

[1716] David Geer. Hackers get to the root of the problem. Computer, 39 (5):17–19, May 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2006:NBe

[1717] Linda Daily Paulson. News brief. Computer, 39(5):20–22, May 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2006/05/r5020.pdf.

Nezhad:2006:WSI

[1718] Hamid R. Motahari Nezhad, Boualem Benatallah, Fabio Casati, and Farouk Toumani. Web services interoperability specifications. Computer, 39 (5):24–32, May 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:2006:PT

[1719] Edward A. Lee. The problem with threads. Computer, 39(5):33–42, May 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tanenbaum:2006:CWM

[1720] Andrew S. Tanenbaum, Jorrit N. Herder, and Herbert Bos. Can we make operating systems reliable and secure? Computer, 39(5):44–51, May 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Buttazzo:2006:ASR

[1721] Giorgio Buttazzo. Achieving scalability in real-time systems. Computer, 39(5):54–59, May 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rae:2006:FES

[1722] Andrew Rae, Colin Fidge, and Luke Wildman. Fault evaluation for securitycritical communications devices. Computer, 39(5):61–68, May 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2006:CSA

[1723] Bob Ward. Computer Society announces summer and fall conferences. Computer, 39(5):69–73, May 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/05/r5069.
pdf.

Anonymous:2006:CCe

[1724] Anonymous. Call and calendar. Computer, 39(5):74–76, May 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/05/r5074.pdf.

Anonymous:2006:Bd

[1725] Anonymous. Bookshelf. Computer, 39(5):77, May 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
05/r5077.pdf.

Anonymous:2006:COd

[1726] Anonymous. Career opportunities.

Computer, 39(5):78–82, May 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2006/05/r5078.pdf.

Williams:2006:DNS

[1727] Laurie Williams. Debunking the nerd stereotype with pair programming. Computer, 39(5):83–85, May 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Crowston:2006:AHO

[1728] Kevin Crowston and James Howison. Assessing the health of open source communities. Computer, 39 (5):89–91, May 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Doran:2006:IPS

[1729] Teresa Doran. IEEE 1220: For practical systems engineering. Computer, 39(5):92–94, May 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lindahl:2006:DSE

[1730] Michael Lindahl. The device software engineer’s best friend. Computer, 39 (5):95–97, May 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:2006:SAS

[1731] Keith Smith. Simplifying Ajax-style Web development. Computer, 39(5): 98–101, May 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Milutinovic:2006:OPN

[1732] Veljko Milutinovic. Our profession needs a reminder. Computer, 39(5):

104, 102–103, May 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//bell.computer.org/dlcomments/;
http://csdl.computer.org/comp/
mags/co/2006/05/r5104.pdf.

Anonymous:2006:ASf

[1733] Anonymous. Article summaries. Computer, 39(6):4, June 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/06/r6004.pdf.

Anonymous:2006:Ld

[1734] Anonymous. Letters. Computer, 39 (6):5–6, June 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
06/r6005.pdf.

Grier:2006:CI

[1735] David Alan Grier. The captured imagination. Computer, 39(6):7–9, June 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/06/r6007.
pdf.

Anonymous:2006:YAf

[1736] Anonymous. 32 & 16 years ago. Computer, 39(6):10–11, June 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/06/r6010.pdf.

Lawton:2006:MVR

[1737] George Lawton. Making virtual reality more accessible. Computer, 39 (6):12–15, June 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
06/r6012.pdf.

Geer:2006:SOA

[1738] David Geer. Silicon optics aims to combine the best of both worlds. Computer, 39(6):16–19, June 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/06/r6016.pdf.

Paulson:2006:WTL

[1739] Linda Dailey Paulson. Wireless technology leaps the last mile. Computer, 39(6):20–22, June 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/06/r6020.pdf.

Dedeke:2006:QUC

[1740] Adenekan (Nick) Dedeke and Benjamin Lieberman. Qualifying use case diagram associations. Computer, 39 (6):23–29, June 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2006:GEI

[1741] Michael Zyda. Guest Editor’s introduction: Educating the next generation of game developers. Computer, 39(6):30–34, June 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/06/r6030.pdf.

Fullerton:2006:PCG

[1742] Tracy Fullerton. Play-centric games education. Computer, 39(6):36–42, June 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Murray:2006:GDE

[1743] Janet Murray, Ian Bogost, Michael Mateas, and Michael Nitsche. Game design education: Integrating computation and culture. Computer, 39 (6):43–51, June 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Argent:2006:BGD

[1744] Lawrence Argent, Bill Depper, Rafael Fajardo, Sarah Gjertson, Scott T. Leutenegger, Mario A. Lopez, and Jeff Rutenbeck. Building a game development program. Computer, 39 (6):52–60, June 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Horswill:2006:EAT

[1745] Ian Horswill and Marlena Novak.

Evolving the artist-technologist. Computer, 39(6):62–69, June 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2006:WBH

[1746] Anonymous. Weyuker and Barbacci honored with Society awards. Computer, 39(6):73–74, June 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/06/r6073.pdf.

Anonymous:2006:CCf

[1747] Anonymous. Call and calendar. Computer, 39(6):75–76, June 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2006/06/r6075.pdf.

Anonymous:2006:Be

[1748] Anonymous. Bookshelf. Computer, 39(6):77, June 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2006/
                                        06/r6077.pdf.

Anonymous:2006:COe

[1749] Anonymous. Career opportunities. Computer, 39(6):78–81, June 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2006/06/r6078.pdf.

Macedonia:2006:ETW

[1750] Michael Macedonia. It’s the end of the tube as we know it. Computer, 39(6):83–85, June 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hawking:2006:WSEa

[1751] David Hawking. Web search engines: Part 1. Computer, 39(6):86–88, June 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kagal:2006:SPC

[1752] Lalana Kagal, Tim Finin, Anupam Joshi, and Sol Greenspan. Security and privacy challenges in open and dynamic environments. Computer, 39 (6):89–91, June 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vonAhn:2006:GP

[1753] Luis von Ahn. Games with a purpose. Computer, 39(6):92–94, June 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holzmann:2006:PRD

[1754] Gerard J. Holzmann. The power of 10: Rules for developing safety-critical code. Computer, 39(6):95–97, June 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2006:DDS

[1755] Neville Holmes. The data doughnut and the software hole. Computer, 39 (6):100, 98–99, June 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

bell.computer.org/dlcomments/.

Anonymous:2006:ASg

[1756] Anonymous. Article summaries. Computer, 39(7):4, July 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/07/r7004.pdf.

Anonymous:2006:Le

[1757] Anonymous. Letters. Computer, 39 (7):5, 7, July 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
07/r7005.pdf.

Grier:2006:AGD

[1758] David Alan Grier. Across the great divide. Computer, 39(7):8–10, July 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2006:YAg

[1759] Anonymous. 32 & 16 years ago. Computer, 39(7):11–12, July 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/07/r7011.pdf.

Geer:2006:USE

[1760] David Geer. UWB standardization effort ends in controversy. Computer, 39(7):13–16, July 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
07/r7013.pdf.

Ortiz:2006:HSR

[1761] Sixto Ortiz, Jr. How secure is RFID? Computer, 39(7):17–19, July 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/07/r7017.
pdf.

Paulson:2006:NBf

[1762] Linda Dailey Paulson. News briefs. Computer, 39(7):20–22, July 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/07/r7020.pdf.

Meyer:2006:CVE

[1763] Bertrand Meyer and Karine Arnout. Componentization: The visitor example. Computer, 39(7):23–30, July 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mann:2006:TPE

[1764] Zoltán Adám Mann. Three public enemies: Cut, copy, and paste. Computer, 39(7):31–35, July 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dedrick:2006:PPK

[1765] Jason Dedrick and Kenneth L. Kraemer. Is production pulling knowledge work to China? A study of the notebook PC industry. Computer, 39 (7):36–42, July 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Konana:2006:CIS

[1766] Prabhudev Konana. Can Indian software firms compete with the global giants? Computer, 39(7):43–47, July 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Andel:2006:CMS

[1767] Todd R. Andel and Alec Yasinac.

On the credibility of Manet simulations. Computer, 39(7):48–54, July 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tseng:2006:WSN

[1768] Yu-Chee Tseng, Meng-Shiuan Pan, and Yuen-Yung Tsai. Wireless sensor networks for emergency navigation. Computer, 39(7):55–62, July 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2006:CSH

[1769] Bob Ward. Computer Society honors Oklahoma teen at Intel Science Fair. Computer, 39(7):67–70, July 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/07/r7067.
pdf.

Anonymous:2006:CCg

[1770] Anonymous. Call and calendar. Computer, 39(7):71–74, July 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/07/r7071.pdf.

Anonymous:2006:Bf

[1771] Anonymous. Bookshelf. Computer, 39(7):75, July 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
07/r7075.pdf.

Anonymous:2006:COf

[1772] Anonymous. Career opportunities. Computer, 39(7):76–80, July 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/07/r7076.pdf.

Anonymous:2006:Pd

[1773] Anonymous. Products. Computer, 39(7):81, July 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
07/r7081.pdf.

Robson:2006:GFS

[1774] Robby Robson. Globalization and the future of standardization. Computer, 39(7):82–84, July 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Narasimhan:2006:AET

[1775] Priya Narasimhan. Assistive embedded technologies. Computer, 39(7): 85–87, July 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jansen:2006:PS

[1776] Bernard J. Jansen. Paid search. Computer, 39(7):88–90, July 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brodine:2006:MSS

[1777] Dick Brodine. Mathematical server sizing. Computer, 39(7):91–93, July 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mahmoud:2006:MSA

[1778] Qusay H. Mahmoud and Leslie Yu.

Making software agents user-friendly. Computer, 39(7):96, 94–95, July 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://bell.computer.
org/dlcomments/.

Anonymous:2006:ASh

[1779] Anonymous. Article summaries. Computer, 39(8):4, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/08/r8004.pdf.

Anonymous:2006:Lf

[1780] Anonymous. Letters. Computer, 39 (8):6–7, August 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
08/r8006.pdf.

Grier:2006:PH

[1781] David Alan Grier. The print of the hand. Computer, 39(8):8–10, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/08/r8008.
pdf.

Anonymous:2006:YAh

[1782] Anonymous. 32 & 16 years ago. Computer, 39(8):11–12, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/08/r8011.pdf.

Ward:2006:NST

[1783] Bob Ward. NC State team wins $20,000 CSIDC prize. Computer, 39 (8):13–16, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/08/r8013.pdf.

Paulson:2006:SSN

[1784] Linda Dailey Paulson. Services science: A new field for today’s economy. Computer, 39(8):18–21, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/08/r8018.
pdf.

Geer:2006:VUT

[1785] David Geer. Vendors upgrade their physics processing to improve gaming. Computer, 39(8):22–24, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/08/r8022.
pdf.

Paulson:2006:NBg

[1786] Linda Dailey Paulson. News briefs.

Computer, 39(8):25–27, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/08/r8025.pdf.

Bimber:2006:GEI

[1787] Oliver Bimber. Guest Editor’s introduction: Computational photography— the next big step. Computer, 39 (8):28–29, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/08/r8028.pdf.

Nayar:2006:CCR

[1788] Shree K. Nayar. Computational cameras: Redefining the image. Computer, 39(8):30–38, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cohen:2006:MC

[1789] Michael F. Cohen and Richard Szeliski.

The moment camera. Computer, 39 (8):40–45, August 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Levoy:2006:LFC

[1790] Marc Levoy. Light fields and computational imaging. Computer, 39(8): 46–55, August 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Debevec:2006:VCR

[1791] Paul Debevec. Virtual cinematography: Relighting through computation. Computer, 39(8):57–65, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kong:2006:CSF

[1792] Joseph S. Kong, Behnam A. Rezaei, Nima Sarshar, Vwani P. Roychowdhury, and P. Oscar Boykin. Collaborative spam filtering using E-mail networks. Computer, 39(8):67–73, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2006:CSC

[1793] Bob Ward. Computer Society connection. Computer, 39(8):77–79, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2006/08/r8077.
                                        pdf.

Anonymous:2006:CCh

[1794] Anonymous. Call and calendar. Computer, 39(8):80–81, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2006/08/r8080.pdf.

Anonymous:2006:COg

[1795] Anonymous. Career opportunities. Computer, 39(8):82–85, August 2006.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/08/r8082.pdf.

Anonymous:2006:Pe

[1796] Anonymous. Products. Computer, 39 (8):86, August 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
08/r8086.pdf.

Anonymous:2006:Bg

[1797] Anonymous. Bookshelf. Computer, 39(8):87, August 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
08/r8087.pdf.

Hawking:2006:WSEb

[1798] David Hawking. Web search engines:

Part 2. Computer, 39(8):88–90, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Macedonia:2006:WGP

[1799] Michael Macedonia. Why graphics power is revolutionizing physics. Computer, 39(8):91–92, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Veeramani:2006:SCV

[1800] Narayan Veeramani. Smart clients versus Web forms. Computer, 39(8): 93–95, August 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stoneburner:2006:TUS

[1801] Gary Stoneburner. Toward a unified security-safety model. Computer, 39 (8):96–97, August 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tanenbaum:2006:TSN

[1802] Andrew S. Tanenbaum, Chandana Gamage, and Bruno Crispo. Taking sensor networks from the lab to the jungle. Computer, 39(8):98–100, August 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2006:QES

[1803] Neville Holmes. Quasireform of English spelling. Computer, 39(8):104, 102–103, August 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://bell.

computer.org/dlcomments/.

Anonymous:2006:ASi

[1804] Anonymous. Article summaries. Computer, 39(9):4, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/09/r9004.pdf.

Anonymous:2006:Lg

[1805] Anonymous. Letters. Computer, 39 (9):5, 7–8, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/09/r9005.pdf.

Grier:2006:EB

[1806] David Alan Grier. The empty box. Computer, 39(9):9–13, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/09/r9009.
pdf.

Anonymous:2006:YAi

[1807] Anonymous. 32 & 16 years ago.

Computer, 39(9):12–13, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/09/r9012.
pdf.

Lawton:2006:MBI

[1808] George Lawton. Making business intelligence more useful. Computer, 39 (9):14–16, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/09/r9014.pdf.

Leavitt:2006:BW

[1809] Neal Leavitt. Browsing the 3D Web.

Computer, 39(9):18–21, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/09/r9018.
pdf.

Paulson:2006:NBh

[1810] Linda Dailey Paulson. News briefs.

Computer, 39(9):22–24, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/09/r9022.
pdf.

Chin:2006:DCB

[1811] George Chin, Jr., Eric G. Stephan, Debbie K. Gracio, Olga A. Kuchar, Paul D. Whitney, and Karen L. Schuchardt. Developing concept-based user interfaces for scientific computing. Computer, 39(9):26–34, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shklovski:2006:GEI

[1812] Irina Shklovski and Michele F. Chang. Guest Editors’ introduction: Urban computing—navigating space and context. Computer, 39(9):36–37, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/09/r9036.
pdf.

Williams:2006:ICC

[1813] Amanda Williams and Paul Dourish. Imagining the city: The cultural dimensions of urban computing. Computer, 39(9):38–43, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Foth:2006:FSN

[1814] Marcus Foth. Facilitating social networking in inner-city neighborhoods. Computer, 39(9):44–50, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kostakos:2006:DUP

[1815] Vassilis Kostakos, Eamonn O’Neill, and Alan Penn. Designing urban pervasive systems. Computer, 39(9):52– 59, September 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kjeldskov:2006:PPC

[1816] Jesper Kjeldskov and Jeni Paay. Public pervasive computing: Making the invisible visible. Computer, 39(9):60– 65, September 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Davis:2006:SUP

[1817] Janet Davis, Peyina Lin, Alan Borning, Batya Friedman, Peter H. Kahn, Jr., and Paul A. Waddell. Simulations for urban planning: Designing for human values. Computer, 39(9): 66–72, September 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2006:ICS

[1818] Anonymous. IEEE Computer Society election. Computer, 39(9):73– 80, September 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
09/r9073.pdf.

Anonymous:2006:IPE

[1819] Anonymous. IEEE President-Elect candidates respond to questions on governance issues. Computer, 39(9): 81–86, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/09/r9081.pdf.

Anonymous:2006:COh

[1820] Anonymous. Career opportunities.

Computer, 39(9):88–93, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2006/09/r9088.
                                        pdf.

Anonymous:2006:Pf

[1821] Anonymous. Products. Computer, 39(9):94, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2006/09/r9094.pdf.

Anonymous:2006:Bh

[1822] Anonymous. Bookshelf. Computer, 39(9):95, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2006/09/r9095.pdf.

Anonymous:2006:CCi

[1823] Anonymous. Call and calendar. Computer, 39(9):99–101, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2006/09/r9099.
                                        pdf.

Anonymous:2006:GDP

[1824] Anonymous. GENI design principles. Computer, 39(9):102–105, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moore:2006:CSS

[1825] James W. Moore. Converging software and systems engineering standards. Computer, 39(9):106–108, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wolf:2006:HMS

[1826] Wayne Wolf. A half-million strong at least. Computer, 39(9):109–110, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Elhanany:2006:TRF

[1827] Itamar Elhanany and Kurt Busch. Is time ripe for fabric on a chip? Computer, 39(9):112–113, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Olsen:2006:CIF

[1828] Kai A. Olsen. Computer intelligence and formalization. Computer, 39(9):116, 114–115, September 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://bell.computer.
org/dlcomments/.

Anonymous:2006:ASj

[1829] Anonymous. Article summaries. Computer, 39(10):4, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/10/rx004.pdf.

Anonymous:2006:Lh

[1830] Anonymous. Letters. Computer, 39 (10):6, October 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
10/rx006.pdf.

Grier:2006:LBL

[1831] David Alan Grier. The language of bad love. Computer, 39(10):7–9, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/10/rx007.
pdf.

Anonymous:2006:YAj

[1832] Anonymous. 32 & 16 years ago. Computer, 39(10):10–11, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/10/rx010.pdf.

Ortiz:2006:PCG

[1833] Sixto Ortiz, Jr. Phone companies get into the TV business. Computer, 39 (10):12–15, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/10/rx012.pdf.

Vaughan-Nichols:2006:WNW

[1834] Steven J. Vaughan-Nichols. Will the new wi-fi fly? Computer, 39(10): 16–18, October 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
10/rx016.pdf.

Paulson:2006:NBi

[1835] Linda Dailey Paulson. News briefs. Computer, 39(10):19–21, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/10/rx019.pdf.

Carver:2006:SII

[1836] Doris Carver, Ron Hoelzeman, Jim Aylor, and Mike Hinchey. Special issue introduction: The IEEE Computer Society’s 60th anniversary. Computer, 39(10):22–25, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bishop:2006:CPD

[1837] Judith Bishop and Nigel Horspool.

Cross-platform development: Software that lasts. Computer, 39(10):26– 35, October 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Baresi:2006:TOW

[1838] Luciano Baresi, Elisabetta Di Nitto, and Carlo Ghezzi. Toward openworld software: Issue and challenges. Computer, 39(10):36–43, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Margaria:2006:SEL

[1839] Tiziana Margaria and Bernhard Steffen. Service engineering: Linking business and IT. Computer, 39(10): 45–55, October 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woodcock:2006:FSV

[1840] Jim Woodcock. First steps in the Verified Software Grand Challenge. Computer, 39(10):57–64, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jackson:2006:WCW

[1841] Michael Jackson. What can we expect from program verification? Computer, 39(10):65–71, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Broy:2006:GCI

[1842] Manfred Broy. The ’Grand Challenge’ in informatics: Engineering softwareintensive systems. Computer, 39(10): 72–80, October 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gries:2006:WWL

[1843] David Gries. What have we not learned about teaching programming? Computer, 39(10):81–82, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2006:COi

[1844] Anonymous. Career opportunities. Computer, 39(10):84–91, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2006/10/rx084.pdf.

Anonymous:2006:Bi

[1845] Anonymous. Bookshelf. Computer, 39(10):92, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2006/10/rx092.pdf.

Anonymous:2006:Pg

[1846] Anonymous. Products. Computer, 39(10):93, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2006/10/rx093.pdf.

Anonymous:2006:CCj

[1847] Anonymous. Call and calendar. Computer, 39(10):97–99, October 2006.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/10/rx097.pdf.

Patterson:2006:TUP

[1848] Laurie J. Patterson. The technology underlying podcasts. Computer, 39 (10):103–105, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Arazi:2006:ESN

[1849] Benjamin Arazi. Enhancing security with nanotechnology. Computer, 39 (10):106–107, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Naaman:2006:EW

[1850] Mor Naaman. Eyes on the world. Computer, 39(10):108–111, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Menzies:2006:MSR

[1851] Tim Menzies and Julian Richardson.

Making sense of requirements, sooner. Computer, 39(10):112–114, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Macedonia:2006:MBH

[1852] Michael Macedonia. Moving beyond HD. Computer, 39(10):115–117, October 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2006:TDS

[1853] Neville Holmes. The technology of data stashing. Computer, 39(10):120, 118–119, October 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://bell.

computer.org/dlcomments/.

Anonymous:2006:ASk

[1854] Anonymous. Article summaries. Computer, 39(11):4, November 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2006/11/ry004.pdf.

Grier:2006:BBD

[1855] David Alan Grier. The benefits of being different. Computer, 39(11): 6–8, November 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
11/ry006.pdf.

Anonymous:2006:YAk

[1856] Anonymous. 32 & 16 years ago. Computer, 39(11):9–10, November 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/11/ry009.
pdf.

Vaughan-Nichols:2006:NAV

[1857] Stephen J. Vaughan-Nichols. New approach to virtualization is a lightweight. Computer, 39(11):12– 14, November 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
11/ry012.pdf.

Paulson:2006:SRM

[1858] Linda Daily Paulson. Speech recognition moves from software to hardware. Computer, 39(11):15–18, November 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/11/ry015.
pdf.

Paulson:2006:NBj

[1859] Linda Dailey Paulson. News briefs.

Computer, 39(11):19–21, November 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/11/ry019.
pdf.

Yi:2006:FSF

[1860] Joshua J. Yi, Lieven Eeckhout, David J. Lilja, Brad Calder, Lizy K. John, and James E. Smith. The future of simulation: A field of dreams. Computer, 39(11):22–29, November 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Foster:2006:SSL

[1861] Ian Foster and Carl Kesselman. Scaling system-level science: Scientific exploration and IT implications. Computer, 39(11):31–39, November 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sloot:2006:MMD

[1862] Peter M. A. Sloot, Alfredo Tirado-Ramos, Ilkay Altintas, Marian Bubak, and Charles Boucher. From molecule to man: Decision support in individualized E-health. Computer, 39(11):

40–46, November 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hunter:2006:MMP

[1863] Peter J. Hunter, Wilfred W. Li, Andrew D. McCulloch, and Denis Noble. Multiscale modeling: Physiome project standards, tools, and databases. Computer, 39(11):48–54, November 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Plale:2006:CLA

[1864] Beth Plale, Dennis Gannon, Jerry Brotzge, Kelvin Droegemeier, Jim Kurose, David McLaughlin, Robert Wilhelmson, Sara Graves, Mohan Ramamurthy, Richard D. Clark, Sepi Yalda, Daniel A. Reed, Everette Joseph, and V. Chandrasekar. CASA and LEAD: Adaptive cyberinfrastructure for real-time multiscale weather forecasting. Computer, 39(11):56– 64, November 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:2006:DIA

[1865] Ying-Dar Lin, Chih-Wei Jan, Po-Ching Lin, and Yuan-Cheng Lai. Designing an integrated architecture for network content security gateways. Computer, 39(11):66–72, November 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2006:Ph

[1866] Anonymous. Products. Computer, 39(11):85, November 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/11/ry085.pdf.

Anonymous:2006:Bj

[1867] Anonymous. Bookshelf. Computer, 39(11):86, November 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/11/ry086.pdf.

Ward:2006:CSM

[1868] Bob Ward. Computer Society magazines plan 2007 coverage. Computer, 39 (11):87–90, November 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/11/ry087.pdf.

Anonymous:2006:CCk

[1869] Anonymous. Call and calendar.

Computer, 39(11):91–93, November 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/11/ry091.
pdf.

Beckwith:2006:GHW

[1870] Laura Beckwith, Margaret Burnett, Valentina Grigoreanu, and Susan Wiedenbeck. Gender HCI: What about the software? Computer, 39(11):97– 101, November 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Trcek:2006:SMR

[1871] Denis Trcek. Security models: Refocusing on the human factor. Computer, 39(11):103–104, November 2006.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Land:2006:UIS

[1872] Susan K. (Kathy) Land and Maj. Bret Wilson. Using IEEE Standards to support America’s Army gaming development. Computer, 39(11):105– 107, November 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tanurhan:2006:PFQ

[1873] Yankin Tanurhan. Processors and FP-GAs quo vadis? Computer, 39(11): 108–110, November 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kim:2006:SGS

[1874] Jong-Weon Kim, Kyu-Tae Kim, and Jong-Uk Choi. Securing E-government services. Computer, 39(11):111–112, November 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jenkins:2006:CI

[1875] Stephen Jenkins. Concerning interruptions. Computer, 39(11):116, 114– 115, November 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://bell.

                                        computer.org/dlcomments/.

Anonymous:2006:ASl

[1876] Anonymous. Article summaries. Computer, 39(12):4, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2006/12/rz004.pdf.

Anonymous:2006:Li

[1877] Anonymous. Letters. Computer, 39(12):6, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/12/rz006.pdf.

Carver:2006:EG

[1878] Doris L. Carver. An expression of gratitude. Computer, 39(12):7, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/12/rz007.
pdf.

Cooper:2006:RSO

[1879] Deborah M. Cooper. Reaching out to shape our future. Computer, 39 (12):8–10, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/12/rz008.pdf.

Grier:2006:SCJ

[1880] David Alan Grier. Songs of comfort and joy. Computer, 39(12):11– 13, December 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
12/rz011.pdf.

Anonymous:2006:YAl

[1881] Anonymous. 32 & 16 years ago.

Computer, 39(12):14–15, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/12/rz014.
pdf.

Vaughan-Nichols:2006:RMW

[1882] Steven J. Vaughan-Nichols. Researchers make Web searches more intelligent. Computer, 39(12):16–18, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2006/12/rz016.
                                        pdf.

Lawton:2006:WTT

[1883] George Lawton. Working today on tomorrow’s storage technology. Computer, 39(12):19–22, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2006/12/rz019.
                                        pdf.

Paulson:2006:NBk

[1884] Linda Dailey Paulson. News briefs. Computer, 39(12):23–25, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2006/12/rz023.
                                        pdf.

Seetharaman:2006:UVC

[1885] Guna Seetharaman, Arun Lakhotia, and Erik Philip Blasch. Unmanned vehicles come of age: The DARPA Grand Challenge. Computer, 39(12): 26–29, December 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2006/
                                        12/rz026.pdf.

Bertozzi:2006:VEV

[1886] Massimo Bertozzi, Alberto Broggi, and Alessandra Fascioli. VisLab and the evolution of vision-based UGVs. Computer, 39(12):31–38, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Touchton:2006:PPA

[1887] Bob Touchton, Tom Galluzzo, Danny Kent, and Carl Crane. Perception and planning architecture for autonomous ground vehicles. Computer, 39(12): 40–47, December 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Urmson:2006:TDS

[1888] Chris Urmson, William ‘Red’ Whittaker, Sam Harbaugh, Michael Clark, and Phillip Koon. Testing driver skill for high-speed autonomous vehicles. Computer, 39(12):48–51, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:2006:DH

[1889] Isaac Miller, Ephrahim Garcia, and Mark Campbell. To drive is human. Computer, 39(12):52–56, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lombardi:2006:IBC

[1890] Paolo Lombardi, Bertrand Zavidovique, and Michael Talbert. On the importance of being contextual. Computer, 39(12):57–61, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McDowell:2006:MBS

[1891] Patrick McDowell, Brian S. Bourgeois, Donald A. Sofge, and S. S. Iyengar.

Memory-based in situ learning for unmanned vehicles. Computer, 39(12): 62–66, December 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Srini:2006:VSA

[1892] Vason P. Srini. A vision for supporting autonomous navigation in urban environments. Computer, 39(12): 68–77, December 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2006:CTE

[1893] Bob Ward. Computer thanks its expert reviewers. Computer, 39(12):78– 80, December 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2006/
12/rz078.pdf.

Ward:2006:KVC

[1894] Bob Ward. Kasturi voted 2007 Computer Society President-Elect: Board of Governors members and new Vice Presidents also named. Computer, 39 (12):81–82, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/12/rz081.pdf.

Anonymous:2006:CCl

[1895] Anonymous. Call and calendar. Computer, 39(12):83–85, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/12/rz083.
pdf.

Anonymous:2006:CAI

[1896] Anonymous. Computer 2006 annual index, volume 39. Computer, 39(12): 90–101, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/12/rz090.pdf.

Anonymous:2006:COj

[1897] Anonymous. Career opportunities.

Computer, 39(12):102–122, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2006/12/rz102.
pdf.

Anonymous:2006:Bk

[1898] Anonymous. Bookshelf. Computer, 39(12):124, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/12/rz124.pdf.

Anonymous:2006:Pi

[1899] Anonymous. Products. Computer, 39(12):125, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2006/12/rz125.pdf.

Gonzalez-Escribano:2006:SP

[1900] Arturo González-Escribano and Diego R.

Llanos. Speculative parallelization. Computer, 39(12):126–128, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gershon:2006:GTD

[1901] Nahum Gershon, Ben Sawyer, and J. R. (Jim) Parker. Games and technology: Developing synergy. Computer, 39(12):129–130, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neill:2006:PDD

[1902] Colin J. Neill and Philip A. Laplante. Paying down design debt with strategic refactoring. Computer, 39(12):131– 134, December 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Covell:2006:DAV

[1903] Michele Covell, Shumeet Baluja, and Michael Fink. Detecting ads in video streams using acoustic and visual cues. Computer, 39(12):135–137, December 2006. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Santini:2006:SWT

[1904] Simone Santini. Standards: What are they good for? Computer, 39(12): 140, 139, December 2006. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://bell.

                                        computer.org/dlcomments/.

Anonymous:2007:ASa

[1905] Anonymous. Article summaries. Computer, 40(1):4, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2007/01/r1004.pdf.

Anonymous:2007:La

[1906] Anonymous. Letters. Computer, 40 (1):6, January 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
01/r1006.pdf.

Chang:2007:MVC

[1907] Carl K. Chang. My vision for computer. Computer, 40(1):7–8, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/01/r1007.
pdf.

Williams:2007:YD

[1908] Michael R. Williams. A year of decision. Computer, 40(1):9–11, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/01/r1009.
pdf.

Grier:2007:WF

[1909] David Alan Grier. The wave of the future. Computer, 40(1):12–14, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/01/r1012.
pdf.

Anonymous:2007:YAa

[1910] Anonymous. 32 & 16 years ago.

Computer, 40(1):15–16, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/01/r1015.pdf.

Ortiz:2007:BCI

[1911] Sixto Ortiz, Jr. Brain-computer interfaces: Where human and machine meet. Computer, 40(1):17–21, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/01/r1017.
pdf.

Paulson:2007:NBa

[1912] Linda Dailey Paulson. News briefs. Computer, 40(1):22–24, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/01/r1022.pdf.

Bahar:2007:ASN

[1913] R. Iris Bahar, Dan Hammerstrom, Justin Harlow, William H. Joyner, Jr., Clifford Lau, Diana Marculescu, Alex Orailoglu, and Massoud Pedram. Architectures for silicon nanoelectronics and beyond. Computer, 40(1): 25–33, January 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fiadeiro:2007:DSS

[1914] José Luiz Fiadeiro. Designing for software’s social complexity. Computer, 40 (1):34–39, January 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parameswaran:2007:RIB

[1915] Manoj Parameswaran, Xia Zhao, Andrew B. Whinston, and Fang Fang. Reengineering the Internet for better security. Computer, 40(1):40–44, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dudek:2007:AAA

[1916] Gregory Dudek, Philippe Giguere, Chris Prahacs, Shane Saunderson, Junaed Sattar, Luz-Abril Torres-Mendez, Michael Jenkin, Andrew German, Andrew Hogue, Arlene Ripsman, Jim Zacher, Evangelos Milios, Hui Liu, Pifu Zhang, Marti Buehler, and Christina Georgiades. AQUA: An amphibious autonomous robot. Computer, 40(1): 46–53, January 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Menzies:2007:STA

[1917] Tim Menzies, David Owen, and Julian Richardson. The strangest thing about software. Computer, 40(1): 54–60, January 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cybenko:2007:PQS

[1918] George Cybenko and Vincent H. Berk.

Process query systems. Computer, 40 (1):62–70, January 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Spohrer:2007:STS

[1919] Jim Spohrer, Paul P. Maglio, John Bailey, and Daniel Gruhl. Steps toward a science of service systems. Computer, 40(1):71–77, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2007:CCa

[1920] Anonymous. Call and calendar. Computer, 40(1):81–83, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2007/01/r1081.pdf.

Ward:2007:CSCa

[1921] Bob Ward. Computer Society connection. Computer, 40(1):84–93, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2007/01/r1084.
                                        pdf.

Anonymous:2007:COa

[1922] Anonymous. Career opportunities. Computer, 40(1):94–100, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2007/01/r1094.pdf.

Anonymous:2007:Ba

[1923] Anonymous. Bookshelf. Computer, 40(1):102, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2007/01/r1102.pdf.

Anonymous:2007:Pa

[1924] Anonymous. Products. Computer, 40(1):103, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2007/01/r1103.pdf.

Cerf:2007:IA

[1925] Vinton G. Cerf. An information avalanche. Computer, 40(1):104–105, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vanGenuchten:2007:ISG

[1926] Michiel van Genuchten. The impact of software growth on the electronics industry. Computer, 40(1):106– 108, January 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Narasimhan:2007:FTC

[1927] Priya Narasimhan. Fault-tolerant CORBA: From specification to reality. Computer, 40(1):110–112, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2007:CPH

[1928] Neville Holmes. The computing profession and higher education. Computer, 40(1):116, 114–115, January 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://bell.computer.
org/dlcomments/.

Anonymous:2007:ASb

[1929] Anonymous. Article summaries. Computer, 40(2):4, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2007/02/r2004.pdf.

Grier:2007:DE

[1930] David Alan Grier. Dirty electricity. Computer, 40(2):6–8, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/02/r2006.
pdf.

Anonymous:2007:YAb

[1931] Anonymous. 32 & 16 years ago. Computer, 40(2):10–11, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2007/02/r2010.pdf.

Paulson:2007:DSD

[1932] Linda Dailey Paulson. Developers shift to dynamic programming languages. Computer, 40(2):12–15, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2007:PCI

[1933] George Lawton. Powering down the computing infrastructure. Computer, 40(2):16–19, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2007/02/r2016.pdf.

Paulson:2007:NBb

[1934] Linda Dailey Paulson. News briefs. Computer, 40(2):20–22, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2007/02/r2020.pdf.

Klingsheim:2007:CSN

[1935] André N. Klingsheim, Veborn Moen, and Kjell J. Hole. Challenges in securing networked J2ME applications. Computer, 40(2):24–30, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schuff:2007:MMO

[1936] David Schuff, Ozgur Turetken, John D’Arcy, and David Croson. Managing E-mail overload: Solutions and future challenges. Computer, 40(2): 31–36, February 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kaps:2007:CSD

[1937] Jens-Peter Kaps, Gunnar Gaubatz, and Berk Sunar. Cryptography on a speck of dust. Computer, 40(2): 38–44, February 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Li:2007:TMD

[1938] Huaizhi Li and Mukesh Singhal. Trust management in distributed systems. Computer, 40(2):45–53, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vijairaghavan:2007:MTI

[1939] Vaarun Vijairaghavan, Darshak Shah, Pallavi Galgali, Amit Shah, Nikhil Shah, Venkatesh Srinivasan, and Lokesh Bhatia. Marking technique to isolate boundary router and attacker. Computer, 40(2):54–58, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tian:2007:AQM

[1940] Min Tian, Andreas Gramm, Hartmut Ritter, Jochen H. Schiller, and Thiemo Voigt. Adaptive QoS for mobile Web services through cross-layer communication. Computer, 40(2):59– 63, February 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hatton:2007:HAD

[1941] Les Hatton. How accurately do engineers predict software maintenance tasks? Computer, 40(2):64–69, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2007:HSN

[1942] Anonymous. Hosaka and Spielberg named winners of 2006 Computer Pioneer Award. Computer, 40(2):73– 77, February 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2007/
                                        02/r2073.pdf.

Anonymous:2007:CCb

[1943] Anonymous. Call and calendar. Computer, 40(2):78–80, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2007/02/r2078.pdf.

Anonymous:2007:COb

[1944] Anonymous. Career opportunities. Computer, 40(2):82–85, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2007/02/r2082.pdf.

Anonymous:2007:Bb

[1945] Anonymous. Bookshelf. Computer, 40(2):87, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2007/02/r2087.pdf.

Anonymous:2007:Pb

[1946] Anonymous. Products. Computer, 40(2):88, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2007/02/r2088.pdf.

Siek:2007:PDT

[1947] Katie A. Siek, Kay Connelly, Suzanne Menzel, and Laura Hopkins. Propagating diversity through active dissemination. Computer, 40(2):89–92, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Peek:2007:CEM

[1948] Daniel Peek and Jason Flinn. Consumer electronics meets distributed storage. Computer, 40(2):93–95, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Luebke:2007:HGW

[1949] David Luebke and Greg Humphreys.

How GPUs work. Computer, 40(2): 96–100, February 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Macedonia:2007:FAF

[1950] Michael Macedonia. The future arrives … finally. Computer, 40(2):101– 103, February 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tripunitara:2007:RMP

[1951] Mahesh Tripunitara and Tom Messerges.

Resolving the micropayment problem. Computer, 40(2):104–106, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grottke:2007:FBR

[1952] Michael Grottke and Kishor S. Trivedi. Fighting bugs: Remove, retry, replicate, and rejuvenate. Computer, 40(2): 107–109, February 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Loka:2007:SDW

[1953] Raghavendra Rao Loka. Software development: What is the problem? Computer, 40(2):112, 110–111, February 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://bell.computer.
org/dlcomments/.

Anonymous:2007:ASc

[1954] Anonymous. Article summaries. Computer, 40(3):4, March 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2007/03/r3004.pdf.

Anonymous:2007:Lb

[1955] Anonymous. Letters. Computer, 40 (3):6–7, March 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
03/r3006.pdf.

Grier:2007:O

[1956] David Alan Grier. Outposts. Computer, 40(3):8–10, March 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/03/r3008.pdf.

Anonymous:2007:YAc

[1957] Anonymous. 32 & 16 years ago.

Computer, 40(3):11–12, March 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/03/r3011.pdf.

Laird:2007:THL

[1958] Cameron Laird. Taking a hard-line approach to encryption. Computer, 40 (3):13–15, March 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2007:VFS

[1959] Neal Leavitt. Vendors fight spam’s sudden rise. Computer, 40(3):16–19, March 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/03/r3016.
pdf.

Paulson:2007:NBc

[1960] Linda Dailey Paulson. News briefs.

Computer, 40(3):20–22, March 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/03/r3020.pdf.

Buell:2007:GEI

[1961] Duncan Buell, Tarek El-Ghazawi, Kris Gaj, and Volodymyr Kindratenko. Guest Editors’ introduction: Highperformance reconfigurable computing. Computer, 40(3):23–27, March 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/03/r3023.
pdf.

Tripp:2007:THL

[1962] Justin L. Tripp, Maya B. Gokhale, and Kristopher D. Peterson. Trident: From high-level language to hardware circuitry. Computer, 40(3):28–37, March 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moore:2007:VEF

[1963] Nicholas Moore, Albert Conti, Miriam Leeser, and Laurie Smith King. Vforce: An extensible framework for reconfigurable supercomputing. Computer, 40 (3):39–49, March 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Herbordt:2007:AHP

[1964] Martin C. Herbordt, Tom Van-Court, Yongfeng Gu, Bharat Sukhwani, Al Conti, Josh Model, and Doug DiSabello. Achieving high performance with FPGA-based computing. Computer, 40(3):50–57, March 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Morris:2007:SMC

[1965] Gerald R. Morris and Viktor K. Prasanna. Sparse matrix computations on reconfigurable hardware. Computer, 40 (3):58–64, March 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Alam:2007:UFD

[1966] Sadaf R. Alam, Pratul K. Agarwal, Melissa C. Smith, Jeffrey S. Vetter, and David Caliga. Using FPGA devices to accelerate biomolecular simulations. Computer, 40(3):66–73, March 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2007:CSCb

[1967] Bob Ward. Computer Society connection. Computer, 40(3):77–79, March 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/03/r3077.
pdf.

Anonymous:2007:CCc

[1968] Anonymous. Call and calendar.

Computer, 40(3):80–82, March 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/03/r3080.pdf.

Anonymous:2007:Pc

[1969] Anonymous. Products. Computer, 40(3):83, March 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
03/r3083.pdf.

Anonymous:2007:COc

[1970] Anonymous. Career opportunities.

Computer, 40(3):84–88, March 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/03/r3084.pdf.

Anonymous:2007:Bc

[1971] Anonymous. Bookshelf. Computer, 40(3):89, March 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
03/r3089.pdf.

Sztipanovits:2007:ENT

[1972] Janos Sztipanovits, John Bay, Larry Rohrbough, Shankar Sastry, Douglas C. Schmidt, Norm Whitaker, Don Wilson, and Don Winter. Escher: A new technology transitioning model. Computer, 40(3):90–92, March 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bussler:2007:FNW

[1973] Christoph Bussler. The fractal nature of Web services. Computer, 40 (3):93–95, March 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hardaway:2007:RPS

[1974] Don Hardaway. Replacing proprietary software on the desktop. Computer, 40 (3):96–97, March 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2007:DTS

[1975] Neville Holmes. Digital technology and the skills shortage. Computer, 40(3): 100, 98–99, March 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://bell.

                                        computer.org/dlcomments/.

Anonymous:2007:ASd

[1976] Anonymous. Article summaries. Computer, 40(4):4, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2007/04/r4004.pdf.

Anonymous:2007:Lc

[1977] Anonymous. Letters. Computer, 40 (4):6–7, April 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
04/r4006.pdf.

Grier:2007:BDT

[1978] David Alan Grier. The best deal in town. Computer, 40(4):8–11, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/04/r4008.
pdf.

Anonymous:2007:YAd

[1979] Anonymous. 32 & 16 years ago.

Computer, 40(4):12–13, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/04/r4012.pdf.

Ortiz:2007:GBE

[1980] Sixto Ortiz, Jr. Getting on board the enterprise service bus. Computer, 40 (4):15–17, April 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2007:NBT

[1981] George Lawton. The next big thing in chipmaking. Computer, 40(4):18–20, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/04/r4018.
pdf.

Paulson:2007:NBd

[1982] Linda Dailey Paulson. News briefs.

Computer, 40(4):21–23, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2007/04/r4021.pdf.

Riehle:2007:EMO

[1983] Dirk Riehle. The economic motivation of open source software: Stakeholder perspectives. Computer, 40 (4):25–32, April 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Budgen:2007:DIB

[1984] David Budgen, Michael Rigby, Pearl Brereton, and Mark Turner. A data integration broker for healthcare systems. Computer, 40(4):34–41, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Aiken:2007:MDM

[1985] Peter Aiken, M. David Allen, Burt Parker, and Angela Mattia. Measuring data management practice maturity: A community’s self-assessment. Computer, 40(4):42–50, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhang:2007:PPD

[1986] Nan Zhang and Wei Zhao. Privacypreserving data mining systems. Computer, 40(4):52–58, April 2007. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cao:2007:DCC

[1987] Jiannong Cao, Yang Zhang, Guohong Cao, and Li Xie. Data consistency for cooperative caching in mobile environments. Computer, 40(4):60–66, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gebis:2007:EEC

[1988] Joe Gebis and David Patterson. Embracing and extending 20th-Century instruction set architectures. Computer, 40(4):68–75, April 2007. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mayer:2007:BDS

[1989] Albrecht Mayer, Harry Siebert, and Klaus D. McDonald-Maier. Boosting debugging support for complex systems on chip. Computer, 40(4):76–81, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2007:COd

[1990] Anonymous. Career opportunities.

Computer, 40(4):82–88, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/04/r4082.pdf.

Ward:2007:CSCc

[1991] Bob Ward. Computer Society connection. Computer, 40(4):89–91, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/04/r4089.
pdf.

Anonymous:2007:CCd

[1992] Anonymous. Call and calendar.

Computer, 40(4):95–97, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/04/r4095.pdf.

Anonymous:2007:Pd

[1993] Anonymous. Products. Computer, 40(4):98, April 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2007/
                                        04/r4098.pdf.

vanLent:2007:GS

[1994] Michael van Lent. Game smarts. Computer, 40(4):99–101, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rich:2007:ATS

[1995] Donald Rich. Authentication in transient storage device attachments. Computer, 40(4):102–104, April 2007. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Krumm:2007:PWD

[1996] John Krumm and Eric Horvitz. Predestination: Where do you want to go today? Computer, 40(4):105–107, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leeper:2007:WFN

[1997] David G. Leeper. Wi-fi—the nimble musician in your laptop. Computer, 40 (4):108–110, April 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hinchey:2007:SSI

[1998] Michael G. Hinchey, Roy Sterritt, and Chris Rouff. Swarms and swarm intelligence. Computer, 40(4):111–113, April 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Harkiolakis:2007:IVE

[1999] Nicholas Harkiolakis. Incorporating a variable-expertise-level system in IT course modules. Computer, 40(4):116, 114–115, April 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://bell.

computer.org/dlcomments/.

Anonymous:2007:ASe

[2000] Anonymous. Article summaries. Computer, 40(5):4, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2007/05/r5004.pdf.

Anonymous:2007:Ld

[2001] Anonymous. Letters. Computer, 40 (5):6–7, May 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
05/r5006.pdf.

Grier:2007:WCH

[2002] David Alan Grier. Working class hero. Computer, 40(5):8–10, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2007:YAe

[2003] Anonymous. 32 & 16 years ago. Computer, 40(5):11–12, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2007:SDN

[2004] George Lawton. Stronger Domain Name System thwarts root-server attacks. Computer, 40(5):14–17, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Heyman:2007:NAT

[2005] Karen Heyman. New attack tricks antivirus software. Computer, 40 (5):18–20, May 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2007:NBe

[2006] Linda Dailey Paulson. News briefs. Computer, 40(5):21–23, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/05/r5021.pdf.

Hatton:2007:ETO

[2007] Les Hatton. Empirical test observations in client-server systems. Computer, 40(5):24–29, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jaimes:2007:GEI

[2008] Alejandro Jaimes, Daniel Gatica-Perez, Nicu Sebe, and Thomas S. Huang. Guest Editors’ introduction: Humancentered computing—toward a human revolution. Computer, 40(5):30–34, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/05/r5030.
pdf.

Alm:2007:CSS

[2009] Norman Alm, Richard Dye, Gary Gowans, Jim Campbell, Arlene Astell, and Maggie Ellis. A communication support system for older people with dementia. Computer, 40(5):35–41, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Guerra-Filho:2007:LHA

[2010] Gutemberg Guerra-Filho and Yiannis Aloimonos. A language for human action. Computer, 40(5):42–51, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

deSilva:2007:IMD

[2011] Gamhewage C. de Silva, Toshihiko Yamasaki, and Kiyoharu Aizawa. An interactive multimedia diary for the home. Computer, 40(5):52–59, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Trivedi:2007:HSA

[2012] Mohan M. Trivedi and Shinko Y.

Cheng. Holistic sensing and active displays for intelligent driver support systems. Computer, 40(5):60–68, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yamagiwa:2007:CNS

[2013] Shinichi Yamagiwa and Leonel Sousa.

Caravela: A novel stream-based distributed computing environment. Computer, 40(5):70–77, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2007:CSCd

[2014] Bob Ward. Computer Society connection. Computer, 40(5):85–88, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/05/r5085.
pdf.

Anonymous:2007:CCe

[2015] Anonymous. Call and calendar. Computer, 40(5):89–91, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2007/05/r5089.pdf.

Anonymous:2007:Pe

[2016] Anonymous. Products. Computer, 40(5):92, May 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2007/
                                        05/r5092.pdf.

Anonymous:2007:Bd

[2017] Anonymous. Bookshelf. Computer, 40(5):93, May 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2007/
                                        05/r5093.pdf.

Burgstahler:2007:IPP

[2018] Sheryl E. Burgstahler and Richard E. Ladner. Increasing the participation of people with disabilities in computing fields. Computer, 40(5):94–97, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Clacy:2007:SMD

[2019] Ben Clacy and Brian Jennings. Service management: Driving the future of IT. Computer, 40(5):98–100, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:2007:BW

[2020] Kwei-Jay Lin. Building Web 2.0. Computer, 40(5):101–102, May 2007.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Joannou:2007:ESS

[2021] Paul Joannou. Enterprise, systems, and software engineering—the need for integration. Computer, 40(5): 103–105, May 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woodfill:2007:VDE

[2022] John Iselin Woodfill, Ron Buck, Dave Jurasek, Gaile Gordon, and Terrance Brown. 3D vision: Developing an embedded stereo-vision system. Computer, 40(5):106–108, May 2007. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:2007:CED

[2023] Jiming Liu. Computing as an evolving discipline: 10 observations. Computer, 40(5):112, 110–111, May 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://bell.computer.
org/dlcomments/.

Anonymous:2007:ASf

[2024] Anonymous. Article summaries. Computer, 40(6):4, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2007/06/r6004.pdf.

Grier:2007:CR

[2025] David Alan Grier. The Camino Real. Computer, 40(6):6–8, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2007:YAf

[2026] Anonymous. 32 & 16 years ago. Computer, 40(6):9–10, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2007:SVW

[2027] Sixto Ortiz, Jr. Searching the Visual Web. Computer, 40(6):12–14, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2007:NBf

[2028] Linda Dailey Paulson and George Lawton. News briefs. Computer, 40 (6):15–18, June 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2007/
                                        06/r6015.pdf.

Keblawi:2007:CFN

[2029] Feisal Keblawi and Dick Sullivan. The case for flexible NIST security standards. Computer, 40(6):19–26, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cleland-Huang:2007:BPA

[2030] Jane Cleland-Huang, Brian Berenbach, Stephen Clark, Raffaella Settimi, and Eli Romanova. Best practices for automated traceability. Computer, 40 (6):27–35, June 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gschwind:2007:OSE

[2031] Michael Gschwind, David Erb, Sid Manning, and Mark Nutter. An open source environment for cell broadband engine system software. Computer, 40 (6):37–47, June 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Aggarwal:2007:ICM

[2032] Nidhi Aggarwal, Parthasarathy Ranganathan, Norman P. Jouppi, and James E. Smith. Isolation in commodity multicore processors. Computer, 40 (6):49–59, June 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tseng:2007:IIM

[2033] Yu-Chee Tseng, You-Chiun Wang, Kai-Yang Cheng, and Yao-Yu Hsieh. iMouse: An integrated mobile surveillance and wireless sensor system. Computer, 40(6):60–66, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chakrabarti:2007:PBA

[2034] Saikat Chakrabarti and Mukesh Singhal. Password-based authentication: Preventing dictionary attacks. Computer, 40(6):68–74, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2007:COe

[2035] Anonymous. Career opportunities.

Computer, 40(6):76–80, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/06/r6076.pdf.

Anonymous:2007:CCf

[2036] Anonymous. Call and calendar. Computer, 40(6):81–83, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2007/06/r6081.pdf.

Ward:2007:CSCe

[2037] Bob Ward. Computer Society connection. Computer, 40(6):84–85, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.computer.
                                        org/comp/mags/co/2007/06/r6084.
                                        pdf.

Anonymous:2007:Be

[2038] Anonymous. Bookshelf. Computer, 40(6):89, June 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2007/
                                        06/r6089.pdf.

Holmes:2007:BA

[2039] Neville Holmes. Binary arithmetic. Computer, 40(6):90–93, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Macedonia:2007:ITT

[2040] Michael Macedonia. iPhones target the tech elite. Computer, 40(6):94–95, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lau:2007:SSW

[2041] Tessa Lau. Social scripting for the Web. Computer, 40(6):96–98, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bohner:2007:ECT

[2042] Shawn Bohner. An era of changetolerant systems. Computer, 40(6):

100–102, June 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hars:2007:DIH

[2043] Laszlo Hars. Discryption: Internal hard-disk encryption for secure storage. Computer, 40(6):103–105, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Price:2007:SRC

[2044] Sean M. Price. Supporting resourceconstrained collaboration environments. Computer, 40(6):108, 106–107, June 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://bell.computer.
org/dlcomments/.

Anonymous:2007:ASg

[2045] Anonymous. Article summaries. Computer, 40(7):4, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2007/07/r7004.pdf.

Grier:2007:BT

[2046] David Alan Grier. The boundaries of time. Computer, 40(7):5–7, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2007:YAg

[2047] Anonymous. 32 & 16 years ago. Computer, 40(7):8–9, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2007:YFW

[2048] George Lawton. These are not your father’s widgets. Computer, 40(7):

10–13, July 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2007:WUF

[2049] Neal Leavitt. For wireless USB, the future starts now. Computer, 40 (7):14–16, July 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2007:NBg

[2050] Linda Dailey Paulson. News briefs. Computer, 40(7):17–19, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/07/r7017.pdf.

Vasconcelos:2007:PSS

[2051] Nuno Vasconcelos. From pixels to semantic spaces: Advances in contentbased image retrieval. Computer, 40 (7):20–26, July 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Treleaven:2007:BSH

[2052] Philip Treleaven and Jonathan Wells. 3D body scanning and healthcare applications. Computer, 40(7):28–34, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bowman:2007:VRH

[2053] Doug A. Bowman and Ryan P. McMahan. Virtual reality: How much immersion is enough? Computer, 40 (7):36–43, July 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shahabi:2007:IAF

[2054] Cyrus Shahabi, Tim Marsh, Kiyoung Yang, Hyunjin Yoon, Albert A. Rizzo, Margaret McLaughlin, and Minyoung Mun. Immersidata analysis: Four case studies. Computer, 40(7):45–52, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nayar:2007:DUP

[2055] Shree K. Nayar and Vijay N. Anand.

3D display using passive optical scatterers. Computer, 40(7):54–63, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ciolfi:2007:SPI

[2056] Luigina Ciolfi, Mikael Fernström, Liam J. Bannon, Parag Deshpande, Paul Gallagher, Colm McGettrick, Nicola Quinn, and Stephen Shirley. The Shannon Portal installation: Interaction design for public places. Computer, 40(7):64–71, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2007:COf

[2057] Anonymous. Career opportunities.

Computer, 40(7):72–73, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/07/r7072.pdf.

Ward:2007:CSCf

[2058] Bob Ward. Computer Society connection. Computer, 40(7):77–80, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/07/r7077.
pdf.

Anonymous:2007:CCg

[2059] Anonymous. Call and calendar. Computer, 40(7):81–83, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2007/07/r7081.pdf.

Anonymous:2007:Bf

[2060] Anonymous. Bookshelf. Computer, 40(7):84, July 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2007/
                                        07/r7084.pdf.

Jansen:2007:CF

[2061] Bernard J. Jansen. Click fraud. Computer, 40(7):85–86, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carlin:2007:IAC

[2062] Anna Carlin and Frederick Gallegos. IT audit: A critical business process. Computer, 40(7):87–89, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nath:2007:SWA

[2063] Suman Nath, Jie Liu, and Feng Zhao. SensorMap for wide-area sensor Webs. Computer, 40(7):90–93, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2007:SCH

[2064] Jeffrey M. Voas and Phillip A. Laplante. Standards confusion and harmonization. Computer, 40(7):94–96, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2007:CC

[2065] Neville Holmes. Consciousness and computers. Computer, 40(7):100, 98– 99, July 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://bell.computer.
org/dlcomments/.

Anonymous:2007:ASh

[2066] Anonymous. Article summaries. Computer, 40(8):4, August 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2007/08/r8004.pdf.

Grier:2007:OTA

[2067] David Alan Grier. In our time: Annie and the boys. Computer, 40 (8):6–9, August 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2007:YAa

[2068] Neville Holmes. 32 & 16 years ago.

Computer, 40(8):10–11, August 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nichols:2007:NIT

[2069] Steven J. Vaughan Nichols. New interfaces at the touch of a fingertip. Computer, 40(8):12–15, August 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2007:PNC

[2070] Sixto Ortiz, Jr. Protecting networks by controlling access. Computer, 40 (8):16–19, August 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2007:NBh

[2071] Linda Dailey Paulson. News briefs. Computer, 40(8):20–22, August 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/08/r8020.pdf.

Soukup:2007:ICG

[2072] Jiri Soukup and Martin Soukup. The inevitable cycle: Graphical tools and programming paradigms. Computer, 40 (8):24–30, August 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ramakrishnan:2007:AES

[2073] Naren Ramakrishnan. From the Area Editor: Search—the new incarnations. Computer, 40(8):31–32, August 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/08/r8031.
pdf.

Joachims:2007:SEL

[2074] Thorsten Joachims and Filip Radlinski. Search engines that learn from implicit feedback. Computer, 40(8): 34–40, August 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smyth:2007:CBA

[2075] Barry Smyth. A community-based approach to personalizing Web search. Computer, 40(8):42–50, August 2007.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jansen:2007:SSM

[2076] Bernard J. Jansen and Amanda Spink.

Sponsored search: Is money a motivator for providing relevant results? Computer, 40(8):52–57, August 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kamvar:2007:DTM

[2077] Maryam Kamvar and Shumeet Baluja.

Deciphering trends in mobile search. Computer, 40(8):58–62, August 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ramakrishnan:2007:TP

[2078] Raghu Ramakrishnan and Andrew Tomkins. Toward a PeopleWeb. Computer, 40(8):63–72, August 2007. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2007:Bg

[2079] Anonymous. Bookshelf. Computer, 40(8):73, August 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
08/r8073.pdf.

Ward:2007:CSCg

[2080] Bob Ward. Computer Society connection. Computer, 40(8):77–79, August 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/08/r8077.
pdf.

Anonymous:2007:CCh

[2081] Anonymous. Call and calendar. Computer, 40(8):80–81, August 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2007/08/r8080.pdf.

Anonymous:2007:COg

[2082] Anonymous. Career opportunities. Computer, 40(8):82–84, August 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2007/08/r8082.pdf.

Sangwan:2007:HBG

[2083] Raghvinder S. Sangwan and Colin J. Neill. How business goals drive architectural design. Computer, 40(8): 85–87, August 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ross:2007:MES

[2084] Ron Ross. Managing enterprise security risk with NIST Standards. Computer, 40(8):88–91, August 2007. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Epstein:2007:EV

[2085] Jeremy Epstein. Electronic voting. Computer, 40(8):92–95, August 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Torrey:2007:HWP

[2086] Cristen Torrey and David W. McDonald. How-to Web pages. Computer, 40 (8):96–97, August 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vanLent:2007:GOM

[2087] Michael van Lent and William Swartout. Games: Once more, with feeling. Computer, 40(8):98–100, August 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hatton:2007:CSQ

[2088] Les Hatton. The chimera of software quality. Computer, 40(8):104, 102–103, August 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://bell.

computer.org/dlcomments/.

Anonymous:2007:ASi

[2089] Anonymous. Article summaries.

Computer, 40(9):4, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/09/r9004.pdf.

Anonymous:2007:Le

[2090] Anonymous. Letters. Computer, 40(9):6, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2007/09/r9006.pdf.

Grier:2007:CC

[2091] David Alan Grier. Controlling the conversation. Computer, 40(9):7– 9, September 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2007:YAb

[2092] Neville Holmes. 32 & 16 years ago. Computer, 40(9):10–11, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

King:2007:MCM

[2093] Brett Allan King and Linda Dailey Paulson. Motion capture moves into new realms. Computer, 40(9):13– 16, September 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geer:2007:PMC

[2094] David Geer. For programmers, multicore chips mean multiple challenges. Computer, 40(9):17–19, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2007:NBi

[2095] Linda Dailey Paulson. News briefs. Computer, 40(9):20–22, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/09/r9020.
pdf.

Pieper:2007:NEP

[2096] Sean M. Pieper, JoAnn M. Paul, and Michael J. Schulte. A new era of performance evaluation. Computer, 40(9): 23–30, September 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Prey:2007:GEI

[2097] Jane Prey and Alf Weaver. Guest Editors’ introduction: Tablet PC technology—the next generation. Computer, 40(9):32–33, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/09/r9032.pdf.

Davis:2007:MPS

[2098] Randall Davis. Magic Paper: Sketch— understanding research. Computer, 40 (9):34–41, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roschelle:2007:III

[2099] Jeremy Roschelle, Deborah Tatar, S. Raj Chaudhury, Yannis Dimitriadis, Charles Patton, and Chris DiGiano. Ink, improvisation, and interactive engagement: Learning with tablets. Computer, 40(9):42–48, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pittman:2007:HRT

[2100] James A. Pittman. Handwriting recognition: Tablet PC text input. Computer, 40(9):49–54, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anderson:2007:CPE

[2101] Richard Anderson, Ruth Anderson, Peter Davis, Natalie Linnell, Craig Prince, Valentin Razmov, and Fred Videon. Classroom presenter: Enhancing interactive education with digital ink. Computer, 40(9):56–61, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tront:2007:FPP

[2102] Joseph G. Tront. Facilitating pedagogical practices through a large-scale tablet PC deployment. Computer, 40 (9):62–68, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2007:ICS

[2103] Anonymous. IEEE Computer Society election. Computer, 40(9):69– 76, September 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
09/r9069.pdf.

Anonymous:2007:CSC

[2104] Anonymous. Computer Society connection: IEEE president-elect candidates address Computer Society concerns. Computer, 40(9):77–83, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/09/r9077.
pdf.

Anonymous:2007:CCi

[2105] Anonymous. Call and calendar. Computer, 40(9):87–89, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.computer.
org/comp/mags/co/2007/09/r9087.
pdf.

Anonymous:2007:COCa

[2106] Anonymous. Career opportunities: Classified advertising. Computer, 40 (9):90–93, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2007/09/r9090.pdf.

Anonymous:2007:Bh

[2107] Anonymous. Bookshelf. Computer, 40(9):95, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2007/09/r9095.pdf.

Watson:2007:CSB

[2108] Hugh J. Watson and Barbara H.

Wixom. The current state of business intelligence. Computer, 40(9):96– 99, September 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coallier:2007:SAE

[2109] François Coallier. Standards, agility, and engineering. Computer, 40(9):100– 102, September 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kohavi:2007:OEL

[2110] Ron Kohavi and Roger Longbotham.

Online experiments: Lessons learned. Computer, 40(9):103–105, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vahid:2007:TSC

[2111] Frank Vahid. It’s time to stop calling circuits “hardware”. Computer, 40 (9):106–108, September 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2007:PCK

[2112] Neville Holmes. The profession as a culture killer. Computer, 40(9):112, 110– 111, September 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://bell.

                                        computer.org/dlcomments/.

Anonymous:2007:ASj

[2113] Anonymous. Article summaries. Computer, 40(10):4, October 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2007/10/mco2007100004.pdf.

Anonymous:2007:Lf

[2114] Anonymous. Letters. Computer, 40(10):6–7, October 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2007/10/mco2007100006.pdf.

Grier:2007:MA

[2115] David Alan Grier. E-mailing from Armenia. Computer, 40(10):8–10, October 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2007:YAh

[2116] Anonymous. 32 & 16 years ago. Computer, 40(10):11–12, October 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2007:WCS

[2117] George Lawton. Web 2.0 creates security challenges. Computer, 40(10): 13–16, October 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2007:NBj

[2118] Linda Dailey Paulson. News briefs. Computer, 40(10):17–20, October 2007.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/10/mco2007100017.
pdf.

Oktaba:2007:SPI

[2119] Hanna Oktaba, Félix García, Mario Piattini, Francisco Ruiz, Francisco J. Pino, and Claudia Alquicira. Software process improvement: The Competisoft Project. Computer, 40(10): 21–28, October 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wolf:2007:GEI

[2120] Wayne Wolf. Guest Editor’s introduction: The embedded systems landscape. Computer, 40(10):29– 31, October 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
10/mco2007100029.pdf.

Henzinger:2007:DES

[2121] Thomas A. Henzinger and Joseph Sifakis. The discipline of embedded systems design. Computer, 40(10): 32–40, October 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sangiovanni-Vincentelli:2007:ESD

[2122] Alberto Sangiovanni-Vincentelli and Marco Di Natale. Embedded system design for automotive applications. Computer, 40(10):42–51, October 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Talla:2007:UDT

[2123] Deepu Talla and Jeremiah Golston. Using DaVinci technology for digital video devices. Computer, 40(10): 53–61, October 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ramacher:2007:SDR

[2124] Ulrich Ramacher. Software-defined radio prospects for multistandard mobile phones. Computer, 40(10):62– 69, October 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2007:CSCh

[2125] Bob Ward. Computer Society connection. Computer, 40(10):73–75, October 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2007/
                                        10/mco2007100073.pdf.

Anonymous:2007:CCj

[2126] Anonymous. Call and calendar. Computer, 40(10):76–77, October 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2007/10/mco2007100076.
                                        pdf.

Anonymous:2007:COh

[2127] Anonymous. Career opportunities. Computer, 40(10):78–86, October 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2007/
                                        10/mco2007100078.pdf.

Anonymous:2007:Bi

[2128] Anonymous. Bookshelf. Computer, 40(10):88, October 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2007/10/mco2007100088.pdf.

Yong:2007:RLS

[2129] Jianming Yong and Elisa Bertino. Replacing lost or stolen E-passports. Computer, 40(10):89–91, October 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Knight:2007:GC

[2130] John Knight. The glass cockpit.

Computer, 40(10):92–95, October 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Griswold:2007:FEM

[2131] William G. Griswold. Five enablers for Mobile 2.0. Computer, 40(10): 96–98, October 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Macedonia:2007:GLV

[2132] Mike Macedonia. Generation 3D: Living in virtual worlds. Computer, 40 (10):99–101, October 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Casanova:2007:DCS

[2133] Marco A. Casanova, Karin K. Breitman, Daniela F. Brauner, and André L. A. Marins. Database conceptual schema matching. Computer, 40(10): 102–104, October 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vanGenuchten:2007:GRC

[2134] Michiel van Genuchten and Doug Vogel. Getting real in the classroom. Computer, 40(10):108, 106–107, October 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://bell.computer.
org/dlcomments/.

Anonymous:2007:ASk

[2135] Anonymous. Article summaries. Computer, 40(11):4, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/11/mco2007110004.
pdf.

Anonymous:2007:Lg

[2136] Anonymous. Letters. Computer, 40 (11):6–7, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2007/11/mco2007110006.pdf.

Grier:2007:CB

[2137] David Alan Grier. Counting beans. Computer, 40(11):8–10, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2007:YAi

[2138] Anonymous. 32 & 16 years ago. Computer, 40(11):12–13, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geer:2007:IDA

[2139] David Geer. Improving data accessibility with file area networks. Computer, 40(11):14–17, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2007:WBT

[2140] Sixto Ortiz, Jr. 4G wireless begins to take shape. Computer, 40(11):18– 21, November 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2007:NBk

[2141] Linda Dailey Paulson. News briefs.

Computer, 40(11):22–24, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
11/mco2007110022.pdf.

Petrenko:2007:TSE

[2142] Maksym Petrenko, Denys Poshyvanyk, Václav Rajlich, and Joseph Buchta. Teaching software evolution in open source. Computer, 40(11):25–31, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Margaria:2007:SEB

[2143] Tiziana Margaria. Service is in the eyes of the beholder. Computer, 40 (11):33–37, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2007/11/mco2007110033.pdf.

Papazoglou:2007:SOC

[2144] Michael P. Papazoglou, Paolo Traverso, Schahram Dustdar, and Frank Leymann. Service-oriented computing: State of the art and research challenges. Computer, 40(11):38–45, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Magedanz:2007:ESC

[2145] Thomas Magedanz, Niklas Blum, and Simon Dutkowski. Evolution of SOA concepts in telecommunications. Computer, 40(11):46–50, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bosch:2007:SOE

[2146] Jan Bosch, Stefan Friedrichs, Stefan Jung, Johannes Helbig, and Alexander Scherdin. Service orientation in the enterprise. Computer, 40(11):51– 56, November 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kaiser:2007:TRP

[2147] Matthias Kaiser. Toward the realization of policy-oriented enterprise management. Computer, 40(11):57– 63, November 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Steffen:2007:FLC

[2148] Bernhard Steffen and Prakash Narayan. Full life-cycle support for end-to-end processes. Computer, 40(11):64–73, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Curbera:2007:CCS

[2149] Francisco Curbera. Component contracts in service-oriented architectures. Computer, 40(11):74–80, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2007:CSCi

[2150] Bob Ward. Computer Society connection. Computer, 40(11):81–85, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
11/mco2007110081.pdf.

Anonymous:2007:CCk

[2151] Anonymous. Call and calendar.

Computer, 40(11):86–87, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
11/mco2007110086.pdf.

Anonymous:2007:COi

[2152] Anonymous. Career opportunities.

Computer, 40(11):88–101, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
11/mco2007110088.pdf.

Anonymous:2007:Bj

[2153] Anonymous. Bookshelf. Computer, 40(11):103, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2007/11/mco2007110103.pdf.

Wolf:2007:GNB

[2154] Wayne Wolf. The good news and the bad news. Computer, 40(11):104– 105, November 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Thomas:2007:US

[2155] Martyn Thomas. Unsafe standardization. Computer, 40(11):109–111, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mikroyannidis:2007:TSS

[2156] Alexander Mikroyannidis. Toward a social Semantic Web. Computer, 40(11): 113–115, November 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Davidson:2007:ESW

[2157] Mary Ann Davidson and Elad Yoran. Enterprise security for Web 2.0. Computer, 40(11):117–119, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sharkey:2007:AKC

[2158] Noel Sharkey. Automated killers and the computing profession. Computer, 40(11):124, 122–123, November 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://bell.computer.
                                        org/dlcomments/.

Anonymous:2007:ASl

[2159] Anonymous. Article summaries. Computer, 40(12):4, December 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2007/12/mco2007120004.
pdf.

Williams:2007:IY

[2160] Michael R. Williams. An interesting year. Computer, 40(12):6– 7, December 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
12/mco2007120006.pdf.

Grier:2007:FN

[2161] David Alan Grier. A force of nature. Computer, 40(12):8–10, December 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2007:YAj

[2162] Anonymous. 32 & 16 years ago. Computer, 40(12):11–12, December 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2007:CWO

[2163] Neal Leavitt. The changing world of outsourcing. Computer, 40(12):13– 16, December 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Heyman:2007:NVP

[2164] Karen Heyman. A new virtual private network for today’s mobile world. Computer, 40(12):17–19, December 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2007:NBl

[2165] Linda Dailey Paulson. News briefs.

Computer, 40(12):20–22, December 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
12/mco2007120020.pdf.

Gil:2007:ECS

[2166] Yolanda Gil, Ewa Deelman, Mark Ellisman, Thomas Fahringer, Geoffrey Fox, Dennis Gannon, Carole Goble, Miron Livny, Luc Moreau, and Jim Myers. Examining the challenges of scientific workflows. Computer, 40(12): 24–32, December 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barroso:2007:CEP

[2167] Luiz André Barroso and Urs Hölzle. The case for energy-proportional computing. Computer, 40(12):33–37, December 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rivoire:2007:MME

[2168] Suzanne Rivoire, Mehul A. Shah, Parthasarathy Ranganathan, Christos Kozyrakis, and Justin Meza. Models and metrics to enable energy-efficiency optimizations. Computer, 40(12):39– 48, December 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Feng:2007:GLE

[2169] Wu chun Feng and Kirk Cameron. The Green500 list: Encouraging sustainable supercomputing. Computer, 40(12): 50–55, December 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Oliver:2007:LCA

[2170] John Y. Oliver, Rajeevan Amirtharajah, Venkatesh Akella, Roland Geyer, and Frederic T. Chong. Life cycle aware computing: Reusing silicon technology. Computer, 40(12): 56–61, December 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2007:LVC

[2171] Bob Ward. Land voted 2008 Computer Society President-Elect. Computer, 40 (12):62–63, December 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2007/12/mco2007120062.pdf.

Chang:2007:CRE

[2172] Carl Chang. Computer recognizes expert reviewers. Computer, 40(12):64– 65, December 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
12/mco2007120064.pdf.

Anonymous:2007:CCl

[2173] Anonymous. Call and calendar.

Computer, 40(12):66–67, December 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2007/
12/mco2007120066.pdf.

Anonymous:2007:CAI

[2174] Anonymous. Computer annual index, volume 40, 2007. Computer, 40 (12):72–84, December 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2007/12/mco2007120072.pdf.

Anonymous:2007:COj

[2175] Anonymous. Career opportunities. Computer, 40(12):86–101, December 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2007/
                                        12/mco2007120086.pdf.

Anonymous:2007:Bk

[2176] Anonymous. Bookshelf. Computer, 40(12):102, December 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2007/12/mco2007120102.pdf.

Stone:2007:NLP

[2177] Allen Stone. Natural-language processing for intrusion detection. Computer, 40(12):103–105, December 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brown:2007:SSM

[2178] Jeff Brown, Bill Shipman, and Ron Vetter. SMS: The Short Message Service. Computer, 40(12):106–110, December 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holzmann:2007:CC

[2179] Gerard J. Holzmann. Conquering complexity. Computer, 40(12):111– 113, December 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Subramanya:2007:EUE

[2180] S. R. Subramanya and Byung K. Yi.

Enhancing the user experience in mobile phones. Computer, 40(12):114– 117, December 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vincent:2007:TOM

[2181] Luc Vincent. Taking online maps down to street level. Computer, 40(12):118– 120, December 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Santini:2007:MCD

[2182] Simone Santini. Making computers do more with less. Computer, 40(12):124, 122–123, December 2007. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://bell.

computer.org/dlcomments/.

Anonymous:2008:ASa

[2183] Anonymous. Article summaries. Computer, 41(1):4, January 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2008/01/mco2008010004.pdf.

Anonymous:2008:YAa

[2184] Anonymous. 32 & 16 years ago.

Computer, 41(1):6–7, January 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kasturi:2008:YR

[2185] Rangachar Kasturi. 2008: The year of revitalization. Computer, 41(1): 8–9, January 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
01/mco2008010008.pdf.

Chang:2008:C

[2186] Carl K. Chang. Computer in 2008. Computer, 41(1):10–11, January 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
01/mco2008010010.pdf.

Ortiz:2008:IRL

[2187] Sixto Ortiz, Jr. Internet researchers look to wipe the slate clean. Computer, 41(1):12–16, January 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2008:NBa

[2188] Linda Dailey Paulson. News briefs. Computer, 41(1):17–19, January 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2008/01/mco2008010017.
pdf.

Cusumano:2008:CSB

[2189] Michael A. Cusumano. The changing software business: Moving from products to services. Computer, 41(1): 20–27, January 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Harel:2008:CPL

[2190] David Harel. Can programming be liberated, period? Computer, 41(1): 28–37, January 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhirnov:2008:AID

[2191] Victor Zhirnov, Ralph Cavin, Greg Leeming, and Kosmas Galatsis. An assessment of integrated digital cellular automata architectures. Computer, 41 (1):38–44, January 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Greenspan:2008:TCP

[2192] Michael Greenspan, Joseph Lam, Marc Godard, Imran Zaidi, Sam Jordan, Will Leckie, Ken Anderson, and Donna Dupuis. Toward a competitive poolplaying robot. Computer, 41(1):46– 53, January 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McKinley:2008:HDE

[2193] Philip McKinley, Betty H. C. Cheng, Charles Ofria, David Knoester, Benjamin Beckmann, and Heather Goldsby. Harnessing digital evolution. Computer, 41(1):54–63, January 2008. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reynolds:2008:MSF

[2194] Robert G. Reynolds, Mostafa Ali, and Thaer Jayyousi. Mining the social fabric of archaic urban centers with cultural algorithms. Computer, 41(1): 64–72, January 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:COa

[2195] Anonymous. Career opportunities. Computer, 41(1):74–81, January 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2008/
                                        01/mco2008010074.pdf.

Ward:2008:CSCa

[2196] Bob Ward. Computer Society connection. Computer, 41(1):82–86, January 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2008/
                                        01/mco2008010082.pdf.

Anonymous:2008:CCa

[2197] Anonymous. Call and calendar. Computer, 41(1):87–89, January 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2008/01/mco2008010087.
                                        pdf.

Anonymous:2008:Ba

[2198] Anonymous. Bookshelf. Computer, 41(1):90, January 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2008/01/mco2008010090.pdf.

Grier:2008:PE

[2199] David Alan Grier. The public eye. Computer, 41(1):91–93, January 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sosnowski:2008:SIM

[2200] Janusz Sosnowski, Dawid Trawczyński, and Janusz Zalewski. Safety issues in modern bus standards. Computer, 41 (1):97–99, January 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lacity:2008:IOC

[2201] Mary C. Lacity and Joseph W.

Rottman. The impact of outsourcing on client project managers. Computer, 41(1):100–102, January 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Serpanos:2008:DSW

[2202] Dimitrios Serpanos and Jörg Henkel.

Dependability and security will change embedded computing. Computer, 41 (1):103–105, January 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hendler:2008:WCF

[2203] Jim Hendler. Web 3.0: Chicken farms on the Semantic Web3. Computer, 41 (1):106–108, January 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2008:HCP

[2204] Neville Holmes. The history of the computing profession. Computer, 41(1): 112, 110–111, January 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

bell.computer.org/dlcomments/.

Anonymous:2008:ASb

[2205] Anonymous. Article summaries. Computer, 41(2):4, February 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2008/02/mco2008020004.pdf.

Anonymous:2008:LEa

[2206] Anonymous. Letters to the editor. Computer, 41(2):6–9, February 2008. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
02/mco2008020006.pdf.

Grier:2008:DRS

[2207] David Alan Grier. Derek the rocket scientist. Computer, 41(2):10–12, February 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:YAb

[2208] Anonymous. 32 & 16 years ago. Computer, 41(2):13–14, February 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
02/mco2008020013.pdf.

Lawton:2008:UCP

[2209] George Lawton. US cell phone industry faces an open future. Computer, 41 (2):15–18, February 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:2008:SWB

[2210] Brad Smith. A storm (Worm) is brewing. Computer, 41(2):20–22, February 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2008:NBb

[2211] Linda Dailey Paulson. News briefs. Computer, 41(2):23–25, February 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2008/02/mco2008020023.
pdf.

McArthur:2008:NSD

[2212] David McArthur. National Science Digital Library: Shaping education’s cyberinfrastructure. Computer, 41(2): 26–32, February 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

OLeary:2008:WEA

[2213] Daniel E. O’Leary. Wikis: ’From Each According to His Knowledge’. Computer, 41(2):34–41, February 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jain:2008:EDH

[2214] Ramesh Jain. EventWeb: Developing a human-centered computing system. Computer, 41(2):42–50, February 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Beroggi:2008:SEI

[2215] Giampiero E. G. Beroggi. Secure and easy Internet voting. Computer, 41(2): 52–56, February 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nair:2008:TTS

[2216] Srijith K. Nair, Ron Gerrits, Bruno Crispo, and Andrew S. Tanenbaum. Turning teenagers into stores. Computer, 41(2):58–62, February 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Matyas:2008:ACP

[2217] Václav (Vashek) Matyáš, Dan Cvrcek, Jan Krhovják, and Marek Kumpost. Authorizing card payments with PINs. Computer, 41(2):64–68, February 2008.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

El-Ghazawi:2008:PHP

[2218] Tarek El-Ghazawi, Esam El-Araby, Miaoqing Huang, Kris Gaj, Volodymyr Kindratenko, and Duncan Buell. The promise of high-performance reconfigurable computing. Computer, 41(2): 69–76, February 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:COAa

[2219] Anonymous. Career opportunities advertisements. Computer, 41(2):78– 80, February 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
02/mco2008020078.pdf.

Anonymous:2008:Bb

[2220] Anonymous. Bookshelf. Computer, 41(2):81, February 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2008/02/mco2008020081.pdf.

Ward:2008:CSCb

[2221] Bob Ward. Computer Society connection. Computer, 41(2):82–87, February 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
02/mco2008020082.pdf.

Anonymous:2008:CCb

[2222] Anonymous. Call and calendar. Computer, 41(2):88–90, February 2008.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2008/02/mco2008020088.
pdf.

Maibaum:2008:PFA

[2223] Tom Maibaum and Alan Wassyng.

A product-focused approach to software certification. Computer, 41(2): 91–93, February 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weaver:2008:SN

[2224] Alfred C. Weaver and Benjamin B.

Morrison. Social networking. Computer, 41(2):97–100, February 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vanLent:2008:BF

[2225] Michael van Lent. The business of fun. Computer, 41(2):101–103, February 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Trevor:2008:DMM

[2226] Jonathan Trevor. Doing the mobile mash. Computer, 41(2):104– 106, February 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Natarajan:2008:AC

[2227] Raja Natarajan. On attending conferences. Computer, 41(2):108,107, February 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://bell.computer.
org/dlcomments/.

Anonymous:2008:ASc

[2228] Anonymous. Article summaries. Computer, 41(3):4, March 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2008/03/mco2008030004.pdf.

Anonymous:2008:LEb

[2229] Anonymous. Letters to the editor. Computer, 41(3):5–6, March 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2008/03/mco2008030005.
                                        pdf.

Grier:2008:LJS

[2230] David Alan Grier. Long John Software and the Digital Jolly Roger. Computer, 41(3):7–9, March 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:YAc

[2231] Anonymous. 32 & 16 years ago. Computer, 41(3):10–11, March 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2008:WIF

[2232] Neal Leavitt. Will IEEE 802.1X finally take off in 2008? Computer, 41 (3):12–15, March 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2008:MW

[2233] George Lawton. Moving the OS to the Web. Computer, 41(3):16–19, March 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2008:NBc

[2234] Linda Dailey Paulson. News briefs.

Computer, 41(3):20–22, March 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2008/03/mco2008030020.
pdf.

Moreno-Ger:2008:CCD

[2235] Pablo Moreno-Ger, Iván Martínez-Ortiz, José Luis Sierra, and Baltasar Fernández-Manjón. A content-centric development process model. Computer, 41(3):24–30, March 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boehm:2008:MDS

[2236] Barry Boehm. Making a difference in the software century. Computer, 41 (3):32–38, March 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Osterweil:2008:DIS

[2237] Leon J. Osterweil, Carlo Ghezzi, Jeff Kramer, and Alexander L. Wolf. Determining the impact of software engineering research on practice. Computer, 41 (3):39–49, March 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hochstein:2008:AAP

[2238] Lorin Hochstein and Victor R. Basili.

The ASC—Alliance Projects: A case study of large-scale parallel scientific code development. Computer, 41(3): 50–58, March 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Oppliger:2008:STS

[2239] Rolf Oppliger, Ralf Hauser, and David Basin. SSL/TLS session-aware user authentication. Computer, 41(3):59–65, March 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Xu:2008:MPF

[2240] Cheng-Zhong Xu, Bojin Liu, and Jianbin Wei. Model predictive feedback control for QoS assurance in Webservers. Computer, 41(3):66–72, March 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:COb

[2241] Anonymous. Career opportunities. Computer, 41(3):74–78, March 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2008/03/mco2008030074.
                                        pdf.

Anonymous:2008:Bc

[2242] Anonymous. Bookshelf. Computer, 41(3):79, March 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2008/
                                        03/mco2008030079.pdf.

Ward:2008:CSCc

[2243] Bob Ward. Computer Society connection. Computer, 41(3):83–85, March 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2008/
                                        03/mco2008030083.pdf.

Anonymous:2008:CCc

[2244] Anonymous. Call and calendar.

Computer, 41(3):86–88, March 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2008/03/mco2008030086.
pdf.

Morreale:2008:TCC

[2245] Patricia Morreale, George Chang, and Lee Wittenberg. Transitioning from a community college to a four-year university. Computer, 41(3):89–91, March 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carey:2008:SW

[2246] Michael J. Carey. SOA what? Computer, 41(3):92–94, March 2008. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Klein:2008:CSE

[2247] Gary Klein, Steven Deal, and Sterling Wiggins. Cognitive systems engineering: The hype and the hope. Computer, 41(3):95–97, March 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laporte:2008:DIS

[2248] Claude Y. Laporte, Simon Alexandre, and Alain Renault. Developing international standards for very small enterprises. Computer, 41(3):98–101, March 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2008:SAB

[2249] Neville Holmes. Supporting acquisition of basic skills. Computer, 41(3):104, 102–103, March 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://bell.computer.
org/dlcomments/;   http://csdl.
computer.org/comp/mags/co/2008/
03/mco2008030104.pdf.

Anonymous:2008:ASd

[2250] Anonymous. Article summaries. Computer, 41(4):4, April 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2008/04/mco2008040004.pdf.

Anonymous:2008:LEc

[2251] Anonymous. Letters to the editor. Computer, 41(4):5–6, April 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2008/04/mco2008040005.
pdf.

Grier:2008:TLA

[2252] David Alan Grier. Thinking locally, acting globally. Computer, 41 (4):7–10, April 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:X

[2253] Anonymous. 32 & 16. Computer, 41 (4):11–12, April 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hayman:2008:MMS

[2254] Karen Hayman. The move to make social data portable. Computer, 41 (4):13–15, April 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2008:PTR

[2255] Sixto Ortiz, Jr. Proponents try to rehabilitate peer-to-peer technology. Computer, 41(4):16–19, April 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2008:PSL

[2256] Linda Dailey Paulson. E-paper soon to be in living color. Computer, 41 (4):20–22, April 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
04/mco2008040020.pdf.

Lin:2008:USM

[2257] Po-Ching Lin, Ying-Dar Lin, Yuan-Cheng Lai, and Tsern-Huei Lee. Using string matching for deep packet inspection. Computer, 41(4):23–28, April 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gorton:2008:DIC

[2258] Ian Gorton, Paul Greenfield, Alex Szalay, and Roy Williams. Dataintensive computing in the 21st century. Computer, 41(4):30–32, April 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
04/mco2008040030.pdf.

Xue:2008:QRG

[2259] Yong Xue, Wei Wan, Yingjie Li, Jie Guang, Linyian Bai, Ying Wang, and Jianwen Ai. Quantitative retrieval of geophysical parameters using satellite data. Computer, 41(4):33–40, April 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Villa:2008:ART

[2260] Oreste Villa, Daniele Paolo Scarpazza, and Fabrizio Petrini. Accelerating real-time string searching with multicore processors. Computer, 41(4): 42–50, April 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kumar:2008:ASQ

[2261] Vijay S. Kumar, Sivaramakrishnan Narayanan, Tahsin Kurc, Jun Kong, Metin N. Gurcan, and Joel H. Saltz. Analysis and semantic querying in large biomedical image datasets. Computer, 41(4):52–59, April 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gokhale:2008:HTH

[2262] Maya Gokhale, Jonathan Cohen, Andy Yoo, W. Marcus Miller, Arpith Jacob, Craig Ulmer, and Roger Pearce. Hardware technologies for high-performance data-intensive computing. Computer, 41(4):60–68, April 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shahabi:2008:PEE

[2263] Cyrus Shahabi, Mehrdad Jahangiri, and Farnoush Banaei-Kashani. ProDA: An end-to-end wavelet-based OLAP system for massive datasets. Computer, 41(4):69–77, April 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:CA

[2264] Anonymous. Classified ads. Computer, 41(4):82–85, April 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2008/04/mco2008040082.pdf.

Anonymous:2008:Bd

[2265] Anonymous. Bookshelf. Computer, 41(4):86, April 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
04/mco2008040086.pdf.

Ward:2008:CSE

[2266] Bob Ward. Computer science enrollments drop. Computer, 41(4): 87–89, April 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
04/mco2008040087.pdf.

Ward:2008:CCa

[2267] Bob Ward. Call and calendar.

Computer, 41(4):90–92, April 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2008/04/mco2008040090.
pdf.

Hallsteinsen:2008:DSP

[2268] Svein Hallsteinsen, Mike Hinchey, Sooyong Park, and Klaus Schmid. Dynamic software product lines. Computer, 41(4):93–95, April 2008. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jurik:2008:RMM

[2269] Andrew D. Jurik and Alfred C. Weaver. Remote medical monitoring. Computer, 41(4):96–99, April 2008. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vanLent:2008:MMS

[2270] Michael van Lent. Massive media shift. Computer, 41(4):100–101, April 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yang:2008:ARD

[2271] Jeonghwa Yang, Bill N. Schilit, and David W. McDonald. Activity recognition for the digital home. Computer, 41 (4):102–104, April 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Olsen:2008:KE

[2272] Kai A. Olsen. The $100,000 keying error. Computer, 41(4):108, 106–107, April 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://bell.computer.
                                        org/dlcomments/.

Anonymous:2008:ASe

[2273] Anonymous. Article summaries. Computer, 41(5):4, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2008/05/mco2008050004.pdf.

Grier:2008:KYA

[2274] David Alan Grier. Know your audience. Computer, 41(5):6–8, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:YAd

[2275] Anonymous. 32 & 16 years ago. Computer, 41(5):9–10, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2008:IFT

[2276] George Lawton. Is it finally time to worry about mobile malware? Computer, 41(5):12–14, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:2008:AIB

[2277] Brad Smith. ARM and Intel battle over the mobile chip’s future. Computer, 41(5):15–18, May 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2008:NBd

[2278] Linda Dailey Paulson. News briefs.

Computer, 41(5):19–21, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Spinellis:2008:AMC

[2279] Diomidis Spinellis. The Antikythera mechanism: A computer science perspective. Computer, 41(5):22–27, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hutchby:2008:ENM

[2280] James A. Hutchby, Ralph Cavin, Victor Zhirnov, Joe E. Brewer, and George Bourianoff. Emerging nanoscale memory and logic devices: A critical assessment. Computer, 41(5):28–32, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cavin:2008:ERA

[2281] Ralph Cavin, James A. Hutchby, Victor Zhirnov, Joe E. Brewer, and George Bourianoff. Emerging research architectures. Computer, 41(5):33–37, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bourianoff:2008:BLA

[2282] George Bourianoff, Joe E. Brewer, Ralph Cavin, James A. Hutchby, and Victor Zhirnov. Boolean logic and alternative information-processing devices. Computer, 41(5):38–46, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Corso:2008:PPP

[2283] Jason J. Corso, Guangqi Ye, Darius Burschka, and Gregory D. Hager. A practical paradigm and platform for video-based human-computer interaction. Computer, 41(5):48–55, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Varshney:2008:IWH

[2284] Upkar Varshney. Improving wireless health monitoring using incentivebased router cooperation. Computer, 41(5):56–62, May 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:Be

[2285] Anonymous. Bookshelf. Computer, 41(5):63, May 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
05/mco2008050063.pdf.

Anonymous:2008:COc

[2286] Anonymous. Career opportunities.

Computer, 41(5):64–66, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2008/05/mco2008050064.
pdf.

Kasturi:2008:IAT

[2287] Rangachar Kasturi and Carl K. Chang.

Introducing the Advanced Technology Executive Forum. Computer, 41 (5):67–69, May 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2008:CSR

[2288] Bob Ward. Computer Society releases Q3 & 4 conference slate. Computer, 41(5):70–73, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2008/05/mco2008050070.pdf.

Anonymous:2008:CCd

[2289] Anonymous. Call and calendar. Computer, 41(5):77–80, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2008/05/mco2008050077.pdf.

Hardaway:2008:NMI

[2290] Don Hardaway, Richard G. Mathieu, and Richard Will. A new mission for the information systems discipline. Computer, 41(5):81–83, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sanchirico:2008:VBN

[2291] Paul Sanchirico. Video 2.0: Building next-generation video networks. Computer, 41(5):84–86, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Seidman:2008:SEC

[2292] Stephen Seidman. Software engineering certification schemes. Computer, 41(5):87–89, May 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2008:CP

[2293] Neville Holmes. The craft of programming. Computer, 41(5):92, 90–91, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://bell.computer.
                                        org/dlcomments/.

Anonymous:2008:ASf

[2294] Anonymous. Article summaries. Computer, 41(6):4, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2008/06/mco2008060004.pdf.

Anonymous:2008:LEd

[2295] Anonymous. Letters to the editor. Computer, 41(6):6–7, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://csdl.computer.org/
                                        comp/mags/co/2008/06/mco2008060006.
                                        pdf.

Grier:2008:TB

[2296] David Alan Grier. The theosophists’ bathroom. Computer, 41(6):8–10, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:YAe

[2297] Anonymous. 32 & 16 years ago. Computer, 41(6):11–12, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2008:DSO

[2298] George Lawton. Developing software online with platform-as-a-service technology. Computer, 41(6):13–15, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaughan-Nichols:2008:MWN

[2299] Steven J. Vaughan-Nichols. Mobile WiMax: The next wireless battle ground. Computer, 41(6):16–18, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2008:NWF

[2300] Linda Dailey Paulson. A new wi-fi for peer-to-peer communications. Computer, 41(6):19–21, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Toyama:2008:ICT

[2301] Kentaro Toyama and M. Bernardine Dias. Information and communication technologies for development. Computer, 41(6):22–25, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2008/06/mco2008060022.pdf.

Heeks:2008:INP

[2302] Richard Heeks. ICT4D 2.0: The next phase of applying ICT for international development. Computer, 41 (6):26–33, June 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:2008:SDT

[2303] Jonathan Donner, Rikin Gandhi, Paul Javid, Indrani Medhi, Aishwarya Ratan, Kentaro Toyama, and Rajesh Veeraraghavan. Stages of design in technology for global development. Computer, 41(6):34–41, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marsden:2008:TED

[2304] Gary Marsden. Toward empowered design. Computer, 41(6):42–46, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Surana:2008:DRW

[2305] Sonesh Surana, Rabin Patra, Sergiu Nedevschi, and Eric Brewer. Deploying a rural wireless telemedicine system: Experiences in sustainability. Computer, 41(6):48–56, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Breitman:2008:IRE

[2306] Karin Breitman and Mike Hinchey. ICT research experience. Computer, 41(6):58, June 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Medeiros:2008:GRC

[2307] Claudia Bauzer Medeiros. Grand research challenges in computer science in Brazil. Computer, 41(6):59–65, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ryan:2008:EIS

[2308] Kevin Ryan. Engineering the Irish software tiger. Computer, 41(6):66–71, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:COd

[2309] Anonymous. Career opportunities.

Computer, 41(6):72–76, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2008/06/mco2008060072.
pdf.

Ward:2008:SLC

[2310] Bob Ward. Society launches Computing Now site. Computer, 41(6): 77–79, June 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
06/mco2008060077.pdf.

Ward:2008:CCb

[2311] Bob Ward. Call and calendar. Computer, 41(6):80–82, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2008/06/mco2008060080.pdf.

Anonymous:2008:Bf

[2312] Anonymous. Bookshelf. Computer, 41(6):83, June 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
06/mco2008060083.pdf.

Magerko:2008:ADG

[2313] Brian Magerko. Adaptation in digital games. Computer, 41(6):87–89, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hazen:2008:TBH

[2314] Mark E. Hazen. The technology behind HomePlug AV powerline communications. Computer, 41(6):90–92, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Li:2008:WLC

[2315] Yangi Li and James A. Landay. Into the wild: Low-cost Ubicomp prototype testing. Computer, 41(6):94–97, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lindvall:2008:BSA

[2316] Mikael Lindvall and Dirk Muthig. Bridging the software architecture gap. Computer, 41(6):98–101, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Martin:2008:KMI

[2317] Luther Martin. Key-management infrastructure for protecting stored data. Computer, 41(6):103–104, June 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mahmoud:2008:MDI

[2318] Qusay H. Mahmoud and Allan Dyer. Mobile devices in an introductory programming course. Computer, 41(6): 108, 106–107, June 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://bell.

                                        computer.org/dlcomments/.

Anonymous:2008:ASg

[2319] Anonymous. Article summaries. Computer, 41(7):4, July 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2008/07/mco2008070004.pdf.

Grier:2008:HG

[2320] David Alan Grier. Harry and Georgie.

Computer, 41(7):5–8, July 2008. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:YAf

[2321] Anonymous. 32 & 16 years ago. Computer, 41(7):9–10, July 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2008:AAM

[2322] Neal Leavitt. Application awareness makes storage more useful. Computer, 41(7):11–13, July 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2008:WIB

[2323] Sixto Ortiz, Jr. The wireless industry begins to embrace femtocells. Computer, 41(7):14–17, July 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2008:NBe

[2324] Linda Dailey Paulson. News briefs.

Computer, 41(7):18–20, July 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2008/07/mco2008070018.
pdf.

Loui:2008:PSR

[2325] Ronald P. Loui. In praise of scripting: Real programming pragmatism. Computer, 41(7):22–26, July 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dally:2008:EEC

[2326] William J. Dally, James Balfour, David Black-Shaffer, James Chen, R. Curtis Harting, Vishal Parikh, Jongsoo Park, and David Sheffield. Efficient embedded computing. Computer, 41 (7):27–32, July 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hill:2008:ALM

[2327] Mark D. Hill and Michael R. Marty. Amdahl’s Law in the multicore era. Computer, 41(7):33–38, July 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.cs.wisc.
edu/multifacet/amdahl/;   http:/
/www.cs.wisc.edu/multifacet/papers/
ieeecomputer08_amdahl_multicore.
pdf.

Vahid:2008:WPD

[2328] Frank Vahid, Greg Stitt, and Roman Lysecky. Warp processing: Dynamic translation of binaries to FPGA circuits. Computer, 41(7):40–46, July 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chang:2008:APD

[2329] Kai-Hui Chang, Igor L. Markov, and Valeria Bertacco. Automating postsilicon debugging and repair. Computer, 41(7):47–54, July 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wang:2008:SMS

[2330] Kun Wang, Zhonghai Yin, Lihua Zhou, Feng Yuan, and Zengxin Li. SI-ISAM: A model for secure inhomogeneous information systems. Computer, 41(7):56–61, July 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:COe

[2331] Anonymous. Career opportunities.

Computer, 41(7):62–64, July 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2008/07/mco2008070062.
pdf.

Ward:2008:CSCd

[2332] Bob Ward. Computer Society connection. Computer, 41(7):65–71, July 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
07/mco2008070065.pdf.

Ward:2008:CCc

[2333] Bob Ward. Call and calendar. Computer, 41(7):72–75, July 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2008/07/mco2008070072.pdf.

Anonymous:2008:Bg

[2334] Anonymous. Bookshelf. Computer, 41(7):79, July 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2008/
                                        07/mco2008070079.pdf.

Arenas:2008:CTS

[2335] Alvaro Arenas and Michael Wilson. Contracts as trust substitutes in collaborative business. Computer, 41 (7):80–83, July 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sarji:2008:HAT

[2336] David K. Sarji. HandTalk: Assistive technology for the deaf. Computer, 41 (7):84–86, July 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kelly:2008:ISI

[2337] Diane Kelly. Innovative standards for innovative software. Computer, 41 (7):88–89, July 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mathieson:2008:MEE

[2338] Kieran Mathieson. Making ethics easier. Computer, 41(7):91–93, July 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rutkowski:2008:NMR

[2339] Anne-Françoise Rutkowski and Michiel van Genuchten. No more reply-to-all. Computer, 41(7):96, 94–95, July 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://bell.computer.
                                        org/dlcomments/.

Anonymous:2008:ASh

[2340] Anonymous. Article summaries. Computer, 41(8):4, August 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/csdl.computer.org/comp/mags/co/
2008/08/mco2008080004.pdf.

Grier:2008:FD

[2341] David Alan Grier. Figures in the distance. Computer, 41(8):5–7, August 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:YAg

[2342] Anonymous. 32 & 16 years ago.

Computer, 41(8):8–9, August 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2008:NWB

[2343] George Lawton. New ways to build rich Internet applications. Computer, 41 (8):10–12, August 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaughan-Nichols:2008:VSS

[2344] Steven J. Vaughan-Nichols. Virtualization sparks security concerns. Computer, 41(8):13–15, August 2008. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2008:NBBa

[2345] Linda Dailey Paulson. News briefs:

Billboard-viewership technology sparks privacy concerns; adding uncertainty to improve chips; vendors develops 3D technology for handhelds. Computer, 41(8):16–18, August 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carin:2008:CSQ

[2346] Lawrence Carin, George Cybenko, and Jeff Hughes. Cybersecurity strategies: The QuERIES methodology. Computer, 41(8):20–26, August 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dong:2008:HAS

[2347] Jing Dong, Raymond A. Paul, and Liang-Jie Zhang. High-assurance service-oriented architectures. Computer, 41(8):27–28, August 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2008/08/mco2008080027.
pdf.

Townend:2008:CCH

[2348] Paul Townend, Jinpeng Huai, Jie Xu, Nik Looker, Dacheng Zhang, Jianxin Li, and Liang Zhong. CROWN-C: A high-assurance service-oriented Grid middleware system. Computer, 41(8): 30–38, August 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tsai:2008:TES

[2349] Wei-Tek Tsai, Xinyu Zhou, Yinong Chen, and Xiaoying Bai. On testing and evaluating service-oriented software. Computer, 41(8):40–46, August 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yen:2008:QRW

[2350] I-Ling Yen, Hui Ma, Farokh B. Bastani, and Hong Mei. QoS-reconfigurable Web services and compositions for high-assurance systems. Computer, 41 (8):48–55, August 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sloan:2008:ATS

[2351] John C. Sloan, Taghi M. Khoshgoftaar, and Venkat Raghav. Assuring timeliness in an e-science serviceoriented architecture. Computer, 41 (8):56–62, August 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Atkinson:2008:SHA

[2352] Colin Atkinson, Daniel Brenner, Giovanni Falcone, and Monika Juhasz. Specifying high-assurance services. Computer, 41(8):64–71, August 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2008:IPE

[2353] Bob Ward. IEEE President-Elect candidates address Computer Society concerns. Computer, 41(8):72– 77, August 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
08/mco2008080072.pdf.

Ward:2008:CSCe

[2354] Bob Ward. Computer Society connection. Computer, 41(8):78–79, August 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

computer.org/comp/mags/co/2008/
08/mco2008080078.pdf.

Ward:2008:CCd

[2355] Bob Ward. Call and calendar. Computer, 41(8):80–82, August 2008. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:Bh

[2356] Anonymous. Bookshelf. Computer, 41(8):83, August 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2008/
                                        08/mco2008080083.pdf.

vanLent:2008:FVH

[2357] Michael van Lent. The future is virtually here. Computer, 41(8):87–89, August 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Watts:2008:ANR

[2358] Lloyd Watts. Advanced noise reduction for mobile telephony. Computer, 41 (8):90–92, August 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Martin:2008:IBE

[2359] Luther Martin. Identity-based encryption comes of age. Computer, 41(8): 93–95, August 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Meyer:2008:SPS

[2360] Bertrand Meyer. Seven principles of software testing. Computer, 41(8): 99–101, August 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mankoff:2008:SCS

[2361] Jennifer Mankoff, Robin Kravets, and Eli Blevis. Some computer science issues in creating a sustainable world. Computer, 41(8):102–105, August 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2008:EUS

[2362] Neville Holmes. The European Union and the Semantic Web. Computer, 41 (8):106, 107–108, August 2008. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:ASi

[2363] Anonymous. Article summaries.

Computer, 41(9):4, September 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://csdl.computer.org/
comp/mags/co/2008/09/mco2008090004.
pdf.

Grier:2008:KWC

[2364] David Alan Grier. The known world:

Click here to empty trash. Computer, 41(9):6–8, September 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:YAh

[2365] Anonymous. 32 and 16 years ago. Computer, 41(9):9–10, September 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2008:ITSa

[2366] Sixto Ortiz, Jr. Industry trends:

Stitching computing systems into a single fabric. Computer, 41(9):11– 13, September 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2008:TNN

[2367] George Lawton. Technology news: New technology prevents data leakage. Computer, 41(9):14–17, September 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2008:NBG

[2368] Linda Dailey Paulson. News briefs: Group releases first 3-D chip standard; supercomputers get energy-efficient; company puts a new spin on musicediting software; attack hijacks domain names of Internet-addressing organization. Computer, 41(9):18–20, September 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sheng:2008:ENG

[2369] Quan Z. Sheng, Xue Li, and Sherali Zeadally. Enabling next-generation RFID applications: Solutions and challenges. Computer, 41(9):21–28, September 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lam:2008:XDP

[2370] Tak Cheung (Brian) Lam, Jianxun Jason Ding, and Jyh-Charn Liu. XML document parsing: Operational and performance characteristics. Computer, 41(9):30–37, September 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Symborski:2008:SUC

[2371] Carl Symborski. Scalable user content distribution for massively multiplayer online worlds. Computer, 41(9): 38–44, September 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kumar:2008:SLN

[2372] Sanjeev Kumar, Jatin Chhugani, Changkyu Kim, Daehyun Kim, Anthony Nguyen, Pradeep Dubey, Christian Bienia, and Youngmin Kim. Second life and the new generation of virtual worlds. Computer, 41(9):46– 53, September 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:2008:RFE

[2373] Ken Smith, Len Seligman, and Vipin Swarup. Research feature: Everybody share: The challenge of datasharing systems. Computer, 41(9):54– 61, September 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lysne:2008:INA

[2374] Olav Lysne, Sven-Arne Reinemo, Tor Skeie, Ashild Gronstad Solheim, Thomas Sodring, Lars Paul Huse, and Bjorn Dag Johnsen. Interconnection networks: Architectural challenges for utility computing data centers. Computer, 41(9):62–69, September 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:RMI

[2375] Anonymous. Report to members: IEEE Computer Society elections. Computer, 41(9):71–79, September 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2008/
                                        09/mco2008090071.pdf.

Ward:2008:CCe

[2376] Bob Ward. Call and calendar. Computer, 41(9):84–86, September 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://csdl.

                                        computer.org/comp/mags/co/2008/
                                        09/mco2008090084.pdf.

Anonymous:2008:Bi

[2377] Anonymous. Bookshelf. Computer, 41(9):87, September 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /csdl.computer.org/comp/mags/co/
                                        2008/09/mco2008090087.pdf.

Chi:2008:WTS

[2378] Ed H. Chi. Web technologies: The Social Web: Research and opportunities. Computer, 41(9):88–91, September 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeMarle:2008:SIV

[2379] Ann DeMarle. Standards: Innovation and value. Computer, 41(9):92– 94, September 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geller:2008:ISP

[2380] James Geller, Soon Ae Chun, and Yoo Jung. IT systems perspectives: Toward the Semantic Deep Web. Computer, 41 (9):95–97, September 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gorton:2008:PXD

[2381] Ian Gorton. The profession: XML does real programmers a service. Computer, 41(9):100, 98–99, September 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://bell.computer.org/
dlcomments/;   http://www.adobe.
com/products/acrobat/readstep2.
html.

Anonymous:2008:DASa

[2382] Anonymous. Departments: Article summaries. Computer, 41(10):4, October 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:La

[2383] Anonymous. Letters. Computer, 41 (10):6–7, October 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2008:KWD

[2384] David Alan Grier. The known world:

Drinking with dinosaurs. Computer, 41 (10):8–10, October 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:YAYa

[2385] Anonymous. 32 and 16 years ago: 32 and 16 years ago. Computer, 41(10): 11–12, October 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2008:ITH

[2386] Neal Leavitt. Industry trends: Highstakes battle rages in graphics-chip marketplace. Computer, 41(10):13– 15, October 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geer:2008:TNT

[2387] David Geer. Technology news: Technology news. Computer, 41(10):16– 19, October 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2008:NBNa

[2388] Linda Dailey Paulson. News briefs: News briefs. Computer, 41(10):20– 22, October 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riva:2008:CPC

[2389] Oriana Riva and Jaakko Kangasharju. Computing practices: Challenges and lessons in developing middleware on smart phones. Computer, 41(10): 23–31, October 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bimber:2008:GEI

[2390] Oliver Bimber. Guest Editor’s introduction: Total recall. Computer, 41 (10):32–33, October 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cichocki:2008:CFN

[2391] Andrzej Cichocki, Yoshikazu Washizawa, Tomasz Rutkowski, Hovagim Bakardjian, Anh-Huy Phan, Seungjin Choi, Hyekyoung Lee, Qibin Zhao, Liqing Zhang, and Yuanqing Li. Cover feature: Noninvasive BCIs: Multiway signalprocessing array decompositions. Computer, 41(10):34–42, October 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schlogl:2008:BFO

[2392] Alois Schlögl and Clemens Brunner.

BioSig: A free and open source software library for BCI research. Computer, 41(10):44–50, October 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McFarland:2008:BCI

[2393] Dennis J. McFarland and Jonathan R.

Wolpaw. Brain-computer interface operation of robotic and prosthetic devices. Computer, 41(10):52–56, October 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pfurtscheller:2008:RBC

[2394] Gert Pfurtscheller, Gernot R. M?ller-Putz, Reinhold Scherer, and Christa Neuper. Rehabilitation with braincomputer interface systems. Computer, 41(10):58–65, October 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lcuyer:2008:BCI

[2395] Anatole L’écuyer, Fabien Lotte, Richard B. Reilly, Robert Leeb, Michitaka Hirose, and Mel Slater. Braincomputer interfaces, virtual reality, and videogames. Computer, 41(10): 66–72, October 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2008:CSCf

[2396] Bob Ward. Computer Society connection: Computer Society honors top experts. Computer, 41(10):82–84, October 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2008:CCf

[2397] Bob Ward. Call and calendar. Computer, 41(10):85–87, October 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:Bj

[2398] Anonymous. Bookshelf. Computer, 41 (10):91, October 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stankovic:2008:HTW

[2399] John A. Stankovic. How things work: Wireless sensor networks. Computer, 41(10):92–95, October 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sangwan:2008:STS

[2400] Raghvinder S. Sangwan, Li-Ping Lin, and Colin J. Neill. Software technologies: Structural complexity in architecture-centric software evolution. Computer, 41(10):96–99, October 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Raman:2008:ICE

[2401] T. V. Raman and Charles L. Chen. Invisible computing: Eyes-free user interfaces. Computer, 41(10):100–101, October 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Higa:2008:ECW

[2402] Darold Higa. Entertainment computing: Walled gardens versus the Wild West. Computer, 41(10):102– 105, October 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2008:PDH

[2403] Neville Holmes. The profession:

The dea[r]th of human understanding. Computer, 41(10):108, 106–107, October 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:DASb

[2404] Anonymous. Departments: Article summaries. Computer, 41(11): 4, November 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2008:KWEa

[2405] David Alan Grier. The known world:

Edward Elgar’s Facebook page. Computer, 41(11):6–8, November 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:YAYb

[2406] Anonymous. 32 and 16 years ago: 32 and 16 years ago. Computer, 41(11): 9–10, November 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2008:ITSb

[2407] Sixto Ortiz Jr. Industry trends: SRM manages the exploding volume of storage. Computer, 41(11):11–14, November 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaughan-Nichols:2008:TNM

[2408] Steven J. Vaughan-Nichols. Technology news: The Mobile Web comes of age. Computer, 41(11):15–17, November 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2008:NBNb

[2409] Linda Dailey Paulson. News briefs: News briefs. Computer, 41(11):18– 20, November 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Price:2008:PPS

[2410] Michael Price. Perspectives: The paradox of security in virtual environments. Computer, 41(11):22– 28, November 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Getov:2008:GEI

[2411] Vladimir Getov. Guest Editor’s introduction: e-Science: The added value for modern discovery. Computer, 41 (11):30–31, November 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wilkins-Diehr:2008:CFT

[2412] Nancy Wilkins-Diehr, Dennis Gannon, Gerhard Klimeck, Scott Oster, and Sudhakar Pamidighantam. Cover feature: TeraGrid science gateways and their impact on science. Computer, 41 (11):32–41, November 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Matsuoka:2008:CSS

[2413] Satoshi Matsuoka, Kazushige Saga, and Mutsumi Aoyagi. Coupledsimulation e-Science support in the NAREGI Grid. Computer, 41(11):42– 49, November 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:2008:SBC

[2414] Craig Lee and George Percivall.

Standards-based computing capabilities for distributed geospatial applications. Computer, 41(11):50–57, November 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Saltz:2008:SCT

[2415] Joel Saltz, Tahsin Kurc, Shannon Hastings, Stephen Langella, Scott Oster, David Ervin, Ashish Sharma, Tony Pan, Metin Gurcan, Justin Permar, Renato Ferreira, Philip Payne, Umit Catalyurek, Enrico Caserta, Gustavo Leone, Michael C. Ostrowski, Ravi Madduri, Ian Foster, Subhashree Madhavan, Kenneth H. Buetow, Krishnakant Shanbhag, and Eliot Siegel. e-Science, caGrid, and translational biomedical research. Computer, 41(11): 58–66, November 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Perrott:2008:SID

[2416] Ron Perrott, Terry Harmer, and Rhys Lewis. e-Science infrastructure for digital media broadcasting. Computer, 41 (11):67–72, November 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:Bk

[2417] Anonymous. Bookshelf. Computer, 41 (11):73, November 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2008:CSCg

[2418] Bob Ward. Computer Society connection: Society magazines plan coverage for 2009. Computer, 41(11): 89–92, November 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2008:CCg

[2419] Bob Ward. Call and calendar. Computer, 41(11):93–95, November 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Walker:2008:SEM

[2420] Alastair J. Walker. Standards: Enterprise maturity models: Have we lost the plot? Computer, 41(11): 96–98, November 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gal-On:2008:ECM

[2421] Shay Gal-On and Markus Levy. Embedded computing: Measuring multicore performance. Computer, 41(11): 99–102, November 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Huberman:2008:WTC

[2422] Bernardo A. Huberman. Web technologies: Crowdsourcing and attention.

Computer, 41(11):103–105, November 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cellary:2008:PCC

[2423] Wojciech Cellary. The profession:

Content communities on the Internet. Computer, 41(11):106–108, November 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:DASc

[2424] Anonymous. Departments: Article summaries. Computer, 41(12):4, December 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:Lb

[2425] Anonymous. Letters. Computer, 41 (12):5–6, December 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kasturi:2008:PME

[2426] Rangachar Kasturi. President’s message: Enhancements to Society’s products and services: Report of progress made in 2008. Computer, 41(12): 7–9, December 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2008:KWEb

[2427] David Alan Grier. The known world:

Evolutionary fervor. Computer, 41(12): 10–12, December 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:YAYc

[2428] Anonymous. 32 and 16 years ago: 32 and 16 years ago. Computer, 41(12):

13–14, December 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geer:2008:ITR

[2429] David Geer. Industry trends: Reducing the storage burden via data deduplication. Computer, 41(12):15– 17, December 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:TNW

[2430] Anonymous. Technology news: Wireless HD video heats up. Computer, 41 (12):18–20, December 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2008:NNB

[2431] Linda Dailey Paulson. News: News briefs. Computer, 41(12):21–23, December 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woo:2008:PEA

[2432] Dong Hyuk Woo and Hsien-Hsin S. Lee. Perspectives: Extending Amdahl’s Law for energy-efficient computing in the many-core era. Computer, 41(12): 24–31, December 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hinchey:2008:GEI

[2433] Mike Hinchey, Paul A. Curto, and Scott Hamilton. Guest Editors’ introduction: Computer celebrates NASA’s 50th anniversary. Computer, 41(12): 32–33, December 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hamilton:2008:CFU

[2434] Margaret H. Hamilton and William R.

Hackler. Cover feature: Universal systems language: Lessons learned from Apollo. Computer, 41(12):34– 43, December 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bajracharya:2008:AMR

[2435] Max Bajracharya, Mark W. Maimone, and Daniel Helmick. Autonomy for Mars Rovers: Past, present, and future. Computer, 41(12):44–50, December 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kumar:2008:IHM

[2436] Sujay Kumar, Christa Peters-Lidard, Yudong Tian, Rolf Reichle, James Geiger, Charles Alonge, John Eylander, and Paul Houser. An integrated hydrologic modeling and data assimilation framework. Computer, 41(12): 52–59, December 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pinninghoff:2008:RFU

[2437] Maria Pinninghoff, Ricardo Contreras, and John Atkinson. Research feature: Using genetic algorithms to model road networks. Computer, 41(12):60– 67, December 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:Bl

[2438] Anonymous. Bookshelf. Computer, 41 (12):71, December 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2008:AI

[2439] Anonymous. Annual index. Computer, 41(12):84–96, December 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2008:RMI

[2440] Bob Ward. Report to members: Isaak named 2009 Computer Society President-Elect. Computer, 41(12):97– 98, December 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2008:CSCh

[2441] Bob Ward. Computer Society connection: Computer recognizes expert reviewers and departing editors. Computer, 41(12):99–100, December 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2008:CCh

[2442] Bob Ward. Call and calendar. Computer, 41(12):101–103, December 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jackson:2008:BPC

[2443] Jerlando F. L. Jackson and Elizabeth M. O’Callaghan. Broadening participation in computing: Computer science faculty diversity during the past decade. Computer, 41(12):104– 107, December 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McGraw:2008:HTW

[2444] Gary McGraw. How things work: Automated code review tools for security. Computer, 41(12):108–111, December 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kornecki:2008:STS

[2445] Andrew Kornecki, Brian Butka, and Janusz Zalewski. Software technologies: Software tools for safety-critical systems according to DO-254. Computer, 41(12):112–115, December 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jensen:2008:ICU

[2446] Christian S. Jensen, Carmen Ruiz Vicente, and Rico Wind. Invisible computing: User-generated content: The case for mobile services. Computer, 41 (12):116–118, December 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Funge:2008:ECL

[2447] John Funge. Entertainment computing: Let us entertain you. Computer, 41 (12):120–122, December 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cole:2008:SNS

[2448] John L. Cole. Security: Network science from an IT and US Army perspective. Computer, 41(12):123– 125, December 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Santini:2008:PMB

[2449] Simone Santini. The profession:

Madrid’s brilliant idea. Computer, 41(12):128, 126–127, December 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:ECCa

[2450] Anonymous. Elsewhere in the CS: Computer highlights society magazines. Computer, 42(1):4–5, January 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Land:2009:PML

[2451] Susan K. (Kathy) Land. President’s message: A leadership challenge in 2009. Computer, 42(1):6–8, January 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2009:KWP

[2452] David Alan Grier. The known world: Politics in play. Computer, 42(1): 9–12, January 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:YAa

[2453] Anonymous. 32 and 16 years ago. Computer, 42(1):13–14, January 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2009:TNC

[2454] Neal Leavitt. Technology news: Is cloud computing really ready for prime time? Computer, 42(1):15–20, January 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2009:NBSa

[2455] Linda Dailey Paulson. News briefs: Start-up unveils flexible supercomputing approach. Computer, 42(1):21– 24, January 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kouzes:2009:CFC

[2456] Richard T. Kouzes, Gordon A. Anderson, Stephen T. Elbert, Ian Gorton, and Deborah K. Gracio. Cover feature: The changing paradigm of dataintensive computing. Computer, 42(1): 26–34, January 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bird:2009:OMG

[2457] Ian Bird, Bob Jones, and Kerk F. Kee.

The organization and management of grid infrastructures. Computer, 42(1): 36–46, January 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ramakrishnan:2009:CP

[2458] Naren Ramakrishnan, Upinder S.

Bhalla, and John J. Tyson. Computing with proteins. Computer, 42(1): 47–56, January 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hanson:2009:BAS

[2459] Mark A. Hanson, Harry C. Powell, Jr., Adam T. Barth, Kyle Ringgenberg, Benton H. Calhoun, James H. Aylor, and John Lach. Body area sensor networks: Challenges and opportunities. Computer, 42(1):58–65, January 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gupta:2009:HKF

[2460] Amar Gupta. The 24-hour knowledge factory: Can it replace the graveyard shift? Computer, 42(1):66–73, January 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berenbach:2009:PED

[2461] Brian Berenbach and Manfred Broy. Professional and ethical dilemmas in software engineering. Computer, 42(1): 74–80, January 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CSCa

[2462] Bob Ward. Computer Society connection: Board of Governors votes to amend bylaws. Computer, 42(1): 91–98, January 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CCa

[2463] Bob Ward. Call and calendar. Computer, 42(1):99–100, January 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cameron:2009:GIG

[2464] Kirk W. Cameron. Green IT: Green introspection. Computer, 42(1):101– 103, January 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Al-Masri:2009:ISP

[2465] Eyhab Al-Masri and Qusay H. Mahmoud. IT systems perspectives: Web service discovery and client goals. Computer, 42(1):104–107, January 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Alexander:2009:SRS

[2466] Perry Alexander. Standards: Rosetta: Standardization at the system level. Computer, 42(1):108–110, January 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hendler:2009:WTW

[2467] Jim Hendler. Web technologies: Web 3.0 emerging. Computer, 42(1):111– 113, January 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2009:PCC

[2468] Neville Holmes. The profession: The credit crunch and the digital bite. Computer, 42(1):116, 114–115, January 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:ECCb

[2469] Anonymous. Elsewhere in the CS:

Computer highlights society magazines. Computer, 42(2):4–5, February 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:DLa

[2470] Anonymous. Departments: Letters.

Computer, 42(2):6–7, February 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2009:KWS

[2471] David Alan Grier. The known world:

Scanning the horizon. Computer, 42 (2):8–11, February 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:YAb

[2472] Anonymous. 32 and 16 years ago.

Computer, 42(2):12–13, February 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaughan-Nichols:2009:ITW

[2473] Steven J. Vaughan-Nichols. Industry trends: Will mobile computing’s future be location, location, location? Computer, 42(2):14–17, February 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2009:TNU

[2474] George Lawton. Technology news: Users take a close look at visual analytics. Computer, 42(2):19–22, February 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2009:NBW

[2475] Linda Dailey Paulson. News briefs: W3C adopts Web-accessibility specifications. Computer, 42(2):23–26, February 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grimshaw:2009:CPO

[2476] Andrew Grimshaw, Mark Morgan, Duane Merrill, Hiro Kishimoto, Andreas Savva, David Snelling, Chris Smith, and Dave Berry. Computing practices: An open grid services architecture primer. Computer, 42(2):27– 34, February 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Caicedo:2009:PIS

[2477] Carlos E. Caicedo, James B. D. Joshi, and Summit R. Tuladhar. Perspectives: IPv6 security challenges. Computer, 42 (2):36–42, February 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Blaze:2009:CFD

[2478] Matt Blaze, Sampath Kannan, Insup Lee, Oleg Sokolsky, Jonathan M. Smith, Angelos D. Keromytis, and Wenke Lee. Cover feature: Dynamic trust management. Computer, 42(2): 44–52, February 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Skogsrud:2009:MTN

[2479] Halvard Skogsrud, Hamid R. Motahari-Nezhad, Boualem Benatallah, and Fabio Casati. Modeling trust negotiation for Web services. Computer, 42 (2):54–61, February 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Beyls:2009:RFR

[2480] Kristof Beyls and Erik H. D’Hollander.

Research feature: Refactoring for data locality. Computer, 42(2):62–71, February 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liesaputra:2009:CRR

[2481] Veronica Liesaputra, Ian H. Witten, and David Bainbridge. Creating and reading realistic electronic books. Computer, 42(2):72–81, February 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:Ba

[2482] Anonymous. Bookshelf. Computer, 42 (2):85, February 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CSCb

[2483] Bob Ward. Computer Society connection: Computer Society awards honor top researchers. Computer, 42(2):89– 94, February 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CCb

[2484] Bob Ward. Call and calendar. Computer, 42(2):95–96, February 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Adcock:2009:ICW

[2485] Matt Adcock, Jae-Woo Chung, and Chris Schmandt. Invisible computing: Are we there yet? user-centered temporal awareness. Computer, 42(2): 97–99, February 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:2009:STS

[2486] Dong hyun Lee, Hoh Peter In, Keun Lee, Sooyong Park, and Mike Hinchey. Software technologies: A survival kit: Adaptive hardware/software codesign life-cycle model. Computer, 42(2):100– 102, February 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Swain:2009:ECW

[2487] Chris Swain. Entertainment computing: Who needs a publisher. Computer, 42(2):103–105, February 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mowshowitz:2009:PTT

[2488] Abbe Mowshowitz and Nanda Kumar.

The profession: And then there were three. Computer, 42(2):108, 106– 107, February 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:ECCc

[2489] Anonymous. Elsewhere in the CS:

Computer highlights society magazines. Computer, 42(3):4–5, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:DLE

[2490] Anonymous. Departments: Letters to the editor. Computer, 42(3):6–7, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2009:KWV

[2491] David Alan Grier. The known world:

Virtual walls. Computer, 42(3):8–10, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:YAc

[2492] Anonymous. 32 and 16 years ago.

Computer, 42(3):11–12, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2009:ITW

[2493] George Lawton. Industry trends:

WDM-PONs hold promise for the long haul. Computer, 42(3):13–16, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geer:2009:TNN

[2494] David Geer. Technology news: Networks on processors improve on-chip communications. Computer, 42(3): 17–20, March 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2009:NBN

[2495] Linda Dailey Paulson. News briefs: News briefs. Computer, 42(3):21–24, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parhami:2009:CPP

[2496] Behrooz Parhami. Computing practices: Puzzling problems in computer engineering. Computer, 42(3): 26–29, March 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marchionini:2009:GEI

[2497] Gary Marchionini and Ryen W. White. Guest Editors’ introduction: Information-seeking support systems. Computer, 42(3):30–32, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pirolli:2009:CFP

[2498] Peter Pirolli. Cover feature: Powers of 10: Modeling complex informationseeking systems at multiple scales. Computer, 42(3):33–40, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chi:2009:ISC

[2499] Ed H. Chi. Information seeking can be social. Computer, 42(3):42–46, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Golovchinsky:2009:CIS

[2500] Gene Golovchinsky, Pernilla Qvarfordt, and Jeremy Pickens. Collaborative information seeking. Computer, 42(3): 47–51, March 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

mcschraefel:2009:BKW

[2501] m. c. schraefel. Building knowledge: What’s beyond keyword search? Computer, 42(3):52–59, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kelly:2009:ECD

[2502] Diane Kelly, Susan Dumais, and Jan O.

Pedersen. Evaluation challenges and directions for information-seeking support systems. Computer, 42(3):60–66, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Russell:2009:IAR

[2503] Daniel M. Russell. Industry-academic relationships. Computer, 42(3):67–68, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CSCc

[2504] Bob Ward. Computer Society connection: Computer Society sponsors prizes at Intel Science Fair. Computer, 42 (3):69–71, March 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CCc

[2505] Bob Ward. Call and calendar. Computer, 42(3):72–74, March 2009. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:Bb

[2506] Anonymous. Bookshelf. Computer, 42(3):75, March 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Baggi:2009:SIM

[2507] Denis Baggi and Goffredo Haus. Standards: IEEE 1599: Music encoding and interaction. Computer, 42(3): 84–87, March 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wolf:2009:ECC

[2508] Wayne Wolf. Embedded computing: Cyber-physical systems. Computer, 42 (3):88–89, March 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Arel:2009:ISP

[2509] Itamar Arel and Scott Livingston. IT systems perspectives: Beyond the Turing Test. Computer, 42(3):90–91, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Samtani:2009:WTO

[2510] Rajan Samtani. Web technologies: Ongoing innovation in digital watermarking. Computer, 42(3):92–94, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lange:2009:GII

[2511] Klaus-Dieter Lange. Green IT: Identifying shades of green: The SPECpower benchmarks. Computer, 42(3):95–97, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2009:PAO

[2512] Neville Holmes. The profession: The automation of originality. Computer, 42(3):100, 98–99, March 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:DEC

[2513] Bob Ward. Departments: Elsewhere in the CS. Computer, 42(4):4–5, April 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:La

[2514] Anonymous. Letters. Computer, 42 (4):6–7, April 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2009:KWT

[2515] David Alan Grier. The known world:

Top of the news. Computer, 42(4): 10–13, April 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:YAd

[2516] Anonymous. 32 and 16 years ago.

Computer, 42(4):14–15, April 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2009:ITC

[2517] Neal Leavitt. Industry trends: Complex-event processing poised for growth. Computer, 42(4):17–20, April 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2009:TNS

[2518] Sixto Ortiz Jr. Technology news: Securing the Internet’s routing infrastructure. Computer, 42(4):21–23, April 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2009:NBT

[2519] Linda Dailey Paulson. News briefs: Technique makes strong encryption easier to use. Computer, 42(4):24–27, April 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hwang:2009:PHV

[2520] Seong Oun Hwang. Perspectives: How viable is digital rights management? Computer, 42(4):28–34, April 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Walker:2009:CPR

[2521] Edward Walker. Computing practices: The real cost of a CPU hour. Computer, 42(4):35–41, April 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ebert:2009:CFE

[2522] Christof Ebert and Capers Jones. Cover feature: Embedded software: Facts, figures, and future. Computer, 42(4):42–52, April 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shokry:2009:MBV

[2523] Hesham Shokry and Mike Hinchey.

Model-based verification of embedded software. Computer, 42(4):53–59, April 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chockler:2009:RFR

[2524] Gregory Chockler, Rachid Guerraoui, Idit Keidar, and Marko Vukolic. Research feature: Reliable distributed storage. Computer, 42(4):60–67, April 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Figueroa:2009:SDA

[2525] Alejandro Figueroa, Günter Neumann, and John Atkinson. Searching for definitional answers on the Web using surface patterns. Computer, 42 (4):68–76, April 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:Bc

[2526] Anonymous. Bookshelf. Computer, 42(4):81, April 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CSCd

[2527] Bob Ward. Computer Society connection: CRA Survey counts new PhDs. Computer, 42(4):87–92, April 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CCd

[2528] Bob Ward. Call and calendar. Computer, 42(4):93–94, April 2009. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bongard:2009:ARB

[2529] Josh Bongard. AI redux: Biologically inspired computing. Computer, 42(4):95–98, April 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Balan:2009:ICD

[2530] Rajesh Krishna Balan and Narayan Ramasubbu. Invisible computing: The digital wallet: Opportunities and prototypes. Computer, 42(4):100–102, April 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2009:ECT

[2531] Michael Zyda, Marc Spraragen, and Balakrishnan Ranganathan. Entertainment computing: Testing behavioral models with an online game. Computer, 42(4):103–105, April 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Buckley:2009:STR

[2532] Jim Buckley. Software technologies: Requirements-based visualization tools for software maintenance and evolution. Computer, 42(4):106–108, April 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Doerschuk:2009:BPC

[2533] Peggy Doerschuk, Jiangjiang Liu, and Judith Mann. Broadening participation in computing: Increasing participation of females and underrepresented minorities in computing. Computer, 42 (4):110–113, April 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Olsen:2009:PHP

[2534] Kai A. Olsen. The profession: Inhouse programming is not passé: Automating originality. Computer, 42(4): 116, 114–115, April 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:DECa

[2535] Anonymous. Departments: Elsewhere in the CS. Computer, 42(5):4–5, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2009:KWD

[2536] David Alan Grier. The known world:

The dictator and the Web design. Computer, 42(5):6–8, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:YAe

[2537] Anonymous. 32 and 16 years ago. Computer, 42(5):9–10, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaughn-Nichols:2009:ITV

[2538] Steven J. Vaughn-Nichols. Industry trends: Vendors draw up a new graphics-hardware approach. Computer, 42(5):11–13, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2009:TNC

[2539] George Lawton. Technology news: The changing face of open source. Computer, 42(5):14–17, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2009:NBC

[2540] Linda Dailey Paulson. News briefs: The changing world of outsourcing. Computer, 42(5):18–21, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paul:2009:CPT

[2541] Nathanael Paul and Andrew S. Tanenbaum. Computing practices: Trustworthy voting: From machine to system. Computer, 42(5):23–29, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pacyna:2009:GEI

[2542] Piotr Pacyna, Anthony Rutkowski, Amardeo Sarma, and Kenji Takahashi. Guest Editors’ introduction: Trusted identity for all: Toward interoperable trusted identity management systems. Computer, 42(5):30–32, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chadwick:2009:CFA

[2543] David W. Chadwick and George Inman. Cover feature: Attribute aggregation in federated identity management. Computer, 42(5):33–40, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Madsen:2009:CSF

[2544] Paul Madsen and Hiroki Itoh. Challenges to supporting federated assurance. Computer, 42(5):42–49, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paci:2009:IAM

[2545] Federica Paci, Rodolfo Ferrini, Andrea Musci, Kevin Steuer, Jr., and Elisa Bertino. An interoperable approach to multifactor identity verification. Computer, 42(5):50–57, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

LApez:2009:STI

[2546] Gabriel López, Cánovas, Antonio F.

Gómez-Skarmeta, and Joao Girao. A SWIFT take on identity management. Computer, 42(5):58–65, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barranco:2009:RFB

[2547] Manuel Barranco, Julián Proenza, and Luís Almeida. Research feature: Boosting the robustness of controller area networks: CANcentrate and ReCANcentrate. Computer, 42(5):66–73, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CSCe

[2548] Bob Ward. Computer Society connection. Computer, 42(5):79–83, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:Bd

[2549] Anonymous. Bookshelf. Computer, 42(5):86, May 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cameron:2009:GIR

[2550] Kirk W. Cameron. Green IT: The road to greener IT pastures. Computer, 42 (5):87–89, May 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Birman:2009:WTE

[2551] Ken Birman, Jared Cantwell, Daniel Freedman, Qi Huang, Petko Nikolov, and Krzysztof Ostrowski. Web technologies: Edge mashups for serviceoriented collaboration. Computer, 42 (5):90–94, May 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kateeb:2009:ECH

[2552] Ali El Kateeb. Embedded computing: Hardware reconfiguration capability for third-generation sensor nodes. Computer, 42(5):95–97, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jansen:2009:ISP

[2553] Bernard J. Jansen, Theresa B. Flaherty, Ricardo Baeza-Yates, Lee Hunter, Brendan Kitts, and Jamie Murphy. IT systems perspectives: The components and impact of sponsored search. Computer, 42(5):98–101, May 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2009:PDG

[2554] Neville Holmes. The profession: The design of government. Computer, 42 (5):102–104, May 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:ECCd

[2555] Anonymous. Elsewhere in the CS: Computer highlights society magazines. Computer, 42(6):4–5, June 2009.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:DLb

[2556] Anonymous. Departments: Letters.

Computer, 42(6):6–7, June 2009. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2009:KWM

[2557] David Alan Grier. The known world:

Marking the fall of sparrows. Computer, 42(6):9–12, June 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:YAf

[2558] Anonymous. 32 and 16 years ago. Computer, 42(6):13–14, June 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geer:2009:ITF

[2559] David Geer. Industry trends: The future of mobile VoIP in the enterprise. Computer, 42(6):15–18, June 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2009:TNT

[2560] George Lawton. Technology news: On the trail of the Conficker Worm. Computer, 42(6):19–22, June 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2009:NBM

[2561] Linda Dailey Paulson. News briefs:

Materials breakthrough could eliminate bootups. Computer, 42(6):23–25, June 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Oppliger:2009:CPI

[2562] Rolf Oppliger, Ruedi Rytz, and Thomas Holderegger. Computing practices: Internet banking: Client-side attacks and protection mechanisms. Computer, 42(6):27–33, June 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rashid:2009:GEI

[2563] Awais Rashid, John Weckert, and Richard Lucas. Guest Editors’ introduction: Software engineering ethics in a digital world. Computer, 42 (6):34–41, June 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karahasanovic:2009:CFE

[2564] Amela Karahasanovic, Petter Bae Brandtzæg, Jeroen Vanattenhoven, Bram Lievens, Karen Torben Nielsen, and Jo Pierson. Cover feature: Ensuring trust, privacy, and etiquette in Web 2.0 applications. Computer, 42 (6):42–49, June 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pieters:2009:PPW

[2565] Wolter Pieters and André van Cleeff. The precautionary principle in a world of digital dependencies. Computer, 42 (6):50–56, June 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tamai:2009:SII

[2566] Tetsuo Tamai. Social impact of information system failures. Computer, 42 (6):58–65, June 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gotterbarn:2009:PPM

[2567] Donald Gotterbarn and Keith W.

Miller. The public is the priority: Making decisions using the software engineering code of ethics. Computer, 42 (6):66–73, June 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CSCf

[2568] Bob Ward. Computer Society connection: IEEE Foundation gives away more than $300,000. Computer, 42 (6):78–79, June 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CCe

[2569] Bob Ward. Call and calendar. Computer, 42(6):80–81, June 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:Be

[2570] Anonymous. Bookshelf. Computer, 42(6):85, June 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Consolvo:2009:ICD

[2571] Sunny Consolvo, James A. Landay, and David W. McDonald. Invisible computing: Designing for behavior change in everyday life. Computer, 42(6):86–89, June 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vassev:2009:STA

[2572] Emil Vassev and Mike Hinchey. Software technologies: ASSL: A software engineering approach to autonomic computing. Computer, 42(6): 90–93, June 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yates:2009:ARE

[2573] Alexander Yates. AI redux: Extracting world knowledge from the Web. Computer, 42(6):94–97, June 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ryoo:2009:SSA

[2574] Jungwoo Ryoo, Phil Laplante, and Rick Kazman. Security: In search of architectural patterns for software security. Computer, 42(6):98–100, June 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Radenski:2009:PEC

[2575] Atanas Radenski. The profession: Expanding computing as an academic service discipline. Computer, 42(6):104, 101–103, June 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:C

[2576] Anonymous. Cover. Computer, 42(7): 3, July 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:DECb

[2577] Anonymous. Departments: Elsewhere in the CS. Computer, 42(7):4–5, July 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2009:KWC

[2578] David Alan Grier. The known world: The character of play. Computer, 42(7):7–9, July 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:YAg

[2579] Anonymous. 32 and 16 years ago. Computer, 42(7):10–11, July 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ortiz:2009:TNI

[2580] Sixto Ortiz Jr. Technology news: IEEE 802.11n: The road ahead. Computer, 42(7):13–15, July 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2009:NBSb

[2581] Linda Dailey Paulson. News briefs:

Software-based networking approach could help Internet research. Computer, 42(7):16–19, July 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chang:2009:EMR

[2582] Carl K. Chang. EIC’s message: Reversing the bipolar drift—introducing the IAB. Computer, 42(7):20–24, July 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Babar:2009:CPS

[2583] Muhammad Ali Babar and Ian Gorton. Computing practices: Software architecture review: The state of practice. Computer, 42(7):26–32, July 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bimber:2009:GEI

[2584] Oliver Bimber. Guest Editor’s introduction: Visual effects and beyond. Computer, 42(7):33–34, July 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roble:2009:CFD

[2585] Doug Roble and Nafees Bin Zafar. Cover feature: Don’t trust your eyes: Cutting-edge visual effects. Computer, 42(7):35–41, July 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Thomas:2009:VGB

[2586] Graham Thomas and Oliver Grau. Virtual graphics for broadcast production. Computer, 42(7):42–47, July 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yang:2009:VEC

[2587] Xubo Yang, Milo Yip, and Xiaoyue Xu. Visual effects in computer games. Computer, 42(7):48–56, July 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Trowbridge:2009:MBB

[2588] Scott Trowbridge and Christopher Stapleton. Melting the boundaries between fantasy and reality. Computer, 42(7):57–62, July 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Whitworth:2009:RFC

[2589] Brian Whitworth and Tong Liu. Research feature: Channel E-mail: A sociotechnical response to spam. Computer, 42(7):63–72, July 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:Bf

[2590] Anonymous. Bookshelf. Computer, 42(7):77, July 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CSCg

[2591] Bob Ward. Computer Society connection: Computer Society presents technical awards. Computer, 42(7): 81–84, July 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CCf

[2592] Bob Ward. Call and calendar. Computer, 42(7):85, July 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lane:2009:SBP

[2593] Gerald T. Lane. Standards: Best practices for standards communities. Computer, 42(7):86–87, July 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Subrahmanian:2009:WTM

[2594] V. S. Subrahmanian. Web technologies:

Mining online opinions. Computer, 42 (7):88–90, July 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shah:2009:GIA

[2595] Amip Shah, Tom Christian, Chandrakant Patel, Cullen Bash, and Ratnesh Sharma. Green IT: Assessing ICT’s environmental impact. Computer, 42(7):91–93, July 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schaffer:2009:ISP

[2596] Henry E. Schaffer, Samuel F. Averitt, Marc I. Hoit, Aaron Peeler, Eric D. Sills, and Mladen A. Vouk. IT systems perspectives: NCSU’s virtual computing lab: A cloud computing solution. Computer, 42(7):94–97, July 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2009:PAR

[2597] Neville Holmes. The profession: Agility and respect. Computer, 42(7):100, 98–99, July 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:EC

[2598] Anonymous. Elsewhere in the CS. Computer, 42(8):4–5, August 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:Lb

[2599] Anonymous. Letters. Computer, 42 (8):6–7, August 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2009:WF

[2600] David Alan Grier. Welcome to the family. Computer, 42(8):8–10, August 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:YAh

[2601] Anonymous. 32 and 16 years ago. Computer, 42(8):11–12, August 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaughan-Nichols:2009:GCM

[2602] Steven J. Vaughan-Nichols. Gameconsole makers battle over motionsensitive controllers. Computer, 42(8): 13–15, August 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2009:CSC

[2603] Linda Dailey Paulson. Caching scheme could boost developing world’s Internet access. Computer, 42(8):16–18, August 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rich:2009:BTB

[2604] Charles Rich. Building task-based user interfaces with ANSI/CEA-2018. Computer, 42(8):20–27, August 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ramakrishnan:2009:PDM

[2605] Naren Ramakrishnan. The pervasiveness of data mining and machine learning. Computer, 42(8):28–29, August 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Koren:2009:MFT

[2606] Yehuda Koren, Robert Bell, and Chris Volinsky. Matrix factorization techniques for recommender systems. Computer, 42(8):30–37, August 2009. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Krause:2009:OSW

[2607] Andreas Krause and Carlos Guestrin.

Optimizing sensing: From water to the Web. Computer, 42(8):38–45, August 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

King:2009:RSA

[2608] Ross D. King, Jem Rowland, Wayne Aubrey, Maria Liakata, Magdalena Markham, Larisa N. Soldatova, Ken E. Whelan, Amanda Clare, Mike Young, Andrew Sparkes, Stephen G. Oliver, and Pinar Pir. The robot scientist Adam. Computer, 42(8):46–54, August 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Xie:2009:DMS

[2609] Tao Xie, Suresh Thummalapenta, David Lo, and Chao Liu. Data mining for software engineering. Computer, 42 (8):55–62, August 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lach:2009:ASP

[2610] John Lach and Vinu Vijay Kumar. Application-specific product generics. Computer, 42(8):64–74, August 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:IPE

[2611] Anonymous. IEEE President-Elect candidates address Computer Society concerns. Computer, 42(8):79–84, August 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:BWW

[2612] Bob Ward. Benjamin Wah wins 2008 Kanai Award. Computer, 42(8): 85–89, August 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CCg

[2613] Bob Ward. Call and calendar. Computer, 42(8):90–92, August 2009. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:Bg

[2614] Anonymous. Bookshelf. Computer, 42(8):93, August 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kuhn:2009:CST

[2615] Rick Kuhn, Raghu Kacker, Yu Lei, and Justin Hunter. Combinatorial software testing. Computer, 42(8): 94–96, August 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ramakrishnan:2009:TPD

[2616] Naren Ramakrishnan, Debprakash Patnaik, and Vandana Sreedharan. Temporal process discovery in many guises. Computer, 42(8):97–101, August 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reisman:2009:ULO

[2617] Sorel Reisman. Using learning objects to affect educational outcomes. Computer, 42(8):102–104, August 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Viega:2009:CCC

[2618] John Viega. Cloud computing and the common man. Computer, 42(8): 106–108, August 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chu:2009:NII

[2619] Maurice Chu. Natural and implicit interaction systems. Computer, 42(8): 109–111, August 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sharkey:2009:RAL

[2620] Noel Sharkey. The robot arm of the law grows longer. Computer, 42(9):116, 113–115, September 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:CHSa

[2621] Anonymous. Computer highlights society magazines. Computer, 42(9): 4–5, September 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:LEa

[2622] Anonymous. Letters to the Editor. Computer, 42(9):6–7, September 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2009:DN

[2623] David Alan Grier. Data of the night. Computer, 42(9):8–11, September 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:YAi

[2624] Anonymous. 32 and 16 years ago. Computer, 42(9):12–13, September 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:2009:RFD

[2625] George Lawton. Routing faces dramatic changes. Computer, 42(9):15– 17, September 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2009:HAB

[2626] Linda Dailey Paulson. HP announces breakthrough in affordable, flexible displays. Computer, 42(9):18–21, September 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moncrieff:2009:DPP

[2627] Simon Moncrieff, Svetha Venkatesh, and Geoff A. W. West. Dynamic privacy in public surveillance. Computer, 42(9):22–28, September 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abrial:2009:FSY

[2628] Jean-Raymond Abrial. Faultless systems: Yes we can! Computer, 42(9): 30–36, September 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Black:2009:FVA

[2629] Sue Black, Paul P. Boca, Jonathan P.

Bowen, Jason Gorman, and Mike Hinchey. Formal versus agile: Survival of the fittest. Computer, 42(9): 37–45, September 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Meyer:2009:PTT

[2630] Bertrand Meyer, Arno Fiva, Ilinca Ciupa, Andreas Leitner, Yi Wei, and Emmanuel Stapf. Programs that test themselves. Computer, 42(9):46– 55, September 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cooke:2009:TPP

[2631] Daniel E. Cooke and J. Nelson Rushton. Taking Parnas’s principles to the next level: Declarative language design. Computer, 42(9):56–63, September 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nuseibeh:2009:SSR

[2632] Bashar Nuseibeh, Charles B. Haley, and Craig Foster. Securing the skies: In requirements we trust. Computer, 42(9):64–72, September 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:ICS

[2633] Bob Ward. IEEE Computer Society election. Computer, 42(9):73– 80, September 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CCh

[2634] Bob Ward. Call and calendar. Computer, 42(9):81–82, September 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:Bh

[2635] Anonymous. Bookshelf. Computer, 42 (9):83, September 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Helal:2009:UIS

[2636] Sumi Helal. University-industry synergy: An eclectic dialog for revised relationships. Computer, 42(9):88– 89, September 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pendyala:2009:WUC

[2637] Vishnu S. Pendyala and Simon S. Y.

Shim. The Web as the ubiquitous computer. Computer, 42(9):90– 92, September 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shukla:2009:MDE

[2638] Sandeep K. Shukla. Model-driven engineering and safety-critical embedded software. Computer, 42(9):93– 95, September 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vogel:2009:ITB

[2639] Doug Vogel, Michiel van Genuchten, and Carol Saunders. IS innovation: Thinking beyond the silo. Computer, 42(9):96–98, September 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kant:2009:TSP

[2640] Krishna Kant. Toward a science of power management. Computer, 42(9): 99–101, September 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2009:MSD

[2641] Neville Holmes. Music, sociality, and digital technology. Computer, 42(10):

104, 102–103, October 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:CHSb

[2642] Anonymous. Computer highlights society magazines. Computer, 42(10): 4–5, October 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:Lc

[2643] Anonymous. Letters. Computer, 42 (10):6–7, October 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2009:MI

[2644] David Alan Grier. The marketplace of ideas. Computer, 42(10):8–11, October 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:YAj

[2645] Anonymous. 32 and 16 years ago. Computer, 42(10):12–13, October 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geer:2009:FBN

[2646] David Geer. The OS faces a brave new world. Computer, 42(10):15– 17, October 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:RDA

[2647] Anonymous. Researchers develop application to identify ancient documents. Computer, 42(10):18–20, October 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Blair:2009:MRT

[2648] Gordon Blair, Nelly Bencomo, and Robert B. France. Models@ run.time. Computer, 42(10):22–27, October 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maoz:2009:UMB

[2649] Shahar Maoz. Using model-based traces as runtime models. Computer, 42(10):28–36, October 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cetina:2009:ACT

[2650] Carlos Cetina, Pau Giner, Joan Fons, and Vicente Pelechano. Autonomic computing through reuse of variability models at runtime: The case of smart homes. Computer, 42(10):37– 43, October 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Morin:2009:MRT

[2651] Brice Morin, Olivier Barais, Jean-Marc Jézéquel, Franck Fleurey, and Arnor Solberg. Models@ run.time to support dynamic adaptation. Computer, 42(10):44–51, October 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Georgas:2009:UAM

[2652] John C. Georgas, André van der Hoek, and Richard N. Taylor. Using architectural models to manage and visualize runtime adaptation. Computer, 42(10): 52–60, October 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gjerlufsen:2009:MMS

[2653] Tony Gjerlufsen, Mads Ingstrup, and Jesper Wolff Olsen. Mirrors of meaning: Supporting inspectable runtime models. Computer, 42(10):61–68, October 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoste:2009:MAC

[2654] Kenneth Hoste and Lieven Eeckhout. A methodology for analyzing commercial processor performance numbers. Computer, 42(10):70–76, October 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abbasi:2009:CTD

[2655] Ahmed Abbasi and Hsinchun Chen. A comparison of tools for detecting fake websites. Computer, 42(10):78– 86, October 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:Bi

[2656] Anonymous. Bookshelf. Computer, 42 (10):92, October 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CSA

[2657] Bob Ward. Computer Society awards recognize excellence. Computer, 42 (10):93–95, October 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:CCi

[2658] Bob Ward. Call and calendar. Computer, 42(10):96, October 2009. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anokwa:2009:OSD

[2659] Yaw Anokwa, Carl Hartung, Waylon Brunette, Gaetano Borriello, and Adam Lerer. Open source data collection in the developing world. Computer, 42(10):97–99, October 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Henderson:2009:UCT

[2660] Peter B. Henderson. Ubiquitous computational thinking. Computer, 42(10): 100–102, October 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Petrakos:2009:CTA

[2661] Nikolaos Petrakos, George W. Dinolt, James Bret Michael, and Pantelimon Stanica. Cube-type algebraic attacks on Wireless Encryption Protocols. Computer, 42(10):103–105, October 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Margaria:2009:CMD

[2662] Tiziana Margaria and Bernhard Steffen. Continuous model-driven engineering. Computer, 42(10):106–109, October 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fern:2009:PSC

[2663] Luis Fernández-Sanz. Personal skills for computing professionals. Computer, 42(11):112, 110–111, November 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:CHSc

[2664] Anonymous. Computer highlights society magazines. Computer, 42(11): 4–5, November 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:Ld

[2665] Anonymous. Letters. Computer, 42 (11):6–7, November 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2009:BA

[2666] David Alan Grier. Bad alignment. Computer, 42(11):8–11, November 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:YAk

[2667] Anonymous. 32 and 16 years ago. Computer, 42(11):12–13, November 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leavitt:2009:ATT

[2668] Neal Leavitt. Anonymization technology takes a high profile. Computer, 42 (11):15–18, November 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2009:STN

[2669] Linda Dailey Paulson. System takes new approach to speech search. Computer, 42(11):20–23, November 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoisie:2009:ESC

[2670] Adolfy Hoisie and Vladimir Getov. Extreme-scale computing. Computer, 42(11):24–26, November 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Torrellas:2009:AES

[2671] Josep Torrellas. Architectures for extreme-scale computing. Computer, 42(11):28–35, November 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ajima:2009:TMT

[2672] Yuichiro Ajima, Shinji Sumimoto, and Toshiyuki Shimizu. Tofu: A 6D mesh/ torus interconnect for exascale computers. Computer, 42(11):36–40, November 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barker:2009:UPM

[2673] Kevin J. Barker, Kei Davis, Adolfy Hoisie, Darren J. Kerbyson, Michael Lang, Scott Pakin, and José Carlos Sancho. Using performance modeling to design large-scale systems. Computer, 42(11):42–49, November 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wilde:2009:PSA

[2674] Michael Wilde, Ian Foster, Kamil Iskra, Pete Beckman, Zhao Zhang, Allan Espinosa, Mihael Hategan, Ben Clifford, and Ioan Raicu. Parallel scripting for applications at the petascale and beyond. Computer, 42(11):50– 60, November 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donofrio:2009:EEC

[2675] David Donofrio, Leonid Oliker, John Shalf, Michael F. Wehner, Chris Rowen, Jens Krueger, Shoaib Kamil, and Marghoob Mohiyuddin. Energyefficient computing for extreme-scale science. Computer, 42(11):62–71, November 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Singh:2009:CBS

[2676] Munindar P. Singh, Amit K. Chopra, and Nirmit Desai. Commitment-based service-oriented architecture. Computer, 42(11):72–79, November 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:SSS

[2677] Bob Ward. Society supports software engineering PE examination. Computer, 42(11):87–89, November 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:Bj

[2678] Anonymous. Bookshelf. Computer, 42 (11):92, November 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Finkelstein:2009:MI

[2679] Shel Finkelstein. Managing interns. Computer, 42(11):93–95, November 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carrero:2009:IWE

[2680] Miguel Carrero. Innovation for the Web 2.0 era. Computer, 42(11):96– 98, November 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cameron:2009:MIC

[2681] Kirk W. Cameron. My IT carbon footprint. Computer, 42(11):99– 101, November 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reif:2009:GTC

[2682] Harry L. Reif and Richard G. Mathieu. Global trends in computing accreditation. Computer, 42(11):102– 104, November 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:2009:TBC

[2683] Neville Holmes. Truth and breadth, clarity and depth in algebra. Computer, 42(12):108, 106–107, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:CHSd

[2684] Anonymous. Computer highlights society magazines. Computer, 42(12): 4–5, December 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:LEb

[2685] Anonymous. Letters to the Editor.

Computer, 42(12):6–8, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Land:2009:IOS

[2686] Susan K. (Kathy) Land. Implementing organizational strategies: Progress made in 2009. Computer, 42(12):9– 11, December 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2009:HAT

[2687] David Alan Grier. Honor among thieves. Computer, 42(12):12–15, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:YAl

[2688] Anonymous. 32 and 16 years ago. Computer, 42(12):16–17, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaughan-Nichols:2009:ARN

[2689] Steven J. Vaughan-Nichols. Augmented reality: No longer a novelty? Computer, 42(12):19–22, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2009:PLE

[2690] Linda Dailey Paulson. Programming language enables creative visualization. Computer, 42(12):23–25, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Feng:2009:TEM

[2691] Wu chun Feng and Pavan Balaji. Tools and environments for multicore and many-core architectures. Computer, 42 (12):26–27, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schneider:2009:PME

[2692] Scott Schneider, Jae-Seung Yeom, and Dimitrios S. Nikolopoulos. Programming multiprocessors with explicitly managed memory hierarchies. Computer, 42(12):28–34, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dave:2009:CSS

[2693] Chirag Dave, Hansang Bae, Seung-Jai Min, Seyong Lee, Rudolf Eigenmann, and Samuel Midkiff. Cetus: A source-to-source compiler infrastructure for multicores. Computer, 42(12): 36–42, December 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tallent:2009:IPB

[2694] Nathan R. Tallent and John M. Mellor-Crummey. Identifying performance bottlenecks in work-stealing computations. Computer, 42(12):44–50, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kelly:2009:ECB

[2695] Terence Kelly, Yin Wang, Stéphane Lafortune, and Scott Mahlke. Eliminating concurrency bugs with control engineering. Computer, 42(12): 52–60, December 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Toral:2009:EMP

[2696] Sergio L. Toral, Manuel Vargas, and Federico Barrero. Embedded multimedia processors for road-traffic parameter estimation. Computer, 42(12):

61–68, December 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:RVC

[2697] Bob Ward. Reisman voted 2010 Computer Society President-Elect. Computer, 42(12):69–70, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2009:MWC

[2698] Bob Ward. Miura wins Cray Award. Computer, 42(12):71–73, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:Bk

[2699] Anonymous. Bookshelf. Computer, 42 (12):82, December 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Apolloni:2009:CEC

[2700] Andrea Apolloni, V. S. Anil Kumar, Madhav V. Marathe, and Samarth Swarup. Computational epidemiology in a connected world. Computer, 42 (12):83–86, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Li:2009:BPF

[2701] Yang Li. Beyond pinch and flick: Enriching mobile gesture interaction. Computer, 42(12):87–89, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Beecroft:2009:PFN

[2702] Alexander J. Beecroft and James Bret Michael. Passive fingerprinting of network reconnaissance tools. Computer, 42(12):91–93, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hislop:2009:ITO

[2703] Gregory W. Hislop. The inevitability of teaching online. Computer, 42(12): 94–96, December 2009. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Broy:2009:THS

[2704] Manfred Broy, Mario Gleirscher, Stefano Merenda, Doris Wild, Peter Kluge, and Wolfgang Krenzer. Toward a holistic and standardized automotive architecture description. Computer, 42(12):98–101, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Santini:2009:SH

[2705] Simone Santini. Simplifying humans.

Computer, 42(12):104, 102–103, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2009:AI

[2706] Anonymous. Annual index. Computer, 42(12):??, December 2009. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).